A Complete Bibliography of Publications in Ecological Modelling (1990–1999)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

05 October 2023 Version 1.03

Title word cross-reference

$129.50 [Jør97f]. $169 [Ano99e]. $195.00/Dfl [Mau93]. 2 [Gal90]. $21.50 [Jør93a]. $225.50/DFl [Bos93b]. $236 [Jør94f]. $24 [Jør93c]. $27.50 [Ste93]. $29.95 [Ano96o, Jør95c]. 3 [EBA^{+}99, TZE96, WU96]. $32.50 [Jør94d]. $35.00 [Ano99d]. $40.50 [Ano99d]. $49.95 [Jør95e]. $52.95 [Jør92a]. $57.00 [Jør93a]. $60.00 [Ano99d]. $67.50 [Vas94]. $69.00 [Ano99d]. $69.96 [Nie95a]. $70 [Ano96q]. $74.00 [Leg92]. $75 [Jør93b]. $80.00 [Ano99b]. ^{1} [Gra98c, TSR97]. ^{14} [FK99]. ^{90} [Mon98a]. _{2} [CCKO94, CC96, CHM96, Chi98b, Com97, HH98, Jør95d, KBL96, KV90, LFM97, MJ99, Tay96]. _{3} [CHM96, HH98, NMS95]. _{4} [CCKO94, CHM96]. a [MCTB95, Zon98a]. i [MC93a]. L [WF99]. σ [Per98]. T [PD99].

-coordinate [Per98]. -D [Gal90, TZE96, WU96]. -response [Com97]. -state [MC93a]. -tilapia [FG94].

/Graham [Jør94e].

0 [Bos93a, Bos93b, HS94, Jør91a, Jør91b, Jør91c, Jør93b, Jør94a, Jør94b, Jør94e, Jør94c, Jør94f, Jør94g, Jør94d, Jør95a, Jør95b, Koh91, Leg91b, Mar99a, McD93, Mit94, Nie95b, Nie95c, Nih91, Sil95a, Ste93, Str91, Ush91, Zuc91]. 0-08-036198 [Jør95a]. 0-13-084989-8 [Vas94]. 0-13-161506-8 [Sil95a]. 0-13-928201-7 [Str92]. 0-231-06919-7 [Jør94d]. 0-231-10502-9 [Ano99d]. 0-231-10503-7 [Ano99d]. 0-306-43455-5 [Jør91b]. 0-306-43766-X [Ste93]. 0-309-03645-3 [Ush91]. 0-387-51603-4 [Ano91i]. 0-387-94309-9 [Wu95]. 0-412-04051-4 [Jør97e]. 0-412-29830-9 [Leg93]. 0-412-29840-6 [Leg93]. 0-415-06187-3 [Nih91]. 0-442-00805-8 [Nie95c]. 0-444-81478-7 [Nie95b]. 0-444-81572-4 [Ulg95]. 0-444-81578-3 [Ulg95]. 0-444-81904-5 [Ano96r]. 0-444-87348-1 [Ano96p]. 0-444-87349-X [Ano96p]. 0-444-87409-7 [Mau93]. 0-444-88030-5 [Str91]. 0-444-88469-6 [Bos93a]. 0-444-88532-3 [Bos93a]. 0-444-88599-4 [Bos93b]. 0-444-88609-5 [Koh91]. 0-47-1-95017 [Jør97h]. 0-471-183 [Ano99c]. 0-471-59515-2 [Ano96o]. 0-471-61315-0 [Leg91b]. 0-471-62559-0 [Zuc91]. 0-471-92193-9 [Jør91a]. 0-471-94263-4 [HS94]. 0-471-95463-2 [Ano96q]. 0-521-452323-6 [Jør95b]. 0-521-45967-2 [Jør95c]. 0-521-56342-9 [Ano99f]. 0-521-56523-5 [Ano99f]. 0-7923-1370-4 [Jør94g]. 0-7923-1386-0 [Jør95f]. 0-7923-1667-3 [Jør94e]. 0-7923-1805-6 [Jør95d]. 0-7923-1919-2 [Jør94a]. 0-7923-1985-0 [Mit94]. 0-7923-2149-9 [Jør94b]. 0-7923-2321-1 [Jør94f]. 0-7923-2377-7 [Jør94c]. 0-7923-2935-X [Jør96a]. 0-7923-2936-8 [Jør96a]. 0-7923-3767-0 [Jør97f]. 0-7923-4523-1 [Mar99a]. 0-7923-4524-X [Mar99a]. 0-7923-4872-9 [Ano99e]. 0-7923-4921-0 [Jør99b]. 0-8176-3421-5 [Jør91c]. 0-8493-4246-5 [Nie95a]. 0-88192-202-1 [Jør95g]. 0-dimensional [Log97]. 036198 [Jør95a]. 0722-0723 [Jør93b]. 0722-0723/ISBN [Jør93b]. 0723/ISBN [Jør93b].

1 [Jør94f, Kom94, McD93, Met98, Nih93, Svi91a]. 1-56881-033-4 [Kom97]. 1-85312-146-0 [McD93]. 1-85742-034-9 [Kom94]. 1/2in [Wu95]. 12-627860-1 [Jør95e]. 137 [BS95b, BS95c]. 139/US [Jør93b]. 159.95 [Str92]. 1992 [Jør94f]. 1993 [Ano95w]. 1995 [Jør99a].

2 [Bos91a, Jør94a]. 2-7380-0775-9 [Jør99a]. 2.0 [VZF96]. 22 [Jør97g]. 240.00 [Ulg95]. 2A [Ano96p]. 2B [Ano96p, Ano96p]. 2DLEAF [PA96]. 2in [Wu95]. 2nd [Hug91, Jør95c, Nie95c].

3 [All91, Ano91i, Bos91a, Bos93a, Hug91, Jør93c, Jør93b, Jør94e, Leg91a, Nih91, Nih93, Svi91b, Svi91a, Ush91]. 3-05-500475-2 [Bos91a]. 3-05-500761-1 [Svi91a]. 3-527-28705-1 [Jør97d]. 3-528-05419-0 [Kom97]. 3-540-18090-7 [All91]. 3-540-19465-7 [Svi91b]. 3-540-51129-6 [Leg91a]. 3-540-51603-4 [Ano91i]. 3-540-51603-4/0-387-51603-4 [Ano91i]. 3-540-52116-X [Hug91]. 3-540-58017-4 [Ano99b]. 3-540-94287-4 [Wu95]. 3-540-97488-1 [Nih93]. 3-923381-22-0 [Jør93b]. 3-923381-24-7 [Jør93c]. 3-population [TGM95]. 38.00 [Mar93]. 38.50/US [Jør93c]. 3rd [Jør99a].

4 [Ano91i, Bos93b, HS94, Jør94g]. 4/0 [Ano91i].

5 [Jør91b, Jør91c, Koh91, Str91]. 521 [Leg92]. 521-36138-9 [Leg92].

6 [Jør95b, Leg91a]. 65§ [Jør97h].

7 [All91, Jør93c, Jør94c, Jør94d, Nie95b, Svi91b].

8 [Bos91b, Nie95c, Sil95a]. 88lysis [Ano96w].

9 [Jør91a, Jør94b, Kom94, Leg92]. 90 [Bos91b]. 90-220-0937-8 [Bos91b]. 91 [CT96].

= [Ano99c].

abatement [Jør95f]. able [Jør92c]. Above [WMRK93, KW99, vWB99]. Above-ground [WMRK93]. absorption [AW97, Bar98]. abstract [Dur96]. abundance [Aok95a, BMHT99, BGTL99, Kaw97, LADL99, MDT93, WCDL96]. abundant [Osh96]. abuse [Kra95a]. Acacia [WJW99]. Académiai [Vas94]. Academic [Ano99e, Jør94a, Jør94b, Jør94f, Jør94g, Jør95d, Jør95e, Jør95f, Jør97c, Jør97f, Mar99a, Mit94]. Academics [Jør94c]. Academy [Ush91]. Acanthaster [RGBM90]. Acari [BM99]. acarine [BM99, CHC91]. access [GG92]. accompanied [Kom97]. according [FFS96, SIU97]. account [Jør92c]. Accounting [HF96]. accumulation [YH96]. Accuracy [YWMM98]. achieve [Tay96]. acid [WM95, dVKvdS95]. acidification [MI95, Sog93]. acidified [BL90]. acidifying [RH96b]. acknowledgement [Ano96-61, Ano97-58]. Acknowledgements [vG95]. acquisition [BSB99, Buw91, TB93]. across [LUC^{+}93a, WC96]. activated [VV91, VV92]. active [SN99b]. activity [FK99, HH99, XY95]. actual [RKM97, VdKK^{+}99]. Adaptation [GU97, HWB96, Hul96, Hul98]. adaptationist [Hul92]. Adapting [LFM97]. Adaptive [TRR96, Wil96a, Omo95, Omo97, Ste95a, WFB96]. Added [MA93b]. Addendum [HKV^{+}95]. Adelaide [GHH^{+}96]. adequacy [Luc93b]. adiabaticity [TBM96]. Adirondack [Pat97]. Adriatic [Tom96]. advantage [Ula95]. advection [MA93b, YWMM98]. advective [SFGH97]. advice [HS92a]. Aedes [RM95]. aerea [FSC^{+}97]. aerosols [Bot98]. affect [Fah98, Leg97b]. affected [BGF99, Jõg98, SvdG99, WK97b]. affinity [SSK^{+}91]. Africa [GMI95, SLP93, TB93, TBG93]. African [GYNE99, HSPM96, LLM99]. Africanized [MRSW93]. after [CD99, MJMN97, MMJ^{+}98]. against [CBP95, MRSK96, SY90]. Agardh [HCdVM98]. agardhii [MVS^{+}95]. Age [HF93, Kar93, AM98b, BLS98, BPTM^{+}97, Got98, HV99, HLF99, IK95, Jen96a, Jør95c, KJ94, PRK97, YAKS97]. age-0 [Got98]. Age-dependent [Kar93]. age-structured [AM98b, Jen96a, Jør95c, YAKS97]. aged [Bar91, GPT95, KBV99, Val99]. Agelaius [ÖM97b]. aggregate [WC96]. aggregates [DB98]. aggregating [Log97, YO95]. Aggregation [Aly91, Mar91, SCM93, YH93]. agricultural [Axe94, Gra95, LSB96, RSKS91, SS98, WHN96]. agro [SHJ95, TB93, TBG93]. agro-ecosystem [TB93, TBG93]. agro-ecosystems [SHJ95]. Agroecosystem [Ano95w, DSKM95, MKA^{+}95, SM95, Sep99, WTL91]. agroecosystems [GHQ99]. agrometeorological [LK91]. AGROSIM [SM95]. AGROSIM-Winter-wheat [SM95]. aid [AR96, Sil93a]. Ailuropoda [CALW99, iWSG96]. Aimophila [SGML91]. Aims [Wis92a]. Air [Mah98b, AL99, BJM92, CFS93, CTL98, FG92, KdH99, Kre91, SH92, Yin96]. air-temperature [CFS93]. air-tree-soil [CTL98]. Akademie [Bos91a, Svi91a]. Akademie-Verlag [Bos91a, Svi91a]. Al [WM95]. Al-solubility [WM95]. Alain [Jør97f]. Alaska [FFR99, SN99b, SR98]. alba [Smi97b]. Albany [Jør97e]. Alberta [HAW97]. albicaulis [KABT90]. alert [Jør91b]. ALEX [LBA^{+}95]. Alfred [Jør94e]. algae [ANFM99, PL95]. algal [AB97, Dua95, MCTB95, MCTB96, PASG97, RFHY97, SK91, SABG93, TKO91, US97]. algorithm [BMHC90, BC90, NZ92, SCM93, Yin96]. algorithmic [AH99, HCdVM98]. algorithms [Ces97a, PC99a, Wil96a]. alien [HR96]. alkaline [GYYL96]. Allee [McC97, WLW99]. Allen [Jør94d]. allocation [BSB99, KJ94, SDS98, SR92, TB93, TH90, Val99, VdKK^{+}99]. Alock [MT93]. along [MPNJ97, Por98, PRK97, Sal92a]. Alpine [TTT98, GF90, HCL96]. altaica [HAW97]. alternative [BJM92, Bar96b, BHL96, EFCM99, MP99, RC93, SGNS91, SG97, TDM96, Cas96]. altitudinal [Rya94]. ALWAYS [BEB99]. Amazon [Kan90]. Amazonian [WFJ96]. ambient [CCKO94, CC96, JTS92, WK97b]. America [Jes99, SFGF94]. American [MGCR98, Pat97]. americanus [Pat97]. Ammodytes [KKNY91]. ammonium [Par93a, RRJ90]. among [Bon93, DDA99, EAG97, HH98, Woo98]. amplitude [AK93, MT96]. ampullacuem [Mar91]. Amsterdam [Ano96p, Ano96r, Bos93a, Bos93b, Jør97g, Koh91, Mau93, Nie95b, Str91, Ulg95]. Anabaena [MDB98]. anadromous [Lee91]. anaerobic [FV97, MLB96, SEJ93]. Analyses [SR95b, BLRV98]. analysing [JTS92]. Analysis [Ano99b, Axe93, Bon93, Cas96, Col98, Com97, GRV93, MPNJ97, TB93, TBG93, Whi99, WJW99, ADF^{+}93, Ano91i, Aok92, BCH^{+}97, BS98a, BWW99, Böl90, BM96, CTDC94, CTL98, CHLT96, Chi98a, Cra98, Dup98, DGWK97, EH91, Els92, FP99, Foo99, GLD97, Gra98a, HRWC91, Han95, Hen95, HK91, HW98, Hug91, Jam93, Jõg98, Jør92a, KO95, Kan90, Kin95, Kle97, KN95, KI98, LCL93, Li95, LV99, LCH95, MDO99, MG98, McK96, Mil92, MHB^{+}96, MP99, Nih93, OH97, OA98, PTM91, PYE94, PH95, Per96, PM96, PK92, PC99b, PT99, Rec94, RL99, RPM^{+}98, RB98, STTR93, SH97, SN95, Sil99, SEJ93, SR95a, Smi97a, SF94, TD94, TM99, UOB94, VVDF95, VK99, VMSP97, VZF96, WG91b, YH99, ZMH96, vT95]. Analytic [SN99b].

Analytical [Han91, Kas93, Kim94, Coo95, Liu96, Mar99b, cZdZ96]. analyze [MRCL92]. analyzing [PP95, RS96]. Anantha [Jør94b]. Andhra [MHA95]. Andrew [CD99]. Angara [KCP^{+}98]. angular [LS98a]. Animal [BF98, AN90, Bla97, CF99, Dur96, WVB93, YH93]. animals [Omo95]. animat [KB97b]. Announcement [Ano91a, Ano94a, Ano95a, Ano96a, Ano96b, Ano96c, Ano96d, Ano93-30]. Announcements [Ano91b, Ano91c, Ano93a, Ano94b]. annual [Bon93, Bos91b, DSNN96, GdC99, LGM99, PR95a]. annuals [KJ94]. ant [KG95, SFGF94]. antartic [Böl90]. anthropogenic [PP95, TTT98]. ANU [Tay96]. ANU-BACE [Tay96]. any [Mar92]. Aphelinus [LRM93]. aphid [DBB94, KD96, KS91, RL99, SKŞ95, Wil91]. Aphididae [DBB94, SKŞ95]. aphids [KR92]. Aphis [WH91]. Apis [Mar98]. Apodemus [BS98b]. apple [ASTL91, DBB94]. Applicability [SM97]. Application [AN90, BS98a, BW96, CBB97, EBA^{+}99, FWCJ95, GKvdSdV95, HCdVM98, JWS95, Ker95, KSD^{+}99, Kra95a, Krä95b, KBSC98, LDB^{+}96, LB97b, MM96, MEBZ96, PSK98, PDAS95, RSKS91, Sch95, Sog94, Tor96, VMB95, VWN^{+}99, vD95, ALB91, BZ95, BL92, Ces97b, CP96, ET95b, Gam98, GET98, Gob93, GSW92, HTW96, Hul98, JA96, KEAP97, KCP^{+}98, LRK96, MPA97, NNR98, PA96, PYE94, PS90, Per98, SH97, SD96, TDM96, VP93, VF96a, Wan95, WW92, WGP96, WE96, WBH^{+}98, XHW^{+}99, ZH97, vOÅ95, EFCM99, MDT93, MFP99, MvdV95, NF96, OA97, PAGL96, Ano99d]. Applications [Foo99, Sep99, Ush91, AWS98, BCB97, CLCG99, DCP99, EMHB99, Jør95b, Jør97c, LK91, MLDH90, OYA95, RH96a, SG98, UDML97]. Applied [McD93, Cac90, CB97, DSKM95, Kin95, KH94b, SH91, TTT98, VVDF95, YWMM98, vHY95, Bos91a, Svi91b, Svi91a]. Applying [GTF99, Dro96]. apportionment [LVS^{+}99]. approach [Aok94, ACS96, Bat96b, Bat98, CALW99, CBC^{+}98, CBB97, CB97, CDDC96, DGS91, DFP^{+}93, DT97, Gao96, GET98, GFS99, GPG^{+}99, HCL94, Jør94b, Kai96, KD99, KMS99, Kle95, KS96, LK99, LW94, Łom99, LMG96, MHA95, MC93a, MvdZ96, MC93b, OS95, ÖÖ99, PT94, RSS91, RFHY97, RPM^{+}98, SD97, SADZ95, SvB96, Tis97, Tor96, VBL90, VS96b, Wae93, WFJ96, WS97, WH91, WHJ^{+}92, WL97, iWHS97, Ziv98]. Approaches [LH92, LLM96, Mül92, THR96, Zon98b]. approximate [Tro90]. approximately [Ano99c]. Approximative [AL91]. apricot [SC92]. April [Ano90l, Ano92z, Ano93z, Ano94w, Ano95-36, Ano96-52, Ano97-44, Ano97-53, Ano98-39, Ano99-36]. Aquaculture [CP96, Cac90, HM94, KUI94, LP91]. aquatic [AG93, CC96, Fer95, Gob93, Lho94, MM97, Mil92, PW93, SGNS91, Str92, TCC^{+}97]. arbitrary [VEA99]. arboreal [ML98]. ARCASIM [MRSK96]. Architectural [Chi98a]. architecture [EBA^{+}99, VWN^{+}99]. Ardea [Smi97b]. area [CO99, Chi98a, De 98, HL97, Mat99, RU97, TTT98, WV91]. area-based [WV91]. areal [Szy96]. areas [CPPB99]. Areawide [LAB^{+}98]. Arenosol [Sog94]. arid [Dun99, GHQ99, KR96, WL96, iWLSL96, ZH98]. armigera [MHA95]. arrangement [WVB93]. arrays [KO95]. arrow [iWSG96]. art [ÅB90, Jør92a, Jør95h, Jør99d, Mül97]. artedii [Jen97]. artefacts [Nal97]. arthropod [CW95]. arthropod-dominated [CW95]. article [Ano98j]. Artificial [LG99, RFHY97, ALGB99, BGF99, BGTL99, CPMC96, GTF99, MDB98, ÖÖ99, PT97, PL95, TS96, TAM^{+}99, WS97, vWB99]. ascendency [Chr94, Pat95, Sal92a]. ascendens [Hua98]. ascendens-intercrop [Hua98]. Ashe [FSG96]. Ashgate [Kom94]. Asia [MJY96]. Asimina [dLNT98]. aspects [Bat96a, Jør93d, Kle97, RH96a, vdHvEH96]. aspen [Hog99]. assemblage [JCC^{+}95]. assemblages [LG91]. assembly [HM97b]. assess [BGTL99]. assessed [HLF99]. Assessing [GRL92, Kit97, BLK^{+}99, BW96, CBC^{+}98, JMPT99, KB99, XJT99]. Assessment [Ano99e, Dye95, AM98a, BG92, BWW99, CS96, DR96, DPSB99, GGP97, JLS98, PJ96, Rec94, RS98a, SADZ95, TVS^{+}94, VJR^{+}98, VK97]. assessments [FP93, LL96a]. assimilation [HC97a]. assist [SMWF98]. associated [FFR99, SWK93]. assumptions [Mah98a, SG97, Sog93]. asychis [LRM93]. asymmetric [Ces97a, Ces97b, Dam99]. asymmetrical [Hul96]. asymmetry [Ume95]. asynchronous [GS98b]. Athens [DCCSL99]. Atlantic [BMHT99, BSN^{+}98, ALGB99, Got98]. Atmosphere [PSL^{+}93, vGDT95, AO97b, BJ95, MJMN97, OA97, SGS93, TvG95]. Atmosphere-terrestrial [PSL^{+}93]. atmospheric [AOD97, CB96, HB93, NNR98, TZE96, Tay96]. Atrazine [AK96, Per93]. attack [SR93]. attempt [HSU96]. attenuation [Liu96]. attractors [BH96, BH99]. attributes [Rob96a, Rob96b]. Attwater [PGS98]. Auerbach [Sil95a]. augmentation [MRSK96]. augmentative [LAB^{+}98]. August [Ano91-27, Ano92-27, Ano93y, Ano94v, Ano96-53, Ano97-43, Ano97-52, Ano98-31, Ano99-38, Ano99-49]. Australasian [BK98]. Australia [BBR99, Dun99, GGJ96, Hil96, HWB96, JTS92, Kin96, MDB98, PJ99, SFS97]. Australian [RH96b, VF96a]. Austrian [SM97]. Author [Ano90a, Ano90b, Ano90c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano94c, Ano94d, Ano94e, Ano94f, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano96k, Ano96l, Ano96m, Ano96n, Ano97g, Ano97h, Ano97f, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, PS91]. autocorrelation [Hen95]. automata [BBK98, Chi97, Dun99, GS98b, KT97, LSB99, RS98b, SB97, Spe97]. Automated [ZBDN96]. Automatic [MN93]. automaton [Rux96]. autonomous [Bye91]. availability [HBP93, WMRK93]. avenae [SKŞ95, SPH97]. avermectins [SMWF98]. avian [BEM97]. avoidance [Loe97b]. AWFUL [Pat98].

B [Bos93b, HS94, Jør93b, Jør96a, Jør97d, Koh91, Wu95]. BACE [Tay96]. Bachman [Ano99e]. back [Ano99f, Löf96]. bacteria [vdHvEH96]. bacterial [VV92]. Baikal [Sem99, SGS^{+}95, Sil99]. baited [OM97a]. Bakema [Bos91b]. Bakuzis [Lea97]. balance [BJ95, ET95a, ET95b, EMHB99, GBBV99, HSU96, KBL96, KR96, LK91, MK95, PL95, SM95, Sog93, TVS^{+}94, Ulr92, VMSP97]. balances [SHJ95]. Balancing [BKT95, CP92]. Balaton [SK91]. balsam [HLF99, RY91]. Baltic [HW98, Sem99, VK97]. bamboo [CALW99, KI93, KI98, iWSG96]. Bambusoideae [CALW99]. Banded [Dun99]. Bangladesh [Mah98b]. Barbara [Jør94e]. bare [Bre95]. bark [Bye96]. barley [Gra94, MMJ^{+}98]. barrier [SSY95]. basal [Chi98a]. based [AR96, AUS96, BM99, Bas98, BEB99, BS98b, BSB99, BKG97, CF99, CS96, CBC^{+}98, CR97a, Chi98a, Chr94, CR97b, DSNN96, DGS91, DL99, FSKC97, GS96, GFS99, GS98a, GGC98, GWAU99, Gri99a, GS98b, GYNE99, Håk96a, HRV^{+}96, HBM^{+}97, HAW97, HCL96, JRB97, Kei97, KD99, KS95, KKNY91, KH98, KTM95, KI98, Lav95, LC97a, LR97, LSB99, LA98, LL96a, Łom99, LS99, LGSF92, MRSW93, MG98, MCGLJ99, MMRR92, MHB^{+}96, MB96, MPA97, RLHD99, RF97, RRF91, RCD93, Sal92b, SV96, SR93, SSW91, SSM^{+}93, SMP96, Spe97, TTT98, Tre91, Uch99, Ulr92, VJR^{+}98, VdKK^{+}99, WV91, WL97, Zon98a, vdBDJ98, vdHvEH96, SS98]. Basel [Jør91c, Jør97d]. Basic [Ano92f, Leg91b, SSH98]. Basin [SN99b, KH94b, Lee91]. Basis [Jør94f, Alb92a, Bos92, Col97, Mor99, SBD^{+}95]. bass [DGS91, JRB97, Tre91, TCC^{+}97]. Bavaria [LH92]. Bay [BBR99, Boc99, CD99, Mar97a, And98, Got98, NF96]. Bayesian [AJK95, DCM92, GFS99, PT94, SBT97, TD94, VK97, VK99, VWN^{+}99]. be [AOE99, PW93]. bean [BM99, TBG93]. bear [Pat97]. beaver [Stu98]. Beck [HS94]. bed [HSN96, OKS99]. bee [Mar98]. BEEAMATE [DHRLE90]. beech [BS98b, Hof95, KI93, KI98, Wis92b]. beech-dwarf [KI98]. beef [BBH92]. bees [MRSW93]. beetle [Bye96, PL90]. beginning [RJN99]. behalf [Jør91a]. Behavior [BC90, Bak96, BS95a, BS95b, BS95c, RC93, Van94]. behavioral [MGCR98]. behaviors [MRSW93]. behaviour [DL99, EP93, HBM^{+}97, KW95, Mon98a, PRF97, Pol96, Rux96, SR93, SM95, Sou99]. being [LT96]. benchmarks [Hae91]. benefits [Lyo96b]. benthic [CFC91, Cla99, Dup98, PBO^{+}96, PL95]. Berge [Bos91b]. Berlin [All91, Ano91i, Bos91a, Hug91, Jør91c, Leg91a, Svi91b, Svi91a]. Berlin/New [Ano91i]. Bernard [Sil95a]. between [ALGB99, Ben99, CHM96, Chi98a, CB97, DCCSL99, Dup98, Gra94, Gra98d, HC97b, JCS^{+}97, KR92, LVS^{+}99, M.a93a, Mar99b, Mat99, NN95, PS93b, PR95c, RU97, SSW91, Ste95b, SGS93, SvB96, TBG93, Ume95, Wu90, iWSG96]. between-generation [Mat99]. between-shoot [Ste95b]. BGC/CHESS [KSD^{+}99]. Bias [WCS95, OM97a, WF99]. Biased [HM97a]. biennials [KJ94]. bifurcations [PRF97, RS93]. big [BSB97]. bimodally [Sel95]. bio [Sem99]. bio-optical [Sem99]. Bioaccumulation [NJ96, Gob93, LB97b]. biochemical [NMS95]. bioconversion [RPM^{+}98]. Biodiversity [Ano99e, Bat98, MPNJ97]. bioeconomic [CH97, HK91]. bioenergetic [Cac90]. bioenergetics [SG97]. biofilm [Hju96]. Biogeochemical [HW98, NF96, ADF^{+}93, AM98a, Log97, TDM96]. Biogeochemistry [Jør99a, KW95]. biogeography [BBN98]. bioindication [SBW^{+}99]. Biological [SGNS91, TKO91, BLS98, BK96, DR96, HvCC94, HSXP98, Il’96, Jør95b, KT94, KS96, LCMR96, LAB^{+}98, MG96b, RS98b, SG98, SS96, Spe94, TS93, Vin96, dGG98, Vas94]. biologists [BWW99]. biology [Mar97a, SOM97]. biomanipulation [JA96]. biomass [AKH99, BD99, BCB97, Chi98a, Cou91, Gae92, KEAP97, KKNY91, SFGH97, SM95, VS96c]. biomass-based [KKNY91]. Biomathematics [Leg91a, Svi91b]. biome [BBKV96]. biophysical [BSN^{+}98, BPN99, NMS95]. bioprocesses [Jør97d]. Biosphere [LWMT97, Ano93s, BGS97, FSKC97, Kir92, KV90, Kra96, SADZ95, SSH97, SV98, SSH98, LRK96, Svi97b]. biostabilisation [CSCD99]. biotic [Aok95a]. biphenyls [LB97b]. birch [Jõg98, LB97a]. bird [FFR99, HTW96, MDO99, Wol94]. birds [CMG98, HS99, NCNN97]. Birkhäuser [Jør91c]. birth [Par93b]. Biscayne [CD99]. bisexual [ACS96, Tuz91]. bison [PGD91]. Bite [CPM99]. bivalves [Dow97]. Biwa [Aok90]. black [HS91, MJGW94, Pat97, RM95, BSL97]. blackbird [ÖM97b]. Blackwell [Jør97b]. bleaching [WFB96]. Bleiswijkse [JA96]. blocks [OP96]. bloom [CV99, Lav95]. blooming [AB97, MVS^{+}95, PASG97]. blooms [MMJN97, RFHY97, Tom96]. blue [HS99, MMJ^{+}98]. bluegill [TCC^{+}97]. bluegill-largemouth [TCC^{+}97]. Board [Ano90f, Ano90g, Ano91o, Ano91p, Ano91q, Ano91r, Ano91s, Ano92n, Ano92o, Ano92p, Ano92q, Ano92r, Ano93m, Ano93n, Ano93o, Ano93p, Ano93q, Ano93r, Ano94l, Ano94m, Ano94n, Ano94o, Ano95p, Ano95q, Ano95r, Ano95s, Ano95t, Ano95u, Ano95v, Ano96-29, Ano96-30, Ano96-31, Ano96-32, Ano96-33, Ano96-34, Ano96-35, Ano96-36, Ano96-37, Ano96-38, Ano97s, Ano97t, Ano97u, Ano97v, Ano97w, Ano97x, Ano97y, Ano97z]. Boca [Nie95a]. BOD [PTM91, PT94, ST92, TD94]. bodies [PU97]. Bohemian [SM97]. boll [LCMR96, LAB^{+}98, MRSK96]. Bonn [Jør97e]. Book [Ano97i, Ano98a, Ano99a]. Boophilus [TMG96, TMGS97]. boosted [GCT96]. Bootstrap [PTM91]. boreal [KV97, KW99, KN95, Kol98, KL99, PA99, SVK93]. borer [ASTL91, HvCC94]. borne [Lin98]. borrowed [Ste93, Ste93]. Bossel [Kom97]. Boston [Jør91c, Jør97c, Jør97e, Jør97f, Mar99a]. Boston/Basel/Berlin [Jør91c]. Bostrichidae [ASTL91]. bottom [MS97]. bottomed [Tho97]. bottomland [KG96]. Bouma [Jør97h]. bound [Ano99c, Jør97b]. boundary [Bot98]. boundary-layer [Bot98]. bounding [Nor96]. bounds [MBF96]. BOWET [MWK95]. box [KdH99, NF96, Tay96]. box-diffusion [Tay96]. Braat [Nih91]. brackish [Szy96]. brant [MJGW94]. brassicae [Axe93]. Braunschweig [Ano95w]. Brazil [CV99]. Brazilian [Kan90]. Brebbia [Jør96a, McD93]. breeder [Smi97a]. breeding [Fah98, ÖM97b, SGML91]. Bremia [SvB93]. Brevoortia [Got98]. Brian [Jør91a, Jør93b, Jør95c]. Briand [Leg91a]. Brisbane [Jør97h]. British [Ano99c]. Brockenbleck [Jør93c]. Brockenblick [Jør93b]. brook [CR97b, PP95, Pow96]. brown [ALGB99, Jes98, VJR^{+}98]. brucellosis [PGD91]. bubbling [Van97]. bucket [ET95a, ET95b]. budget [BKT95, Chi98b, KN95, OKS99, SFGH97]. budgets [HKMRS95, Jør95b, SGML91]. budworm [RY91, SY90]. buffer [dLNT98]. build [Ano92g]. building [Kom97]. Bulgarian [KRD96]. Burns [Leg92]. bush [UDVESM^{+}95]. Byrnes [Spe94].

C [Ano96o, HCdVM98, Jør95a, Jør95e, Jør96a, Jør97g, Leg91a, Mau93, McD93, Nih91, Sil95a, CCKO94, CHM96, CPB97, FK99, HH98, MJMN97, MMJ^{+}98, NMS95, RRNM97]. C-mecoprop [FK99]. cabbage [SMP96]. Cabo [CV99]. Calanus [Mar97a]. calculated [Bou95]. Calculating [LLF97, BM97b, CP92, Liu96, UDVESM^{+}95]. calculation [KT94, Löf96, VKH^{+}94]. calculations [Löf96, Oel95, TVS^{+}94]. calculus [Wen92]. calibrated [Bot95]. Calibration [JH95, Bot95, KH94a, Kle97, SH97, VJR^{+}98, VP93]. calicivirus [BK98]. California [MU99, Spe94]. Call [BK97, Ano91a, Ano93-30]. calling [RC93]. Calluna [HB93]. Calvi [NF96]. Camargue [TAM^{+}99]. Cambic [Sog94]. Cambridge [Ano99f, Jør95b, Jør95c, Jør97d, Leg92]. campaign [FSC^{+}97]. campestris [HAW97]. Canada [HAW97, Hog99, SADZ95, Yin96]. Canadian [BL90]. Canal [Spe94]. Canberra [JTS92]. CANDY [FOS95]. cannot [GA96]. canonical [Mar97b, VS96b]. canopies [Ces97a, Ces97b, CA98, Liu97, Mar99b, Tro90, dCF99]. Canopy [Sto92, Ano94k, CLCG99, EBA^{+}99, HH98, Liu96, MPA97]. capacity [CALW99, EMHB99, KRD96]. capital [PC99b]. carabids [TIH98]. carbo [HE96]. carbohydrate [BSB99]. Carbon [VVK97, Ano96q, BKT95, BHBM96, Buw91, CB96, CFC91, Col97, CTKW98, CG93, Dup98, EGJ95, Ema96, FK99, GDS98, GSS94, Gra95, Hju96, KH96, Kir99, KN95, NF96, Par96, RGRF98, SGS93, SvB97, SvB98, TB93, TZE96, Tay96, Val99, WBH^{+}98, Zon98a, vDvB95, vWB99, vdPV99]. carbon-allocation [Val99]. carinata [MMJN97]. Carlo [Ann99, Jør97f, Ann97, DCM92, HC97a, NCNN97, Yoo99]. Carolina [TM99]. carp [KH94b, SGNS91, Spe94]. Carraro [Jør97f]. carrying [CALW99]. cascade [PG98]. case [BSN^{+}98, CCK92, FG92, FSG96, Gae92, HSB99, IPS99, LB97a, MDO99, PT95, RCS97, SH99, SS98, SRW97, TKG97, Ush91, VK97, VFC98, VVC99]. cassava [GYNE99]. Cassin [SGML91]. cassinii [SGML91]. catastrophe [iT96]. catch [McC95a, RS96]. catch-effort [RS96]. catchment [BHBM96, HL97, KSD^{+}99, SM95, SS98, Var93]. catchments [MWK95, Mon98b, PJ99, VS96a]. categorical [MDT93]. Catena [Jør93c, Jør93b]. catfish [MGNR91]. cations [WFJ96]. Catolaccus [LCMR96, MRSK96]. Catolaccus-boll [LCMR96]. cattle [BBH92, ST99, TMG96, TMGS97]. Caucasus [TC98]. Caulerpa [AH99, HCdVM98]. cause [RA96]. caused [AK93, CMG98, SNU96]. causing [SvB93]. CD [Jør99a]. cell [Zon98a]. cell-based [Zon98a]. cells [Bel98]. Cellular [BBK98, Dun99, GS98b, KT97, LSB99, Rux96, RS98b, SB97, Spe97]. censoring [PD99]. census [WQ97]. center [BL96]. central [BSE95, FSA99, SFS97, TC98, WFJ96]. century [BL93, GPO97]. CenW [Kir99]. Cercopithecus [SLP93, SL96]. CERES [AR96, CKK^{+}98, Mah98b]. CERES-Maize [CKK^{+}98]. CERES-Rice [Mah98b]. cerifera [SSY95]. Cervus [DDA99, SDBA98]. Ch [Ano96q]. Chaetomorpha [FSC^{+}97]. chain [CW95, HKV^{+}95, HC97a, Hea95, KSR97, KVAW91, PW97, RS93, VG98, WG91b]. chains [RPT93, vdB98]. Challenges [VB99]. Change [Jør99b, ADF^{+}93, BG93, BH94, BH99, CHM96, DSB^{+}99, DH95, FS97, HS94, IPS99, KV97, Lin98, LL96a, LCH95, MWY92, PSC93, SD97, VF96c, VW97, VdKK^{+}99, WHN96]. changes [AKH99, BBKV96, BU98, BEM97, CBC^{+}98, HC96, HC97c, Jør92c, Nie94, PA99, PR95b, SJ96, SNU96, SSY95, SY99, TS96, VS96a]. Changing [Ano99f, GF90, Il’96, SD96, SDF96, SVK93, VV92]. channel [MGNR91]. Chao [XJTL99]. Chaohu [Xu97]. Chaos [GAP93, All90, Pat97, RS93, RM93, Suá99, Van97]. Chaotic [WFK^{+}91, PG98]. chaparral [MWY92]. Chapman [Jør97e, Leg93, Mar93]. char [CR97b]. Characterisation [SEJ93]. characteristic [Zei91]. characteristics [CP92, FS97, GYYL96, GPW97, PT97, TBvH95]. Characterization [SRW97, PW91]. Characterizing [Pow96, Aok95b]. charge [Sog93]. Charlesworth [Jør95c]. Chemical [Jør95f, ADF^{+}93, BW96, DB98, KHL97, LG91, Sog93]. chemicals [FFS96, Gob93, HTW96, KdH99]. chemistry [MEBZ96, SA97]. chenopod [Dun99]. Chesapeake [And98, Got98]. CHESS [KSD^{+}99, PDAS95]. chew [CPM99]. Chichester [Jør97h, Zuc91]. chicken [PGS98]. Chile [Mar97a]. chilensis [Mar97a]. China [BKT95, HM94, HSXP98, JPA^{+}99, LWMT97]. chinensis [MT95, MT96]. Chinese [JAZ^{+}99, MT95, MT96, XJTL99]. chloride [MPA96]. chlorophyll [MCTB95, Rec93, Zon98a]. chlorophyll-nutrient [Rec93]. Choice [DFPG91, DG93, HRRP99]. cholinesterase [CMG98]. chopped [MMJ^{+}98]. chromium [AS94]. chronic [Gra98a]. chubs [Jen97]. Cichla [FG94]. Cincinnati [Jør97e]. cinnabarinus [SSW91]. circle [Van94]. circuit [Smi97b]. citrus [MPA96, YAKS97, FFR99]. City [Jør97e, DCCSL99]. Clark [Jør91c]. clarkii [ANM99]. class [MY97]. classical [GQB91, MDH96]. classification [BGS97, BGL99, HF93, LT96, Met98, Sil97]. Classified [SL91a, Mor99]. Classifying [LKS96]. clay [VGDvdZ97]. cleared [CBS97]. Cliffs [Sil95a]. Climate [Lin98, SvB97, Wae93, ADF^{+}93, BG93, BBKV96, BGS97, CHM96, ET95b, FS97, FSA99, GF90, HMC99, IPS99, Jes99, KV97, Kre91, LL96a, MU99, Oel95, PA99, PSC93, RS91, SD97, SVK93, SvB96, SvB98, VS96a]. climate-biosphere [BGS97]. climate-dependent [Kre91]. climate-induced [Jes99]. Climate-soil [Wae93]. climate-vegetation [BGS97, SvB98]. climates [Gra94]. climatic [AKH99, Dye95, GPO97, Hog99, MWY92, SVK93, UDVESM^{+}95, VF96a]. climatic/soil [GPO97]. clipping [Cou91]. Clonal [WK97a, SDS98]. closed [LLWP95]. closure [Kat90]. cloth [Ano99d]. cloud [Bot98]. cloud-topped [Bot98]. clouds [Bot98]. clover [Hil96]. CLUE [VF96c, VF96b]. CLUE-CR [VF96c]. cluster [Böl90]. clustering [Bar98, FFS96, MM96, dCF99]. CN/CHESS [PDAS95]. CO [Jør95d, LFM97, Jør94b, MR91b, CCKO94, CC96, CHM96, Chi98b, Com97, HH98, KV90, MJ99, Tay96, Jør94b]. co-existence [MR91b]. Co-operation [Jør94b]. Co-ordination [Jør94b]. coal [CSCD99]. coast [Por98]. Coastal [Ano91i, Ano96p, BGF99, CPH99, CCK92, CFC^{+}96, Jør97c, Leg97b, MR91a, MP99, ÖM97b, RCD98, TM99, vdBDJ98]. coccifera [RPT93]. Coccinella [SPH97]. coded [CS96]. coefficient [Håk97, IV91, Rec93]. coefficients [SABG93]. Coexistence [Ebe93, Hua98, Jen97, Kei97, Mar91, MMRR92, SG98, TSMC96]. Cohen [Leg91a]. Cole [Nih93]. Coleoptera [ASTL91, BBL97]. colepes [ŠŠZ90]. collaborative [SRV95]. Collecting [Ano92g]. collection [Kom97]. collective [Mon98a, Mon98b]. collembolan [LADL99]. collisions [Gav95]. colonies [Mar98]. colonization [SFGF94]. colony [DHRLE90, HTW96, KG95, PT99, Wol94]. colony-growth [KG95]. color [GTF99, Nih93]. coloradensis [MGD98]. colour [Ano99b]. Columbia [Ano99d, Jør93a, Jør94d, Lee91]. combinations [vdSdVK96]. Combined [KHL97, CBC^{+}98, SNU96, TKG97]. combining [PSN98]. comment [Day90, Lyo96a]. Comments [Ann99, Yoo99, Cra98]. commercial [CD99]. Commission [Ush91]. Committee [Ush91]. common [BK97]. commonly [Sog93]. communication [Hae91]. communities [CPMC96, Gal90, KR96, Rod99, Svi91a, TC98, Wu90, YKH98]. Community [HM97b, All91, BP93, CW95, CCR98, DH95, DH96, Dup98, Foo99, HLG96, MM97, RPT93, TS96, VV91, WFK^{+}91, WK97a, Ziv98, Svi91a, Leg91a]. community-level [Ziv98]. Comparative [Aok97, Aok94, Jør94a, Mah98b, VMSP97, Nih93]. compared [KW99, VVK97]. Comparing [MDO99, ML92b, SPE95, Gra98d]. Comparison [CHM96, CCR98, EBBW98, LSB99, RJN99, TDM96, vGDT95, BG93, Ben99, DSKM95, FRS^{+}95, JD99, KdH99, LCL93, LVS^{+}99, OR99, PW97, PT97, Per93, RRNM97, THR96, VW97, VGDvdZ97, Zon98b]. comparisons [CU93, PR95c, VB99]. compartment [Log97, PR95b]. compartmental [GMM91, Lee91]. compartments [Aok95a]. compensatory [Jen93]. compete [TSMC96]. competing [MR91b, TKO91]. competition [ALB92b, Bon93, BP93, BH94, CB97, Dam99, EAG97, Gra94, Gri99b, GQB91, Han95, Il’96, Jen96b, Kaw97, KTM95, Mau90, McC99, McC92, McC95b, MMRR92, Pie95, TKO91, UH99, Van96, Wil96b, YKH98, YKH99]. competitive [ALB92b, Cha96, MY97, RSC96, WLW99]. complement [CHC91]. completely [Hug91]. completeness [PC99a]. complex [BK96, JPS92, LG91, Liu93b, Mar97b, TDM96, Wen92, Sil95a]. complexity [Kei97]. complicated [CMH99]. component [LCMR96, Vin96]. components [HH98, SM95]. Composite [SL91b, ŠŠZ90]. composition [Jør92c, MRCL92, PS93b, SJ96]. composting [Kai96]. Comprehensive [CT96, HK91]. computation [PSK98]. Computational [McD93, Lav96, OS95, Jør96a]. computations [KW99]. Computer [dC95, Axe93, Axe94, Jør92a, LBA^{+}95, SM99, ST92, WJW99, Wil96a]. computers [Hae91, PFP^{+}95, VNLB96]. computing [SS96, Leg91b]. concentration [CB96, CPH99, DCCSL99, JTS92, LG91]. concentrations [Ahn99, Fom97, FK99, RMT92, ST92, vdSdVK96]. concept [SS96, SSH98]. Concepts [AO97b, Ush91, Ano92f, GSW92, Jør97c]. Conceptual [MvdZ96, HSU96, HS96, PG97, PJ99, VF96b, WL97]. concerning [Aok95a, Han97a]. Concise [Jør95a]. conclusions [Ano92m, MWJ92]. concurrent [Lav95]. condensation [Mar97b]. Conditioned [RM93]. conditions [Bar96b, BLRV98, Kat90, KV97, MAY^{+}98, MvdV95, RKM97, SNU96, SK91, SVK93, VEA99]. conductance [Fri95, HC97c]. Conference [Ano99e, Jør94f, Jør99a, Jør99b]. confidence [Håk96a]. configuration [MC93a, Mar99b]. configurations [Was95]. confinement [Nal97]. confusum [OM90, Tuz91]. conical [Tho97]. coniferous [KW99, vDvB95, vWB99]. conifers [Ste95b]. conjunction [SVK93]. connectance [ABK^{+}92]. connectionist [AH99]. connectivity [HF96, SL96, WVB93]. consensus [vdBDJ98]. consensus-based [vdBDJ98]. consequences [Gra98a, HBM^{+}97, KR92, PACS96, RS91, Sil99, Koh91]. consequent [Was95]. Conservation [GGJ96, ASP93, Bat98, Dub97a, Jet92, Liu93b, SDS96, PSJ97, Ano96o]. consider [dLNT98]. considerations [dCF99]. constant [Hut91]. Constraints [SV98, GBBV99, GT97, Jør92e]. constructing [PR95a]. construction [DGWK97, KD99]. consumption [CB96, CPH99, VV92, Woo98]. contagion [RS96]. contagious [LA96]. containing [SvB98]. contaminant [ML97]. contaminated [LT96, UMJ^{+}97, Jør99a]. contamination [MvdZ96, VBL90]. Contemporary [AO97a]. content [FK99, Håk95a]. Contents [Ano90e, Ano90d, Ano91j, Ano91k, Ano91m, Ano91l, Ano91n, Ano92h, Ano92i, Ano92l, Ano92k, Ano92j, Ano93k, Ano93l, Ano93h, Ano93i, Ano93j, Ano94g, Ano94h, Ano94i, Ano94j, Ano95i, Ano95j, Ano95k, Ano95l, Ano95m, Ano95n, Ano95o, Ano96w, Ano96s, Ano96t, Ano96u, Ano96v, Ano96x, Ano96y, Ano96z, Ano96-27, Ano96-28, Ano97k, Ano97l, Ano97o, Ano97p, Ano97m, Ano97n, Ano97j, Ano97-67, Ano97-68]. context [HS92c]. contexts [Hul96]. Continental [SADZ95, BHWP91]. continuation [MD97]. continuous [BGS97, Jen94, NN95, Par96]. continuously [Hut91]. contrasting [Gra98b, Kin95, KR96]. contributing [All90]. contribution [SS98]. control [ALS97, DCP99, DT97, FG92, FG94, GA96, HvCC94, Hon92, KH96, LCMR96, LAB^{+}98, MG96b, MFP99, Mos92, Par93b, PC99b, RPJK94, Rip95, SGNS91, SG98, SM99, Sep99, Smi97a, SMY^{+}97, Spe94, TGM^{+}99, TRR96]. controlled [GS97, MT93]. Controlling [BHL96, HLF99, Kin96, THR96]. controls [HW98, KW99, WM95]. controversy [HSPM96]. conventional [RSKS91, SKW91]. convergence [BFK96]. Conversion [VF96b]. coordinate [Per98]. Coordinator [Jør93c]. Cootes [TPS98]. copepods [GRBF97, Omo97]. copper [LL96c, RMT92]. coral [AK93, McC92, McC95a, RGBM90, WFB96]. core [SH91]. Coregonus [Jen97]. cormorants [HE96]. corn [EH91, RSKS91]. correction [EFCM99]. correlated [BPTM^{+}97, Els92]. Correlation [NN95, HF96, KW99, LCL93]. Correspondence [Ano97q, Ano98b, Ano98c, Ano98d, Ano98e]. correspondences [PT99]. corridors [TIH98]. corroboration [CR97b]. Corvus [PT99]. Cost [WGP96, TBM96]. Costa [BSN^{+}98, VF96c, VZF96]. Costanza [Jør93a]. costs [Rux95]. cotton [LAB^{+}98, LC97b, MLDH90, NHWS91, SCEZ^{+}94]. could [Gri99a]. Council [Ush91]. count [WQ97]. Country [DSB^{+}99]. counts [WCDL96]. coupled [HRV^{+}96, HMW96, PG98, PSL^{+}93, Van94]. Coupling [NMS95, WU96, YH99]. courtroom [Swa96]. cover [All91, Ano99b, BEM97, Kom94, Leg93, PJL^{+}96, Leg93, Vas94]. covered [FS97]. cowpea [TB93, TBG93]. CR [VF96c]. Craig [Jør97c]. crane [LWMT97]. crappie [AM98b]. crayfish [ANM99, AFM99, ANFM99, NAFM99]. CRC [Nie95a]. created [MM97]. creating [CF99]. Creation [Ano94k]. Creek [PJL^{+}96, Var93]. Cremlingen [Jør93b]. Cremlinger [Jør93c]. Crete [VPFM97]. crickets [RC93]. crisis [Jør91b]. CRISP [ANM99, AFM99, ANFM99, NAFM99]. CRISP-crayfish [NAFM99]. criteria [AM98a, MTCL96, NJ96]. criterion [SR92]. critical [JMN98, Kat90, LL96a, Sch96a, SSH98, dVPO^{+}95, vT95]. criticality [RAM99]. crop [AR97a, Ben99, BSB99, BILL98, CKK^{+}98, GGZ99, HCL94, Jol96, KEAP97, LC97b, LK91, MHJ96, RKM97, SDS96, SHJ95, TAM^{+}99, TKG97, VVDF95, Whi95, ZH97, ZH98]. cropping [JO95]. crops [CB96, JO95, SR92, Bos91b]. cross [FZ99, RSC96]. cross-diffusivity [RSC96]. cross-validated [FZ99]. crowded [WWJ99]. crown [LWMT97, RU97, Ume95]. crowns [Ces97a, Ces97b, KS95, WW92]. Crustacea [MMJN97]. Cryptomeria [Chi98b]. Cs [BS95b, BS95c]. Cs-137 [BS95b, BS95c]. cubic [MY97]. cultivation [GMI95]. cultural [GT97]. culture [MLGV99, SR95b, Szy96, TN98]. cultured [Dow97]. Cummins [Jør97g]. cumulative [Pow96]. cuniculi [Sey92]. current [Kra96, MAY^{+}98]. curtain [PASG97]. curve [SEJ93]. Cushing [Jør97g]. cyanobacteria [MDB98]. cyanobacterial [Bel98]. Cyathura [MMJN97]. cycle [Bár96a, BLS98, GSS94, GdC99, KGD97, NF96, OM97a, Omo95, Rip95, SvB97, SvB98, Tay96, Wis92b, Ano96q]. cycles [ADF^{+}93, ALB91, DPSB99, GU97, Kir99, Log97, PHB90, RM92]. cyclical [ACS96, Coq95]. cycling [Cha95, CFC^{+}96, Com97, Ema96, GS98a, Leg93, SA97, dVKvdS95, vDvB95]. cyst [And98]. Czech [KSD^{+}99].

D [Bos91b, Jør95f, Kom97, Leg93, EBA^{+}99, Gal90, TZE96, WU96]. DaD [Log97]. DaD-}stability [Log97]. Daily [CLCG99, Ano96w, Ant98, BILL98, Fri98, Han91, Liu96, MT93, PJ99, Szy96, Tro90]. DAISY [MJMN97, MMJ^{+}98, SHJ95]. Dale [Sey92]. dam [BM96, GG99, SMY^{+}97]. damage [BHL96, TAM^{+}99]. dammed [Sal97]. dampening [BMHT99]. Danish [Nie94]. Danube [KH94b]. Daphnia [HE96, KHL97, MB96]. Dasyneura [Axe93]. Data [vT95, AO97a, AOE99, Ano92g, Ano96w, Ant98, BKT95, CHM96, Cra98, DFP^{+}93, DSKM95, EB93, FP93, Foo99, GFS99, GPO97, HM97a, KBSC98, LRK96, MCGLJ99, ML98, MKA^{+}95, MM96, MDT93, Met98, MWK95, MFP99, Mor99, NNR98, PTM91, PC99a, RS96, STTR93, SPE95, SM95, SFS97, TBvH95, TD94, WL96, DGWK97]. database [KLFH96, MJ99]. dataset [TBvH95]. date [RJN99]. David [Ano99d, Jør95e, Jør95g]. day [FRS^{+}95, SD97]. daylength [FRS^{+}95]. DC [Ush91]. dealing [MM96]. DeAngelis [Leg93]. death [Par93b]. decade [BL93]. December [Ano91u, Ano91y, Ano93-28, Ano95-38, Ano96-54, Ano97-45, Ano97-40, Ano97-50, Ano97-42, Ano98-33, Ano99-41, Ano99-52]. deciduous [Bos93b, VB99]. Decision [HS92a, KL99, AR96, GA97a, HS92c, McK96, Rec94, SH92, VK99, WGP96, WHJ^{+}92]. decision-theoretic [McK96]. decisions [HRRP99, VKT94]. decline [HLF99]. Decomposing [PR95a, VVK97]. decomposition [JHM97, Kin95, VNLB96, YH96]. decoupled [Bye91]. decoupling [Mos92]. decrease [CB96]. deduced [Van94]. deep [GRV93]. deer [BW94, DDA99, RPH99, SDBA98, XHW^{+}99]. DeerMOM [XHW^{+}99]. define [Håk96a]. Defining [Hil96]. definition [Bas98, Cao95]. defoliation [Hog99, SCH92]. Deforestation [Kan90, Kan91, RS91]. degradation [HSPM96]. degraded [SDS96]. degree [SD97]. degress [Mar91]. del [Mar97a]. DELAQUA [RPJK94]. delay [SCEZ^{+}94]. Delia [SMP96]. demand [GG92, SG98]. demographic [Akç91, GRL92, SLP93, TB93]. demography [McC96]. Dendroctonus [BBL97]. Dendrolimus [Ned99]. denitrification [HSN96]. Denmark [BCB97, FKNM^{+}97, SS98]. Density [Jen96a, KJ94, Bar98, DGS91, DH96, GRL92, Gra98a, JCC^{+}95, Jen95, KL98a, LBV96, Mat99, Mau97, ML98, Par93b, Sen90, SN95, Smi97a, SLP93, iWHS97]. Density-dependent [Jen96a, KJ94, DGS91, Jen95, LBV96, Sen90, SN95]. depend [HS97b]. dependence [DH96, Gra98a, KL98a, LLF97, Mau97, ND97, Smi97a]. dependent [BBKV96, DGS91, Jen95, Jen96a, Kar93, KJ94, Kre91, LBV96, Sen90, SN95]. depletion [Dub97a, Var93]. deposit [ML97]. deposition [AOD97, AL99, HB93, SA97, TBvH95, dVKvdS95]. depth [FK99]. Derivation [KRD96, LLF97, WF99]. derivative [Sil93a]. derived [CBP95]. describe [JHM97]. Describing [UB97, Bac93, BHBM96, CWPD95, SSW91, SMDSZSXF93, SPH97, vdHvEH96]. Description [FKNM^{+}97, Bár96a, Bos92, CR97b, CPPB99, HS97a, JZV92, VJR^{+}98]. descriptors [Böl90, DCCSL99]. desert [Day90, HC96, WJW99]. desertification [VPFM97]. desertification-threatened [VPFM97]. Design [Jør92a, MJ99, Lav95, LC97b, MCGLJ99, MM97, RB98, ŠŠZ90, XHW^{+}99]. Designing [AR97a]. despite [Der98]. destocking [HSPM96]. detail [Sto92]. detect [HC97c, Rod99]. detecting [MEBZ96, Jør97c]. Detection [Day90, BC90, Day91, TD94]. determination [Omo97, SABG93]. determine [LSB96]. Determining [HvCC94, Håk99, Kat90]. Deterministic [SY99, Nor96, VU93a, VVDF95]. deterministic/stochastic [VU93a]. detritivore [VVK97]. Detroit [Jør97e]. developed [HM97a]. Developing [SH99, Bat96a, MCU98]. Development [Ano96r, DCM92, FDC^{+}96, Jør92c, JLS98, KMWB98, PJ96, Str91, AH99, Bos92, FK99, Gav95, HCdVM98, Hof95, How96a, How96b, Jør94b, Jør97a, Jør99d, Lyo96a, MD97, PA96, STTR93, SMP96, SDBA98, WMRK93, Wil96b, Xu97, Nih91, Jør94a]. Developments [Koh91, Ulg95]. Dfl [Ano96p, Ano96r, Bos93a, Koh91, Nie95b, Str91, Ulg95]. Dfl.100/US$57 [Bos91b]. Dicks [Jør91a]. Diego [Jør95e, Jør97c]. diet [BK91a]. dietary [CMG98, MGNR91]. difference [Rom96]. differences [Rom96]. different [Ben99, BWW99, BLRV98, DA95, DSKM95, FSA99, Gon97, Gra94, Håk97, Jen93, KEAP97, KH94a, LLWP95, Nie97, RJN99, RÅ99, Sto92, VP93, VF96a, Wil91]. differential [HC97b, WVCS94]. differentiated [PMV96]. differently [WBH^{+}98]. differing [TC98]. difficulties [MCTB95]. diffusion [Bla97, Coo95, CTKW98, Cri94a, HKV^{+}95, JHM95, Kal99, KVAW91, PL90, Sen90, SN95, Tay96, YWMM98]. diffusive [Jes99, KH94b]. diffusivity [RSC96]. diflubenzuron [HL96]. digester [SEJ93]. digestion [MLB96]. digital [AL99, LAB95]. dimension [Pay99]. dimensional [BCH^{+}97, BS95a, BS95b, BS95c, CWC95, Coo95, CTKW98, GS96, HBM^{+}97, Log97, Mar92, MA96b, Per98, SvB97, SvB98, WCC95, Wil96b, WTL91]. dimensions [KS96]. dinoflagellate [And98]. dioxide [CB96, CTKW98, Jør95f, Kre91, WBH^{+}98]. Dipodomys [Mau90]. Diptera [CS96]. direct [BLS98]. direction [AL99]. discharge [Håk96b]. discharged [PT94]. discontinuous [Ces97a, Ces97b]. Discount [How96a, How96b, Lyo96a, Lyo96b]. Discounting [Liu93a]. discovery [TDK98]. Discrete [dVZ93, Jen94, NN95, Par96, SSS^{+}91, SB97]. Discrete-event [dVZ93]. discriminant [MDO99]. discrimination [ALGB99, BJM92]. Discussion [Ano92m]. disease [BK98, MHJ96, Sch96b]. diseases [Hug91]. diskette [Kom97, Leg91b, Jør97d]. diskettes [Wu95]. Dislodgement [NHWS91]. dispar [SC96]. Dispersal [MA96a, DFHR96, GAP93, HMA90, HP94, JWEB92, JB97, LGM99, Mal96, McC99, YWMM98]. dispersion [IV91, LW96, Per98, SH92]. display [Ume95]. Disposal [Jør94e, SADZ95]. Dissipation [SJP99]. dissipative [CTDC94]. Dissolved [SF94, VU93a, BHBM96, CFC91, CP96, FS97, GH94, LG91, Mon98b, Tun90, VU93b]. distance [McC99, RDFZ96]. distributed [BKT95, Cla99, KMWB98, PM96, Sel95, SCEZ^{+}94, Sil93a, TDM96]. distribution [AL94, Bac93, BPTM^{+}97, CA98, GG99, GRBF97, GA97b, HAW97, HS97b, IK95, Ken92, KV90, LS98a, LWMT97, MR97, Oga96, PAGL96, SIU97, TC98, WHA98, WWJ99, Yam97]. distributional [PC99a, PBO^{+}96]. distributions [Gae92, HL97, HCL96, IPS99, JCC^{+}95, MDO99, TKO91, WC96]. District [GGJ96]. Disturbance [Van96, And96, And95, GS97, LA96, LTMP97, MJGW94, PA99, SN99a]. disturbance-controlled [GS97]. disturbances [BPTM^{+}97, IK95, dVZ93]. disturbed [Ema96]. ditch [PL95]. diurnal [CFS93, Liu96, NCNN97]. diverse [VB99]. Diversity [Aok95a, De 98, SSH97, BM97b, GPT95, Han98, JSC91, JD99, Kan90, Kan91, Kol98, LADL99, LBV96, LW94, MA96a, Mal96, Ziv98]. DM [All91, Ano99b, Bos91a, Jør93c, Jør93b, Jør97d, Nih93]. DM148.00 [Leg91a]. DM160.00 [Hug91]. DM42.00 [Svi91a]. DM88.00 [Ano91i]. DM98 [Ano97-68, Kom97]. Dmitrii [Nie95a]. do [BL93, FSA99, PT94]. documentation [BK97, HGB98, MKA^{+}95]. does [Fah98, HS97b]. doing [VK99]. domain [VNLB96]. Domestic [Ano99d]. dominated [CW95, PU97]. Dordrecht [Jør94a, Jør94b, Jør94e, Jør94c, Jør94f, Jør94g, Jør95d, Jør95f, Jør96a, Jør97f, Jør99b, Mar99a, Mit94]. doubling [CB96, RS93]. Douglas [CFS93]. Douglas-fir [CFS93]. down [TGM^{+}99]. downscaling [Oel95]. downstream [Cri94a]. downy [SvB93]. dreissenid [PACS96]. driven [DFHR96, DH95]. Droop [LC97a]. drought [GBBV99]. dry [EFGE98, HB93, RGRF98, WK97a, Ano96p]. dry-grassland [WK97a]. drying [KS96]. due [NHWS91]. dump [CSCD99]. Duncan [Jør94a]. dune [dC95]. dung [SMWF98]. Dunn [Jør97d]. duration [CR97b, GBBV99, Håk96c]. during [BU98, Col97, FSC^{+}97, SGML91, Whi95]. Durraiappah [Jør94b]. Dwarf [Cou91, KI93, KI98]. dwelling [KD96]. Dynamic [EP93, Gal90, Gao96, PW97, Par93a, SBT97, AJK95, AL91, Bac93, Bak96, BG92, BGF99, BJ97, CWPD95, CBB97, GYYL96, GMI95, Gra98d, HW98, Jør97a, Jør99d, Kir92, LCMR96, LVS^{+}99, Mar97b, MVS^{+}95, Nie92, Nie95d, NZ92, PR95a, Pat97, PC99b, RA96, SH91, SGS93, Svi99, YS98, cZdZ96, dZcZ96, Jør95b, Wu95]. Dynamical [Ter93, KK97, Nie94]. Dynamics [CW95, DJ98, Gra95, Jen93, Jen94, NJ96, PL90, AUS96, Aly91, ANM99, ANFM99, AK93, Axe93, BBK98, BCB97, BS98b, BMV99, BP93, BK91b, BC96, Bou95, Cac90, CALW99, CFC91, CK97, Chi97, CHC91, CC94, CL90, CU93, Com97, Cou91, CG93, CDDC96, DGS91, DLN^{+}95, DH99, DNR95, DF97, EGJ95, EH91, ESD98, FSC^{+}97, FV97, FOS95, FSKC97, GLD97, GW98, GH94, GA97b, Gor98, GSS94, GT97, HKV^{+}95, HMA90, Han95, Hea95, HPB93, HBP93, IK95, Jac96, Jar95, Jes98, Jõg98, JHM95, JB97, JN94, Kar93, KEAP97, KI93, KABT90, KV97, Ker95, KD96, KH94b, KS91, KCP^{+}98, KI98, KVAW91, LSC95, Lea97, LD97, LSB99, LRS^{+}93, LA96, LVH96, LA98, LJ94, MHJ96, Mal95, MC93a, MGCR98, Mar97a, MGD98, MMJN97, MRL94, MC93b]. dynamics [McC96, McC95b, MLDH90, MLGV99, MB96, NN95, Ned99, NAFM99, OS95, Osh96, PS93a, PHB90, PGD91, Pow96, PMP97, RPT93, RL99, Rob96b, Sal97, SML98, SBW^{+}99, SC96, Shu98, SPH97, SBD^{+}97, SFS97, SDBA98, Stu98, TMG96, TMGS97, TSR97, Tuz91, UB97, VU93a, VU93b, Van96, WL96, WLW99, WJW99, WVCS94, Wil99, Wil91, WFK^{+}91, WK97a, WH91, WV91, WVB93, iWHS97, WWJ99, YKH98, YKH99, YH99, dC95, vdPV99, Jør97e, Jør97d, Leg93]. Dysaphis [DBB94]. early [FSC^{+}97]. earth [Ano96q, Ano96r, Ste93, Ano99b]. eastern [Hil96, IPS99, VB99]. eco [BM94, Ano96r]. eco-engineering [Ano96r]. eco-ethological [BM94]. ecodistricts [BW96]. ecodynamical [Sem99]. ecohydrological [Dro96]. ECOLECON [Liu93b]. Ecological [All91, Ano93h, Ano93i, Gra98d, Hug91, Jør90, Jør93a, Jør97i, Jør99c, Kol98, Leg97a, Sil97, Ush91, XJT99, ZMH96, AL91, Ano90e, Ano91k, Ano91m, Ano91l, Ano91n, Ano92k, Ano93j, Ano94h, Ano94i, Ano95i, Aok93, BMHC90, BLK^{+}99, Bir94, BC96, BC90, BH96, Chi98b, CG98, CDDC96, DFPG91, Der98, DG93, FP98, Fer95, FSG96, Gam98, GT97, GWAU99, GS98b, Gue91, HL96, JZV92, Jør92b, Jør92e, Jør93d, JNM95, Jør95h, Jør97c, Jør97b, Jør99d, Kaw97, KH94a, KLFH96, Kra93, LG99, LW96, LCL93, LL96b, LS98b, LS99, Luc93b, LCH95, MPP98, ML92a, Mül98, PA96, PWU93, PG97, PG99b, PS93a, PH95, Pat95, Pay99, Pol96, Pow93, RB96, RO97, Rod99, Ryk96, Sal92b, SNUM91, SN99a, Sol99, SVK93, iT96, UMJ^{+}97, Ulg95]. ecological [UDML97, Van97, VUDM^{+}99, Web95, Wen92, Wis92a, Woo98, Xu97, XJTL99, Zei91, Liu93b, Jør97i, Ano92l, Ano93k, Ano93l, Ano94j, Ano97k, Ano97l, Ano90d, Ano91j, Ano92h, Ano92i, Ano92j, Ano94g, Ano99c, Ano99d, Jør91a, Jør93a, Ush91, Zuc91]. ecological-economic [Woo98, Liu93b]. ecologically [JCC^{+}95]. Ecology [Gra98c, ÅB90, AN90, All91, Bre92, Fle91, GSW92, Gri99a, HGB98, Jør91c, Jør92d, JPS92, Jør94c, Jør95e, Jør97b, Jør99c, Ken95, Leg91b, LDB^{+}96, Lho94, Łom99, Mar92, PYE94, RF97, Sil95a, SBT97, ŠŠZ90, Suá99, SL91a, Pat98, Ano91i, Bos91a, Jør94d, Nie95a, Svi91b]. ECOMONT [CBC^{+}98]. econetworks [Han96, Han97b]. economic [BLRV98, BSN^{+}98, BPN99, CG98, Gra98d, Jør94b, Woo98, Liu93b]. economics [Jør93a, Jør94c]. economists [Lyo96b]. ECOPATH [CP92]. ecophysiological [Bos91b]. ecoscreening [Sol99]. ecosys [Gra98b]. Ecosystem [AYO99, Chr95, HM94, Jør92b, JMN98, Mar99a, Was95, AM98a, Axe94, BGF99, BLK^{+}99, BK92b, BKG97, CV99, CFC91, Che90, CP92, CH97, CM91, CP96, DPSB99, FZ99, FDC^{+}96, FSKC97, Fri98, Gae92, GPO97, GGC98, Håk96b, Håk97, Håk99, HL96, HKMRS95, HMC99, HW98, HM97b, HLF99, JSC91, Kai96, KV97, KMS99, Kir92, Kra96, KBSC98, LAB95, LUC^{+}93a, Lav96, LRS^{+}93, Loe97a, LMG96, MUM98, McC95a, MS97, MAY^{+}98, MCTB95, Mit94, Mül92, MWJ92, Mül97, MP99, NZ92, OS95, PW93, PHB90, PEAS99, PSL^{+}93, PK92, RA97, Rip95, ST96, SGS^{+}95, Sil99, Sil93b, Str92, TB93, TBG93, Tor96, UAA97, Ulr92, UDVESM^{+}95, Whi99, WE96, iWLSL96, Xu97, XJT99, XJTL99, YS98, cZdZ96, dZcZ96, Sil95a]. ecosystem-fisheries [McC95a]. ecosystem-scale [HKMRS95]. ecosystemic [Sol99]. Ecosystems [Ano99f, Bos93b, JPS92, JPS99, Jør99b, Mau93, PSJ97, SJP99, AOD97, Aly91, Ano96p, Aok95b, Aok97, AW97, BK92a, BJ97, Bos92, Bre92, Bre95, CFC^{+}96, Com97, EFGE98, Fer95, FDC^{+}96, Gao96, GYYL96, GDS98, GSS94, Gra95, Han97a, Han98, HRV^{+}96, HPB93, HBP93, JPA^{+}99, Jør97a, Jør97g, LUC^{+}93a, Leg92, LH92, LL96c, MRCL92, MCU98, Mes92, Mil92, MG97, Nih93, PT97, PA99, PMP97, SBW^{+}99, Str94a, SGS93, SHJ95, Tho97, UB98, VMSP97, WL96, WBH^{+}98, YH96, vDvB95, Jør95g, Mau93]. Ecotechnological [Str94b]. ecotechnology [Zuc91]. ecotoxicity [VVK97]. ecotoxicological [AHHK97, FFS96]. ecotoxicology [Jør95b]. ECOWIN [Fer95]. ectoparasitic [Mar98]. Ed [Jør99a, Jør95c, Nie95c]. eddy [KW99]. edge [BC90, CFS93]. Edited [Ano99c, Ano99b, Ano99d, Ano99e, Ano99f, Jør97c, Jør97h, Jør97e, Jør99b, Jør97g]. Edition [Jør99a, Mar99a, Ano99f, Hug91, Jør94d]. Editor [Ano96p, Ano96r, Hug91, Jør91a, Jør93a, Jør93b, Jør94g, Jør95a, Jør95f, Kom94, Nie95b, Str92, Vas94, Ess90, Koh90, WG91a]. Editorial [Ano97r, Ano98f, Ano98g, Ano98h, Ano99g, Ano99h, Ano99i, HS92b, HSJM96, Jør96b, Ric95, Ano90f, Ano90g, Ano91o, Ano91p, Ano91q, Ano91r, Ano91s, Ano92n, Ano92o, Ano92p, Ano92q, Ano92r, Ano93m, Ano93n, Ano93o, Ano93p, Ano93q, Ano93r, Ano94l, Ano94m, Ano94n, Ano94o, Ano95p, Ano95q, Ano95r, Ano95s, Ano95t, Ano95u, Ano95v, Ano96-29, Ano96-30, Ano96-31, Ano96-32, Ano96-33, Ano96-34, Ano96-35, Ano96-36, Ano96-37, Ano96-38, Ano97s, Ano97t, Ano97u, Ano97v, Ano97w, Ano97x, Ano97y, Ano97z]. Editors [All91, Ano91i, Ano96q, Bos93a, Bos93b, HS94, Jør94a, Jør94e, Jør94f, Jør95d, Jør96a, Koh91, Leg92, Mau93, McD93, Nih91, Nih93, Sil95a, Str91, Svi91b, Zuc91]. Eds [Jør97b]. edulis [HS99]. Edwards [FSG96]. eelgrass [Bac93, BCB97]. Effect [Cha96, RSC96, SD96, SDF96, SDS96, WG98, And98, BDD97, BPTM^{+}97, DBB94, DH96, HS99, JB97, KI93, KL99, LS98a, MvdZ96, McC97, MT93, MT95, MT96, Mil92, Ned97, PG99a, PL95, SKŞ95, Sel95, SMDSZSXF93, Str96, SL96, TKO91, VDV95, WLW99, Ziv98, dCF99]. Effective [CMS90, WE96, CL99, Kaw97]. Effects [Dub97a, Han95, LGM99, LA96, Mal96, McC99, Rux96, VF96b, WVB93, YKH98, YKH99, vHY95, ADF^{+}93, AB97, Bar98, CBC^{+}98, CWC95, CR97b, CD99, De 98, Dua95, FS97, Fle91, Gar90, HTW96, HL96, HSXP98, JRB97, KHL97, KH98, LG91, LL96a, MS98, MGD98, MMJN97, MGNR91, McC96, MI95, ML97, MLGV99, ND97, Pat95, PSC93, PASG97, RRNM97, RÅ99, Sch96a, SNU96, SCH92, Smi95b, SA97, Spe97, Ver96, WU96, WCC95, WHA98, XJTL99]. efficiency [BCD93, GPG^{+}99, Omo97, RRF91, SM99, SCM93, SSK^{+}91]. Efficient [JTS92, HCL94]. effluent [RCD98, ST92]. effluents [SKW91]. effort [RS96]. egg [Mau97, STTR93, SY90]. eggs [Jen93, NHWS91]. Egrets [Smi97b]. eight [GPO97]. Eijsackers [Jør94f]. Eisenbach [MWK95]. elaphus [DDA99, SDBA98]. elasticity [GG92]. electric [GG99]. element [Kal99]. elementary [PMP97]. Elements [Jør99a, Bár96a, GPW97, Wen92]. elephant [DPSB99]. elevated [CCKO94, CC96, HH98, MJ99]. elevation [AL99, RCD98]. Ellis [Str92, Vas94]. Elsevier [Ano96p, Ano96r, Bos93a, Bos93b, Jør97g, Koh91, Mau93, McD93, Nie95b, Str91, Ulg95]. embayment [BGF99]. embodied [SV98]. Emergence [FP98]. Emergent [KB97b, BG92, OS95, Phi95]. Emerging [RB99, JPS92, JPS99, PSJ97, SJP99]. Emergy [BM97a, BM96, Chr94, JNM95, UOB94, TBM96, Pat92, Pat95]. Emergy-based [Chr94]. Emergy/exergy [BM97a]. Emission [Jør95f, Koh91]. emissions [HWB96, PRK97, Tay96, WHJ^{+}92]. emphasis [Hon92, Jør99d]. empirical [CFS93, CHC91, GPW97, LHAJ99, MCGLJ99, SH99]. empirically [MHB^{+}96]. Empiricism [Rut93]. Encephalitis [TGM95, Lin98]. enclosures [Boc99]. encounter [Bye96]. encyclopedia [Jør95a]. endangered [McK96, PGS98]. energetics [MGCR98, ZMH96]. Energy [HSU96, KB99, Pat92, BHWP91, Gra95, Han97b, HKMRS95, HPB93, HBP93, Jør95b, Kir99, MGNR91, Mau90, OP96, RGRF98, Rip95, SGML91, Smi97b, Spe97, SV98, Ter93, THR96, Wu90]. energy-circuit [Smi97b]. energy-flow [BHWP91]. energy-processing [Mau90]. energy-transport-reaction [Rip95]. energy/protein [MGNR91]. engineering [HSXP98, XJTL99, Zuc91, Ano96r]. England [HS94]. Englewood [Sil95a]. Engman [Mar93]. enhancing [dLNT98]. enrichment [CHM96]. entropy [Aok90, BU98, PT99]. envelope [Löf96]. environ [JNM95]. environment [AL94, DCP99, Dur96, Gao96, GET98, HC97c, Il’96, Mar99b, MS95, PT95, Per98, Ste95a, Ume95, Vir96, WW92, Wu90, AN90]. Environmental [Jør95f, Koh91, Kom94, RKM97, Str91, TBM96, Ulg95, VP93, Bos93a, BH94, Cle98, DD̆Á97, DFP^{+}93, DT97, EMHB99, FF91, HS94, HSHS96, JK99, Jør95a, Jør97d, Jør97f, JLS98, KW99, KHL97, MGD98, MHB^{+}96, Nor96, PFP^{+}95, Ped98, Pow93, RH96a, Sam96, SNU96, SIU97, SSY95, UOB94, Ush91, VK97, VK99, Vas94, WHN96, Koh91]. environmentally [SDS96]. environments [De 98, Ebe93, Liu93a, Ula95, Ano99f]. environs [Pat92]. EOLE [Sal97]. epidemic [Sch96b, TC96]. epidemics [Sch96b, Hug91]. epidemiology [Mor92]. epilimnetic [OA98]. epiphytic [TSMC96]. equation [HC97b, Jen96a, M.a93a, MRLC96, Sil95b, TDK98, WVCS94, YWMM98]. equations [ALB92b, Chi97, GQB91, HRV^{+}96, Kim94, Liu96, Mor92, Sog93]. Equilibrium [Kin95, NMS95, Sog93, UB97]. Equivalence [Hul92]. era [GRJP96]. Erie [ÖM97b]. Erik [Mar99a]. Erithacus [RB99]. erosion [BLRV98, CBP95, Jør93b]. erratic [AM98b]. Erratum [Ano94p, Ano96-39, Ano97-27, Ano97-28, Ano98i]. error [Hen95, SY99, WG98]. errors [PM96]. Escape [Ned97]. establishment [Böl90]. Estes [All91]. Estimate [BSV91, BILL98, CBB97, MD97, NCNN97]. estimates [Els92, Sto92]. Estimating [AOE99, AL99, FZ99, GH94, Spe94, Col97, DA95, GFS99, Han91, MvdZ96, MR91a]. Estimation [BJM92, KL98b, PL90, Val99, VF96a, CBP95, DFP^{+}93, HC97a, KB97a, KV90, Kra96, ML92a, Mow97, OR99, SV96, SPP^{+}97, VPFM97]. estuaries [FKNM^{+}97]. Estuarine [Ano91i, FKN97, MPNJ97, PSK98, PAGL96, WHA97, WHA98]. estuary [MMJN97, LVH96]. ethanol [RPM^{+}98]. ethological [BM94]. etiquette [BK97]. ETR [Rip95]. eucalypt [WW92]. Eucalyptus [BS98a, Kin96, SRB99]. Eulerian [JK99]. Euphausia [MS98]. Eurocourses [Jør95f]. Europe [KdH99]. European [Jør94f, AH98, MRSW93, RB99, Sey92, Wis92b, vWB99]. EUROSEM [BLRV98]. EUROSEM-GRIDSEM [BLRV98]. EUROSOL [Jør94f]. eutrophic [AB97, JA96, Szy96, XJTL99, YS98]. eutrophication [Aok95b, FKN97, GRV93, Håk99, Leg97b, MPNJ97, Sal92a, SMY^{+}97]. eutrophied [MMJN97]. evaluate [AS94, GBBV99, Kin96, MRSK96]. Evaluating [AWS98, Tun90, VS97, WHA98, Loe97a, PMV96]. Evaluation [Böl90, FTH99, HL96, JHM97, KW95, MTCL96, MMJ^{+}98, OKS99, RB98, SMY^{+}97, CT96, GA97a, HB93, HSHS96, MKA^{+}95, Met98, NCNN97, OP96, PL90, SGML91, SM99, SMWF98, Smi95b, VVDF95, Ver96, vT95]. evanescent [Pol96]. evaporation [VF96a, VPFM97]. evapotranspiration [LRG98, M.a99, MWK95, RKM97]. even [Bar91, Val99]. even-aged [Bar91, Val99]. Event [GS98b, LCL93, dVZ93]. Event-based [GS98b]. events [Rux96, SB97]. Everglades [Jam93, VFC98, WSGR96]. evolution [Coq95, KJ94, MAY^{+}98, MCTB95, MCTB96, RMT92, Jør95c]. evolutionary [FAF98, MG98, MC93b, Vin96]. Examination [Nie97, Sal92a]. examine [SJ96]. examining [CALW99]. example [Bot95, CS96, DD̆Á97, FP93, GGP97, McC92, MB96, PFP^{+}95, Rut93]. Examples [Ano92s, HE96]. excess [KV90]. Exchange [Kir94, DH99, GDS98, PA96, RSD94, SGS93, WBH^{+}98, Wu90]. Exchanges [Mar99b]. exclusion [ALB92b]. exemplified [Wis92a]. exercises [Jør93c]. Exergy [Aok92, BJ97, Jør92d, Svi97b, Xu97, BM97a, BU98, Jør92e, JNM95, MPNJ97, Nie95d, Nie97, Pat95, Sal92a, SJ96, ZMH96]. existence [MR91b]. existing [FFS96, HWB96]. exotic [SFGF94]. expanded [JK99]. expansion [GQB91]. expected [MCU98]. experience [RC97, TTT98]. experiences [DSB^{+}99]. experiment [HSXP98, TIH98, VDV95]. Experimental [Jør92a, CHM96, Gra95, MPA96]. experimentation [HSHS96, PG97]. experiments [Cas96, Gra98a, Jør93b, LL97, MJ99, Nal97]. expert [CF99, Dro96, HV98, RPJK94, ŠŠZ90, THR96]. explain [Gon97]. explicit [And96, CF99, Cha99, KdH99, LA98, TGM^{+}99, VdKK^{+}99, WU96, WJW99]. exploitation [BMHT99, Jen91]. Exploitative [Kaw97, Han95]. exploited [MVBP93]. Exploiting [BH99, PMV96, SW99]. exploration [BSN^{+}98]. Exploring [Gor98]. exponential [Coo95]. exposure [CMG98, GCT96, MvdZ96]. extended [ESD98, SMP96, Whi99]. extension [GC91, KBV99]. external [Sil99, VEA99]. Extinction [McC96, GA96, GH97, Liu93a, MBF96, SW99]. extra [WCDL96]. Extrapolating [Sou99, BL93]. extremal [AO97a, Pat95].

F [Ano96o, Bos91a, Bos91b, Jør94d, Jør96a, Leg91a, Leg91b, SKŞ95]. F. [HAW97]. facilitating [KD99]. factor [PS90, WRHW94]. factorial [HSHS96]. Factors [All90, HLF99, Håk99, Kin96, RB98, SIU97, TTT98, Woo98]. FAGUS [Hof95]. Falco [PT99]. falcon [PT99]. Falconer [Kom94]. fall [PS90, SC92]. fangiana [iWSG96]. far [UB97]. farm [Tho97]. farming [Jol96, RH96b]. farmland [ZH98]. fast [RM92]. fat [Cac90]. fate [AK96, BW96, JMA97, KdH99, SRW97, vdHvEH96]. feasible [Tom96]. feature [Foo99]. February [Ano93-27, Ano94y, Ano95-39, Ano97-41, Ano97-51, Ano98-32, Ano99-34, Ano99-44]. feed [MLGV99]. Feedback [Par93b, All90, Mos92, Ula95]. Feedbacks [Dup98, DSB^{+}99, Woo98]. feeding [AG93, AK93, Boc99, DBB94, MT93, ML97, MLGV99, RY91, RFB99]. females [HS97b]. Fenhann [Koh91]. feral [Hon92]. fertility [Kin95]. Fertilization [Cla99, KEAP97]. fertilizer [KBL96]. fertilizers [ZH98]. fescue [HAW97]. Festuca [HAW97]. fever [TMG96, TMGS97]. ff [Jør99a]. fiber [SCEZ^{+}94]. Field [Ano91i, BMV99, Bre95, CBC^{+}98, FP93, FSC^{+}97, GPW97, JMA97, Kat90, LK99, MJMN97, MMJ^{+}98, Per96, RC93, STTR93, SPE95, SN99b, cZdZ96, dZcZ96]. fields [Sem99]. figures [Ano99b, Nih91]. filter [ST92, TD94]. filtering [ESD98]. finding [McC97, MG96a]. Findlay [Nih93]. fine [iWHS97]. finite [Kal99]. Finland [Vir92]. fir [CFS93, HLF99, RY91]. Fire [WSGR96, BPTM^{+}97, FSG96, KO95, KT97, KG95, LTMP97, MU99, PA99, PD99, RPT93, SFGF94, Tap93, dVZ93]. fire-return [PD99]. fire-sensitive [FSG96]. First [Jør94f, PH95, PK92, HRWC91, SPP^{+}97]. fish [BMHT99, Bey98, BGTL99, Cac90, FSA99, Gam98, Jen91, Jen93, LLM99, LR97, MG98, MVBP93, Por98, RLHD99, RSD94, RFB99, Rom96, SG97, SNU96, SIU97, Sen90, SN95, SHEM95, SSK^{+}91, WCS95]. fisheries [AM98b, GG92, MS98, McC95a, MS97, PMV96]. fishermen [DL99]. fishery [GC95, Got98, Gra98d, MCT97, PC99b, RS96, WFK^{+}91]. fishes [DS99]. fishing [McC95a]. Fishman [Jør91b]. fit [RS98a, SR95a]. fitness [ZH97]. Fitting [AJK95]. five [Per93, vdSdVK96]. FIWALD [Sch95]. fixation [Kas93]. fixation-leaching [Kas93]. Fixed [LB97a, HSN96, RCD93]. Fjord [FKNM^{+}97, SS98]. FL [Nie95a]. flaming [RAM99]. flamingos [TAM^{+}99]. flatwoods [LRG98]. flavicollis [BS98b]. flea [Sey92]. flexibility [MA93b]. flocks [Rux95]. flooded [RRJ90]. floodplain [WFJ96]. Florida [Smi97b, Ahn99, CG93, CD99, Jam93, LRG98, MLDH90, Por98, RM95]. Flow [Aok95b, Mal95, AH98, Aok97, BHWP91, FP99, GGJ96, HSU96, Han97a, HW98, JCS^{+}97, MA96b, Nak92, PH95, PK92, Rip95, SH91, Tor96, VJR^{+}98, VNLB96, VFC98, VVC99, Whi99]. Flow- [Mal95]. Flow-indices [Aok95b, Aok97]. flow-thru [PK92]. flower [TBG93]. flowering [RJN99]. flows [BSE95, PH95, THR96]. fluctuating [dGG98]. fluctuations [AK93, TAB^{+}93]. fluid [BU98]. Flux [SV96, GDS98, Jam93]. Flux-based [SV96]. fluxes [DA95, HW98, KW99, RGRF98, SS98, SV98, VVK97, vWB99]. fly [JWEB92, SMP96]. flying [MGCR98]. focus [Håk99]. foliage [Bar98]. foliar [VB99]. following [RCD93]. fontinalis [PP95]. Food [Hae91, CW95, Col98, Gob93, GYNE99, HKV^{+}95, HSU96, Han95, Jac96, Jen96b, Kei97, KSR97, Kle95, KVAW91, Leg91a, Leg93, PACS96, PW97, RS93, Spe97, VG98, WG91b, vdB98, Jør97e]. food-chain [KSR97]. Food-web [Hae91]. food-webs [Gob93]. foodweb [CL99]. footed [PT99]. forage [BBH92, BBH92]. Foraging [Rux95, BF98, CF99, DuB97b, MS98, MGCR98, RRF91, TWRW93]. forcing [Bot98]. forecast [KMS99]. forecasting [Alb92a, AH99]. Forest [Ano99e, IK95, Jør99b, LRS^{+}93, MvdV95, Ulr92, vGDT95, AOD97, AUS96, AO97b, Bar98, Bat96a, BS95b, BD99, BS98b, BK91b, Bos96, BJ95, BPTM^{+}97, BFK96, Cha99, CFS93, CU93, Com97, DDA99, DSB^{+}99, Dye95, GET98, GFS99, GPT95, GBBV99, GGP97, HMA90, HKMRS95, HMC99, HMB99, Jes99, JPA^{+}99, KO95, KT97, KI93, KV97, KW99, KMS99, Kir99, KL99, KSD^{+}99, Kre91, KI98, LH92, LSB99, LA96, LTMP97, LB97a, Liu93b, Liu97, LA98, LMG96, MA96a, Mal96, MUM98, MU99, MPA97, Ned99, OYA95, OA97, Osh91, Osh96, PA99, PRK97, PSC93, RM95, Rob96b, dLNT98, RÅ99, SDBA98, SM97, TvG95, TIH98, Tor96, Tro90, VS97, Van96, VS96b, VS96c, WU96, WM95, WW92, WCC95, Wil96b, Wis92b, WV91, dVZ93, dVKvdS95, vDvB95, vHY95, vHvGT95, vdSdVK96]. Forest [Jør99b]. forest-soil [vdSdVK96]. Forest-Soil-Atmosphere [vGDT95, BJ95, TvG95]. forested [ASP93, Jør95g, RCD98]. forestry [Gra98d, Löf96, Ano99e]. forests [ADF^{+}93, AYO99, BS95a, BS95b, BS95c, BvSH91, Cha99, Chi98b, JAZ^{+}99, KABT90, KN95, KBV99, Kol98, KL99, LB97a, LL96a, Mow97, VB99, vWB99, Bos93b]. foreword [Sil95a]. FORGRO [MvdV95]. forhyd [Bou95]. form [CFC^{+}96, Par96, Smi97a, VV92, OYA95]. ForM-S [OYA95]. formal [AN90]. Formalisation [Dro96]. format [Kom97]. Formation [CTDC94, GPW97, Mal95, Van97]. formed [ST99]. former [GPO97]. formerly [Ano91i]. Formind [KH98]. forming [And98]. FORMOSAIC [LA98]. forms [Nie97]. formula [BM97b, Tro90]. formulae [Han96]. formulation [EH91, RL99, UDML97]. formulations [Bar96b, GPT95, MP99]. FORSOL [VMB95]. FORSUM [Krä95b]. ForSVA [AO97b, OA97]. Fossil [Tay96]. fossorial [And96]. fouling [Svi91a]. four [BG93, Per96, SH91, VF96a]. four-year [SH91]. fox [ALS97, SM99]. Fractal [Ken92, DB98, LLM96]. fractals [LL96b]. fractional [HSHS96]. fragmentation [Fah98, HMA90, JMPT99]. fragmented [MA96a]. framework [CC94, GGZ99, GGP97, LJ94, Loe97a, MB96, PGD91, PP95, WL97]. France [ST96, Sal97, SvdG99, TS93, TAM^{+}99]. Francisco [Jør97e]. Frankfurt [LRK96]. Frederick [Ano91i]. free [Cla99, GRBF97, HH99]. free-ranging [HH99]. free-spawners [Cla99]. freely [GSG95]. Frequency [HL97, DD̆Á97, MT93]. Fresco [Jør97h]. freshwater [MM97, MR91a, Rya94, Sog93, VCGG96, WHA98, XJT99]. fringed [HSB99]. Frio [CV99]. frugilegus [PT99]. fruit [DBB94, YAKS97]. fuel [SADZ95, Tap93, Tay96]. fuelbed [KO95]. fugacity [BW96]. Full [Ano98j, Nih93]. function [BJ97, Dam99, FRS^{+}95, HE96, Mor99, UAA97, VEA99, Whi99]. Functional [LCMR96, BDD97, BD97, HCL94, MRL94, MRLC96, PWU93, PT97]. functioning [PSN98]. functions [Bas98, MWY92, MPP98, Vol98]. fundamental [MWY92]. Fundamentals [Ulg95]. fungal [YAKS97]. fungus [SvB93]. further [Gav95, GdC99]. future [And95, Gri99a, IPS99, JPS92, Jør97h, Lyo96b]. Fuzzy [CHLT96, Foo96, FFS96, LT96, Met98, Sal92b, Bár96a, BS98b, Cao95, CS96, DSNN96, Dro96, MM96, PT99, Rob96a, Rob96b, Sil97, THR96, WGP96, iWLSL96, ZBDN96].

G [All91, Ano91i, Ano96q, Jør94a, Jør96a, Jør97g, Jør99b, Koh91, Nie95c, Ste93, Svi91b]. gain [Hua98]. gains [Löf96]. gall [Axe93]. Gal’perin [Svi91a]. game [BSB97]. gap [BFK96, CL90, IK95, WU96, JPA^{+}99]. garrigue [RPT93]. Gary [Jør97e, Nih93]. gas [HWB96, PA96, PRK97, SR94, SR95b, WBH^{+}98]. gases [Jør95f]. Gaussian [Mow97]. general [Col98, FDC^{+}96, Håk96a, Håk96b, Svi97a, TKG97, Yin96]. generalist [PR95c]. Generalists [VV91]. generalizability [Bre92]. generalization [Vol98]. generalized [IT97]. generating [SVK93]. Generation [SSM^{+}93, Mat99, MN93, SR98]. generations [Jen94]. Generative [VUDM^{+}99]. generator [Fri98, Oel95]. generic [CR97a, DT97, GGZ99, LBA^{+}95, Mar99b, SOM97, LFM97]. genericness [RA97]. genetic [Kir92, Löf96]. genetically [vdHvEH96]. genetics [DHRLE90]. genotypes [Wil91]. genotypic [GdC99]. Geographic [KBV99, Hil96, Ken92, KG95, PC99a, RPH99]. geologic [LL96b]. George [MEBZ96]. GePSi [CR97a, LFM97]. German [KH94b, KK97, SM95, SHJ95]. Germany [Ano95w, Jør93c, Jør93b, BSE95, FWCJ95, MKA^{+}95, MWK95, PDAS95, SPE95, Smi95a, VMB95, WM95]. germination [And98, SvB93]. Giant [CALW99, iWSG96]. Gilbert [Nih91]. Gillman [Jør97b]. gillnet [PMV96]. gills [SSK^{+}91]. GIS [CPPB99, EB93, Mow97, SS98, TTT98]. GIS-based [SS98]. glacial [HL97]. glaciation [SADZ95]. Global [Jør99b, KBL96, MY97, TC96, YH99, Ano96r, BBKV96, DSB^{+}99, Fri98, Kra93, Log97, LRK96, ML92b, Oel95, RS91, SvB97, SvB98, Svi99, Yin96, Jør91b, Jør94b, Nie95b]. globulus [BS98a, SRB99]. goal [Bas98, BJ97]. goals [BSB97]. Gomatam [Sam96]. goodness [RS98a, SR95a]. goodness-of-fit [RS98a, SR95a]. Gosselink [Nie95c]. GPA [RL99]. GPA-Phenodynamics [RL99]. GPSF [GGZ99]. gradient [FKN97, GPO97, MPNJ97, PRK97, Sal92a]. Gradients [Mül98, BvSH91, CFS93]. Graham [Jør94e]. grain [SKŞ95, Tho97]. graminoid [Cou91]. grammars [VUDM^{+}99]. Grande [GHQ99]. grandis [MRSK96]. grapes [ASTL91]. graphs [Nie95a]. grass [CCKO94, CHM96, Hut91, MMJ^{+}98, SGNS91, Spe94, TSMC96, cZdZ96, dZcZ96]. grasses [DCCSL99]. grasshopper [HC97b]. grasshoppers [BWW99]. grassland [ASP93, CL90, CU93, GYYL96, GLD97, GPO97, SCH92, TS96, WK97a, WBH^{+}98, YH96]. grasslands [GHQ99]. Gray [Jør96a]. grazed [Hut91]. grazer [BDD97]. grazers [McC92]. Grazing [ASP93, BK91a, CH97, HWB96, ST99, Str96, TMGS97, iWLSL96]. Great [Smi97b]. greater [TAM^{+}99]. Greece [DCCSL99, PS93a]. green [KG96, MMJN97, RL99]. green-tree [KG96]. greenhouse [CHLT96, HS92a, HWB96]. GRIDSEM [BLRV98]. Gromiec [Str91]. Gross [Svi91b, Liu96]. ground [WMRK93]. groundwater [Bat96b, GGJ96, Per93, Var93, VBL90, WTL91]. group [Han96]. grouped [Cra98]. grouping [KH98, YO95]. grouse [HBM^{+}97]. Grove [Jør97e, MPA96]. growing [Gav95, MT96, Whi95]. grown [RSKS91, SR94, SR95b]. Growth [Gam98, AFM99, ACS96, Bac93, BCD93, Ben99, BSE95, Boc99, Bon93, BSB99, Bos91b, CFS93, CS97, Dam99, DBB94, DDA99, Dow97, FSC^{+}97, GS96, GFS99, Gor98, Gra98b, GGP97, HV98, HF93, HKMRS95, HMW96, Hof95, IT97, JLS98, KEAP97, KS95, KIPK99, KG95, Kin96, KG96, Kir99, Kit97, KN95, KBV99, Kre91, LA91, LC97a, LR97, LBV96, LB97a, LLWP95, MGNR91, MT93, MT95, MvdV95, MI95, MLGV99, Mos92, Mow97, PR95a, PL90, RÅ99, SJ96, Sch96a, SG97, SRV95, SS96, SSS^{+}91, SK91, SM97, SABG93, SL91b, TDK98, VEA99, VS97, VS96b, VS96c, WMRK93, WRHW94, ZH98, Zon98b, vHY95]. guide [WHJ^{+}92]. Guinea [GMI95]. Gulf [PS93a, ML97]. Gurney [Mar93]. Gypsy [STTR93, CWC95, SC96, WVCS94, Wil99]. Gyrodinium [And98].

H [Ano91i, Ano99b, Ano99f, Bos91b, Jør94e, Jør94d, Jør97h, Koh91, Kom97, Mau93]. Habitat [RPH99, RRF91, BBL97, Fah98, FSA99, GA97a, HF96, HSB99, HC97b, Jes98, Jør95g, LGM99, LADL99, LW94, NG96, ÖM97b, ÖÖ99, SD96, SDF96, Spe97, SHEM95, SL96, TMG96, TMGS97, VW97, Ziv98]. habitats [Jør97c]. Haematoloechus [MGD98]. Hails [Jør97b]. Hall [Jør97e, Leg93, Mar93, Sil95a]. Hallam [Svi91b]. hallii [HAW97]. Hamers [Jør94f]. Hamilton [Web95]. Hamiltonian [Chi97]. Hannon [Wu95]. Hansen [Bos93a]. Hard [All91, Ano99b, Ano99f, Kom94, Leg93, Vas94]. hardback [Kom97]. Hardbound [Jør91b, Jør91c, Jør95b, Jør97d, Jør97c, Jør97h, Jør97e, Jør97f, Jør97g, Ano99e, HS94, Jør93a, Jør94a, Jør94b, Jør99b]. Hardcover [Jør92a]. hardwood [KG96, KBV99, LBV96]. hardwoods [OA97]. hare [RF97]. harem [MK97]. harvest [Jen96a]. Harvesting [McC95b, Vir96, CD99, HNCS92, KL98a, LA91, MG98, MG96a, MCT97, RÅ99, TAB^{+}93, Vir92]. harvests [BSB97]. hat-trick [Bla95]. hatchery [ALGB99]. Haurie [Jør97f]. Hazardous [Jør94e, Jør94e]. HB [Ulg95, Mar99a]. HD [Kom97]. health [XJT99, XJTL99]. heat [EGJ95, KW99, KS96]. heathlands [Coq95, HB93]. heavy [KCP^{+}98, MVBP93]. heavy-metal [MVBP93]. hedgerows [TIH98]. height [HM97a, Hut91]. held [Ano99e, Jør91a, Jør93a]. helicopter [MJGW94]. Helicoverpa [DFHR96, MHA95, RDFZ96]. Heliothis [NHWS91]. helmeted [McC96]. Hemisphere [Rya94]. herbage [SFS97]. herbicide [Jar95]. herbivore [And96, SR93]. herbivores [BK91a, HLG96, LGSF92]. herbivorous [LW96]. HERBSIM [DNR95]. herd [BW94, ST99]. Herendeen [Ano99c]. Herman [Ano99f]. herring [Jen97]. Herrling [Jør96a]. hervibore [Vir92]. heterogeneity [Gao96, GPW97, LGM99, RRF91, VDV95, WC96, Ziv98]. heterogeneous [AW97, LLM96, Tis97]. Hierarchical [Lav95, LMG96, Mül92, GF90, Kir92, Kir94, Mau90, PW93]. hierarchically [JMPT99]. hierarchies [BG92]. Higashi [Leg92]. High [GA96, ASTL91, MCTB95, MCTB96, RCD93, SSK^{+}91]. higher [ALB91, ALB92b, Rod99]. higher-order [ALB91, ALB92b]. highlands [GMI95]. highly [HW98]. hilly [GLD97]. Himalayan [MDO99]. history [CALW99, CU93, KJ94]. Hoekstra [Jør94d]. holism [PW93]. holistic [Jør94b]. Holland [EBBW98]. Holly [MA93b]. Holly-Preissmann [MA93b]. Holocene [ET95b]. holographic [FZ99]. Hom [DBB94, SKŞ95]. home [MvdZ96]. home-range [MvdZ96]. homogeneous [PS93a, Tap93]. homogenization [ALB92b, FP99]. honey [MRSW93, Mar98]. honeybee [DHRLE90]. honeyeater [McC96]. Hopf [PRF97]. Horwood [Str92, Vas94]. Host [Axe94, BL92, BLS98, Bye96, KS91, LW96, MGD98, MG96b, PGD91, TMG96, TMGS97, WS97]. host-parasite [WS97]. host-parasite-landscape [TMG96, TMGS97]. Host-parasitoid [Axe94, BL92, MG96b]. hosts [MHA95]. hourly [Ant98]. hoyi [Jen97]. humus [Che90]. hunger [OM97a]. hunted [BSV91]. hunting [Col98]. Hurricane [CD99]. Hutchinson [ALB92b]. hyacinths [HSXP98]. Hybrid [RPJK94, KMS99, FSKC97]. Hydrilla [CC96]. hydro [GG99]. hydro-electric [GG99]. hydrobiology [Jør94a]. hydrochemical [WFJ96]. Hydrodynamic [RC97, RCS97, TM99]. hydrodynamics [OZIA96]. hydroecology [Gue91]. hydroelectric [DT97]. hydrogen [BSL97]. hydrography [US97]. hydrologic [PJL^{+}96]. Hydrological [BL90, Bár96a, KMWB98, SvB98, VW97, VS96a]. hydrological/water [KMWB98]. hydrology [MR91a, VWN^{+}99, Mar93]. hydrophobic [Gob93]. hyperbola [Han91]. Hyperbolic [Fle91]. hypolimnetic [OA98]. hypotheses [SC96]. hypothesis [Loe97a, PG97, PG99b, WFB96, Was95]. hysteresis [Smi95b, VDV95, Ver96].

ice [FS97]. ice-covered [FS97]. ichthyological [GSG96]. Ictalurus [MGNR91]. Ideal [GRBF97]. identification [SPP^{+}97]. Identifying [Gae92]. II [BS95b, Cas96, Ces97b, CP92, CG98, DGWK97, MRLC96, OA97, SH97, TBG93, VS96c]. III [BS95c]. illustrate [Woo98]. illustrated [JTS92, Jør91b, Jør91c, TGM^{+}99, Jør97b]. illustrations [Nih93]. Image [VZF96]. imagery [BD99, Foo96]. immune [TC96]. Immunity [GCT96]. Impact [Jør99b, BK98, CTL98, DR96, Kit97, LAB95, McC95b, PGD91, RH96b, SADZ95, SMWF98, Sil97, Spe94, UMJ^{+}97, VSG92, dVKvdS95]. Impacts [Jør99b, BG93, CH97, DT97, HB93, Jør91a, Jør97c, KG96, McC95a, Pow96, SVK93]. Implementation [Lho94, CPB97, HAOR97]. implemented [Gao96]. Implementing [MC93a, SOM97]. implication [Aok97]. Implications [BH96, CU93, LW96, Pie95, Spe94, BM94, Bon93, Bye91, Chi98b, Got98, GPT95, KIPK99, PSL^{+}93, TKO91, Tre91]. importance [DLN^{+}95, LC91]. important [Rux95]. imported [KG95, SFGF94]. impounded [BJ94]. impoundments [Stu98]. improve [SCEZ^{+}94]. Improved [Ano99e, CT96, GPG^{+}99]. improving [SY90]. inbreeding [McC96]. incineration [WGP96]. Inclusion [SCEZ^{+}94]. Inclusive [Aok93]. income [LBV96]. incompatible [DJ98]. Incorporating [Liu96, Gra98a]. incorporation [MRLC96, MJMN97, MMJ^{+}98]. increase [CB96]. increases [WF99]. increasing [ALS97, SCM93]. increment [HM97a]. increments [PR95a]. independently [Bot95]. Index [Ano97-29, Ano97-30, Ano97-31, Ano97-32, Ano97-33, Ano97-34, Ano97-35, Ano97-36, Ano97-37, Ano97-38, Ano98k, Ano98l, Ano98m, Ano98n, Ano98o, Ano98p, Ano98q, Ano98r, Ano98s, Ano98t, Ano98u, Ano98v, Ano98w, Ano98x, Ano98y, Ano98z, Ano98-27, Ano98-28, Ano99j, Ano99k, Ano99l, Ano99m, Ano99n, Ano99o, Ano99p, Ano99q, Ano99r, Ano99s, Ano99t, Ano99u, Ano99v, Ano99w, Ano99x, Ano99y, Ano99z, Ano99-27, Ano99-28, Ano99-29, Ano99-30, Ano99-31, Ano90a, Ano90b, Ano90c, Ano90n, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano94c, Ano94d, Ano94e, Ano94f, Ano94q, Ano94r, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95-30, Ano95x, Ano95y, Ano95z, Ano95-27, Ano95-28, Ano95-29]. index [Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano96k, Ano96l, Ano96m, Ano96n, Ano96-44, Ano96-45, Ano96-46, Ano96-47, Ano96-48, Ano96-49, Ano96-40, Ano96-41, Ano96-42, Ano96-43, Ano97g, Ano97h, Ano97f, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97-65, Ano97-66, Ano97-64, Ano97-59, Ano97-60, Ano97-61, Ano97-62, Ano97-63, Aok93, BM97b, JSC91, LA97, WF99, iWHS97]. India [MHA95, SD96, Shu98]. indicators [Vas94, Xu97, XJT99]. indices [Aok95b, Aok97, BMHT99, CBP95, Hen95]. indirect [MS98, Pat95]. Individual [CR97b, DL99, GWAU99, JRB97, Łom99, Rom96, RCD93, VJR^{+}98, BM99, BL92, BF98, Ben99, CF99, CPB97, DGS91, Gri99a, HBM^{+}97, HCL96, IPS99, Kei97, KH98, KTM95, KI98, Lav95, LR97, LSB99, Lho94, LA98, LS98b, LS99, MG98, MB96, MPA97, PR95c, RLHD99, RB99, RRF91, SBD^{+}95, Sou99, Spe97, Tis97, Tre91, Uch99, WW92, Wol94]. Individual-based [CR97b, DL99, GWAU99, JRB97, Łom99, RCD93, VJR^{+}98, BM99, CF99, DGS91, Gri99a, HBM^{+}97, HCL96, Kei97, KH98, Lav95, LR97, LSB99, LA98, LS99, MG98, MB96, MPA97, RLHD99, RRF91, Spe97, Tre91, Uch99]. individual-oriented [BL92, Lho94, LS98b, Wol94]. individual/landscape [CPB97]. individuals [Gav95, KD96, RCD93, SBD^{+}95, WC96]. induced [BH97, Jes99, Mal95, PJL^{+}96]. induction [Coq95, DGWK97]. inductive [HAW97]. industrial [Jør95f]. industry [Jør91a]. infection [GCT96, KGD97, SvB93]. infective [TC96]. INFER [HV98]. inference [ZBDN96]. inferences [LR97]. infiltration [Str96]. inflow [VSG92, WHA98]. Influence [FK99, RU97, Sam96, SK91, Bel92, Bot95, CFC^{+}96, HV99, JTS92, KL98b, MGCR98, Mar91, TTT98, TMG96, TMGS97]. influenced [BJ94, WMRK93]. influences [GdC99, Sil99]. Information [Aok94, BK92a, Cle98, DFPG91, DG93, Hil96, Kir94, KRD96, LW94, RPH99, SV98, VKT94]. informational [DFP^{+}93, UAA97]. Ingham [Jør97d]. ingredient [MG97]. inhibition [CMG98, Par93a]. initial [Bar96b, HV99, Liu93a]. initiation [RGBM90]. Inland [KKNY91, CMS90]. Innsbruck [Str95]. inorganic [Col97, NF96]. input [PH95, RSKS91, WG98]. input-output [PH95]. INRA [Jør99a]. insect [DJ98, Hog99, MD97, MG96b, SMWF98, YWMM98]. insect-mediated [DJ98]. Insecta [CS96]. insecticide [CMG98]. insects [LW96, Oga96]. insights [GdC99]. instability [FAF98, RA96]. instantaneous [BCH^{+}97]. Institute [Jør91a]. Institution [Kom94]. instream [VJR^{+}98]. instrument [PC99b]. intake [BBH92, BK91a, MGCR98]. Integrated [GT97, Jør94f, ANM99, AFM99, ANFM99, Fri95, HM94, Jol96, NAFM99, RB98, TN98, VVDF95, VF96c, WRHW94, Jør94f]. Integrating [Ano99b, Kre91, Cle98, Jør97h, Jør94c]. Integration [BD99, Jør97e, Pat95, Mar99a, Mit94]. intensities [RÅ99]. intensity [GBBV99, McC95a]. Intensive [MK95, MLGV99]. Inter [Hae91, Rya94, GW98]. Inter-processor [Hae91]. inter-relationships [Rya94]. Inter-site [GW98]. interacting [Dub97a, GS98a, Sam96]. Interaction [Jør94g, MR97, ALB91, BD97, Der98, HV99, HC97b, KR92, ÖÖ99, RBM97, SSW91, Sey92, SvB96, iWSG96]. Interactions [DDA99, Axe94, AHHK97, DH95, Hua98, JCC^{+}95, LUC^{+}93a, MR91b, PU97, PG99b, PGD91, PSL^{+}93, Rod99, SSS^{+}91, TBG93, TMG96, TMGS97, TCC^{+}97, VVK97, WS97, vdPV99]. interactive [RPH99]. interannual [Dun99, Hog99]. interception [Bar98, KS95, LS98a, Liu97, Sto92, dCF99]. interconnection [Bir94]. intercrop [Hua98]. intercropping [Pie95]. Interdisciplinary [RB96, CDDC96]. interest [BL96]. interface [BEM97]. interior [Kar93]. International [Ano99d, Jør91a, Jør93a, Jør99a, Jør99b, Jør94b, Jør94e, Ano95w]. Internet [Leg97a]. interpolation [PM96]. interpretation [Gue91]. interpreted [AOE99]. Interscience [Leg91b]. interspecific [BP93, BH94, GQB91, ÖÖ99]. intertidal [CPPB99, Mau93]. interval [PD99]. intracolonial [MRSW93]. Intraspecific [UH99, BP93]. intratidal [PSK98]. introduced [WCS95]. Introduction [AR97b, Ano92t, FB93, JMW92, Jør97a, LR96, MH96, Rot95, SFK96, Bey98, Jør97b, LG99, Zuc91, Bos93a]. introgression [DCP99]. inundative [SY90]. invading [KG95]. invasion [MS95]. invasions [SG98]. invasive [HCdVM98]. investigate [SLP93]. Investigating [MPA97]. Investigation [Mau97]. investigations [Alb92a, Kra93]. invicta [SFGF94]. invitation [Ano93s]. Irish [Per98]. IRM [WRHW94]. irradiance [VCGG96]. irreversible [Coq95]. irrigated [SRB99]. irrigation [Spe94]. Isarithmic [KO95]. ISBN [All91, Ano91i, Ano96o, Ano96p, Ano96q, Ano96r, Ano99c, Ano99b, Ano99d, Ano99e, Ano99f, Bos91a, Bos91b, Bos93a, Bos93b, HS94, Hug91, Jør91a, Jør91b, Jør91c, Jør93c, Jør93b, Jør94a, Jør94b, Jør94e, Jør94c, Jør94f, Jør94g, Jør94d, Jør95a, Jør95b, Jør95c, Jør95d, Jør95e, Jør95f, Jør95g, Jør96a, Jør97d, Jør97h, Jør97e, Jør97f, Jør99a, Jør99b, Koh91, Kom94, Kom97, Leg91a, Leg91b, Leg92, Leg93, Mar99a, Mau93, McD93, Mit94, Nie95b, Nie95a, Nie95c, Nih91, Nih93, Sil95a, Str91, Str92, Svi91b, Svi91a, Ulg95, Ush91, Vas94, Wu95, Zuc91]. ISEE [Jør93a]. ISEM [GRJP96]. Island [FFR99, SSY95, Wis92a]. islands [WV91]. isolated [De 98, KI98]. Isopoda [MMJN97]. isotopes [vDvB95]. Israel [WJW99]. ISSN [Jør93b]. Italy [BCB97, FKNM^{+}97, RMT92, UOB94]. iterative [KL98b].

J [All91, Ano99b, HS94, Jør94a, Jør94f, Jør95d, Jør97d, Jør97c, Jør97h, Jør99b, Koh91, Leg91a, Leg91b, Mar93, Nie95b, Nie95c, Nih91, Str91, Svi91b, Zuc91]. Jack [Jør91b]. jacobsoni [Mar98]. Jain [Jør92a]. Jakeman [HS94]. Jansen [Bos91b]. January [Ano90j, Ano92x, Ano93v, Ano94u, Ano95-33, Ano96-58, Ano97-56, Ano98-43, Ano99-45]. Japan [Chi98b, KKNY91]. Japanese [TGM95]. japonica [Chi98b]. Jasper [LFM97]. Jeffries [Jør91c]. JERS [Met98]. JERS-1 [Met98]. Jersey [VU93a, VU93b]. John [Ano91i, Ano96o, Ano96q, Ano99c, HS94, Jør92a, Jør97h]. Joint [LS98a, VK97]. Jonathan [Nih93]. Jordan [Ano96o]. Jörg [Jør93c]. Jørgensen [Bos93a, Mar99a, Mit94, Sil95a, Str91, Ulg95, Zuc91]. journal [Jør90, Jør93e]. July [Ano91v, Ano92u, Ano95-40, Ano96-57, Ano98-29, Ano98-38, Ano98-44, Ano99-39, Ano99-50]. June [Ano90k, Ano91z, Ano92y, Ano93x, Ano94x, Ano95-34, Ano96-56, Ano97-47, Ano97-48, Ano98-34, Ano98-42, Ano98-45, Ano99-40, Ano99-51]. juniper [FSG96]. Jürgen [Hug91]. juvenile [Lee91].

Kalish [Jør91b]. Kalman [ESD98, TD94]. Kappa [ML92b]. Keban [SMY^{+}97]. Kenneth [Ano91i]. Keoladeo [SD96, Shu98]. Keulen [Jør97h]. Key [Ano94q, Ano94r, Ano95x, Ano95y, Ano95z, Ano95-27, Ano95-28, Ano95-29, Ano96-40, Ano96-41, Ano96-42, Ano96-43, LV99, WBH^{+}98]. Keyword [Ano95-30, Ano96-44, Ano96-45, Ano96-46, Ano96-47, Ano96-48, Ano96-49]. Kiadó [Vas94]. kinds [MS95]. Kinetic [LV99, SMDSZSXF93, SABG93]. kinetics [Fom97, ML96]. kinship [HBM^{+}97]. Kirk [Jør97e]. kiwifruit [Buw91]. Kluwer [Ano99e, Jør94a, Jør94b, Jør94c, Jør94f, Jør94g, Jør95d, Jør95f, Jør96a, Jør97f, Jør99b, Mar99a, Mit94]. Knowledge [CS96, HAW97, BS98b, DSNN96, Dro96, HV98, Jør97h, KD99, Sal92b, Ush91]. Knowledge-based [HAW97, BS98b, DSNN96, KD99, Sal92b]. known [DPSB99]. Köhl [Ano99e]. Kooijman [Jør95b]. Kovács [Vas94]. Kramer [Jør99b]. Kranz [Hug91]. krill [MS98]. Krummbach [MWK95, SM95]. Kullback [Aok93]. Kuparuk [SN99b]. Kwiatkowska [Jør94e].

L [Ano99d, Jør95b, Jør97h, Leg93, McD93, Nih91, MR91b]. L-V [MR91b]. L. [ALGB99, Bac93, BM99, BCB97, DDA99, Ned99, ÖM97b, RPT93, RJN99, SC96, SMP96, SDBA98, TB93, Yam97]. laboratory [STTR93]. Lac [ST96, TS93]. Lack [HRV^{+}96]. lactuca [SFGH97]. lactucae [SvB93]. lacustrine [Mon98a, PT95]. Ladoga [MAY^{+}98]. lagoon [Cha95, RSdP98, SBD^{+}97, BCB97, CO99, CB97, CS97, FKNM^{+}97, FSC^{+}97, SPP^{+}97]. Lagrangian [YH93]. lags [NJ96, Phi95]. Laguna [Boc99]. lake [AJK95, And95, Aok95b, Aok97, BJ94, BH97, BGTL99, GRV93, Håk95a, Håk96b, HL97, Håk97, Håk99, Jen97, KBSC98, LHAJ99, LB97b, MVS^{+}95, Nie94, OZIA96, OH97, PR95b, SJ96, SF94, VF96a, VPFM97, WFJ96, XJTL99, Aok90, Gob93, HSU96, HSXP98, Jac96, JA96, LB97b, MAY^{+}98, MEBZ96, ÖM97b, PT95, RMT92, Sem99, SGS^{+}95, Sil99, Smi97b, SK91, Xu97, XJTL99]. lake-ecosystems [Aok95b, Aok97]. lake-wide [BH97]. lakes [Aok94, AB97, BLK^{+}99, FS97, FSA99, HL97, HS97a, JA96, JN94, LLM99, OA98, Rec93, RPJK94, Rya94, Sal97, SH97, SHEM95, VF96a, VSG92]. lance [KKNY91]. Land [GGJ96, HSPM96, VF96b, AOD97, BSN^{+}98, CBC^{+}98, Jør97h, LSB96, PJL^{+}96, VF96c, VdKK^{+}99]. land-cover-induced [PJL^{+}96]. land-use [CBC^{+}98]. Landscape [Bak96, HMA90, LW94, Rob96a, WSGR96, AW97, BDD97, CPB97, DG96, EFCM99, EBBW98, HMC99, HMB99, JMPT99, Li95, LTMP97, LA97, LA98, LGSF92, MA96a, SFS97, ŠŠZ90, TMG96, TMGS97, TWRW93, VFC98, VVC99, Ziv98, dVZ93]. landscape-scale [SFS97]. landscapes [BEM97, CBC^{+}98, Ema96, GLD97, Liu93b, PSC93, Tis97]. language [AN90, MC97, SNUM91]. Large [ABK^{+}92, AK93, BH94, HLG96, LGSF92, SBD^{+}95, SD97, TM99, TWRW93, VGDvdZ97]. Large-amplitude [AK93]. Large-scale [ABK^{+}92, BH94, SD97]. largemouth [JRB97, Tre91, TCC^{+}97]. Larsen [Koh91]. larvae [Jen93]. larval [LR97, Por98]. late [ET95b]. latitude [FRS^{+}95]. Latitudinal [Rya94, WU96]. Latitudinal-altitudinal [Rya94]. lattice [HMW96, KI93, Kei97]. lattice-structured [KI93]. launches [GRJP96]. Lausanne [Jør97g]. law [Fle91, Jør94e, Ken95, LC97a, Mos92, Jør94e]. layer [Bot98, HKV^{+}95, KVAW91, PT95, SN99b]. leaching [Kas93, MPA96, PS90, Per93, Per96, Whi95]. lead [CB96, DCCSL99]. leading [RS93]. Leaf [HH98, AW97, CWPD95, Fri95, Kat90, LC91, NMS95, PA96, RU97, SC92, VVK97, WBH^{+}98, Yam97, dCF99]. learn [Gri99a]. learned [Gri99a]. Learning [VK99, DL99, DGWK97]. leaves [NMS95]. Lecture [Ano91i]. legislation [Jør95f]. Leishmania [KGD97]. length [Ano98j]. Lepidoptera [RY91, NHWS91, STTR93]. Leslie [SSW91, SMP96]. lessons [VK99]. Letter [Ess90, Koh90, WG91a]. lettuce [SvB93]. level [AW97, BM97a, BSN^{+}98, CCK92, CU93, GCT96, GPO97, MRSW93, MS98, NMS95, RB99, SvdG99, VB99, Ziv98]. levels [CL99, Kir92, Sto92, Ter93]. Levin [Svi91b]. lichen [Vir92, Vir96]. lichen-reindeer [Vir96]. Liebig [LC97a]. Liénard [Mor92]. Liénard-type [Mor92]. Life [Ush91, CALW99, Cas96, Gra98a, GdC99, Kan90, KJ94, Sel95, SY90, WS97]. life-history [KJ94]. Light [Zon98b, AO97a, CA98, Dua95, Gri99b, KS95, KTM95, Liu96, ND97, PG99a, SK91, WW92, dCF99]. Light-limited [Zon98b]. LIGNUM [PSN98]. likelihood [Dam99]. lima [BM99]. liming [OH97]. limit [BM97b, DPSB99, RM92]. limitation [Sey92]. limited [Com97, RÅ99, WL96, Zon98b]. Limiting [Jen97, PGS98]. limits [Håk96a, Wis92a]. limnological [And95]. Limnology [Str95, Jør94a, Jør95h, Kra95a]. line [Cra98]. linear [BH99, Day91, Pat97, SH91, SBT97, TKG97]. linkages [JCS^{+}97]. linked [Kir99]. Linking [HMC99, LRS^{+}93]. litter [VVK97, YH96]. littoral [BGTL99, HL96, Mau93]. load [MLB96, TH90]. loading [SS98, UOB94]. loads [dVPO^{+}95]. Loam [MK95, DNR95, Jar95]. local [GH97, HC97c, JB97, RSS91, Ume95, YH99]. location [SFGH97]. locomotion [Mal95]. locomotion-induced [Mal95]. locust [HC96]. logarithmic [Fle91]. logic [Hul98, ZBDN96]. Logical [BK92a]. Logical-information [BK92a]. Logistic [MD97, MBF96, HRV^{+}96, IT97, M.a93a, MDO99, PG98, Sil95b]. Logofet [Nie95a]. Loire [MFP99]. London [Jør91a, Jør97c, Jør97e, Jør97f, Leg93, Mar93, Mar99a, Nih91]. Long [DuB97b, SBD^{+}97, CH97, GFS99, LLF97, Mon98a, Ped98, PR95b, Pol96, RDFZ96, WK97a]. long-distance [RDFZ96]. Long-term [DuB97b, CH97, GFS99, Ped98, PR95b, Pol96, WK97a]. longevity [SKŞ95]. longicornis [SSW91]. Loss [FSC^{+}97, SMDSZSXF93]. lotic [CW95, CFC91, SBW^{+}99]. Lotka [GQB91, Jes99, KH94b, M.a93a, RSC96, Sil95b, iWSG96]. Louis [Svi91b]. Louisiana [CCK92]. Lovett [Nih93]. low [Fom97, FK99, GCT96, LV99, RSKS91, SLP93]. low-density [SLP93]. low-input [RSKS91]. lower [BM97b]. Ltd [HS94]. LTER [BL93]. Ludwig [Leg91b]. Lugo [Jør95d]. lumped [GBBV99, PJ99, Sil99]. lunar [Omo95]. lunatus [BM99]. luteum [Mar91]. Lymantria [SC96]. Lymantriidae [STTR93]. lysimeter [VGDvdZ97]. lysimeters [Sog94]. Lysina [KSD^{+}99].

M [Bos91b, HS94, Jør94a, Jør95b, Jør99b, Leg91a, Leg92, Str91, Wu95, FFR99]. M/V [FFR99]. MA [Kom97]. Maarel [Ano96p]. Maastricht [Jør94f]. Mac [Wu95]. machine [DGWK97]. machines [BGL99]. Mackenzie [Koh91]. MACRO [Jar95]. macroalgae [MMJN97, SPP^{+}97]. macroalgal [DF97]. macrophyte [CC96, Nie97, PU97]. macrophytes [AB97, CBS97, MGGSMC^{+}96, Shu98, WHA98]. Macroscopic [Kan91, Aok93]. made [Chi98b, SR94, UB98]. MAGE [GYYL96]. Magela [Var93]. magic [FWCJ95, Sog94]. magna [HE96]. magnesium [Tor96]. mainland [Hil96]. maintaining [GPT95]. maize [BSB99, EBA^{+}99, MMJ^{+}98, CKK^{+}98]. major [CMS90, WFJ96]. make [Rom96]. MAKEDEP [AWS98]. makers [WHJ^{+}92]. making [Rec94, WGP96]. malaria [Mor92, TSR97]. Male [Whi98, HS97b, RC93]. mammal [JD99, Sou99]. man [Chi98b, SR94, UB98]. man-made [Chi98b, SR94, UB98]. managed [WBH^{+}98]. Management [Kom94, Rip95, Bat96b, Bat96a, BHL96, BW94, CPH99, CHLT96, Got98, Gra98c, GHQ99, JCS^{+}97, Jør93a, Ker95, KG96, KUI94, KL99, LAB^{+}98, LW96, PMV96, PGD91, RSKS91, RB98, ST96, SCEZ^{+}94, Str94b, TRR96, TCC^{+}97, VK99, WCC95, XHW^{+}99, vdBDJ98, Bos93a, Jør94a, Jør97f]. managements [Gra94]. Managing [HWB96, Jol96, Vin96]. mangrove [Gra98d, RM95]. Mann [Ano91i]. map [Foo99, HMW96, Kan90, PG98]. maple [LB97a, RU97]. mapping [BD99, Met98]. maps [LAB95, LCL93, Mar92, ML92b, Van94]. March [Ano90i, Ano92-29, Ano93t, Ano94s, Ano96-60, Ano97-46, Ano98-41, Ano98-30, Ano99-33, Ano99-43, RM95]. Mardi [Jør97e]. marginal [vdPV99]. Margit [Vas94]. Mariculture [CPPB99]. marina [Bac93, BCB97, CB97]. marine [AL94, Ano91i, BLS98, CBB97, GMM91, HKV^{+}95, HW98, Jør97a, KVAW91, MGM92, Omo95, Per98, Svi91a, VMSP97]. MARIOLA [UDVESM^{+}95]. Markov [HC97a, MGM92, RPT93]. Markovian [CCR98, Li95, PS93a]. Marquardt [BMHC90]. marsh [CCK92, ÖM97b, TPS98]. marsh-breeding [ÖM97b]. marsupials [ML98]. Martinus [Jør94e]. Marvin [Jør94g]. Masked [FP93]. mass [HSU96, KS96, VMSP97]. mass-balance [VMSP97]. Massif [SM97]. massively [Smi91]. massively-parallel [Smi91]. Mastering [Suá99]. material [THR96, Ulr92]. materials [KS96]. mates [McC97]. Mathematical [AS94, CCKO94, Hug91, Il’96, Kra93, MAY^{+}98, SN95, SGS^{+}95, SR94, Str95, UDML97, ASTL91, BM94, Buw91, CC94, GPT95, GQB91, HS99, HGB98, JO95, Kra95a, Kra96, MGGSMC^{+}96, Nie95a, Per96, SC92, Svi91b, SLP93, WHA97, Jør91c, Str91]. Mathieson [Mau93]. Mating [MK97, ML98, Whi98]. Matrices [Nie95a]. Matrix [Osh91, CCR98, CHC91, GS97, GA97b, Jen95, Jen96a, KBV99, Lea97, LBV96, LB97a, SN99a, WH91]. matter [AL94, Ano96q, CK97, Han97b, LHAJ99, MMJ^{+}98, PAGL96, RGRF98]. Matthias [Jør94c]. maturation [Aok97]. Mature [Omo97]. maturity [Chr95, KJ94, PEAS99]. MAXIMS [RFB99]. maximum [BILL98, Dam99, Was95]. May [Ano90h, Ano91t, Ano94t, Ano95w, Ano95-32, Ano97-55, Ano98-36, Ano99-47, Jør93a, Jør99a, AOE99, Leg97b, PW93, Rux95]. McAleer [HS94]. mean [BL93, Hut91, Ken95, LA97]. meaning [Ryk96]. means [SEJ93]. measure [Aok93, BM97a, PEAS99]. measured [Ahn99]. Measurement [Jør92a, PH95]. measurements [CBC^{+}98, KW99, VF96a]. measures [GA96, Han97a, SV98]. Measuring [CMH99, McD93]. Mechanics [McD93, Mes92]. mechanism [CP96]. mechanisms [BP93, Nih93, WWJ99]. mechanistic [CC96, CHC91, Dua95, GGP97, LK99, SSM^{+}93, SRW97]. mecoprop [FK99]. media [LLM96]. mediated [DJ98, MR91b, Sey92, TSMC96]. Mediterranean [HCdVM98, AH99, Cha95, MS97, UDVESM^{+}95]. Meeting [Jør91a, Str95]. Megalurothrips [TBG93]. meiofauna [PBO^{+}96]. Meissner [Jør96a]. Mejillones [Mar97a]. Mekong [BM96]. melanoleuca [CALW99, iWSG96]. Melbourne [Jør97e]. mellifera [Mar98]. melodia [ST96]. memory [HRRP99]. menhaden [Got98]. Menten [LC97a]. Menten-Monod [LC97a]. mercury [Håk96c]. meromictic [PR95b]. mesoscale [GPW97, KMWB98, Rut93]. mesotrophic [BGTL99, ST96, TS93]. Messages [Str95]. Messara [VPFM97]. Meta [KD99, VK97]. meta-model [VK97]. Meta-modeling [KD99]. metabolic [HAOR97]. metal [MVBP93, PJ96]. metals [BH91, KCP^{+}98]. metapopulation [BH94, GH97, GYNE99, SL96, WVB93]. metapopulations [GAP93, LBA^{+}95, PG98]. meteorological [DDA99, VF96a]. meteorology [JTS92]. methane [Jam93]. methanogenesis [LV99]. Method [WW92, Akç91, DD̆Á97, Håk96a, HC97a, Kal99, KH94a, Löf96, MRCL92, MPP98, Oel95, PW91, PS90, PASG97, Rod99, RCD93, SWK93, SL91a, VNLB96, ZMH96]. methodology [GTF99, WRHW94]. methods [DA95, Jør96a, Jør97b, Kat90, Kra95a, Leg91b, LLF97, PL90, RB96, WQ97, YWMM98, vT95]. metric [UB97]. mexicana [KGD97]. Mexico [Jør97e, GHQ99, ML97, RSdP98]. Michael [Ano99e, Jør97b]. Michaelis [LC97a]. micro [BBL97]. micro-habitat [BBL97]. microalgal [Zon98b]. Microbial [LLM96, BKG97, FK99, Kai96, LG91, LLWP95, LJ94, SR94, VVK97]. microbial-detritivore [VVK97]. microcontaminants [HE96]. microcosm [VVK97]. microorganisms [CFC91, VV91]. microphytes [MGGSMC^{+}96]. micropus [KGD97]. Microsatellites [ALGB99]. microscale [DB98]. microsporidian [LSC95]. microtine [HH99]. middle [LL97, MFP99, Wis92b]. midge [Axe93]. migration [Cri94a, DR96, Dye95, GMM91, HC97b, Kas93, Lee91, MGM92, Mon98b, RSD94, RDFZ96]. Mikhail [Svi91a]. mildew [SvB93]. mill [SKW91]. mine [BH91, CSCD99]. mineralisation [Dup98, FK99]. mineralization [Fom97, Whi95]. minimal [SvB96]. minimise [HWB96]. Minimizing [DT97]. minimodels [Kan91]. Minimum [SR92]. Minshall [Jør97g]. minutum [SY90]. missing [MG97]. mite [BM99, CHLT96, Mar98, YAKS97]. mitis [SLP93, SL96]. Mitsch [Nie95b, Nie95c, Zuc91]. mixed [HKV^{+}95, Jõg98, JO95, Jol96, KVAW91, PT95, PR95c, SR95b]. mixed-crop [Jol96]. mixed-cropping [JO95]. mixed-layer [KVAW91]. mixing [SEJ93]. mixta [Gor98]. mixture [SR94]. mixtures [Wil91]. MOAB [CF99]. mobilising [Jør97h]. mobility [BWW99, PJ96]. Mobilizing [HV98]. mode [YKH98, YKH99]. Model [BFK96, GDS98, HS97a, Jør93d, KV90, LL96a, LRK96, SH97, ST99, SR95a, SPP^{+}97, TC98, ÅB90, AR96, AG93, AJK95, AYO99, ASTL91, AM98b, And96, Ano92f, Ano92g, Ano92-30, AL99, ACS96, AO97b, Bac93, BG93, BBN98, BZ95, BCH^{+}97, BH97, BBH92, Bak96, BM99, BMHC90, BLK^{+}99, Bar96b, BS98a, BBR99, BF98, Bel98, BK97, BS95a, BS95b, BS95c, BEB99, BS98b, BvSH91, BM94, Bon93, BSB99, BW96, BP93, BK91b, Bos96, Bot98, Bot95, BLS98, BH94, BHBM96, BILL98, Buw91, Bye96, Cac90, CW95, CT96, CPH99, CO99, CKK^{+}98, CFC91, Cha95, Cha99, CA98, CFS93, CC96, CHM96, CTL98, CLCG99, Che90, CK97, Chi98a, Cla99, CR97b, CWPD95, CHC91, CS97, CL90, CPB97, CTKW98, Cri94a, Cri94b, CP96, Dam99, DSNN96, De 98, DFP^{+}93, Del92]. model [DJ98, DCM92, DFHR96, Dua95, DF97, DPSB99, Dup98, Dur96, Dye95, EP93, EH91, ET95a, ET95b, EMHB99, Fer95, FWCJ95, FF91, FDC^{+}96, FV97, FK99, FRS^{+}95, FTH99, FOS95, Fri95, FSKC97, Gal90, GC91, GYYL96, Gar90, Gav95, GS96, GFS99, GG99, GCT96, GS98a, GU97, GPO97, GMI95, GGC98, Gob93, GA97b, GPT95, GBBV99, GF90, GMM91, Gra98d, GKvdSdV95, GHH^{+}96, GQB91, GYNE99, Håk96a, Håk96c, HKV^{+}95, HF93, HSU96, HSN96, HL96, HMA90, HRV^{+}96, HKMRS95, HMC99, HMB99, HB93, HS96, HMW96, HBM^{+}97, Hen95, HK91, HS99, Hil96, HCdVM98, HW98, Hju96, Hof95, HNCS92, HAOR97, HC97b, HWB96, HM97b, HSXP98, HCL96, HLF99, Hut91, IT97, Jac96, Jam93, JK99, Jar95, Jen95, Jen96b, JPA^{+}99, Jõg98, Jol96, JWS95, JLS98, KT97, Kas93]. model [KEAP97, KI93, Kei97, KD99, KS95, KW99, Ken95, KGD97, Ker95, KSR97, KG95, KB99, Kin95, Kin96, KG96, Kir99, KKNY91, KUI94, Kle97, KH94b, KHL97, KH98, KN95, KBV99, Kol98, Kom97, KTM95, KJ94, KSD^{+}99, Kra93, Kra96, Krä95b, KW95, KMWB98, KI98, KVAW91, KL98b, Lav95, Lee91, LA91, LAB^{+}98, LC97a, LC97b, LR97, LSB99, LRM93, Li95, LVH96, LWMT97, LVS^{+}99, LBV96, Lin98, LHAJ99, LK99, Liu93b, LLWP95, Liu97, LA98, Loe97a, LGSF92, LCH95, LB97b, MRSW93, MWY92, Mal96, MS98, Mar97a, Mar98, MG98, MGNR91, MVBP93, MGM92, MRL94, McC92, McC95a, MS97, MJ99, MDT93, MJGW94, MU99, MWK95, MvdV95, Mon98a, Mon98b, MVS^{+}95, MN93, MRSK96, MG97, MGGSMC^{+}96, MJMN97, MCT97, MP99, MDH96, Nal97, Nie95d, NMS95, ND97, NHWS91, OM97a]. model [Oel95, OZIA96, OA97, OR99, OM90, Osh91, OH97, OA98, ÖM97b, PA96, PW97, PR95a, PT95, PYE94, PJ96, PL96, PU97, PSK98, PS90, Par96, Pat97, Per98, PSN98, PMV96, Pie95, PS93b, PR95b, PD99, PJ99, Pow96, PC99b, PSC93, PMP97, RSKS91, RS93, Rec93, RS96, RY91, RGBM90, RSD94, RFB99, RGRF98, Rip95, RF97, RL99, RDFZ96, RRF91, RÅ99, RSdP98, RCD93, RS98a, RCD98, ST96, Sal97, SJ96, Sam96, SRB99, SR93, SC92, SH99, SPE95, Sch95, Sch96b, SG97, SS96, SH92, SM95, SNUM91, SIU97, SSW91, Sem99, Sen90, SN95, SN99a, SSM^{+}93, SCEZ^{+}94, Sey92, SC96, SH91, SMDSZSXF93, Sil95b, SR98, ST92, SPH97, SY90, Smi91, Smi97a, Smi97b, Sog94, SA97, SMP96, SFS97, Spe97, SF94, SHEM95, SRW97, SFGF94, SDS98, Stu98]. model [SKW91, SGS93, SHJ95, SvB96, SvB97, SvB98, SLP93, Szy96, TB93, TS96, TGM95, TC96, TTT98, Tay96, TS93, TKO91, TD94, TRR96, Tre91, TCC^{+}97, TWRW93, TDM96, Uch99, UDVESM^{+}95, UDML97, Val99, VEA99, VMB95, VU93a, VU93b, VVK97, VJR^{+}98, VVDF95, VP93, Van97, Var93, VKT94, VK97, VF96c, VF96b, VZF96, VBL90, VGDvdZ97, WG98, WU96, WG91b, WW92, Whi99, WJW99, WVCS94, Wil99, Wil96b, WFK^{+}91, WBH^{+}98, Wol94, Woo98, WHA97, WV91, iWSG96, iWLSL96, WTL91, XHW^{+}99, XY95, YH93, YAKS97, YH96, cZdZ96, dZcZ96, Ziv98, ZH97, ZH98, Zon98a, vD95, vDvB95, vOÅ95, vT95, vdBDJ98, vdHvEH96, vdPV99, Bac93, BS95a, Ces97a, Ces97b, ET95a, ET95b]. Model-based [LL96a]. Model-Theoretical [Jør93d]. modeled [BMHT99]. Modeling [AOD97, BW94, BH91, CWC95, DNR95, Ema96, EBA^{+}99, FFR99, HTW96, IPS99, JA96, LUC^{+}93a, MT93, MT95, MT96, PG99a, Ped98, PBO^{+}96, RBM97, Sch96a, Str91, Ume95, VKT94, Vin96, VS96b, VS96c, XJTL99, Ahn99, BD99, BHL96, BLRV98, CC94, Gao96, GET98, GPG^{+}99, Gor98, HM94, JRB97, Jør91c, KD99, LAB95, LJ94, LFM97, Mah98a, MAY^{+}98, PSL^{+}93, PRK97, Rec94, RCD93, SABG93, Swa96, TM99, WFB96, Wu95, WL97, YS98, Jør92a, Hug91]. modelled [vWB99]. modellers [Ano93s]. Modelling [AH98, AL94, Ano93i, Ant98, AB97, BJ94, BGF99, BBKV96, Bel92, BK92b, BSL97, BCD93, BU98, BK91a, BCB97, Boc99, BBL97, BPTM^{+}97, Ces97a, Ces97b, CBS97, CCK92, CG98, DS99, DLN^{+}95, DH99, Der98, Dub97a, EGJ95, ESD98, FKN97, Fom97, GGJ96, GSS94, HH99, HE96, HP94, Hon92, Hua98, JCS^{+}97, Jes99, JPA^{+}99, JAZ^{+}99, Jør97i, JMA97, Kai96, KABT90, KV97, KIPK99, KMS99, Kle95, Koh91, Kom97, KS96, LW96, LS99, LP91, LL96c, M.a99, MMJN97, McD93, MCTB96, MM97, MG96b, MR91a, MI95, ML97, MPA96, Nie94, Oga96, Osh96, PA99, PAGL96, PD99, Pol96, PJL^{+}96, PASG97, RPT93, Rob96b, RRNM97, RS91, SML98, SBW^{+}99, SKŞ95, Sel95, Shu98, SPP^{+}97, SDBA98, Ste95a, Str94a, Str95, SSK^{+}91, Str96, TTT98, Tho97, Tis97]. Modelling [TDK98, TN98, TPS98, Tuz91, TAB^{+}93, UMJ^{+}97, US97, VCGG96, WL96, WCDL96, WM95, WHN96, Whi95, WCC95, Wil91, Wis92b, WK97b, Wu90, dVPO^{+}95, dVKvdS95, vdSdVK96, Alb92a, And95, Ano90e, Ano91k, Ano91m, Ano91l, Ano91n, Ano92k, Ano93h, Ano93j, Ano94h, Ano94i, Ano95i, AH99, BK92a, BL90, Bre95, BKG97, BH96, CV99, CBC^{+}98, CSCD99, CM91, CMS90, DD̆Á97, DR96, DL99, Dun99, EFCM99, EB93, EFGE98, Foo96, Fri98, Gam98, GLD97, GS97, GWAU99, Gri99a, GS98b, Håk97, HAW97, IV91, Jør90, Jør92b, Jør93d, JNM95, Jør95h, Jør97d, Jør97b, JD99, Jør99c, Jør99d, KH98, Kra95a, KCP^{+}98, KBSC98, Leg97a, LDB^{+}96, LG99, LK91, LBA^{+}95, LSB96, LS98b, MJY96, MDB98, MLB96, Mar97b, MC93b, MCTB95, MFP99, MN93, Nie92, NF96]. modelling [Nor96, NNR98, OYA95, ÖÖ99, PFP^{+}95, PS93a, Pay99, PP95, RPH99, RC97, RCS97, RKM97, RFHY97, RB99, Rob96a, RPM^{+}98, SBD^{+}95, SGML91, SD97, Sep99, Sil93a, Sil93b, SS98, Sol99, SVK93, TZE96, TGM^{+}99, Tor96, Ulg95, VdKK^{+}99, VWN^{+}99, VB99, Wae93, WFJ96, Wen92, WGP96, Wis92a, WHJ^{+}92, WRHW94, Zon98b, ANM99, AFM99, ANFM99, Jør97i, TBG93, Ano93k, Ano93l, Ano97k, Ano97l, Ulg95, Ano92l, Ano94j, Kom94, Ano90d, Ano91j, Ano92h, Ano92i, Ano92j, Ano94g, HS94, Kom97, Nih91]. Models [AUS96, Ano95w, DCP99, Håk95a, HC96, JN94, SMWF98, TSR97, AR97a, AR97b, AM98a, AWS98, Aly91, AL91, AS94, BJM92, BBK98, BGF99, BK98, Bir94, Bla97, BJ95, BILL98, BK96, BFK96, Bye91, CF99, Cas96, Chi97, CCR98, CP92, Cle98, DHRLE90, DSKM95, DCCSL99, DGWK97, Ebe93, EBBW98, Gam98, Gon97, GT97, GGP97, Håk95b, Håk96b, HL97, Håk97, Håk99, HC97a, HM97a, HS92c, HSHS96, HR96, HGB98, JCC^{+}95, JHM97, JH95, Jet92, Jør92c, Jør97a, Jør99d, KH94a, KB97a, KdH99, KLFH96, Koh91, Kom97, Lea97, LCMR96, Leg97b, LADL99, LRS^{+}93, Lho94, LB97a, Log97, Łom99, LS98b, Luc93b, LCH95, LRK96, M.a93a, Mah98b, MC93a, MD97, ML92a, Mar97b, McC97, MKA^{+}95, MG96b, MHB^{+}96, MRLC96, MPA97, NN95]. models [NJ96, NZ92, Par93a, PG98, PT97, PG97, PG99b, PS93a, Per93, Per96, Phi95, PM96, PL90, Pow93, RSS91, RLHD99, RO97, RA97, RM92, RM93, RJN99, RRNM97, Rux95, RS98b, Ryk96, Sal92b, Sch96a, SvB93, SRV95, SOM97, SGS^{+}95, Sil99, Sog93, SBT97, SCM93, SRW97, Sto92, Str94b, SWK93, SY99, Svi99, TvG95, TKG97, VS97, Van97, VK97, VMSP97, VW97, VUDM^{+}99, VFC98, VVC99, Wan95, Web95, WCDL96, Whi98, WE96, YH99, Zei91, vHY95, Jør93c, vGDT95, Jør93b]. modes [HPB93]. Modification [MRLC96]. modified [CKK^{+}98, Lea97, iWSG96, vdHvEH96]. modular [KH96, MC97, Sil93a, TP97]. Modularity [RA97, AR97b, Der98]. Mohren [Jør99b]. moisture [RRNM97]. molecular [Kir92]. molecular-genetic [Kir92]. molting [MJGW94]. Mondego [FKNM^{+}97]. mongolica [WK97b]. Monitoring [UB98, EFGE98, TBvH95]. monkey [SLP93, SL96]. monocarpic [SN99a]. monocultures [Dam99, HMW96, WWJ99]. monoculus [FG94]. monocyclic [MHJ96]. Monod [LC97a]. Monographs [Bos91b]. monomictic [MVS^{+}95]. monoxide [TZE96]. monsoonal [MJY96]. Monte [Ann99, Ano99e, Ann97, DCM92, HC97a, NCNN97, Yoo99]. Monthly [Aok90, SD97, Yin96]. morphogenic [Phi95]. morphogenic/pedogenic [Phi95]. morphology [BKG97]. morphometric [HL97]. Mortality [CJMT96, FFR99, Jen94, KL98a, Mal96, Ped98, RCD93, RB98, SNU96]. mosaic [Coq95, MRL94, Wis92b]. mosaics [CBS97, LA98]. mosquito [RM95]. mosquitoes [LSC95]. mosses [Mar91]. most [LL97, Osh96]. Mostafa [Jør94e]. moth [Ned99, STTR93, SC96, WVCS94, Wil99]. moths [CWC95, DFHR96, RDFZ96]. motivated [RF97]. mountain [CBC^{+}98, PL90, WBH^{+}98]. Mountains [Pat97]. mouse [BS98b]. Movement [RLHD99, BWW99, Bla97, HBP93, KS91, ML98, Sou99, TIH98, VGDvdZ97, WCS95]. Movements [Jør94e, Tis97]. moves [Swa96]. moving [PL96]. MS [Leg91b]. MS-DOS [Leg91b]. mucilage [Tom96]. muelleri [BBR99]. Multi [BCH^{+}97, AJK95, HM94, HLG96, Jol96, KBV99, LBV96, MS95, PSK98, VF96c, WRHW94]. Multi-dimensional [BCH^{+}97]. multi-factor [WRHW94]. multi-lake [AJK95]. multi-patchy [MS95]. multi-regional [Jol96]. multi-scale [VF96c]. multi-species [HM94, HLG96, KBV99, LBV96]. multi-step [PSK98]. multiconstituent [PL96]. multiple [FZ99, GU97, LC97a, VK97, WFK^{+}91]. Multiresolution [JMPT99]. multiscale [LW94]. multivariable [PYE94]. Multivariate [Kle97, BC90, MRCL92, MEBZ96]. Murai [Ano96r]. Murray [MDB98]. mussel [HS99, TN98]. mussels [LD97, PACS96]. mutual [TKO91]. mutualist [FF91]. Myrica [SSY95]. Mysis [Gor98]. Mytilus [HS99]. myxomatosis [Sey92].

N [KBL96, MJMN97, MMJ^{+}98, RRNM97, SPE95, SA97]. N-SIM [SPE95]. name [Jør93e]. NAP [vOÅ95]. natal [McC99, SLP93]. National [SD96, Shu98, SvdG99, Ush91]. native [FG94]. natural [ALGB99, Boc99, BK91b, BLRV98, BC96, CU93, Col97, DDA99, Gra98c, Kir94, SR94, SR95b, TTT98, UB98, VK99, Wil96a, Jør95d]. NDVI [LRK96]. near [ML97, PT99, Smi95a]. necked [BS98b]. need [RS98b]. Needed [UB97]. needs [Lin98]. negentropy [DFP^{+}93]. Negev [WJW99]. neglect [Rux95]. neighborhood [Bon93]. neighbourhood [Gri99b]. nematode [FP93, LVH96]. nematodes [ML97]. Neotoma [KGD97]. nested [CA98]. nesting [HTW96, Sen90, Wol94]. nests [PT99]. Net [AYO99, AO97a, BD99, Han91, HLF99, JPA^{+}99, JAZ^{+}99, KMS99, PA99, Val99, WK97b]. Netherlands [Bos91b, Jør94c, Mit94, EBBW98, JA96, Jør99b, vdSdVK96]. Network [FP98, HPB93, HBP93, Pat95, Ano94k, Aok92, BHWP91, CO99, CPMC96, CP92, CL99, DCCSL99, FP99, Foo99, GTF99, GGP97, HW98, HC97c, Leg92, LSB99, ÖÖ99, PWU93, RFHY97, SH99, TS96, VWN^{+}99, Ano91i]. networks [ALGB99, BGF99, BGTL99, CFC^{+}96, DL99, FZ99, JHM95, LLM99, LDB^{+}96, LG99, LKS96, MDB98, MDO99, MFP99, Mor99, Nak92, PT97, PHB90, PH95, SBW^{+}99, Svi97a, TAM^{+}99, THR96, vWB99]. Neuenkirchen [Smi95a]. Neural [DCCSL99, VWN^{+}99, ALGB99, BGF99, BGTL99, CPMC96, DL99, FZ99, Foo99, GTF99, GGP97, HC97c, LLM99, LDB^{+}96, LG99, LKS96, MDB98, MDO99, MFP99, NNR98, ÖÖ99, PT97, RFHY97, SH99, SBW^{+}99, TS96, TAM^{+}99, THR96, vWB99]. neuro [THR96]. neuro-fuzzy [THR96]. neutral [Van96]. Nevada [MU99]. Newfoundland [HLF99]. newly [MM97]. Newman [Leg91a]. NICCCE [vD95]. NICCE [vDvB95]. niche [Cao95, Han95, MWY92, Wan95, ZH97]. niche-fitness [ZH97]. niches [PR95b]. Nienhuis [Mau93]. Nigeria [Osh91]. Nigerian [Osh96]. Nihoul [Del92]. Nijhoff [Jør94e]. Nile [Boc99]. niloticus [Boc99, FG94]. nine [RRNM97]. nitrate [BMV99, Par93a, WTL91]. nitrification [HSN96]. Nitrogen [CFC^{+}96, LVS^{+}99, AOD97, BSE95, BSB99, Cou91, EGJ95, FOS95, Gra95, Gra98b, HSN96, HB93, HH98, HW98, Hju96, HCL94, Jør95f, KEAP97, KBL96, Ker95, MR97, MLGV99, MPA96, NAFM99, PRK97, RSKS91, RRJ90, RGRF98, RÅ99, SPE95, SM95, SS98, SHJ95, VVK97, VB99, Whi95, WMRK93, WBH^{+}98, WHJ^{+}92, vDvB95, vdPV99]. nitrogen-tillage-residue-management [RSKS91]. NJ [Sil95a]. No [Jør93b]. Noctuidae [NHWS91]. noises [Sam96]. Non [Bye91, PS93a, BH99, KS91, dLNT98, Rux95, Tap93, WCS95]. Non-autonomous [Bye91]. non-homogeneous [Tap93]. non-linear [BH99]. non-preferred [KS91]. non-random [WCS95]. non-spatial [Rux95]. non-timber [dLNT98]. nonbreeding [MGCR98]. nonhomogeneous [Li95]. Nonlinear [KB97a, CTDC94, CDDC96, LDB^{+}96, Mal95]. nonliving [Ano96q]. nonoverlapping [Jen94]. nonstationary [Pat97]. Noord [EBBW98]. Noord-Holland [EBBW98]. Nordmoen [Sog94]. north [AH99, FSA99, MKA^{+}95, SPE95, SM95, GGJ96, HCdVM98, Jes99, MWK95, SFGF94].

North-America [Jes99]. north-central [FSA99]. north-western [AH99, HCdVM98]. northeastern [ADF^{+}93, GHQ99, LH92]. northern [CO99, DT97, DuB97b, KBV99, LBV96, OA97, Spe94, Tom96, Vir92, BSN^{+}98, Rya94]. Northwest [BMHT99]. Norway [BTvH95, Ces97b, FWCJ95, Jõg98, Sog94, US97]. Nosema [OM90]. Nostrand [Nie95c]. note [Ano91-29, Day91, GC91]. Notes [Ano91i]. November [Ano91x, Ano92-28, Ano93u, Ano95-37, Ano96-59, Ano96-55, Ano96-51, Ano97-39, Ano97-49, Ano98-35, Ano99-37, Ano99-48, Jør91a, Jør99b]. NPP [AOE99, PA99, VB99]. NTRM [RSKS91]. nuclear [SADZ95]. nuclide [SADZ95]. NuCM [SA97]. NUCSAM [GKvdSdV95]. null [JCC^{+}95]. number [KJ94, RCD93]. Numeracy [Ano99c]. Numerical [CV99, CSCD99, KUI94, Bot98, Coo95, OZIA96, Por98, VDV95, YWMM98]. Nutrient [BvSH91, Got98, BCB97, Cha95, Com97, DLN^{+}95, DH99, GS98a, KH96, Kir99, Leg93, MR91a, MI95, OKS99, Rec93, SA97, dVKvdS95]. nutrient-cycling [Com97]. nutrient-limited [Com97]. nutrient-periphyton [DLN^{+}95, DH99]. nutrients [HKMRS95, LC97a, ND97].

O [Jør97h, Jør97e, Nie95a, KBL96]. oats [Gra94]. Object [DR96, HAOR97, LC97b, LS98b, SSS^{+}91, Sil93b, AR97a, BL92, BF98, CR97a, Der98, Dur96, Fer95, GGZ99, HMB99, Lav95, Lav96, MRSW93, MC93a, MB96, SNUM91, SSM^{+}93, SOM97, Ziv98]. Object-oriented [DR96, HAOR97, LC97b, LS98b, SSS^{+}91, Sil93b, AR97a, BL92, CR97a, Der98, Fer95, GGZ99, HMB99, Lav95, Lav96, MRSW93, MC93a, MB96, SNUM91, SSM^{+}93, SOM97, Ziv98]. objective [AM98a, BZ95]. objectives [VKT94]. Observations [Szy96]. observe [BWW99]. observed [ALB91]. occupancy [BGTL99]. occurrence [DB98]. ocean [GTF99, BMHT99]. oceanic [AKH99, Kle95]. October [Ano91w, Ano91-28, Ano92w, Ano93-29, Ano95-35, Ano97-54, Ano98-37, Ano98-46, Ano99e, Ano99-35, Ano99-46]. Odiel [PAGL96]. odour [OM97a]. odour-baited [OM97a]. Oedaleus [HC97b]. off [FFR99, Gra98d]. offs [BSN^{+}98, BPN99]. offtake [HLG96]. Ohio [CMS90]. oil [CMS90, Del92, FFR99, Jør91a]. oilseed [MJMN97]. Okeechobee [Smi97b]. Old [Nie95b, CFS93, Mow97]. old-growth [CFS93, Mow97]. oleracea [Yam97]. oligotrophic [Ula95]. One [Coo95, MA96b, CWC95, CTKW98, Mar92]. One-dimensional [Coo95, MA96b, CWC95, CTKW98, Mar92]. Ontario [BW96, Gob93, HSU96, Jac96, LB97b]. Open [Ano93s, DS99, GG92]. open-access [GG92]. Openness [JPS99]. operation [Jør94b]. operational [DT97]. Operations [Jør97f]. operator [MA93b]. optical [EAG97, Sem99]. Optimal [FG92, Håk95b, HS91, Vir92, Bat98, BGL99, Gra98d, HH98, HF96, Jen96a, LA91, MG96a, Par93b, SDS98, SR92, TAB^{+}93]. Optimality [Yam97]. optimised [VEA99]. Optimising [HLG96]. Optimization [BMHC90, KJ94, Nie95d, DuB97b, Gra98d, HSB99, KL98b, NG96, Nie97, Wil96a]. optimize [SCEZ^{+}94]. optimum [Sep99]. option [BW94]. options [Col98, XHW^{+}99, Jør97h]. Opus [BMV99, Smi95a]. orchid [HSB99]. orchids [KB99]. order [ALB91, ALB92b, HRWC91, Rod99]. ordered [MDT93]. ordering [Rux96, SB97]. ordinary [Vol98]. ordination [Jør94b]. Oreochromis [Boc99, FG94]. Øresund [BCB97]. organ [SSS^{+}91]. organic [Ano96q, BHBM96, CFC91, CK97, FK99, Gob93, Håk95a, MMJ^{+}98, RRJ90]. organic-matter [MMJ^{+}98]. organics [LG91]. organising [Foo99]. organism [RSD94, VEA99]. organisms [MvdZ96]. organization [BM97a, BU98, KB97b, Mau90, MC93b, PW93]. organized [Jør91a, RAM99]. Organizing [CL99, JMN98]. orientated [BF98]. orientation [KB97b]. oriented [AR97a, BL92, CR97a, Der98, DR96, Fer95, GGZ99, HMB99, HAOR97, JH95, Lav95, Lav96, LC97b, Lho94, LS98b, MRSW93, MC93a, MB96, SNUM91, SSS^{+}91, SSM^{+}93, SOM97, Sil93b, Sog93, Wol94, Ziv98]. Orta [RMT92, PT95]. orthophosphate [MCTB96]. Oryza [AFM99]. Osbeck [MT95, MT96]. Oscar [Str92]. Oscillating [JWEB92, FV97, GSG95]. oscillations [GS98a, Van94]. Oscillatoria [MVS^{+}95]. Oscillatory [SR93]. Osenberg [Jør97c]. OSIRIS [MB96]. other [GPT95]. otolith [SG97]. otters [All91]. outbreak [RGBM90]. outbreaks [Ned97]. outlier [TD94]. outline [PW93]. output [Omo97, PH95, TN98]. overburden [CSCD99]. overcrowding [FG94]. overexploited [GG92]. overparameterized [RO97]. Overview [BHBM96, Ano90m, EB93, SB97, SSH98]. owl [Kit97]. Oxford [Jør95a, Jør97b, Jør97g]. oxides [Jør95f, WHJ^{+}92]. oxygen [ANFM99, CPH99, Coo95, CP96, FS97, GH94, Kal99, PL95, SF94, Tun90, VU93a, VU93b]. oyster [Whi99]. ozone [CTL98, Jør91b, ML96, WK97b].

P [Bos93a, Jør94f, Leg92, Mau93, Jør92a]. Pacific [Jør97e, MJGW94]. package [Bou95]. pages [Ano99c, Ano99b, Ano99d, Ano99e, Ano99f, Jør97d, Jør99b, Ano90l, Ano90j, Ano90k, Ano90i, Ano90h, Ano91-27, Ano91u, Ano91y, Ano91z, Ano91v, Ano91t, Ano91x, Ano91w, Ano91-28, Ano92z, Ano92-27, Ano92x, Ano92y, Ano92u, Ano92-29, Ano92-28, Ano92w, Ano92v, Ano93z, Ano93y, Ano93-28, Ano93-27, Ano93v, Ano93x, Ano93t, Ano93u, Ano93-29, Ano93w, Ano94w, Ano94v, Ano94y, Ano94u, Ano94x, Ano94s, Ano94t, Ano95-36, Ano95-38, Ano95-39, Ano95-33, Ano95-34, Ano95-40, Ano95-32, Ano95-37, Ano95-35, Ano95-31, Ano96-52, Ano96-53, Ano96-54, Ano96-58, Ano96-56, Ano96-57, Ano96-60, Ano96-59, Ano96-55, Ano96-51, Ano96-50, Ano97-44, Ano97-53, Ano97-43, Ano97-52, Ano97-45, Ano97-40, Ano97-50, Ano97-42, Ano97-41, Ano97-51, Ano97-56, Ano97-47, Ano97-48, Ano97-46, Ano97-55, Ano97-39, Ano97-49, Ano97-54, Ano98-39, Ano98-31, Ano98-33]. Pages [Ano98-32, Ano98-43, Ano98-34, Ano98-42, Ano98-45, Ano98-29, Ano98-38, Ano98-44, Ano98-41, Ano98-30, Ano98-36, Ano98-35, Ano98-37, Ano98-46, Ano98-40, Ano99-36, Ano99-38, Ano99-49, Ano99-41, Ano99-52, Ano99-34, Ano99-44, Ano99-45, Ano99-40, Ano99-51, Ano99-39, Ano99-50, Ano99-33, Ano99-43, Ano99-47, Ano99-37, Ano99-48, Ano99-35, Ano99-46, Ano99-32, Ano99-42, Jør97c, Jør97h, Jør97e, Jør97f, Jør97g]. paleoclimatic [ET95a, ET95b]. Pallas [Pat97]. Pälvinen [Ano99e]. panda [CALW99, iWSG96]. paper [Ann99, Ano99d, Ano99f, Leg93, SKW91, Yoo99]. Paperback [Ush91, Bos93a, Leg91b]. papers [Ano91a, Ano93-30]. parable [BL96]. Paradise [TPS98]. paradox [Gon97]. parallel [CM91, Hae91, PFP^{+}95, Smi91, VNLB96, WSGR96]. Parameter [ML92a, Ann97, Ann99, DPSB99, Els92, KB97a, LB97a, LCH95, PD99, WG98, Yoo99]. Parameterisation [BILL98, Fri98]. parameterizations [Cas96]. Parameters [Jør92e, BWW99, DDA99, DA95, GRL92, GFS99, GH94, HL97, KN95, KRD96, MJ99, MLGV99, PJ99, RKM97, SJ96, SV96, SPP^{+}97, SRW97, iWLSL96]. parametric [OZIA96]. parasite [BLS98, JCC^{+}95, MGD98, PGD91, SY90, TMG96, TMGS97, WS97]. parasites [KR92]. parasitism [LRM93, MG97]. parasitoid [Axe94, BL92, MG96b, MRSK96]. Pareloup [ST96, TS93, Sal97]. Paris [Jør97e, Jør99a]. Park [SD96, Shu98]. Parker [Ano99d]. Part [CG98, SH97, CO99, GSS94, Sil95a, SM97, HS97a, ŠŠZ90]. part-whole [Sil95a]. parthenogens [ACS96]. partially [IK95, Mor99]. particulate [LHAJ99]. partition [Håk97]. partitioning [Ben99, HH98, KB99, VS96c]. pass [DBB94]. passage [PH95]. past [And95]. pastoral [HSPM96, WHN96]. pasture [RGRF98]. pastures [BPN99]. Patagonia [vdBDJ98]. Patch [LRM93, IK95, LA97, RA96, SL96, WVB93, WL97]. patch-based [WL97]. patches [CBS97, DB98, MS95]. patchiness [LR97]. patchy [Liu93a, MS95]. path [And96, Whi99]. pathogen [LSC95, Loe97b, OM90, YAKS97]. Patten [Sil95a]. Pattern [PT99, Van97, Bak96, Day90, JAZ^{+}99, KK97, LS98a, Mal95, MUM98, MU99, ST99, Svi99, TGM^{+}99, VdKK^{+}99, YKH98, YKH99, Mar99a, Mit94]. Patternizing [CPMC96]. Patterns [Nih93, PW91, WFB96, ACS96, BWW99, Cha99, Day91, JWEB92, JMPT99, Jør97e, LTMP97, MEBZ96, PBO^{+}96, SEJ93, Sou99, TTT98, UB98, Ziv98]. Patton [Jør95g]. Patuxent [VVC99]. pawpaw [dLNT98]. PB [Ulg95, Mar99a]. PC [Kom97]. PCB [Jac96, SRW97]. PCBs [JRB97]. PCLake [AJK95]. pea [RJN99]. peach [RL99, SC92]. pedogenic [Phi95]. pelagic [GGC98, Jac96, Lav95, Lav96, PACS96]. Penaeus [MT93, MT95, MT96]. penicillatus [MT93]. Penning [Bos91b]. Penumbra [Ste95b]. perch [TAB^{+}93]. percolation [LLM96]. perennial [CHM96, MHJ96, SCH92, SN99a]. perennial-species [SCH92]. Performance [SRV95, WQ97, EBBW98, Jør92a, KB99, Loe97a, Mah98b, Met98]. performances [BG93]. Pergamon [Jør95a]. Period [RS93, Van97]. Period-doubling [RS93]. periodic [AK93, Ebe93]. periodically [SN99a]. periods [Jen94, LLF97]. periphyton [DLN^{+}95, DH99, SML98]. permafrost [Wae93]. Permanent [MMRR92, DCCSL99, vdB98]. permutation [RS98a]. persimilis [BM99]. persistence [GS97, HL96, Jar95, LGM99, MDH96, Sch96a, SN99a, SL96, Uch99]. personatus [KKNY91]. perspective [BF98, Gra98c, McK96, SG98, Leg92]. perspectives [BM96, VWN^{+}99]. Perturbation [Str92, CMH99]. perturbations [SY99, vdB98]. perturbed [RM93]. Peruvian [Kra96]. pest [DCP99, Hon92, LW96, Ned99, RB98]. Pesticide [Per96, DNR95, MA96b, PS90]. pesticides [Fom97, IV91, JMA97, PMP97, SNU96, SMDSZSXF93, VBL90, VGDvdZ97]. pests [MG96b, WCC95]. Peter [Ano99e, Jør95a]. Peters [Jør96a, Kom97]. Peterson [Ano99d]. Petroleum [Jør91a]. PGEN [Fri95]. pH [MFP99, OH97, PT95]. Phalacrocorax [HE96]. pharmacokinetics [RH96a]. phase [HC96, LCL93, Ned99, RB99, iWHS97]. Phaseolus [BM99]. phases [LLWP95]. Phenodynamics [RL99]. phenology [LRK96, RL99]. Phenomenological [KK97, Alb92a]. phenotypic [DH95]. phenotypically [PR95c]. Philippines [Boc99]. Phillip [BBR99]. philosophy [ŠŠZ90]. phoeniceus [ÖM97b]. phosphorus [Ahn99, Kim94, TVS^{+}94, TPS98, vdPV99]. photochemical [WHJ^{+}92]. photoinhibition [EP93]. photoproduction [Tro90]. photosynthesis [Bel92, CCKO94, CC96, CLCG99, Col97, EP93, Fri95, Han91, HH98, Kir92, Liu96, NMS95, Ste95b, Sto92, TC98]. photosynthetic [BCD93, HH98, IT97, WK97b]. Phrynocephalus [HF93, XY95]. phycotoxin [Bel98]. physical [ADF^{+}93, HW98, Mar99b]. physico [HSXP98]. physico-biological [HSXP98]. physiological [Col97, Kre91]. Physiologically [LGSF92, GYNE99, RF97]. physiologically-based [GYNE99, RF97]. Physiology [Jør99b]. phytoplankton [Bel92, CV99, Col97, JK99, LC97a, Mos92, ND97, Par93a, PR95b, SJ96, SH99, SBT97, Ste95a, TDK98, VCGG96, WE96, Zon98a]. phytoplankton-zooplankton [ND97]. Phytoseiidae [BM99]. Phytoseiulus [BM99]. phytotoxicity [MM96]. Picaninny [PJL^{+}96]. Picoplankton [HKV^{+}95, KVAW91]. pigs [Hon92]. Pinder [Jør96a]. pine [Che90, CG93, GS96, KABT90, KS95, KIPK99, Lea97, LRG98, Ned99, PL90, VS96c]. pini [Ned99]. Pinogram [Ano92g, Ano92s, Ano92-30, Ano92f]. Pinus [CJMT96, IPS99, KABT90, Kin95]. pipe [Chi98a]. Pisum [RJN99]. placement [HF96, HSB99]. plain [TM99]. Plaines [AK96]. plains [ASTL91, GHQ99]. planci [RGBM90]. planetary [Bot98]. planktivores [EAG97]. planktivory [LK99]. plankton [Gae92, Gon97]. planktonic [MDT93]. Planning [Ano99e, Ano96r, Jol96]. Plant [LFM97, AR97b, BSE95, Bon93, BKG97, CA98, CR97a, CCR98, CH97, Dam99, Gal90, GG99, GS97, Gri99b, HMW96, HR96, HSB99, Hug91, HCL96, JHM97, KH96, LW96, MJMN97, RKM97, RA97, SR93, Sch96b, SRV95, SN99a, SSS^{+}91, SSM^{+}93, SCEZ^{+}94, ST92, TC98, Vir92, WL96, Wil91, WK97a, Wu90, YKH98, YH99]. plant-herbivore [SR93]. plant-hervibore [Vir92]. plant/soil [BKG97]. plantaginea [DBB94]. plantation [BS98a, CG93, Lea97, SRB99]. plantations [CJMT96, Kin95, Kin96]. PLANTGRO [HV98]. plants [Ben99, Bon93, CWPD95, DJ98, HH98, KS91, LGM99, NMS95, SDS98, WGP96]. plasma [SSK^{+}91]. Plateau [FSG96]. plateaus [LC91]. platforms [ML97]. Playing [dLNT98]. Plenum [Jør91b]. plot [BEB99, BMV99, Smi95a]. plot-based [BEB99]. plume [HS92c]. PNET [PDAS95, KSD^{+}99]. PnET-BGC [KSD^{+}99]. PnET-BGC/CHESS [KSD^{+}99]. PNET-CN [PDAS95]. PNET-CN/CHESS [PDAS95]. pod [Axe93]. point [CBS97, Day91, WQ97]. points [LL97, SS96]. poisoning [Hon92]. policies [MG96a]. policy [Jør94b, Jør94g]. Polis [Jør97e]. poll [US97]. pollen [DJ98]. pollination [DJ98]. pollutants [KCP^{+}98, PT94, SSK^{+}91]. pollution [AL99, BJM92, Bas98, FG92, Jør91b, Kre91, MVBP93, McD93, NNR98]. POLMOD.PEST [PMP97]. polychlorinated [LB97b]. polyculture [FG94]. Polymorphism [ACS96]. pomi [WH91]. pond [CP96, HM94, MCTB95, MCTB96, Szy96]. ponderosae [BBL97]. ponds [FG94, GH94, LP91, MGGSMC^{+}96]. pondweed [Spe94]. pool [HAOR97]. Population [EH91, GSG95, Gra98a, KD96, Mat99, ML98, RA96, WH91, All90, ANM99, ALS97, BSV91, Bar96b, BDD97, BS98b, CALW99, CHC91, CC94, CD99, DCP99, DDA99, DF97, ESD98, Fah98, FP93, Gam98, GW98, GRL92, GS97, GA97b, HTW96, HBM^{+}97, HS99, Jen91, Jen95, Jes98, Jet92, Jõg98, JB97, Kar93, Kit97, KH94b, KS91, KHL97, KJ94, LGM99, Liu93a, Łom99, MRSW93, Mal95, MC93a, Mar97a, MGD98, Mar98, MMJN97, MLDH90, MB96, NN95, Ned97, Nih91, OM97a, Osh91, Osh96, Par93b, PGD91, PL90, Pow96, RPH99, RL99, Sch96a, SBW^{+}99, SN99a, SC96, SPH97, Smi97b, SBD^{+}97, Sou99, SCM93, SDBA98, SW99, SL91b, SY99, TGM95, TMG96, TMGS97, Tuz91, Uch99, WJW99, WVCS94, Wil99, Wil91, WK97a, YAKS97, ZH97]. Population-dynamic [RA96]. populations [ALGB99, Axe93, Bye96, Cla99, DS99, DG96, Day90, DGS91, Dub97a, FAF98, Gar90, GA96, HE96, Hua98, JWEB92, Jen93, Jen94, Jen96a, Ken92, KD96, LLWP95, MHA95, MK97, MG98, MVBP93, Mau90, Mil92, MRSK96, PMV96, PP95, PR95c, RM95, RCD93, SBD^{+}95, SM99, SMWF98, SLP93, TAB^{+}93, VJR^{+}98, WLW99, YH99, dGG98, Jør95c]. porous [LLM96]. Port [BBR99]. Portland [Jør95g]. portrait [Ned99]. Portugal [FKNM^{+}97]. Portulaca [Yam97]. positive [FP98, Ula95]. Possible [KR92, HBM^{+}97, SR95b]. postdepositional [PJ96]. potato [RL99]. potential [IPS99, Kat90, KG95, LC91, M.a99, PJ96, TIH98, VPFM97]. potentials [Tis97]. pounds [Ano99c]. Power [WHJ^{+}92, GG99, Ken95, Pat95, Sal92a, WF99, iWHS97]. pp [All91, Ano91i, Ano96p, Bos91a, Bos93a, Bos93b, HS94, Hug91, Jør91a, Jør91b, Jør91c, Jør93a, Jør93c, Jør93b, Jør94a, Jør94b, Jør94e, Jør94c, Jør94f, Jør94g, Jør94d, Jør95a, Jør95b, Jør95c, Jør95e, Jør97b, Jør99a, Koh91, Kom94, Kom97, Leg91a, Leg91b, Leg92, Leg93, Mar93, Mar99a, Mau93, McD93, Mit94, Nie95b, Nie95a, Nie95c, Nih91, Nih93, Sil95a, Ste93, Str91, Str92, Svi91b, Svi91a, Ulg95, Ush91, Vas94, Wu95, Zuc91]. pp.

[Ano96o, Ano96p, Ano96q, Ano96r, Jør95d, Jør95f, Jør95g, Jør96a]. practice [Jør97b, TGM^{+}99]. Pradesh [MHA95]. prairie [HSB99, PGS98, RBM97]. precipitation [MGD98, Yin96]. predation [GRBF97, HS99, Kit97, McC92, WFK^{+}91]. Predator [MR91b, TSMC96, ALB91, AHHK97, BD97, FG94, HV99, HNCS92, KH94b, MS95, MDH96, PRF97, PR95c, RM92, RM93, SSW91, Smi91, Van94]. Predator-mediated [TSMC96]. predators [HC97b, KR92, MS98]. predict [And95, BH99, CPH99, Håk95a, Håk96b, Håk96c, LBV96, Tom96, WW92]. predictability [Bre92, HRV^{+}96, Leg97b]. Predicting [Chi97, LLM99, LG91, Mon98a, Mon98b, PACS96, PJ99, RÅ99, TS96, CKK^{+}98, CFS93, Com97, DGS91, Dow97, Gob93, KT97, MJY96, MDO99, ML98, RJN99, ST92, TAM^{+}99, VW97]. Prediction [AKH99, HS97a, PT97, SH97, CP96, EFCM99, HH98, Lin98, LHAJ99, Liu93a, Liu97, OR99, PC99a, RFHY97, Rut93, SWK93, SvdG99, TDK98, WG98, McD93]. predictions [CHM96, Håk96a, HL96, HM97a, PL96, PS90, Szy96]. Predictive [LADL99, FK99, Håk95b, HL97, Håk99, NHWS91, Pow93, VWN^{+}99, Wil99]. predominantly [Hul96]. Preface [Ano95-41, Ano97-57, Frä92, vGDT95, Jør94e]. preference [SIU97]. preferences [AK93]. preferred [KS91]. Preissmann [MA93b]. Preliminary [NCNN97, RSdP98, ASTL91, Cha95, TTT98]. Prenosil [Jør97d]. Prentice [Sil95a]. prescribed [PS90]. presence [Els92, Wae93]. Presentation [OH97, OA98]. Press [Ano99d, Ano99f, Jør93a, Jør94d, Jør95b, Jør95c, Jør95e, Jør95g, Jør97c, Leg92, Nie95a, Ush91]. Prey [LR97, ALB91, AHHK97, BD97, CALW99, HV99, HRRP99, HNCS92, Kit97, KH94b, MDH96, PRF97, RM92, RM93, SSW91, Smi91, Van94]. Price [Jør97d, Kom97, Str92, Ano99d]. primary [AO97a, BGF99, BD99, CCR98, CFC^{+}96, CBB97, Dua95, GPG^{+}99, HLF99, JPA^{+}99, JAZ^{+}99, KMS99, MR91a, PA99, Rut93, SH99, SCH92, Str96, Szy96, Val99]. primer [Leg91b]. principle [AO97a, ALB92b, Web95]. Principles [VKH^{+}94, CR97a, Håk99, Pat95]. Probabilistic [TVS^{+}94, Ken95, VBL90, iWLSL96]. probabilities [Liu93a]. probability [LCL93, MA96a, Tun90]. problem [Håk96c, TBM96, Ush91, VWN^{+}99]. problem-solving [Ush91]. Problems [Alb92a, Ush91, Aly91, Il’96, MWJ92, Nie95a, Sol99]. Procambarus [ANM99]. procedure [JO95, SVK93, VdKK^{+}99]. procedures [PTM91]. Proceedings [Ano99e, Jør91a, Jør93a, Jør94f, Jør99b]. process [AL94, Aok97, BU98, Bos92, CBC^{+}98, Coq95, FSKC97, GFS99, Gue91, JH95, Jen96b, KD99, LUC^{+}93a, LRS^{+}93, MLB96, MGM92, PG99b, SSM^{+}93, TP97]. process-based [CBC^{+}98, FSKC97, SSM^{+}93]. process-oriented [JH95]. processes [BSB99, Bos91b, CBS97, Cri94b, Han98, Jen93, Jør93c, Kre91, MPP98, MR97, Mat99, McC95a, MPA97, Mow97, NMS95, PBO^{+}96, Wae93, Wu90]. processing [Ano92g, Hea95, Mau90]. processor [Hae91]. processors [CM91]. produce [GA96]. producer [CFC^{+}96]. producers [CBB97].

Production [AYO99, AO97a, ANM99, AFM99, ANFM99, Aok90, BGF99, BD99, DSNN96, GPG^{+}99, Gra98d, HCL94, HLF99, KMS99, ML97, NAFM99, RGRF98, Rut93, SH99, SCH92, Str96, SHJ95, Szy96, US97, Woo98]. Productivity [Jõg98, BS98a, Dua95, DF97, JPA^{+}99, JAZ^{+}99, Mah98b, MR91a, PA99, PGS98, SRB99, Val99]. products [dLNT98]. profiles [JMPT99]. profit [SCEZ^{+}94]. profitability [SRB99]. prognaus [SM97]. prognosis [GSG96]. prognostic [ŠŠZ90]. program [Hul92, RH96a, SM99, Wil96a, DSB^{+}99]. programming [GPT95, JO95, MC93a, SNUM91, WSGR96]. Programs [Leg91b, Jør93c, LBA^{+}95]. progress [Löf96]. project [LFM97, SRV95]. projecting [AYO99]. projection [GU97, Jen95, Osh91]. promoters [JLS98]. promotes [Uch99]. Propagating [Mow97]. Propagation [vdB98, PM96]. Proper [Jør94f]. properties [BBN98, Bye91, BH99, EAG97, FFS96, Gae92, HRRP99, Lea97, LL96b, MPNJ97, RB99, Sal92a, Svi97a]. property [SY99]. proposals [BM96]. proposed [BJ94]. Prost [Jør99a]. Protection [Jør94f, CBP95, FF91, Vas94, Jør94f]. Protein [Cac90, MGNR91]. proteins [SSK^{+}91]. prototypes [HS92a]. province [EBBW98]. przewalskii [HF93, XY95]. Publ [Jør94f]. Public [Jør94g]. Publication [Ano90m, Leg91b]. Publications [McD93]. Published [Jør91a]. Publisher [Ano91-29, Ano96-61, Ano97-58]. Publishers [Ano99e, Jør94a, Jør94b, Jør94c, Jør94g, Jør97f, Mar99a, Mit94]. Pudoc [Bos91b]. punctatus [MGNR91]. pure [Bar91, PR95c]. purification [CFC91]. purposes [ET95a, ET95b]. putting [Jør97b]. pyrethroid [MHA95].

QUAL2E [VU93b]. qualify [MTCL96]. Qualitative [Gue91, SvB96, Van94]. Quality [Zei91, BJ94, DCM92, HS97a, HSXP98, Jes98, Jør94a, KIPK99, Kom94, KMWB98, Mah98a, PL96, PU97, PSK98, RC97, RCS97, Rec94, RPJK94, SH97, SBW^{+}99, SH92, SCEZ^{+}94, Str91, Str94b, SKW91, SWK93, TM99, VKT94, VKH^{+}94]. Quantification [PWU93, Chr95, XHW^{+}99]. Quantifying [BSN^{+}98, BPN99, FP99, MCU98, LA97, SWK93]. quantitative [CB97, KUI94, Jør95e]. quantization [Kir92]. quartic [Van97]. quasi [Kle95]. Quebec [DT97, BLK^{+}99]. Quercus [RPT93, WK97b]. queueing [Mau97]. queuing [Bat96b].

R [All91, Ano96q, Jør93b, Jør95g, Jør99a, Mar93, Ver96]. r. [SBD^{+}97]. rabbit [BK98, Sey92]. rabbits [BK98]. rabies [ALS97, SM99, TGM^{+}99]. race [PS93b]. radar [BD99]. Radial [Mor99]. radiata [CJMT96, Kin95]. Radiation [Bar98, Ant98, LS98a, LRS^{+}93, NZ92, Sto92, VCGG96, Yin96]. radiative [Bot98, Ces97a, Ces97b]. radicum [SMP96]. radiocesium [Håk97]. radionuclear [KCP^{+}98]. radionuclide [Kas93]. radionuclides [BS95a, BS95b, BS95c, Per98]. radiosity [CA98]. rain [Cha99, GET98, MPA97, NHWS91, Osh91, Osh96, Van96]. rainbow [CR97b, VJR^{+}98]. rainfall [Ahn99, Ano96w, Dun99, GHH^{+}96, Liu97, PJ99]. rainfall-runoff [PJ99]. rainforest [KH98]. Raj [Jør92a]. RAMAS [LBA^{+}95]. RAMAS/space [LBA^{+}95]. Random [Bla97, Bye96, CBS97, Jen93, LLM96, MRLC96, PT94, Rec93, SR93, SY99, WCS95, YO95]. random-search [MRLC96]. randomness [HS96]. range [FDC^{+}96, KG95, MvdZ96]. rangeland [CH97]. ranging [HH99, Sou99]. rank [Aok95a]. rank-abundance [Aok95a]. rape [Axe93, MJMN97]. rare [WCDL96]. Rasmussen [Koh91]. rat [MLDH90]. rate [BMHT99, BL96, Bye96, Håk97, Kit97, MGCR98, MCTB95, MCTB96, SMDSZSXF93, Tap93, WK97b, WRHW94]. rates [FK99, How96a, How96b, LG91, Lyo96a, Par93b]. ratio [BM97a, MD97, MGNR91, Zon98a]. ration [MT93]. Raton [Nie95a]. Ravera [Str92]. raw [Che90]. RCD [BK98]. reaction [HKV^{+}95, Kas93, KVAW91, PC99b, Rip95]. reaction-diffusion [HKV^{+}95, KVAW91]. reactions [LCH95]. reactor [HSN96]. reaeration [PL95]. Real [Bos92, KT94]. Real-structure [Bos92]. realism [RS98b]. Realistic [DPSB99, GW98]. Realized [MWY92]. really [Gon97]. realms [Jes99]. reasoning [Gue91]. receiving [RCD98, SKW91]. recharge [Var93]. recharge-depletion [Var93]. Recknagel [Bos91a]. recognition [KK97]. recognoscitive [VUDM^{+}99]. Recolonization [CBS97]. reconciled [PW93]. reconsidered [DG93]. Reconstructing [Yin96]. recordings [BC96]. records [SVK93, Yin96]. recovery [Str92]. recruitment [Cri94b, HK91, JCS^{+}97]. rectangular [Han91]. red [DDA99, HBM^{+}97, Lea97, LWMT97, ÖM97b, PT99, SDBA98, SFGF94]. red-crown [LWMT97]. red-footed [PT99]. red-winged [ÖM97b]. redox [PJ96]. redtail [MT93]. reduced [FV97]. reduces [WF99]. reduction [FV97, VSG92]. reductionism [PW93]. reef [DS99, McC92, McC95a, Whi99]. reefal [ABK^{+}92]. reference [PR95b, Svi91a]. reflectance [EBA^{+}99]. reflections [Gra98c]. refuges [TSMC96]. regard [SDBA98]. regenerated [DDA99]. regeneration [MPA97, YKH99]. regimes [CJMT96, KEAP97, PA99]. region [BL93]. regional [ADF^{+}93, AWS98, BSN^{+}98, EB93, Jol96, LAB95, RSD94, SF94]. regionalising [PJ99]. regions [VF96a, ZH98]. regression [FZ99, HL97, LWMT97, MD97, MDO99, PT97, SR95a]. regulated [VKH^{+}94]. regulating [Cri94b]. Regulation [GC95, Mau90, KJ94, ZH98]. reindeer [Vir92, Vir96]. reinforcement [DL99]. Reinhold [Nie95c]. related [PBO^{+}96, SR95a]. relating [HKMRS95]. relation [BMHT99, HH98, MvdV95, OM97a, RPT93, RF97, Sil95a, TCC^{+}97, VCGG96]. relations [FP98]. Relationship [M.a93a, Aok95a, Bla95, Chi98a, PS93b, RU97, Ume95]. relationships [DCCSL99, Jør95g, LDB^{+}96, Mau97, Rec93, Rya94]. relative [RCD98]. release [Bel98, BS95c, Col97]. releases [SY90]. relevance [Jet92]. reliability [Fle91]. Remarks [JZV92]. Remediation [Jør94f]. remote [CLCG99, CBP95, GET98, GPG^{+}99, GTF99, Mar93]. remotely [BD99, Foo96]. removal [Got98, HSN96, Hju96, LL96c]. renewable [Wil96a]. repeated [Ann97, Ann99, Yoo99]. replacement [MN93]. Reply [Ann99, Ver96, How96b]. representation [CS96, HMB99, Lav96, SN99b]. representative [SL91a]. representing [GPO97, MGNR91]. reproduction [RB99]. reproductive [ACS96, GU97, Jen94, Omo95, PGS98, RC93]. Republic [KSD^{+}99]. required [Tay96]. Research [Jør94f, Ush91, ZH98, Ano94k, Jør97f, MWK95, Sal92b, ZMH96]. Reserve [LWMT97]. reservoir [HP94, Jør94a, KG96, MA96b, ST96, SMY^{+}97, Str94b, Str94a, TS93, TM99]. reservoira [JRB97]. reservoirs [BLK^{+}99, DT97, HS97a, JN94, PASG97, RPJK94, SH97, VKH^{+}94]. Residence [Han97b, HP94, LRM93, Whi98]. resident [CR97b]. residue [JHM97, RSKS91]. residues [Whi95]. resistance [MHA95, Sey92]. resolutions [TDM96]. Resource [McC92, Swa96, FP99, GS98a, Gra98c, McC95b, MMRR92, PGD91, SR93, Sey92, SW99, TRR96, Woo98]. resource-based [GS98a, MMRR92, SR93]. Resources [CB97, Boc99, BP93, Dub97a, Jør96a, Leg97a, VPFM97, VK99, Wil96a]. respect [KW95, SRV95]. respiration [Coo95, PG99a, VB99, Was95]. responds [HC97c]. Response [GF90, SvB93, AOD97, AKH99, BDD97, BD97, CKK^{+}98, Cas96, CCK92, Com97, DD̆Á97, DSB^{+}99, EBBW98, GRL92, Gra98a, Gra98b, HMC99, HE96, JK99, Kre91, LCMR96, MWY92, MPP98, MG98, MLDH90, MRLC96, OH97, PS91, Ped98, PA99, PJL^{+}96, SJ96, UDVESM^{+}95, VW97, VS96a, Vol98, WC96, dVZ93]. Responses [CHM96, Cou91, DH95, Dun99, Hog99, Jen91, Kre91, VB99]. Restoration [Ned99, LH92]. results [Bot95, BL93, DSKM95, DSB^{+}99, SH91, TC96, BZ95]. resuspension [BH97, MR91a, SK91]. retention [AS94, Håk97, MR91a, Por98, TN98]. return [PD99]. reuse [KD99]. Revealing [BC96]. reveals [DPSB99]. Review [TvG95, Ano97i, Ano98a, Ano99a, HR96, LBA^{+}95, LL96a, MG96b, RFB99, Sch96a]. Revised [Mar99a, Hug91]. Revisiting [Del92]. Reynolds [Leg91b]. rhino [HS91]. Rica [BSN^{+}98, VF96c, VZF96]. rice [ANM99, AFM99, ANFM99, BG93, BKT95, JMA97, Mah98b, NAFM99, TAM^{+}99, Mah98b]. richness [KL99, NCNN97, WV91]. Richter [Jør93c]. Ridge [LFM97]. rigida [BCD93, CB97, CS97, FSC^{+}97, SPP^{+}97]. Rio [GHQ99]. riparian [HMA90, LADL99, dLNT98]. Ripley [WF99]. rise [CCK92, HS92c]. risk [GRBF97, JLS98, Rec94, RB98, SW99, SR92, Jør94g]. risks [BLK^{+}99, FZ99, MBF96]. Risto [Ano99e]. River [FKNM^{+}97, Lee91, SN99b, BJ94, Håk96b, JHM95, KCP^{+}98, MDO99, SKW91, UMJ^{+}97, VU93a, VU93b, YS98, vdPV99, BM96, CMS90, KH94b, MDB98, MFP99, PAGL96, Jør97g]. riverine [CS96, SS98]. rivers [AH98, BLK^{+}99, JCS^{+}97, SIU97]. Robert [Ano99c, Jør91b, Jør93a]. robin [RB99]. robins [MGCR98]. robust [KH94a, LA97]. Rochebut [Sal97]. rock [ST92]. rock-plant [ST92]. rocky [MS97]. rocky-bottom [MS97]. rodent [Kit97]. rodenticide [MLDH90]. rodents [HH99]. roe [RPH99]. Roger [Kom94]. Röhrig [Bos93b]. role [Ano96q, GG92, LW96, MHA95, Mat99, PG99b, SDS96]. Roles [Nor96, AUS96, MR91a]. Rom [Jør99a]. rook [PT99]. root [BKG97, Gra98b, SMP96, SDS98, TP97, iWHS97]. root-morphology [BKG97]. rootworm [EH91]. Rosemary [Jør97b]. Roskilde [FKNM^{+}97]. rotational [TMGS97]. Rothamsted [Par96]. rotifer [ACS96, GRL92]. rough [HAW97]. routine [WU96]. Routledge [Nih91]. roving [Whi98]. rubecula [RB99]. rule [DGWK97]. rules [Bár96a, Ben99, RLHD99]. run [WTL91]. Running [Bre95, SML98]. runoff [PJ99, Str96]. Russel [Jør97c]. Russian [KS91, KN95]. rust [PS93b, YAKS97]. Ruth [Jør94c, Wu95].

S [Ano96r, Bos93a, Jør92a, Jør95b, Jør99b, Leg92, Mit94, Str91, Tor96, Ulg95, Zuc91, Mar97b, OYA95, SV96, SA97]. S-system [Tor96, Mar97b]. S-systems [SV96]. S. [Mar91]. S.E. [PJ99]. Sabate [Jør99b]. safe [JWS95, AWS98]. safety [Sol99]. sago [Spe94]. Sahel [Str96]. salinity [CKK^{+}98, Met98]. Salmo [ALGB99]. salmon [Bar96b, Cri94a, Cri94b, JCS^{+}97, Mau97, TN98, VK97]. salmonid [LK99]. salmonids [Lee91]. salt [RM95]. Salvelinus [PP95]. samango [SLP93, SL96]. same [DSKM95]. sample [AL99, DFPG91, DG93]. Sampling [DFPG91, DG93, Ann97, Ann99, NCNN97, OM97a, Yoo99]. San [Jør95e, Jør97c, Jør97e]. sand [KKNY91]. sandpile [RAM99]. sandy [GLD97]. SAR [Met98]. Satellite [GPG^{+}99, AO97a, LRK96, RC93, SFS97]. sativa [AFM99]. sativum [RJN99]. saturation [AOD97]. savanna [Rut93]. saving [HS91]. say [AOE99]. Scale [Ano99d, ABK^{+}92, BD97, BH94, BPTM^{+}97, Cle98, HKMRS95, NF96, RBM97, SD97, SS98, SFS97, VF96c]. scales [FDC^{+}96, WC96]. Scaling [AW97, Nal97, CLCG99, Ken92, TGM^{+}99, Jør95e]. scaling-down [TGM^{+}99]. scenario [IPS99, SVK93]. scenarios [BHL96, FSA99, PG97, PG99b, TBvH95, VF96c]. Schaefer [GC91]. schedule [BSB97]. scheme [PSK98, WL97]. schemes [JO95, Jol96, Kim94]. Schmitt [Jør97c]. Schneider [Jør95e]. Schnellnhuber [Ano99b]. schools [Rom96]. Schwarz [VNLB96]. Schweitzer [Ste93]. Science [Ano99b, Jør94g, Jør95f, Jør97b, McD93, Nie95b, Jør93a]. Sciences [Ano99e, Jør99b, Ush91]. Scolothrips [SSW91]. scolytidae [BBL97]. Scots [GS96, KS95, KIPK99, VS96c]. Sea [KKNY91, All91, CCK92, Del92, Leg97b, SNUM91, AH99, BSL97, HCdVM98, HW98, KK97, Per98, Sem99, Tom96]. seagrass [BBR99, CL99, OKS99, WHA97]. search [And96, BSB97, DL99, LW96, MRLC96]. searching [Bye96, MK97]. season [CR97b, SGML91, SCH92, TMG96, TMGS97]. Seasonal [GdC99, KL98a, AH98, And98, Bac93, CPH99, EGJ95, Got98, GHH^{+}96, Håk96b, HC97b, KR96, Smi97a, dZcZ96]. seasonality [WFJ96]. seasonally [RM93, dGG98]. seasons [Whi95]. seaweed [LA91]. Second [Mar99a]. secret [Ula95]. SED [VP93]. Sediment [Jør94f, BH97, PJ96, SK91]. sedimentation [AL94, MR91a]. sediments [And95, Cha95, Håk95a, KK97, UMJ^{+}97]. seedbanks [GA97b]. seedling [KG96]. segment [PL96]. select [SL91a]. Selected [Jør94f]. selecting [JO95, SH92]. selection [BK91a, LK99, Loe97b, McC95a, Omo97, VWN^{+}99]. Selective [MCT97, MG98]. Self [CFC91, TKO91, Bar91, BU98, DJ98, FV97, Foo99, JMN98, KB97b, MC93b, MM97, RAM99, RF97]. Self- [TKO91]. self-design [MM97]. self-incompatible [DJ98]. self-motivated [RF97]. self-organising [Foo99]. self-organization [BU98, KB97b, MC93b]. self-organized [RAM99]. self-organizing [JMN98]. self-oscillating [FV97]. Self-purification [CFC91]. self-thinned [Bar91]. SEM [KBSC98]. Semantics [Wen92]. semelparous [CALW99]. semi [BH96, GHQ99, KR96, MGM92, Omo95, WL96, iWLSL96, ZH98]. semi-arid [GHQ99, KR96, WL96, iWLSL96, ZH98]. semi-lunar [Omo95]. semi-Markov [MGM92]. semi-stable [BH96]. semiarid [CL90, SCH92]. semistable [BH99]. senegalensis [HC97b]. Senegalese [HC97b]. sensed [BD99, Foo96]. sensing [CLCG99, CBP95, GET98, GPG^{+}99, GTF99, Mar93]. sensitive [FSG96, LL97, WBH^{+}98]. Sensitivity [BBN98, Bar96b, Els92, HS92c, HSHS96, Jam93, Nak92, PC99a, WG91b, BS98a, HK91, Kle97, LL97, MBF96, Per96, SH97, SSY95]. sensor [MCGLJ99]. separability [AL91]. September [Ano92v, Ano93w, Ano95-31, Ano96-50, Ano98-40, Ano99-32, Ano99-42, Jør94f]. septempunctata [SPH97]. sequences [MN93]. sequential [Mow97]. serial [RSS91]. series [DFP^{+}93, MDT93, MFP99, NNR98, OYA95, PTM91, TD94, Jør97b]. set [Cao95, DSKM95, MKA^{+}95]. Seto [KKNY91]. sets [GPO97, MWK95, Sil97, WGP96]. settlement [DS99]. several [Ano99b, Bos91b, EBBW98, Han97a, Kar93]. sewage [SKW91]. sexual [Loe97b]. shading [RU97, Ste95b, TKO91]. shallow [AB97, CO99, CBB97, CS97, FKNM^{+}97, GH94, JA96, LP91, Nie94, PL96, SFGH97, SNUM91, XJTL99, YS98]. shallow-sea [SNUM91]. SHALOM [Ziv98]. Shannon [Jør97g, BM97b]. sheep [HWB96]. shelf [BK92b, BHWP91]. shell [Kom97]. shifting [GMI95, Jes99]. shifts [BBKV96]. shoaling [Sen90, SN95]. Shock [MLB96]. Shoot [GS96, Ste95b, SDS98]. Shoot-based [GS96]. shoots [KS95]. shopping [BL96]. Short [Hea95, KW99, Ped98]. short- [Ped98]. Short-term [Hea95, KW99]. shortgrass [RBM97]. shovelers [DuB97b]. shrimp [GMM91, MGM92, MT93, MT95, MT96, MLGV99]. shrimps [TSMC96]. shrub [Cou91, Day90, SSY95, SFS97]. shrublands [Dun99]. Shugart [Ano99f]. Shuttleworth [BILL98]. Shuttleworth-Wallace [BILL98]. Sidney [Jør97c]. Sierra [MU99]. Significance [HRRP99, IV91, PHB90, SFGH97, EAG97, Gri99b, Omo95, Omo97, Ste95b]. significant [Böl90, JCC^{+}95]. silos [Tho97]. silt [Jar95]. silvopastoral [BEB99]. SIM [SPE95]. similar [Ebe93]. similarity [Jen97]. Simon [Svi91b]. simpas [Kom97]. Simple [BK98, Jen95, WWJ99, BM97b, BM94, BHBM96, Bye91, Dam99, Håk96b, Håk96c, HS99, Ker95, KJ94, LA91, LK91, Mar97a, Mar99b, MTCL96, MP99, Pie95, Rux96, SRB99, SBD^{+}95, Van94, Van97, Var93, WFJ96, WG91b]. simple-biology [Mar97a]. Simplest [Svi99, SvB97]. simplification [MN93]. simplifications [AWS98]. Simplified [ML96, SGS93, Cou91, KS96]. Simplifying [Mah98a]. SIMULAT [DNR95]. simulate [Ker95, Per98, VF96c, WU96]. Simulated [FS97, JHM95, MGCR98, RM95, BJ95, GHQ99, Rom96, Stu98, VS96a]. Simulating [CMG98, KH96, Kre91, KR96, LC91, MRL94, RH96b, SM95, SA97, SFGF94, WC96, Akç91, Jes98, LA98, OH97, OA98, RSD94]. Simulation [ALS97, BSE95, BSB99, Bos91b, BK91b, Chi98b, CS97, Coq95, CG93, CD99, DA95, DH95, FSA99, FOS95, FSG96, Gra94, Gra98b, Hog99, Jar95, Jen91, Jør92a, KS91, Kom97, LSC95, LSB96, LRG98, MPP98, MGD98, MG98, MWK95, MLDH90, MLGV99, MK95, OP96, OM90, PGD91, PGS98, PRK97, RC93, ST96, SG97, SCH92, SSY95, SHJ95, TMG96, TMGS97, TS93, VGDvdZ97, WTL91, XY95, vHvGT95, AR96, Ano92s, Ano92-30, ACS96, Axe93, Axe94, BZ95, BG92, Bar98, BL92, BMV99, Bos96, Bou95, BKG97, CCKO94, Cla99, Cle98, CHC91, CL90, DFHR96, DF97, Dup98, EH91, EBA^{+}99, FG94, GLD97, GW98, GGZ99, GF90, Gra95, GT97, Gra98d, Hae91, Han95, HB93, HWB96, HLF99, Jac96, Jam93, JHM97, Kai96, Kat90, KGD97, KMS99, KG96, KUI94]. simulation [Kol98, LD97, LCMR96, LSB99, LRM93, LS98b, LS99, LMG96, MHA95, Mar91, MGNR91, MC97, MTCL96, MJGW94, MVS^{+}95, MB96, MRLC96, Mow97, Nal97, NCNN97, PT95, Per93, PMV96, PSC93, RB96, Rec94, RY91, RGRF98, RL99, RDFZ96, SGNS91, STTR93, SM99, Sem99, SSS^{+}91, SY90, Smi95a, Smi97a, Svi91a, TC98, TIH98, TKG97, TWRW93, WJW99, WK97a, Woo98, WVB93, iWSG96, iWLSL96, WL97, YO95, Ziv98, dC95, dVZ93, vT95, vdBDJ98, NAFM99, Jør97d, Kom97, Bos91b]. simulation-model [Dup98]. Simulations [AHHK97, DSB^{+}99, KH98, Ann97, Ann99, CJMT96, Cou91, MB96, MJMN97, SPE95, SBD^{+}97, SR94, VVK97, WSGR96, Yoo99]. simulator [Bar91, CR97a, SM97, TP97, LFM97]. simultaneous [HSN96, Hju96]. SimuSolv [RH96a]. simzoo [Kom97]. Sinarundinaria [iWSG96]. Singapore [Jør97h, Jør97e]. single [Che90, GC95, HE96, Uch99]. single-species [GC95, HE96]. sink [KH96]. sinks [Jør95d, TZE96]. SIRS [GCT96, TC96]. Site [MK95, SABG93, Ant98, BTvH95, BL93, DNR95, GKvdSdV95, JWS95, Krä95b, KW95, MvdV95, OYA95, RGBM90, Sch95, SH92, VMB95, WL96, dVPO^{+}95, vD95, vHY95, vHvGT95, vOÅ95, TBvH95, GW98]. Site-specific [SABG93, Ant98, SH92]. sites [GPO97, Kin95, LUC^{+}93a, LT96]. siting [CPPB99]. Sitobion [SKŞ95, SPH97]. six [VW97]. sixteen [TvG95]. size [BSV91, Cas96, DFPG91, DG93, Gae92, GGC98, Håk95b, HE96, HS97b, LA97, LK99, MvdZ96, MG98, Omo97, PMV96, Spe97, VV92, WWJ99, YH99]. size- [Cas96]. size-based [GGC98]. size-selection [LK99]. size-selective [MG98]. size-structure [YH99]. size-structured [PMV96]. sizes [Liu93a]. sjostedti [TBG93]. skylarks [DSNN96]. slash [CG93]. slicks [Del92]. slime [Tom96]. slope [CSCD99]. Slovenia [DDA99]. Slow [RM92]. Slow-fast [RM92]. sludge [VV91, VV92]. small [FAF98, HL97, JD99, KS96, MJY96, NF96, PR95b, RBM97, ST99]. small-scale [RBM97]. smallmouth [DGS91]. Smith [Ver96]. smog [WHJ^{+}92]. smolts [Cri94a]. smoothed [HM97a]. Snarpe [Jør97d]. SND [HSN96]. Sneznik [DDA99]. snowmelt [SD97]. social [PRF97, UH99]. societies [Nie97]. Society [Jør93a, Jør94g]. socio [GT97]. socio-cultural [GT97]. sockeye [Cri94a, Cri94b]. Soft [Ano99c, Jør97b]. Softbound [Jør93a, Jør94d]. Software [MCGLJ99, CP92, Jør97d, LS99]. Soil [Bou95, Jør94f, MMJ^{+}98, vGDT95, AS94, And96, AO97b, BBN98, BJ95, BKG97, CTL98, Che90, CK97, Coo95, CTKW98, CBP95, DA95, ET95a, ET95b, EMHB99, Fom97, FK99, GYYL96, GPO97, GSS94, Gra98b, JHM97, Jar95, Jør93c, Kal99, Kas93, KdH99, KRD96, KSD^{+}99, KR96, LAB95, LRS^{+}93, LKS96, MvdZ96, Met98, MWK95, MI95, MK95, MJMN97, OA97, Par96, RKM97, RRJ90, SM95, SH91, SMDSZSXF93, SDS96, Smi95b, Sog93, SA97, TvG95, TP97, VVDF95, VDV95, Ver96, VWN^{+}99, Wae93, WL96, WM95, ZBDN96, vdHvEH96, vdSdVK96, Bot95, Jør93b]. soil-carbon [Par96]. soil-crop [VVDF95]. soil-oriented [Sog93]. soil-plant [RKM97]. soil-plant-atmosphere [MJMN97]. Soils [Jør99a, BS95b, Böl90, KRD96, Per96, PRK97, RH96b, RRNM97, dVKvdS95]. SOILVEG [vHvGT95]. solar [Ant98, NZ92, SV98, VCGG96, WU96, Yin96]. Solenopsis [SFGF94]. solids [TN98, TPS98]. Solling [Bou95, BJ95, BTvH95, EGJ95, FWCJ95, GKvdSdV95, JWS95, Krä95b, KW95, MvdV95, OYA95, PDAS95, Sch95, TBvH95, VMB95, WM95, dVPO^{+}95, vD95, vHY95, vHvGT95, vOÅ95]. solubility [WM95]. solute [SS98, VDV95]. Solution [Wil96a, Han91, Kal99, Kas93, Kim94, Mar99b, SBD^{+}95, VNLB96, vdSdVK96]. solutions [Coo95]. solving [Sol99, Ush91, YWMM98]. somatic [SG97]. Some [Cra98, MCTB95, RH96a, Bat96a, Löf96, Svi97a, vT95]. SOMM [CK97]. Sonntag [Ano96q]. Sons [HS94, Jør97h]. Soons [Jør94e]. sorption [SH91]. sorption/dynamic [SH91]. source [AL99, KH96, LVS^{+}99]. source-sink [KH96]. Sources [Tap93, JTS92, TZE96]. South [MDB98, SLP93, Ahn99, Osh91, HM94, TM99]. south-western [Osh91]. southeast [Por98]. southeastern [JRB97, Kin96]. southern [BW96, GHQ99, HWB96]. southwest [PAGL96]. soybean [VEA99]. sp [TMG96, TMGS97]. sp. [And98]. space [CBB97, EFGE98, Gri99b, Lav96, PW91, UB97, LBA^{+}95]. spacing [Sou99]. Spain [PAGL96]. sparrow [SGML91]. Spatial [BG92, CM91, Hen95, HSB99, Jør95e, LAB95, MHJ96, NG96, PM96, SB97, TM99, Bak96, BSB97, BD97, BvSH91, BPTM^{+}97, BGTL99, Cha99, CLCG99, Chi97, CWC95, Day90, Gao96, GLD97, GMI95, HS97b, HF96, JAZ^{+}99, Kar93, KL98b, LS98a, LWMT97, MR97, MUM98, McC96, MEBZ96, MPA97, Mow97, ÖM97b, ÖÖ99, PBO^{+}96, PL90, RBM97, Rux95, Rux96, ST99, Smi91, TC98, TDM96, Tuz91, VdKK^{+}99, WCC95, WL97, Yam97, YKH98, YKH99]. spatial-temporal [Tuz91]. spatialized [MRL94]. spatially [And96, AW97, BKT95, CF99, Cha99, Cla99, GW98, KdH99, KMWB98, LA98, MvdZ96, PMV96, PM96, TGM^{+}99, TDM96, WU96, WJW99]. spatially-distributed [BKT95]. spatially-explicit [And96, WU96, WJW99]. spatio [Gri99b, LD97, LAB^{+}98, MC97, TSR97]. spatio-temporal [Gri99b, LD97, MC97, TSR97]. spatiotemporal [EFCM99]. spawner [Mau97]. spawners [Cla99]. spawning [CR97b, Tre91]. special [PR95b, SRV95]. specialists [VV91]. speciation [BH91]. Species [JCC^{+}95, AUS96, AHHK97, BEM97, Cha96, De 98, DH95, Dye95, Ebe93, EBBW98, GC95, GS98a, GU97, HM94, HMC99, HMB99, HLG96, HE96, HCdVM98, IPS99, Jet92, Jør92c, Kei97, KSR97, KG95, KH94b, KH98, KBV99, KL99, LBV96, Liu93b, MRCL92, MDO99, MGD98, McK96, MG96a, MPA96, MR91b, MCT97, NCNN97, Osh96, PWU93, PBO^{+}96, PR95b, PC99b, RSC96, RBM97, dLNT98, SJ96, Sam96, Sch96a, SG98, SCH92, SD96, SDF96, SFGF94, VG98, VB99, WCDL96, WFK^{+}91, WK97a, WC96, WBH^{+}98, WV91, Ziv98]. species-area [De 98]. species-level [VB99]. species-specific [PR95b]. specific [Ant98, KR92, PR95b, SH92, SABG93]. specification [AN90]. specifications [VP93]. Specom [Che90]. spectral [BGL99, VCGG96]. SPESS [Bar91]. sphere [BCH^{+}97]. Spider [BM94, BM99, CHLT96, UH99]. spill [CMS90, FFR99]. Spilopsyllus [Sey92]. Splachnum [Mar91]. sponge [CD99]. spore [Mar91, YWMM98]. spp [MDB98, RDFZ96]. spread [CWC95, HR96, Kar93, MHA95, Tap93, TGM^{+}99, WCC95]. spreading [Del92, KT97]. spring [Axe93, Mar97a, ZH98]. Springer [All91, Ano91i, Ano99b, Hug91, Leg91a, Nih93, Svi91b, Wu95]. Springer-Verlag [Wu95]. spruce [Bou95, BJ95, Ces97b, EGJ95, FWCJ95, GKvdSdV95, Jõg98, JWS95, Krä95b, KW95, MvdV95, OYA95, PDAS95, RY91, Sch95, SY90, VMB95, dVPO^{+}95, vD95, vHY95, vHvGT95, vOÅ95, BTvH95]. SPUR [CT96, FTH99]. SPUR-91 [CT96]. spur2.4 [FTH99]. squid [GdC99]. Sr [Mon98a]. St. [FFR99]. stabilisation [Tay96]. Stability [GS98a, HNCS92, Kei97, Li95, Mil92, MDH96, Nie95a, Aly91, BDD97, CCK92, GSW92, MY97, Nak92, NJ96, ND97, Phi95, Pie95, TC96, VMSP97, YH99, Log97]. stabilization [MGGSMC^{+}96]. stable [BH96, FAF98]. stage [Cas96, GU97, Mar97a]. stage-projection [GU97]. stage-structured [Cas96, Mar97a]. stages [LV99, PGS98]. stalk [HvCC94]. stand [Bar98, BSE95, Bou95, BJ95, Ces97b, Che90, EGJ95, FWCJ95, HLF99, KABT90, Kol98, KI98, Lea97, Mal96, PDAS95, SM97]. stand/raw [Che90]. standard [HGB98, KRD96, Tun90, VF96a]. stands [Bar91, GPT95, GBBV99, Hog99, Jõg98, LBV96, RU97, Val99, WK97b, YKH99]. Stanley [Sil95a]. starfish [ALB91, ABK^{+}92, AK93, RGBM90]. starfish-coral [AK93, RGBM90]. starting [MRCL92]. State [Jør95h, Jør99d, LCH95, Mül97, Aok92, Bak96, BFK96, CP92, CTKW98, Han97b, Kim94, Kle95, MC93a, TBM96, UB97, Whi99, Xu97]. State-of-the-art [Jør99d, Mül97]. states [Coq95, SR95b, UB97, IPS99]. Stationarity [BMHT99]. stations [WHJ^{+}92]. statistic [ML92b].

Statistical [Ahn99, EB93, LL96b, MB93, Mes92, PS93b, RS98a, HC97c, Kra95a, LVS^{+}99, MTCL96, Oel95, UDML97, WCDL96, Leg91b]. Statistics [AJK95, MRCL92, MEBZ96]. status [BMHT99, CFC^{+}96, MI95, VSG92]. steady [Aok92, CP92, CTKW98, Han97b, Kim94, Kle95, SR95b, TBM96, Whi99]. steady-state [CP92, CTKW98, Kle95, Whi99]. steep [PL96]. STELLA [CG98]. step [PSK98]. sterilization [BW94]. Stirling [GGJ96]. Stochastic [Bey98, Jet92, TH90, BBR99, BS95a, BS95b, BS95c, BK96, DCP99, De 98, Gar90, GMM91, IV91, Lee91, MDT93, RSS91, SR98, Smi91, VU93a]. stochastically [Bat96a]. stochasticity [Akç91, McC96, PD99]. stock [Cri94b, HK91]. stock-recruitment [HK91]. stocks [ND97]. stoichiometric [PU97]. stomach [Bey98]. stomatal [HC97c, Kat90]. stopping [Bat98]. storage [DLN^{+}95, MJY96, PHB90]. Straskraba [Jør94a]. Strategies [Nie92, FG92, FAF98, HS91, HvCC94, LGM99, MK97, MRSK96, NCNN97, PMV96, SGNS91, SM99, SMY^{+}97, Whi98]. strategy [ADF^{+}93, Jes98, LA91, MK97, TGM^{+}99]. stratification [LP91]. stratified [CPH99]. stratiform [Bot98]. straw [MJMN97, MMJ^{+}98]. stream [CR97b, Jør97g, Kim94, PG99a, RLHD99]. stream-resident [CR97b]. streamflow [PJ99]. streams [DLN^{+}95, DH99, PL96]. stress [CKK^{+}98, RKM97, vHvGT95]. stresses [Ped98, PP95, vHY95]. stressors [KHL97]. stripe [PS93b]. strongly [VKH^{+}94]. Stroosnijder [Jør97h]. structural [Cha99, CBB97, GPT95, JN94, Jør99d, KS95, KIPK99, MDH96, Nie92, Nie94, Nie95d, Xu97, vdHvEH96]. structural-dynamic [Nie92]. structurally [TC98]. Structure [GSG96, AR97a, BS95a, BvSH91, Bos92, BLS98, Ces97a, CTDC94, DH95, DH96, Dup98, ET95a, HV99, Jør97a, Kol98, KV90, LUC^{+}93a, LKS96, LA97, McC96, McC95a, Ned99, PW97, PSN98, RSdP98, RA96, SCEZ^{+}94, SPP^{+}97, Spe97, Ter93, UAA97, VKT94, Whi99, YH99]. structure-ecosystem [LUC^{+}93a]. structured [AM98b, CALW99, Cas96, Gao96, Jen96a, JMPT99, Jør95c, KI93, Mar97a, Mar99b, PMV96, YAKS97]. structures [Coq95, Han98, KK97]. Stuart [Nih93]. students [Jør91c]. Studies [All91, Ano91i, DSB^{+}99, Hug91, BGS97, KV97, Leg92, LFM97, PK92, RCS97, SD97, Ush91]. Study [Cha99, Sal97, SNU96, Aok94, Aok97, Bar98, BvSH91, BSN^{+}98, CCK92, DCCSL99, EAG97, FG92, FSG96, Gae92, HMC99, IPS99, KCP^{+}98, LD97, Mah98b, MDO99, Mar91, OS95, PS93a, PS93b, Por98, SRB99, SC92, SH99, Sey92, SH91, SS98, SRW97, TZE96, Tor96, VU93b, VG98, VF96b, VFC98, VVC99, WK97a, WVB93, iWSG96, iWHS97, YH93]. studying [Kol98, KL99]. sturgeon [GG99]. sub [Håk97]. sub-models [Håk97]. subalpine [Mow97, YKH99]. subfamily [CALW99]. Subject [Ano90n, Ano97-65, Ano97-66, Ano97-64, Ano97-59, Ano97-60, Ano97-61, Ano97-62, Ano97-63, BSB97, WLW99]. subjected [Jen93]. subjectivity [VKT94]. submerged [CC96]. submodel [MMJ^{+}98]. submodels [Str91]. subroutine [WTL91]. subsiding [RCD98]. subspecies [MRSW93]. substances [Cha96, Mil92, Mon98b]. Substrate [VV92]. subsurface [DH99, Fom97, LJ94]. subunits [Böl90]. success [Cla99, Håk99, LW96, PR95c]. successful [RB96]. Succession [KSR97, CL90, CH97, LRS^{+}93, WE96]. successional [Cha99, CCR98]. successive [Whi95]. sugar [RU97]. sugarbeet [Smi95a]. sugarcane [HvCC94, MLDH90]. sugi [Chi98b]. suitability [LSB96]. sulfate [FV97]. sulfate-reduction [FV97]. sulphide [BSL97]. sulphur [Kre91, Jør95f]. Sumatra [GET98]. summer [FSC^{+}97]. Super [SBD^{+}95]. Super-individuals [SBD^{+}95]. superba [MS98]. supercomputers [Smi91]. Supplement [Jør93b]. supply [MJY96, SG98]. supply/demand [SG98]. Support [BGL99, GA97a, HS92a, HS92c, KL99, LS98b, LS99]. suppression [OM97a]. Surface [VFC98, VVC99, DH99, Fom97, MU99, RU97, Rya94, SvB93]. surfactants [JMA97]. survey [AJK95]. surveys [WCS95]. survival [Fah98, Jen93, KG96, LR97, SDF96]. susceptible [Bye96, TC96]. susceptible-infective-immune-susceptible [TC96]. suspected [LT96]. suspended [AL94, LHAJ99, PAGL96, TN98, TPS98]. Sustainability [GHQ99, KI98, BSN^{+}98, BPN99, Com97, Pat97, Pat98, SSH98, UOB94, UB98, Jør93a, Ano99b]. Sustainable [SSH98, dGG98, Ano96r, How96a, How96b, KL98a, Lyo96a, Nih91]. sustaining [Pat98]. Sven [Mar99a, Sil95a]. Svicobians [Log97]. swallow [CPM99]. swamp [BL96]. sward [Hut91]. SWBCM [EMHB99]. Sweden [Håk96c, Lin98]. switching [SS96]. Switzerland [Ano99e]. symbolic [Hul98]. symmetry [Hul96]. sympatric [VJR^{+}98]. sympatry [CR97b]. Symposium [Ano93-30]. Synchronisation [PG98, GdC99, HH99]. synchronized [IK95]. synergism [FP98]. syntax [Wen92]. Synthesis [Pat97, Sol99, GWAU99, UAA97]. synthetic [SVK93]. Syrphidae [CS96]. System [Ano99b, Jes99, ANM99, AFM99, ANFM99, AHHK97, BG92, BM99, BL92, BEB99, Bir94, BLRV98, CF99, Cha96, CTL98, CHLT96, CU93, DSB^{+}99, FG94, FV97, GA97a, GSG95, HS92a, Hil96, KL99, KCP^{+}98, LCMR96, LLWP95, Mar97b, MMRR92, MLGV99, NAFM99, PR95a, PAGL96, Phi95, Pie95, RPH99, RSC96, RPJK94, RB98, SH92, SY90, ŠŠZ90, Tor96, UB97, Vir92, Vir96, YAKS97, dC95, ŠŠZ90]. system-level [CU93]. systems [All90, AL91, Aok92, Aok93, BM97a, Bos91a, BC96, BK96, Bye91, BH99, Cle98, CG98, DD̆Á97, EFCM99, FP98, Gra98c, GS98b, HV98, HSPM96, Hea95, HWB96, Hua98, JZV92, Jør92a, JPS92, Jør95b, JMN98, Jør99c, Kir94, Mau90, ML98, Mon98a, MY97, Mül98, OP96, PRF97, PYE94, Pol96, RSKS91, Rob96a, Rob96b, RH96b, dLNT98, SV96, Smi91, Spe94, TMGS97, THR96, UB97, Vin96, Wae93, Wen92, WHN96, Woo98, HS94, Jør95a, Str91].

T [Bos91b, Leg92, Mar93]. table [Cas96, Gra98a]. tables [Nih91]. taboos [Col98]. Tabu [BSB97]. TACT [AR96]. tactical [AR96]. tactile [EAG97]. taeniorhynchus [RM95]. Taihu [HSXP98]. tailed [XHW^{+}99]. Tampamachoco [RSdP98]. tank [MJY96]. Tarawera [CJMT96]. taxation [GC95, PC99b]. taxifolia [AH99, HCdVM98]. Taxing [GG92]. Taxodium [Hua98]. technique [ESD98, Håk96b, MG96a]. Techniques [Jør92a, DCM92, Dro96, DGWK97, HC97c, JD99, NNR98, OR99, SR95a]. Technology [Jør94g, ÅB90, KCP^{+}98]. teleost [SSK^{+}91]. temperate [JPA^{+}99, Bos93b]. Temperature [ND97, BBL97, CFS93, CWPD95, CP96, Dua95, FOS95, Håk96b, LLF97, LV99, LP91, Mah98b, MGD98, MGNR91, RRNM97, Sch96a, SKŞ95, Sel95, SMDSZSXF93, VB99, Yin96]. temperature- [FOS95]. temperatures [MD97, MT93, OA98, Rya94]. tempo [HBP93]. Temporal [LTMP97, MJMN97, CLCG99, Gri99b, HSB99, Jør95e, LD97, LAB^{+}98, LA96, MC97, MEBZ96, Rux96, SB97, TSR97, TM99, Tuz91]. Ten [Gri99a, Bos91b]. tephritid [JWEB92]. term [CH97, DuB97b, GFS99, Hea95, KW99, Mon98a, Ped98, PR95b, Pol96, SBD^{+}97, WK97a]. terminology [JZV92]. terrain [TDM96]. terrestrial [AO97a, AOE99, FSKC97, Fri98, GDS98, GSS94, KB99, Mes92, NCNN97, PSL^{+}93, SGS93, VZF96, Ano99f, Str92]. territorial [HBM^{+}97]. territories [HS97b]. test [AHHK97, Dam99, KMWB98, LK99, Sog93]. Testing [GPO97, Håk97, HRWC91, Lea97, Ryk96, Ces97b, Loe97a, PS90, PG97, PG99b, SC96]. tests [RS98a, VVK97]. Tetranychidae [BM99]. Tetranychus [BM99, SSW91]. Texas [FSG96, ASTL91, GHQ99]. texture [FK99, KRD96]. Thailand [BM96]. their [ALB92b, BMHT99, Bos92, FFS96, Gam98, Jet92, KB99, MRCL92, Mar99b, Mil92, MN93, PW91, SIU97, Wan95, WCC95, vdPV99]. theorem [Pat98]. theoretic [Bat96b, McK96, PC99b]. Theoretical [Jør93d, PR95c, AG93, Aok94, EAG97, HH98, Löf96, McC97, YH99, dCF99, Leg92]. Theories [Mar99a, Mit94, Nih93]. Theory [ÅB90, Ano99d, Hul98, Jør95b, Ush91, Bac93, Bey98, Cao95, Chi98a, GPW97, GWAU99, Il’96, Jør92b, Jør97b, LW94, Loe97b, Mau97, Mül92, MWJ92, Mül97, Rob96a, Rob96b, SSW91, Sep99, Ulr92, Van96, Wis92a]. Thermaikos [PS93a]. thermal [FSA99, LP91, MT95, Sal97]. thermally [CPH99]. thermic [MT96]. thermodynamic [Bas98, CMH99]. thermodynamics [Jør94c]. thickets [SSY95]. thickness [SN99b]. thinned [Bar91]. Thomas [Ano99d, Jør94d, Svi91b, PBO^{+}96]. threatened [HSB99, VPFM97]. three [BS95a, BS95b, BS95c, FKNM^{+}97, GS96, JHM97, KSR97, MGD98, MR91b, Per98, VG98, Whi95, WVCS94, Wil96b, WTL91]. three-dimensional [BS95a, BS95b, BS95c, GS96, Wil96b, WTL91]. three-species [KSR97]. threshold [FSG96, KB97a]. Thripidae [TBG93]. thrips [TBG93]. thru [PK92]. Thysanoptera [TBG93]. tick [Lin98, TMG96, TMGS97]. tick-borne [Lin98]. tidal [KK97]. tidally [BJ94]. tilapia [Boc99, FG94]. tillage [RSKS91]. Timber [Jør95g, BSB97, dLNT98]. Time [Kir92, Phi95, SGML91, BBKV96, DFP^{+}93, Han97b, HP94, KT94, LRM93, LLF97, MGCR98, Mat99, MDT93, NN95, NJ96, NNR98, PW91, PTM91, RSC96, SS96, SB97, TD94, VU93a, Ste93]. time-dependent [BBKV96]. Time-dynamic [Kir92]. time-series [NNR98]. time-variable [VU93a]. time-varying [RSC96]. times [Sel95]. Timing [Tre91, KL98a].

Timo [Jør94f]. Timothy [Jør94d]. Tokyo [Jør97d, Jør97c, Jør97e, Jør97g]. Tolba [Jør94e]. tolerant [OA97]. tool [BZ95, BL92, GW98, LG99]. toolkit [RPJK94]. tools [ALGB99, JTS92, Kle97]. topographic [Ant98]. topped [Bot98]. Torda [PT99]. Toronto [Jør97c, Jør97h, Jør97e]. Tortricidae [RY91]. total [Ahn99, GDS98]. toxic [Cha96, Mil92, Mon98b]. toxicants [Gar90]. Toxicity [Leg97b, Gra98a]. toxicological [KB97a, RH96a]. Trace [Jør99a, PRK97]. tracer [HP94, SEJ93]. trade [BSN^{+}98, BPN99, Gra98d]. trade-offs [BSN^{+}98, BPN99]. traits [CU93]. trajectories [Rom96]. Transboundary [Jør94e, BHL96]. transect [Cra98, WCS95, WQ97]. transfer [Ces97a, Ces97b, DJ98, KS96]. Transformation [RRJ90, BH91, LG91]. transformations [HL97, RRNM97]. transgene [DCP99]. Transient [Kal99, DLN^{+}95, PSC93, VNLB96]. Transition [GS97, SN99a, CHC91, WH91]. Translating [WC96]. translocation [HS91, KH96]. Transmission [TGM95, Bar98, Loe97b]. transpiration [BILL98, CWPD95, Fri95, LC91, SSY95, Sto92]. transport [AK96, AS94, And98, BMV99, BH91, Del92, FSC^{+}97, IV91, Kim94, LLM96, LJ94, MA96b, RRJ90, Rip95, SFGH97, SADZ95, Smi95b, SRW97, SDS98, VDV95, Ver96, VKH^{+}94, WTL91]. traps [OM97a]. Tree [Jør99b, AUS96, CTL98, Che90, DPSB99, HV98, HM97a, HMB99, IPS99, KD96, KG96, KH98, KTM95, KL98b, LBV96, MI95, Osh91, Osh96, PR95a, Ped98, Sch96a, Tro90, WW92, YKH99]. tree-based [KTM95]. tree-dwelling [KD96]. tree-influence [KL98b]. tree-species [IPS99]. treedyn [Bos96]. trees [Bar98, Bye96, LS98a, PSN98, SC92, WJW99]. trembling [Hog99]. trend [BCH^{+}97, GRV93]. trends [FP93, JTS92, MCU98, vdSdVK96]. trial [CJMT96]. Tribolium [OM90, Tuz91]. Trichogramma [SY90]. trick [Bla95]. triloba [dLNT98]. trip [MS98]. triploid [Spe94]. tritrophic [BM99, GYNE99]. TROLL [Cha99]. Trophic [BHWP91, CO99, PHB90, YS98, CFC^{+}96, CL99, Han98, HPB93, HBP93, RSdP98, Svi97a, Ter93, VMSP97, VSG92]. Trophic-dynamic [YS98]. Tropical [EFGE98, BvSH91, BK91b, BPN99, Cha99, GET98, KH98, MPA97, Osh91, Osh96, PRK97, RS91, Tor96, Var93]. tropospheric [ML96]. Trotman [Jør94e]. trout [ALGB99, CR97b, Jes98, LB97b, PP95, Pow96, VJR^{+}98]. true [PS90]. trutta [ALGB99]. Trybom [TBG93]. tsetse [OM97a]. tucunare [FG94]. Tundisi [Jør94a]. tundra [GF90, Jes99]. Tunicate [Lav95]. turbulence [Bel92]. Turen [GHH^{+}96]. turnover [MJMN97, MMJ^{+}98, Par96]. twig [ASTL91]. Two [dZcZ96, AHHK97, Bot95, Cha96, Dub97a, GS98a, HBM^{+}97, LCL93, MS95, MG96a, MWK95, MCT97, PD99, PC99b, RSC96, Sal97, Sil99, SHJ95, WCC95, Wu95, iWHS97, cZdZ96]. two-dimensional [HBM^{+}97, WCC95]. two-parameter [PD99]. two-phase [LCL93, iWHS97]. two-species [Cha96, MG96a, PC99b, RSC96]. Two-variable [dZcZ96]. twospotted [BM99]. type [BEM97, MRLC96, MR91b, SL91b, Mor92]. types [KH94a]. tyrranus [Got98].

U [Jør92a, Jør96a, Leg92]. U.S. [ADF^{+}93, DSB^{+}99]. U.S.A. [Smi97b]. UFIS [KLFH96]. UK [Ano99f, Kom94, Zuc91, EMHB99]. UK£ [Mit94]. Ulrich [Bos93b]. Ulva [BCD93, CB97, CS97, FSC^{+}97, SFGH97, SBD^{+}97, SPP^{+}97]. Unbias [Smi95b, Ver96]. uncertain [McC95b, Vir96]. uncertainties [SWK93]. Uncertainty [MHB^{+}96, iT96, DCM92, GGP97, HRWC91, KH94a, Kle97, Mow97, OR99, PM96, Rec94, TVS^{+}94, VKT94, WG98, WGP96]. uncrowded [WWJ99]. uncultivated [MHA95]. undergoing [Smi97a]. underpinning [Löf96]. understanding [Bos92, CL99, CDDC96]. understory [KI93]. underwater [Sem99, WQ97]. uneven [GPT95, KBV99]. uneven-aged [GPT95, KBV99]. unfold [Han96]. unfolding [BHWP91]. ungauged [PJ99]. unguiculata [TB93]. ungulates [TWRW93]. unified [CC94, GSG96, Jør94d, LLWP95]. unionid [LD97]. unique [Wen92]. Uniqueness [Bre92]. United [IPS99]. units [EBBW98, PWU93]. University [Ano99d, Ano99f, Jør93a, Jør94d, Jør95b, Jør95c, Leg92]. unknown [AL99]. unlikely [DPSB99]. unmixing [BGL99]. unsaturated [IV91, VNLB96, VGDvdZ97]. unsteady [MA96b]. unthinned [Lea97]. update [BFK96]. updating [RS98b]. upgraded [FTH99]. upland [BHBM96]. upon [HS99, SSW91, VdKK^{+}99]. uptake [BS95c, LC97a, Par93a]. upwelling [BGF99, CV99]. urban [BEM97, DCCSL99, JTS92]. urbanizing [Jes98]. Ursus [Pat97]. urticae [BM99]. US$126 [Jør94g]. US$165.00 [Mit94]. US$169 [Jør99b]. US$24.50 [Jør91b]. US$278 [Jør94e]. US$285.50 [Nie95b]. US$29.95 [Jør91c]. US$40 [Ano99f]. US$55 [Ano99c]. US$59.95 [Wu95]. US$90 [Ano99f]. USA [Kom97, CCK92, FSG96, JRB97, TM99]. Use [AM98a, FG94, HC97c, Kat90, MDB98, SJ96, SRB99, TAM^{+}99, VF96b, iWHS97, AOD97, AO97a, Ano96r, Bár96a, BSN^{+}98, BILL98, BGTL99, BGS97, Bye91, CS96, CBC^{+}98, Cle98, Col98, Gam98, HCL94, Jør97h, KN95, LAB^{+}98, LGSF92, LRK96, Mar92, Mar97b, MG98, MJ99, MN93, RS96, Sil99, TKG97, UOB94, VK97, VF96c, VdKK^{+}99, WW92]. used [AWS98, Ano92f, Ano92g, DD̆Á97, JHM97, KMS99, Sog93]. useful [SL91a]. usefulness [RO97]. user [Bot95]. users [Bot95]. Using [And95, BSB97, DGWK97, PFP^{+}95, Smi91, AR96, Ano92s, AL99, ACS96, BKT95, BGF99, CS96, Cha99, Chi97, CPMC96, CM91, Dam99, DNR95, DA95, Dye95, FP99, FZ99, FOS95, Gae92, GLD97, HSN96, HSHS96, Hen95, Hil96, HLF99, Jar95, KT97, Kle95, KL98b, LLM99, Lea97, LKS96, MPP98, Mau97, MWK95, Mow97, NCNN97, Pow96, PT99, RPH99, RPT93, RH96a, RS98a, Sal97, SGNS91, SBW^{+}99, SM95, SNUM91, SM99, SD97, Sem99, SR95a, Sto92, SHJ95, TTT98, Tay96, VU93b, WSGR96, Ziv98]. USSR [GPO97]. utilisation [Buw91]. utility [HRWC91]. utilization [HPB93, McC92]. Utricularia [Ula95].

V [Ano99b, Ano99d, FFR99, MR91b]. v3.0 [FSKC97]. vaccination [Lin98]. Vahl [HCdVM98]. validated [FZ99, SFS97]. Validating [BK96].

Validation [BZ95, SHEM95, EH91, ET95a, FSA99, Gra95, LRK96, MB93, MKA^{+}95, MFP99, MHB^{+}96, PA96, Pow93, RL99, Ryk96, Wil99]. validations [Håk96a]. Validity [DSKM95, Ano95w]. Valley [VPFM97, CJMT96]. value [KB97b, TBM96]. values [PS90, WG98]. VanBlaricom [All91]. vapour [KW99, WBH^{+}98]. variability [AH98, And98, CPH99, CBB97, Dun99, Håk96b, LAB95, SSM^{+}93]. variable [Bak96, BFK96, GYYL96, HKV^{+}95, KVAW91, LB97a, MD97, MvdZ96, Ste95a, VU93a, WVCS94, dZcZ96]. variable-parameter [LB97a]. variables [AL91, CS96, HL97, Han97a, Kar93, SL91a, WE96, cZdZ96]. variance [Ken95]. variation [BHBM96, De 98, GHH^{+}96, Hog99, Liu96, LP91, MJMN97, Tap93, dZcZ96]. variational [Web95]. variations [AO97a, Aok90, Bac93, Mah98b, PJL^{+}96, UDVESM^{+}95]. variety [PS93b, Wil91]. various [BJ95]. Varroa [Mar98]. varying [Mar91, RSC96]. VC [Pay99]. vector [BGL99]. vegetal [dCF99]. Vegetation [BEM97, CBP95, AO97b, BCH^{+}97, BBK98, BS95c, BGS97, Foo96, GHQ99, KR96, LUC^{+}93a, LRK96, M.a99, MRL94, ML92b, MN93, OA97, Rob96a, SGNS91, Stu98, SvB96, SvB97, SvB98, Svi99, TTT98, TCC^{+}97, VZF96, VW97]. vegetative [WMRK93]. Vejle [SS98]. velocity [Sch96b]. Velzen [Jør95f]. Venice [BCB97, FKNM^{+}97, CO99, CB97, CS97, FSC^{+}97, SBD^{+}97, SPP^{+}97]. venom [BM94]. ventilated [Tho97]. ventilation [SSK^{+}91]. Veracruz [RSdP98]. Verita [Ano99e]. Verlag [Ano99b, Bos91a, Jør93c, Jør93b, Kom97, Svi91a, Wu95]. version [CKK^{+}98, Wu95]. versus [Bre92, LB97a, MWY92, YH99]. vertebrate [Hon92]. vertical [And98, GA97b, Oga96]. vespertinus [PT99]. VHC [Jør97d]. VI [AS94]. via [Mos92]. Viability [DG96, McK96, LBA^{+}95, SLP93]. viable [BW94]. Victoria [BBR99, PJL^{+}96]. Vieweg [Kom97]. Vigna [TB93]. vines [Buw91]. violating [Tun90]. Virginia [SSY95]. virginiana [IPS99]. Virtual [BWW99, PG97, WWJ99]. virulence [Sey92]. visual [AG93, EAG97, WCS95, WQ97]. vital [Rob96a, Rob96b]. Vladimirovich [Svi91a]. Voivodina [PT99]. Vol [Ano91n, Ano91i, Ano92i, Ano92k, Ano93l, Ano93h, Ano93i, Ano94h, Ano94j, Ano95n, Ano96p, Ano97l, Jør95f, Mau93, Svi91b, Ulg95, Ano90e, Ano90d, Ano91j, Ano91k, Ano91m, Ano91l, Ano92h, Ano92l, Ano92j, Ano93k, Ano93j, Ano94g, Ano94i, Ano95j, Ano95k, Ano95l, Ano95m, Ano95o, Ano96w, Ano96s, Ano96t, Ano96u, Ano96v, Ano96x, Ano96y, Ano96z, Ano96-27, Ano96-28, Ano97k, Ano97o, Ano97p, Ano97m, Ano97n]. Volterra [GQB91, Jes99, KH94b, M.a93a, RSC96, Sil95b, iWSG96]. Volterra-System [Jes99]. Volume [Ano97-67, Ano97-68, Ano95i, Ano97g, Ano97h, Ano97f, Ano97-65, Ano97-66, Ano97-64, PR95a]. volumes [Ano96p, Jør96a]. VORTEX [LBA^{+}95]. Vries [Bos91b]. vs [MK97]. vulnerability [VBL90]. Vulpes [SM99]. VULPEST [VBL90].

W [Bos91b, Jør93c, Jør93b, Jør94d, Jør96a, Jør97c, Jør97g, Nie95b, Nie95c, Zuc91]. W-3302 [Jør93c, Jør93b]. Wadden [KK97]. wading [HTW96, Wol94]. Wageningen [Bos91b, Jør99b]. walk [LLM96, YO95]. Wallace [BILL98]. Walp [TB93]. warm [FG94, SCH92]. warm-season [SCH92]. warm-water [FG94]. warming [Dye95, HMC99, Jør94b, KN95]. Washington [Jør97e, Ush91]. waste [HSN96, Jør95f, MGGSMC^{+}96, SADZ95, TH90, WGP96]. wastes [BH91, Jør94e]. wastewater [RCD98]. Water [BJ95, Kom94, MJY96, Rec94, VPFM97, vWB99, ANFM99, BJ94, BSE95, BMV99, Bou95, Bre95, CB96, CO99, CBB97, CS97, Cou91, CBP95, DA95, DCM92, EGJ95, ET95a, ET95b, EMHB99, FG94, FOS95, GBBV99, Gra95, Gra98b, HSN96, HS97a, HSXP98, Jar95, Jør94a, Jør96a, Kat90, KEAP97, KW99, Ker95, Kir99, KdH99, KRD96, KSD^{+}99, KR96, KMWB98, LK91, LHAJ99, LC91, Mah98a, MGNR91, MWK95, MEBZ96, MGGSMC^{+}96, MK95, PL96, PU97, PSK98, RC97, RCS97, RKM97, RPJK94, RGRF98, Rip95, SFGH97, SML98, SH97, SBW^{+}99, SM95, SH91, SA97, Str91, Str94b, SDS98, SKW91, SWK93, SHJ95, Szy96, TM99, VKT94, VNLB96, VKH^{+}94, VFC98, VVC99, VSG92, WL96, WMRK93, WBH^{+}98, ZH98, Kom94, McD93]. water- [FOS95]. water-balance [ET95a, ET95b]. waterfleas [HE96]. waters [CPH99, GGC98]. watershed [Jes98, TVS^{+}94, VVC99]. watersheds [BL90, KMWB98]. Waterstone [Jør94g]. wave [YKH99]. waves [ALB91, ABK^{+}92, Sch96b]. way [SL91a]. weather [Fri98, Oel95, RSS91, SR98, SVK93]. web [GYNE99, Hae91, HSU96, Jac96, Kei97, Kle95, PACS96, PW97, Spe97]. webs [Gob93, Jør97e, vdB98, Leg91a, Leg93]. Wedding [AH99]. weed [GA96, GA97b, Yam97]. weevil [LCMR96, LAB^{+}98, MRSK96]. Weibull [PD99]. weight [MGCR98]. Weighted [LA97]. Weinheim [Jør97d]. well [DPSB99]. Wellesley [Kom97]. Wenzel [Ano99b]. West [GMI95, TB93, TBG93]. Western [GGJ96, AH99, EH91, HSB99, Hog99, Osh91, US97, HCdVM98]. Westerschelde [LVH96]. wetland [LL96c, MR97, MM97, MR91a, RCD98, Shu98, Stu98]. Wetlands [AK96, CS96, Nie95b, ÖM97b, UMJ^{+}97, vdPV99, Nie95c]. whale [MK97]. wheat [AR96, BZ95, BSE95, CWPD95, KEAP97, KT94, KS91, PS93b, SRV95, SS96, SM95, Smi95a, SvdG99, WMRK93, ZH98]. wheat/sugarbeet [Smi95a]. whether [PS90]. which [Gon97]. Whippany [VU93a, VU93b]. White [XHW^{+}99, Hil96, RY91]. White-tailed [XHW^{+}99]. whitebark [KABT90]. whitei [OM90]. whole [PG99a, Sil95a]. whole-stream [PG99a]. wide [BH97, GPO97]. Wiesbaden [Kom97]. wild [DCP99, Gra94]. wildfires [RAM99]. wildland [BEM97]. wildland-urban [BEM97]. wildlife [BSB97, BHL96, GA97a, HF96, Jen96a, LBA^{+}95, NG96, Jør95g]. Wiley [Ano96o, Ano96q, Ano99c, HS94, Jør91a, Jør92a, Jør97h, Leg91b, Zuc91]. Wiley-Interscience [Leg91b]. Will [CB96]. Wind [BH97, DFHR96, NHWS91]. wind-driven [DFHR96]. Windows [Wu95, Gao96]. Winemiller [Jør97e]. winged [ÖM97b]. Winn. [Axe93]. Winter [KEAP97, SM95, Axe93, BZ95, BSE95, CL99, SRV95, SS96, TWRW93, WMRK93]. Wisniewski [Jør95d]. within [ALS97, BEM97, CA98, DCP99, Mat99, Mil92, PWU93, PS90, PGD91, Ste95b, UB97]. within- [Mat99]. within-shoot [Ste95b]. without [Ant98]. Wittum [Jør96a]. wood [KIPK99]. woody [GHQ99]. word [Ano94q, Ano94r, Ano95x, Ano95y, Ano95z, Ano95-27, Ano95-28, Ano95-29, Ano96-40, Ano96-41, Ano96-42, Ano96-43]. workbook [Jør91c]. working [DD̆Á97, PFP^{+}95]. Workshop [Ano95w, GWAU99, Jør93a, MWJ92, vGDT95]. World [Bos93b, Mau93, AYO99, Ano96p, CMH99, Jør97g, McC95b, Nie95b]. Wrobel [McD93]. Wulff [Ano91i].

X [Ano96p, Jør96a, Hug91, Mar99a, Ste93, Jør96a]. XIII [Mau93]. XRMA [KBSC98]. XV [Leg93]. xvi [All91]. xvii [Nih91]. xylose [RPM^{+}98].

Yancheng [LWMT97]. year [FRS^{+}95, GU97, JRB97, SH91]. years [Gri99a]. Yellow [BS98b]. Yellow-necked [BS98b]. Yenisey [KCP^{+}98]. yield [BG93, Jen96a, Kin95, LLM99, SDS96, SvdG99, Vol98, Mah98b]. yields [dGG98]. York [Ano91i, Ano96o, Ano96q, Ano99c, Jør91b, Jør92a, Jør94d, Jør97d, Jør97c, Jør97h, Jør97e, Jør97g, Leg91b, Leg92, Str92, Wu95, MEBZ96, Pat97]. Young [Jør95a, GS96, JRB97, KS95]. young-of-the-year [JRB97]. Yugoslavia [PT99].

zea [NHWS91]. Zepp [Ano96q]. zero [SvB97, SvB98]. zero-dimensional [SvB97, SvB98]. zeros [WCDL96]. Zone [BSN^{+}98, BSL97, BGTL99, Hil96, IV91, vdBDJ98]. zones [DLN^{+}95, Kan90]. zoobenthos [PS93a]. Zoom [JA96]. Zooplankton [GPW97, AKH99, ESD98, ND97]. Zostera [Bac93, BBR99, BCB97, CB97].

References

Aagren:1990:TMA

[ÅB90] Göran I. Ågren and Ernesto Bosatta. Theory and model or art and technology in ecology. Ecological Modelling, 50(1–3):213– 220, March 1990. CODEN ECMODT. ISSN 0304-3800 (print),

              1872-7026 (electronic). URL http://www.sciencedirect.
              com/science/article/pii/030438009090051H.

Asaeda:1997:MEM

[AB97] Takashi Asaeda and Truong Van Bon. Modelling the effects of macrophytes on algal blooming in eutrophic shallow lakes. Ecological Modelling, 104(2–3):261–287, December 8, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001294.

Antonelli:1992:LSS

[ABK^{+}92] P. L. Antonelli, R. H. Bradbury, N. D. Kazarinoff, X. Lin, and R. E. Reichelt. Large-scale starfish waves and reefal connectance. Ecological Modelling, 61(3–4):187–194, June 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380092900179.

Aparici:1996:PBR

[ACS96] Eduardo Aparici, Maria José Carmona, and Manuel Serra. Polymorphism in bisexual reproductive patterns of cyclical parthenogens. A simulation approach using a rotifer growth model. Ecological Modelling, 88(1–3):133–142, July 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095000763.

Aber:1993:SRA

[ADF^{+}93] John D. Aber, Charles Driscoll, C. Anthony Federer, Richard Lathrop, Gary Lovett, Jerry M. Melillo, Paul Steudler, and James Vogelmann. A strategy for the regional analysis of the effects of physical and chemical climate change on biogeochemical cycles in northeastern (U.S.) forests. Ecological Modelling, 67(1):37–47, May 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009390098D.

Anastacio:1999:CCRb

[AFM99] P. M. Anastácio, A. F. Frias, and J. C. Marques. CRISP (crayfish and rice integrated system of production): 1. Modelling rice (Oryza sativa) growth and production. Ecological Modelling, 123(1):17–28, November 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000691.

Aksnes:1993:TMA

[AG93] Dag L. Aksnes and Jarl Giske. A theoretical model of aquatic visual feeding. Ecological Modelling, 67(2–4):233–250, June 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009390007F.

Abrahamsson:1998:MSF

[AH98] Otto Abrahamsson and Lars Håkanson. Modelling seasonal flow variability of European rivers. Ecological Modelling, 114(1):49–58, December 1, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098001173.

Aussem:1999:WCA

[AH99] Alex Aussem and David Hill. Wedding connectionist and algorithmic modelling towards forecasting Caulerpa taxifolia development in the north-western Mediterranean Sea. Ecological Modelling, 120(2–3):225–236, August 17, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099001040.

Axelsen:1997:SPP

[AHHK97] Jørgen Aagaard Axelsen, Niels Holst, Timo Hamers, and Paul Henning Krogh. Simulations of the predator–prey interactions in a two species ecotoxicological test system. Ecological Modelling, 101(1):15–25, August 1, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097019509.

Ahn:1999:SMT

[Ahn99] Hosung Ahn. Statistical modeling of total phosphorus concentrations measured in south Florida rainfall. Ecological Modelling, 116(1):33–44, March 1, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098001598.

Aldenberg:1995:FDM

[AJK95] T. Aldenberg, J. H. Janse, and P. R. G. Kramer. Fitting the dynamic model PCLake to a multi-lake survey through Bayesian statistics. Ecological Modelling, 78(1–2):83–99, March 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094001193.

Antonelli:1993:LAP

[AK93] Peter Antonelli and Vlastimil Kr̆ivan. Large-amplitude periodic fluctuations in starfish-coral dynamics caused by feeding preferences. Ecological Modelling, 67(2–4):251–257, June 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009390008G.

Alvord:1996:AFT

[AK96] H. H. Alvord and R. H. Kadlec. Atrazine fate and transport in the des plaines wetlands. Ecological Modelling, 90(1):97–107, September 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095001506.

Akcakaya:1991:MSD

[Akç91] H. Reşit Akçakaya. A method for simulating demographic stochasticity. Ecological Modelling, 54(1–2):133–136, May 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380091901038.

Aoki:1999:PRZ

[AKH99] Ichiro Aoki, Teruhisa Komatsu, and Kangseok Hwang. Prediction of response of zooplankton biomass to climatic and oceanic changes. Ecological Modelling, 120(2–3):261–270, August 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001076.

Alymkulov:1991:ASV

[AL91] E. D. Alymkulov and N. K. Luckyanov. Approximative separability of variables in dynamic models of ecological systems. Ecological Modelling, 57(3–4):165–172, October 15, 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009190111D.

Abril:1994:MDS

[AL94] J. M. Abril and M. García León. Modelling the distribution of suspended matter and the sedimentation process in a marine environment. Ecological Modelling, 71(4):197–219, February 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094901341.

Antonic:1999:EDU

[AL99] Oleg Antonić and Tarzan Legović. Estimating the direction of an unknown air pollution source using a digital elevation model and a sample of deposition. Ecological Modelling, 124(1):85–95, December 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099001490.

Antonelli:1991:HOP

[ALB91] P. Antonelli, X. Lin, and R. H. Bradbury. A higher-order predator–prey interaction with application to observed starfish waves and cycles. Ecological Modelling, 58(1–4):323–332, November 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190043Z.

Albrecht:1992:PMF

[Alb92a] Karl-Friedrich Albrecht. Problems of modelling and forecasting on the basis of phenomenological investigations. Ecological Modelling, 63(1–4):45–69, September 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290061I.

Antonelli:1992:HCE

[ALB92b] P. Antonelli, X. Lin, and R. H. Bradbury. On Hutchinson’s competition equations and their homogenization: a higher-order principle of competitive exclusion. Ecological Modelling, 60(3–4):309–320, April 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290038G.

Aurelle:1999:MAN

[ALGB99] Didier Aurelle, Sovan Lek, Jean-Luc Giraudel, and Patrick Berrebi. Microsatellites and artificial neural networks: tools for the discrimination between natural and hatchery brown trout (Salmo trutta, L.) in Atlantic populations. Ecological Modelling, 120(2–3):313–324, August 17, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099001118.

Allen:1990:FCC

[All90] J. C. Allen. Factors contributing to chaos in population feedback systems. Ecological Modelling, 51(3–4):281–298, June 1990. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009090072O.

Allen:1991:CES

[All91] K. Radway Allen. The community ecology of sea otters: G. R. VanBlaricom and J. A. Estes (Editors). Ecological Studies, 65. Springer, Berlin, 1988. xvi + 247 pp., Hard cover, DM 158.00. ISBN 3-540-18090-7. Ecological Modelling, 54(1–2):147–149, May 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190110M.

Artois:1997:SRC

[ALS97] Marc Artois, Michel Langlais, and Christelle Suppo. Simulation of rabies control within an increasing fox population. Ecological Modelling, 97(1–2):23–34, April 15, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096000592.

Alymkulov:1991:AME

[Aly91] E. D.-A. Alymkulov. Aggregation in models of ecosystems dynamics and stability problems. Ecological Modelling, 58(1–4):383–386, November 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380091900475.

Alewell:1998:UOC

[AM98a] C. Alewell and B. Manderscheid. Use of objective criteria for the assessment of biogeochemical ecosystem models. Ecological Modelling, 107(2–3):213–224, April 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097002184.

Allen:1998:ASM

[AM98b] Micheal S. Allen and Leandro E. Miranda. An agestructured model for erratic crappie fisheries. Ecological Modelling, 107(2–3):289–303, April 1, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098000064.

Abel:1990:AFS

[AN90] D. E. Abel and B. S. Niven. Application of a formal specification language to animal ecology. I. Environment. Ecological Modelling, 50(1–3):205–212, March 1990. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009090050Q.

Anderson:1995:UPP

[And95] N. John Anderson. Using the past to predict the future: lake sediments and the modelling of limnological disturbance. Ecological Modelling, 78(1–2):149–172, March 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009400124Z.

Andersen:1996:SEM

[And96] Douglas C. Andersen. A spatially-explicit model of search path and soil disturbance by a fossorial herbivore. Ecological Modelling, 89(1–3):99–108, August 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001247.

Anderson:1998:ESV

[And98] Jon Thomas Anderson. The effect of seasonal variability on the germination and vertical transport of a cyst forming dinoflagellate, Gyrodinium sp., in the Chesapeake Bay. Ecological Modelling, 112(2–3):85–109, October 15, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009800074X.

Anastacio:1999:CCRc

[ANFM99] P. M. Anastácio, S. N. Nielsen, A. F. Frias, and J. C. Marques. CRISP (crayfish and rice integrated system of production): 4. Modelling water, algae and oxygen dynamics. Ecological Modelling, 123(1):29–40, November 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099001659.

Anastacio:1999:CCRa

[ANM99] P. M. Anastácio, S. N. Nielsen, and J. C. Marques. CRISP (crayfish and rice integrated system of production): 2. Modelling crayfish (Procambarus clarkii) population dynamics. Ecological Modelling, 123(1):5–16, November 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099001647.

Annan:1997:RPS

[Ann97] J. D. Annan. On repeated parameter sampling in Monte Carlo simulations. Ecological Modelling, 97(1–2):111–115, April 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096000804.

Annan:1999:RCP

[Ann99] J. D. Annan. Reply to “Comments on the paper: On repeated parameter sampling in Monte Carlo simulations”. Ecological Modelling, 124(2–3):255–257, December 13, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099001696.

Anonymous:1990:AIa

[Ano90a] Anonymous. Author index. Ecological Modelling, 49(3–4): 311–312, January 1990. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438009090033D.

Anonymous:1990:AIb

[Ano90b] Anonymous. Author index. Ecological Modelling, 50(4):233– 276, April 1990. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009090036G.

Anonymous:1990:AIc

[Ano90c] Anonymous. Author index. Ecological Modelling, 51(3–4):315– 316, June 1990. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009090074Q.

Anonymous:1990:CEMb

[Ano90d] Anonymous. Contents of Ecological Modelling, vol. 51. Ecological Modelling, 51(3–4):317–318, June 1990. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009090075R.

Anonymous:1990:CEMa

[Ano90e] Anonymous. Contents of ecological modelling, vol. 49. Ecological Modelling, 49(3–4):313–314, January 1990. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009090034E.

Anonymous:1990:EBa

[Ano90f] Anonymous. Editorial Board. Ecological Modelling, 50(1–3): iii, March 1990. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009090038I.

Anonymous:1990:EBb

[Ano90g] Anonymous. Editorial Board. Ecological Modelling, 51(1–2): iii, May 1990. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009090054K.

Anonymous:1990:PMb

[Ano90h] Anonymous. Pages 1–160 (May 1990). Ecological Modelling, 51(1–2):??, May 1990. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1990:PMa

[Ano90i] Anonymous. Pages 1–230 (March 1990). Ecological Modelling, 50(1–3):??, March 1990. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1990:PJa

[Ano90j] Anonymous. Pages 153–314 (January 1990). Ecological Modelling, 49(3–4):??, January 1990. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1990:PJb

[Ano90k] Anonymous. Pages 161–318 (June 1990). Ecological Modelling, 51(3–4):??, June 1990. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1990:PA

[Ano90l] Anonymous. Pages 231–328 (April 1990). Ecological Modelling, 50(4):??, April 1990. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1990:PO

[Ano90m] Anonymous. Publication overview. Ecological Modelling, 50(4): 231, April 1990. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009090035F.

Anonymous:1990:SI

[Ano90n] Anonymous. Subject index. Ecological Modelling, 50(4):277– 328, April 1990. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009090037H.

Anonymous:1991:ACP

[Ano91a] Anonymous. Announcement and call for papers. Ecological Modelling, 54(3–4):321, June 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009190082C.

Anonymous:1991:Aa

[Ano91b] Anonymous. Announcements. Ecological Modelling, 55(1– 2):153–154, July 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438009190074B.

Anonymous:1991:Ab

[Ano91c] Anonymous. Announcements. Ecological Modelling, 57(3–4): 313, October 15, 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438009190121G.

Anonymous:1991:AIa

[Ano91d] Anonymous. Author index. Ecological Modelling, 54(3–4): 322, June 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190083D.

Anonymous:1991:AIb

[Ano91e] Anonymous. Author index. Ecological Modelling, 55(3–4):321, August 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190092F.

Anonymous:1991:AIc

[Ano91f] Anonymous. Author index. Ecological Modelling, 57(3–4):314, October 15, 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190122H.

Anonymous:1991:AId

[Ano91g] Anonymous. Author index. Ecological Modelling, 58(1–4):387, November 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380091900486.

Anonymous:1991:AIe

[Ano91h] Anonymous. Author index. Ecological Modelling, 59(3– 4):301, December 15, 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380091901876.

Anonymous:1991:NAM

[Ano91i] Anonymous. Network analysis in marine ecology: Frederick Wulff, John G. Field and Kenneth H. Mann (Editors). Coastal and Estuarine Studies, formerly: Lecture Notes on Coastal and Estuarine Studies, Vol. 32. Springer, Berlin/New York, 1989. 284 pp., DM88.00. ISBN 3-540-51603-4/0-387-51603-4. Ecological Modelling, 57(3–4):297–299, October 15, 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009190119L.

Anonymous:1991:CEMa

[Ano91j] Anonymous. Contents Ecological Modelling, vol. 54. Ecological Modelling, 54(3–4):323–324, June 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009190084E.

Anonymous:1991:CEMb

[Ano91k] Anonymous. Contents ecological modelling, vol. 55. Ecological Modelling, 55(3–4):322–323, August 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009190093G.

Anonymous:1991:CEMd

[Ano91l] Anonymous. Contents ecological modelling, vol. 58. Ecological Modelling, 58(1–4):389–390, November 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380091900497.

Anonymous:1991:CEMc

[Ano91m] Anonymous. Contents of ecological modelling, vol. 57. Ecological Modelling, 57(3–4):315–316, October 15, 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009190123I.

Anonymous:1991:CEMe

[Ano91n] Anonymous. Contents of ecological modelling, vol. 59. Ecological Modelling, 59(3–4):302, December 15, 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380091901887.

Anonymous:1991:EBa

[Ano91o] Anonymous. Editorial Board. Ecological Modelling, 54(1–2): v, May 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190094H.

Anonymous:1991:EBb

[Ano91p] Anonymous. Editorial Board. Ecological Modelling, 55(1–2): iii, July 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190060E.

Anonymous:1991:EBc

[Ano91q] Anonymous. Editorial Board. Ecological Modelling, 57(1–2):iii, October 1, 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190050B.

Anonymous:1991:EBd

[Ano91r] Anonymous. Editorial Board. Ecological Modelling, 58(1–4):iii, November 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190027X.

Anonymous:1991:EBe

[Ano91s] Anonymous. Editorial Board. Ecological Modelling, 59(1– 2):iii, December 1, 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438009190124J.

Anonymous:1991:PM

[Ano91t] Anonymous. Pages 1–149 (May 1991). Ecological Modelling, 54(1–2):??, May 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1991:PDa

[Ano91u] Anonymous. Pages 1–150 (1 December 1991). Ecological Modelling, 59(1–2):??, December 1, 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1991:PJb

[Ano91v] Anonymous. Pages 1–160 (July 1991). Ecological Modelling, 55(1–2):??, July 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1991:POa

[Ano91w] Anonymous. Pages 1–163 (1 October 1991). Ecological Modelling, 57(1–2):??, October 1, 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1991:PNb

[Ano91x] Anonymous. Pages 1–390 (November 1991). Ecological Modelling, 58(1–4):??, November 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1991:PDb

[Ano91y] Anonymous. Pages 151–302 (15 December 1991). Ecological Modelling, 59(3–4):??, December 15, 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1991:PJa

[Ano91z] Anonymous. Pages 151–324 (June 1991). Ecological Modelling, 54(3–4):??, June 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1991:PA

[Ano91-27] Anonymous. Pages 161–323 (August 1991). Ecological Modelling, 55(3–4):??, August 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1991:POb

[Ano91-28] Anonymous. Pages 165–316 (15 October 1991). Ecological Modelling, 57(3–4):??, October 15, 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1991:PNa

[Ano91-29] Anonymous. Publisher’s note. Ecological Modelling, 57(3– 4):301–312, October 15, 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438009190120P.

Anonymous:1992:AIa

[Ano92a] Anonymous. Author index. Ecological Modelling, 60(3–4):321, April 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290039H.

Anonymous:1992:AIb

[Ano92b] Anonymous. Author index. Ecological Modelling, 61(3–4): 309, June 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380092900249.

Anonymous:1992:AIc

[Ano92c] Anonymous. Author index. Ecological Modelling, 62(4):329, August 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290006Z.

Anonymous:1992:AId

[Ano92d] Anonymous. Author index. Ecological Modelling, 63(1–4):333, September 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290077R.

Anonymous:1992:AIe

[Ano92e] Anonymous. Author index. Ecological Modelling, 64(4):331, November 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290030I.

Anonymous:1992:BCU

[Ano92f] Anonymous. Basic concepts used in the pinogram model. Ecological Modelling, 61(1–2):9–28, May 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290095V.

Anonymous:1992:CPD

[Ano92g] Anonymous. Collecting and processing data used to build the pinogram model. Ecological Modelling, 61(1–2):29–57, May 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009290096W.

Anonymous:1992:CEMa

[Ano92h] Anonymous. Contents of Ecological Modelling, vol. 60. Ecological Modelling, 60(3–4):322–323, April 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290040L.

Anonymous:1992:CEMb

[Ano92i] Anonymous. Contents of Ecological Modelling, vol. 61. Ecological Modelling, 61(1–2):310, May 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290100S.

Anonymous:1992:CEMe

[Ano92j] Anonymous. Contents of Ecological Modelling, vol. 64. Ecological Modelling, 64(4):332–333, November 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380092900319.

Anonymous:1992:CEMd

[Ano92k] Anonymous. Contents of ecological modelling, vol. 62. Ecological Modelling, 62(4):330–331, August 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380092900072.

Anonymous:1992:CEMc

[Ano92l] Anonymous. Contents of ecological modelling vol. 61. Ecological Modelling, 61(3–4):310, June 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290025A.

Anonymous:1992:DC

[Ano92m] Anonymous. Discussion and conclusions. Ecological Modelling, 61(1–2):96–100, May 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438009290098Y.

Anonymous:1992:EBa

[Ano92n] Anonymous. Editorial Board. Ecological Modelling, 60(1):iii, January 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380092900083.

Anonymous:1992:EBb

[Ano92o] Anonymous. Editorial Board. Ecological Modelling, 61(1–2): iii, May 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290093T.

Anonymous:1992:EBc

[Ano92p] Anonymous. Editorial Board. Ecological Modelling, 62(1–3): iii, July 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290078S.

Anonymous:1992:EBd

[Ano92q] Anonymous. Editorial Board. Ecological Modelling, 63(1–4):iii, September 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290055J.

Anonymous:1992:EBe

[Ano92r] Anonymous. Editorial Board. Ecological Modelling, 64(1):iii, October 15, 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290045G.

Anonymous:1992:ESU

[Ano92s] Anonymous. Examples of simulation using pinogram. Ecological Modelling, 61(1–2):101–147, May 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290099Z.

Anonymous:1992:I

[Ano92t] Anonymous. Introduction. Ecological Modelling, 61(1–2):1– 8, May 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290094U.

Anonymous:1992:PJc

[Ano92u] Anonymous. Pages 1–232 (July 1992). Ecological Modelling, 62(1–3):??, July 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1992:PS

[Ano92v] Anonymous. Pages 1–333 (September 1992). Ecological Modelling, 63(1–4):??, September 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1992:PO

[Ano92w] Anonymous. Pages 1–82 (15 October 1992). Ecological Modelling, 64(1):??, October 15, 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1992:PJa

[Ano92x] Anonymous. Pages 1–85 (January 1992). Ecological Modelling, 60(1):??, January 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1992:PJb

[Ano92y] Anonymous. Pages 149–310 (June 1992). Ecological Modelling, 61(3–4):??, June 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1992:PAa

[Ano92z] Anonymous. Pages 167–323 (April 1992). Ecological Modelling, 60(3–4):??, April 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1992:PAb

[Ano92-27] Anonymous. Pages 233–331 (August 1992). Ecological Modelling, 62(4):??, August 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1992:PN

[Ano92-28] Anonymous. Pages 261–333 (November 1992). Ecological Modelling, 64(4):??, November 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1992:PM

[Ano92-29] Anonymous. Pages 95–165 (March 1992). Ecological Modelling, 60(2):??, March 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1992:PSM

[Ano92-30] Anonymous. The pinogram simulation model. Ecological Modelling, 61(1–2):58–95, May 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290097X.

Anonymous:1993:A

[Ano93a] Anonymous. Announcements. Ecological Modelling, 70(3– 4):324, December 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438009390065Z.

Anonymous:1993:AIa

[Ano93b] Anonymous. Author index. Ecological Modelling, 65(3–4):309, February 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380093900927.

Anonymous:1993:AIb

[Ano93c] Anonymous. Author index. Ecological Modelling, 66(3–4):305, April 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380093901218.

Anonymous:1993:AIc

[Ano93d] Anonymous. Author index. Ecological Modelling, 67(2–4): 307, June 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390011G.

Anonymous:1993:AId

[Ano93e] Anonymous. Author index. Ecological Modelling, 68(3–4):303, August 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390024M.

Anonymous:1993:AIe

[Ano93f] Anonymous. Author index. Ecological Modelling, 69(3–4):319, October 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390034P.

Anonymous:1993:AIf

[Ano93g] Anonymous. Author index. Ecological Modelling, 70(3–4):325, December 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380093900662.

Anonymous:1993:CEMc

[Ano93h] Anonymous. Contents of ecological modelling, vol. 68. Ecological Modelling, 68(3–4):304–305, August 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390025N.

Anonymous:1993:CEMd

[Ano93i] Anonymous. Contents of ecological modelling, vol. 69. Ecological Modelling, 69(3–4):320–321, October 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390035Q.

Anonymous:1993:CEMe

[Ano93j] Anonymous. Contents of ecological modelling, vol. 70. Ecological Modelling, 70(3–4):326–327, December 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093900673.

Anonymous:1993:CEMa

[Ano93k] Anonymous. Contents of ecological modelling, vol. 65. Ecological Modelling, 65(3–4):310–311, February 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093900938.

Anonymous:1993:CEMb

[Ano93l] Anonymous. Contents of ecological modelling, vol. 67. Ecological Modelling, 67(2–4):308–309, June 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390012H.

Anonymous:1993:EBa

[Ano93m] Anonymous. Editorial Board. Ecological Modelling, 65(1–2):iii, January 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380093901229.

Anonymous:1993:EBb

[Ano93n] Anonymous. Editorial Board. Ecological Modelling, 66(1–2): iii, March 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390036R.

Anonymous:1993:EBc

[Ano93o] Anonymous. Editorial Board. Ecological Modelling, 67(1): iii, May 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380093900949.

Anonymous:1993:EBd

[Ano93p] Anonymous. Editorial Board. Ecological Modelling, 68(1–2): iii, July 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390101W.

Anonymous:1993:EBe

[Ano93q] Anonymous. Editorial Board. Ecological Modelling, 69(1–2):iii, September 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390044S.

Anonymous:1993:EBf

[Ano93r] Anonymous. Editorial Board. Ecological Modelling, 70(1–2):iii, November 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380093900684.

Anonymous:1993:OIB

[Ano93s] Anonymous. Open invitation to biosphere modellers. Ecological Modelling, 70(3–4):iii, December 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009390055W.

Anonymous:1993:PM

[Ano93t] Anonymous. Pages 1–155 (March 1993). Ecological Modelling, 66(1–2):??, March 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1993:PN

[Ano93u] Anonymous. Pages 1–157 (November 1993). Ecological Modelling, 70(1–2):??, November 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1993:PJa

[Ano93v] Anonymous. Pages 1–159 (January 1993). Ecological Modelling, 65(1–2):??, January 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1993:PS

[Ano93w] Anonymous. Pages 1–161 (September 1993). Ecological Modelling, 69(1–2):??, September 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1993:PJb

[Ano93x] Anonymous. Pages 103–309 (June 1993). Ecological Modelling, 67(2–4):??, June 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1993:PAb

[Ano93y] Anonymous. Pages 119–305 (August 1993). Ecological Modelling, 68(3–4):??, August 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1993:PAa

[Ano93z] Anonymous. Pages 157–305 (April 1993). Ecological Modelling, 66(3–4):??, April 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1993:PF

[Ano93-27] Anonymous. Pages 161–311 (February 1993). Ecological Modelling, 65(3–4):??, February 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1993:PD

[Ano93-28] Anonymous. Pages 161–327 (December 1993). Ecological Modelling, 70(3–4):??, December 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1993:PO

[Ano93-29] Anonymous. Pages 163–321 (October 1993). Ecological Modelling, 69(3–4):??, October 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1993:SAC

[Ano93-30] Anonymous. Symposium announcement and call for papers. Ecological Modelling, 68(1–2):vi, July 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390102X.

Anonymous:1994:Aa

[Ano94a] Anonymous. Announcement. Ecological Modelling, 72(1–2): 152, March 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094901511.

Anonymous:1994:Ab

[Ano94b] Anonymous. Announcements. Ecological Modelling, 74(1–2): 137, July 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094901163.

Anonymous:1994:AIa

[Ano94c] Anonymous. Author index. Ecological Modelling, 71(4):291, February 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094901406.

Anonymous:1994:AIb

[Ano94d] Anonymous. Author index. Ecological Modelling, 72(3–4):283, April 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094900914.

Anonymous:1994:AIc

[Ano94e] Anonymous. Author index. Ecological Modelling, 73(3–4): 332, June 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094900701.

Anonymous:1994:AId

[Ano94f] Anonymous. Author index. Ecological Modelling, 74(3–4):320, August 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094901325.

Anonymous:1994:CEMa

[Ano94g] Anonymous. Contents of Ecological Modelling, vol. 71. Ecological Modelling, 71(4):292–293, February 1994. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094901414.

Anonymous:1994:CEMb

[Ano94h] Anonymous. Contents of ecological modelling, vol. 72. Ecological Modelling, 72(3–4):284–285, April 1994. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094900922.

Anonymous:1994:CEMc

[Ano94i] Anonymous. Contents of ecological modelling, vol. 73. Ecological Modelling, 73(3–4):333–334, June 1994. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009490071X.

Anonymous:1994:CEMd

[Ano94j] Anonymous. Contents of ecological modelling, vol. 74. Ecological Modelling, 74(3–4):321–322, August 1994. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094901333.

Anonymous:1994:CCR

[Ano94k] Anonymous. Creation of the canopy research network. Ecological Modelling, 71(4):289, February 1994. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094901392.

Anonymous:1994:EBa

[Ano94l] Anonymous. Editorial Board. Ecological Modelling, 71(1–3):iii, January 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094900728.

Anonymous:1994:EBb

[Ano94m] Anonymous. Editorial Board. Ecological Modelling, 72(1–2): iii, March 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094901422.

Anonymous:1994:EBc

[Ano94n] Anonymous. Editorial Board. Ecological Modelling, 73(1–2): iii, May 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094900930.

Anonymous:1994:EBd

[Ano94o] Anonymous. Editorial Board. Ecological Modelling, 74(1–2): iii, July 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094901066.

Anonymous:1994:E

[Ano94p] Anonymous. Erratum. Ecological Modelling, 73(1–2):167, May 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094901058.

Anonymous:1994:KWIa

[Ano94q] Anonymous. Key word index. Ecological Modelling, 73(3–4): 331, June 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094900698.

Anonymous:1994:KWIb

[Ano94r] Anonymous. Key word index. Ecological Modelling, 74 (3–4):319, August 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380094901317.

Anonymous:1994:PMa

[Ano94s] Anonymous. Pages 1–152 (March 1994). Ecological Modelling, 72(1–2):??, March 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1994:PMb

[Ano94t] Anonymous. Pages 1–167 (May 1994). Ecological Modelling, 73(1–2):??, May 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1994:PJa

[Ano94u] Anonymous. Pages 1–195 (January 1994). Ecological Modelling, 71(1–3):??, January 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1994:PAb

[Ano94v] Anonymous. Pages 139–322 (August 1994). Ecological Modelling, 74(3–4):??, August 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1994:PAa

[Ano94w] Anonymous. Pages 153–285 (April 1994). Ecological Modelling, 72(3–4):??, April 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1994:PJb

[Ano94x] Anonymous. Pages 169–334 (June 1994). Ecological Modelling, 73(3–4):??, June 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1994:PF

[Ano94y] Anonymous. Pages 197–293 (February 1994). Ecological Modelling, 71(4):??, February 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1995:A

[Ano95a] Anonymous. Announcement. Ecological Modelling, 77(1):88, January 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900357.

Anonymous:1995:AIa

[Ano95b] Anonymous. Author index. Ecological Modelling, 77(2–3):291, February 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900756.

Anonymous:1995:AIb

[Ano95c] Anonymous. Author index. Ecological Modelling, 78(3):290, April 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900810.

Anonymous:1995:AIc

[Ano95d] Anonymous. Author index. Ecological Modelling, 79(1–3): 293, May 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900691.

Anonymous:1995:AId

[Ano95e] Anonymous. Author index. Ecological Modelling, 80(2–3): 304, July 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900454.

Anonymous:1995:AIe

[Ano95f] Anonymous. Author index. Ecological Modelling, 81(1–3):303, August 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900047.

Anonymous:1995:AIf

[Ano95g] Anonymous. Author index. Ecological Modelling, 82(3):314, November 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900152.

Anonymous:1995:AIg

[Ano95h] Anonymous. Author index. Ecological Modelling, 83(3): 406, December 15, 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095900128.

Anonymous:1995:CEM

[Ano95i] Anonymous. Contents of ecological modelling, volume 81. Ecological Modelling, 81(1–3):305–306, August 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095900055.

Anonymous:1995:CVa

[Ano95j] Anonymous. Contents of vol. 77. Ecological Modelling, 77 (2–3):293–294, February 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095900764.

Anonymous:1995:CVb

[Ano95k] Anonymous. Contents of vol. 78. Ecological Modelling, 78 (3):291–292, April 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095900829.

Anonymous:1995:CVc

[Ano95l] Anonymous. Contents of vol. 79. Ecological Modelling, 79 (1–3):295–296, May 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095900705.

Anonymous:1995:CVd

[Ano95m] Anonymous. Contents of vol. 80. Ecological Modelling, 80 (2–3):305–306, July 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095900462.

Anonymous:1995:CVe

[Ano95n] Anonymous. Contents of vol. 82. Ecological Modelling, 82 (3):315–316, November 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095900160.

Anonymous:1995:CVf

[Ano95o] Anonymous. Contents of vol. 83. Ecological Modelling, 83 (3):407–408, December 15, 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095900136.

Anonymous:1995:EBa

[Ano95p] Anonymous. Editorial Board. Ecological Modelling, 77(1):iii, January 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900314.

Anonymous:1995:EBb

[Ano95q] Anonymous. Editorial Board. Ecological Modelling, 78(1–2): iii, March 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900179.

Anonymous:1995:EBc

[Ano95r] Anonymous. Editorial Board. Ecological Modelling, 79(1–3): iii, May 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900640.

Anonymous:1995:EBd

[Ano95s] Anonymous. Editorial Board. Ecological Modelling, 80(1): iii, June 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900403.

Anonymous:1995:EBe

[Ano95t] Anonymous. Editorial Board. Ecological Modelling, 81(1–3): iii, August 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900004.

Anonymous:1995:EBf

[Ano95u] Anonymous. Editorial Board. Ecological Modelling, 82(1):iii, September 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900306.

Anonymous:1995:EBg

[Ano95v] Anonymous. Editorial Board. Ecological Modelling, 83(1–2):iii, December 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900365.

Anonymous:1995:IWV

[Ano95w] Anonymous. International Workshop on the validity of Agroecosystem Models Braunschweig, Germany, 12–14 May 1993. Ecological Modelling, 81(1–3):v, August 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095900012.

Anonymous:1995:KWIa

[Ano95x] Anonymous. Key word index. Ecological Modelling, 77(2– 3):289, February 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095900748.

Anonymous:1995:KWIb

[Ano95y] Anonymous. Key word index. Ecological Modelling, 78(3): 289, April 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900802.

Anonymous:1995:KWIc

[Ano95z] Anonymous. Key word index. Ecological Modelling, 79(1–3): 291, May 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900683.

Anonymous:1995:KWId

[Ano95-27] Anonymous. Key word index. Ecological Modelling, 80(2–3): 303, July 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900446.

Anonymous:1995:KWIe

[Ano95-28] Anonymous. Key word index. Ecological Modelling, 81 (1–3):301, August 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095900039.

Anonymous:1995:KWIf

[Ano95-29] Anonymous. Key word index. Ecological Modelling, 82(3):313, November 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900144.

Anonymous:1995:KI

[Ano95-30] Anonymous. Keyword index. Ecological Modelling, 83(3): 405, December 15, 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438009590011X.

Anonymous:1995:PS

[Ano95-31] Anonymous. Pages 1–108 (September 1995). Ecological Modelling, 82(1):??, September 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1995:PM

[Ano95-32] Anonymous. Pages 1–296 (May 1995). Ecological Modelling, 79(1–3):??, May 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1995:PJa

[Ano95-33] Anonymous. Pages 1–94 (January 1995). Ecological Modelling, 77(1):??, January 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1995:PJb

[Ano95-34] Anonymous. Pages 1–95 (June 1995). Ecological Modelling, 80(1):??, June 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1995:PO

[Ano95-35] Anonymous. Pages 109–210 (October 1995). Ecological Modelling, 82(2):??, October 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1995:PA

[Ano95-36] Anonymous. Pages 173–292 (April 1995). Ecological Modelling, 78(3):??, April 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1995:PN

[Ano95-37] Anonymous. Pages 211–316 (November 1995). Ecological Modelling, 82(3):??, November 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1995:PD

[Ano95-38] Anonymous. Pages 295–408 (15 December 1995). Ecological Modelling, 83(3):??, December 15, 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1995:PF

[Ano95-39] Anonymous. Pages 95–294 (February 1995). Ecological Modelling, 77(2–3):??, February 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1995:PJc

[Ano95-40] Anonymous. Pages 97–306 (July 1995). Ecological Modelling, 80(2–3):??, July 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1995:P

[Ano95-41] Anonymous. Preface. Ecological Modelling, 81(1–3):1, August 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900020.

Anonymous:1996:Aa

[Ano96a] Anonymous. Announcement. Ecological Modelling, 86(2–3): 305, May 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096900241.

Anonymous:1996:Ab

[Ano96b] Anonymous. Announcement. Ecological Modelling, 90(2):185, October 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096900733.

Anonymous:1996:Ac

[Ano96c] Anonymous. Announcement. Ecological Modelling, 90(3): 285, November 1, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096900988.

Anonymous:1996:Ad

[Ano96d] Anonymous. Announcement. Ecological Modelling, 91(1– 3):295, November 15, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096900447.

Anonymous:1996:AIa

[Ano96e] Anonymous. Author index. Ecological Modelling, 84(1–3):329, January 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096900071.

Anonymous:1996:AIb

[Ano96f] Anonymous. Author index. Ecological Modelling, 85(2–3):312, March 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096900605.

Anonymous:1996:AIc

[Ano96g] Anonymous. Author index. Ecological Modelling, 86(2–3):308– 309, May 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096900265.

Anonymous:1996:AId

[Ano96h] Anonymous. Author index. Ecological Modelling, 87(1–3): 305, June 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096900411.

Anonymous:1996:AIe

[Ano96i] Anonymous. Author index. Ecological Modelling, 88(1–3): 310, July 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096900708.

Anonymous:1996:AIf

[Ano96j] Anonymous. Author index. Ecological Modelling, 89(1–3):293, August 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096900563.

Anonymous:1996:AIg

[Ano96k] Anonymous. Author index. Ecological Modelling, 90(3): 295, November 1, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096901003.

Anonymous:1996:AIh

[Ano96l] Anonymous. Author index. Ecological Modelling, 91(1– 3):299, November 15, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096900460.

Anonymous:1996:AIi

[Ano96m] Anonymous. Author index. Ecological Modelling, 92(2–3):291, December 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096901088.

Anonymous:1996:AIj

[Ano96n] Anonymous. Author index. Ecological Modelling, 93(1– 3):295, December 16, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096900927.

Anonymous:1996:CCF

[Ano96o] Anonymous. Conservation: C. F. Jordan. John Wiley, New York, 1995. US $29.95, ISBN 0-471-59515-2, 340 pp. Ecological Modelling, 87(1–3):297–298, June 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380096881671.

Anonymous:1996:DCE

[Ano96p] Anonymous. Dry coastal ecosystems. Ecosystems of the world, volumes 2A and 2B: E. van der Maarel (Editor). Elsevier, Amsterdam, 1993. Vol. 2A: Dfl. 430, ISBN 0-444-87348-1, 600 pp. Vol. 2B: Dfl. 470, ISBN 0-444-87349-X, 664 pp. Ecological Modelling, 87(1–3):300, June 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380096881701.

Anonymous:1996:RNO

[Ano96q] Anonymous. The role of nonliving organic matter in the earth’s carbon cycle: R. G. Zepp and Ch. Sonntag (Editors). John Wiley, New York, 1995. US $70, ISBN 0-471-95463-2, 342 pp. Ecological Modelling, 87(1–3):298, June 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380096881683.

Anonymous:1996:TGP

[Ano96r] Anonymous. Toward global planning of sustainable use of the earth. Development of global eco-engineering: S. Murai (Editor). Elsevier, Amsterdam, 1995. Dfl. 325, ISBN 0-444-81904-5, 434 pp. Ecological Modelling, 87(1–3):299, June 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380096881695.

Anonymous:1996:CVa

[Ano96s] Anonymous. Contents of vol. 84. Ecological Modelling, 84 (1–3):331–332, January 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096900083.

Anonymous:1996:CVb

[Ano96t] Anonymous. Contents of vol. 85. Ecological Modelling, 85(2–3):313–314, March 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096900617.

Anonymous:1996:CVc

[Ano96u] Anonymous. Contents of vol. 86. Ecological Modelling, 86 (2–3):310–311, May 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096900277.

Anonymous:1996:CVd

[Ano96v] Anonymous. Contents of vol. 87. Ecological Modelling, 87(1–3):307–308, June 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096900423.

Anonymous:1996:CVDa

[Ano96w] Anonymous. Contents of vol. 88lysis of daily rainfall data. Ecological Modelling, 88(1–3):311–312, July 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009690071X.

Anonymous:1996:CVe

[Ano96x] Anonymous. Contents of vol. 89. Ecological Modelling, 89 (1–3):295–296, August 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096900575.

Anonymous:1996:CVf

[Ano96y] Anonymous. Contents of vol. 90. Ecological Modelling, 90 (3):297–298, November 1, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096901015.

Anonymous:1996:CVg

[Ano96z] Anonymous. Contents of vol. 91. Ecological Modelling, 91(1– 3):300–301, November 15, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096900472.

Anonymous:1996:CVh

[Ano96-27] Anonymous. Contents of vol. 92. Ecological Modelling, 92 (2–3):293–294, December 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438009690109X.

Anonymous:1996:CVi

[Ano96-28] Anonymous. Contents of vol. 93. Ecological Modelling, 93(1– 3):296–297, December 16, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096900939.

Anonymous:1996:EBa

[Ano96-29] Anonymous. Editorial Board. Ecological Modelling, 84(1–3):iii, January 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096900058.

Anonymous:1996:EBb

[Ano96-30] Anonymous. Editorial Board. Ecological Modelling, 85(1):iii, February 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096900526.

Anonymous:1996:EBc

[Ano96-31] Anonymous. Editorial Board. Ecological Modelling, 86(1):iii, April 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096900046.

Anonymous:1996:EBd

[Ano96-32] Anonymous. Editorial Board. Ecological Modelling, 87(1–3): iii, June 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096900393.

Anonymous:1996:EBe

[Ano96-33] Anonymous. Editorial Board. Ecological Modelling, 88(1–3): iii, July 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438009690068X.

Anonymous:1996:EBf

[Ano96-34] Anonymous. Editorial Board. Ecological Modelling, 89(1–3): iii, August 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438009690054X.

Anonymous:1996:EBg

[Ano96-35] Anonymous. Editorial Board. Ecological Modelling, 90(1):iii, September 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096900289.

Anonymous:1996:EBh

[Ano96-36] Anonymous. Editorial Board. Ecological Modelling, 91 (1–3):iii, November 15, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096900435.

Anonymous:1996:EBi

[Ano96-37] Anonymous. Editorial Board. Ecological Modelling, 92(1): iii, November 21, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096900290.

Anonymous:1996:EBj

[Ano96-38] Anonymous. Editorial Board. Ecological Modelling, 93 (1–3):iii, December 16, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096900885.

Anonymous:1996:E

[Ano96-39] Anonymous. Erratum. Ecological Modelling, 93(1–3):291, December 16, 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096900903.

Anonymous:1996:KWIa

[Ano96-40] Anonymous. Key word index. Ecological Modelling, 84 (1–3):327–328, January 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438009690006X.

Anonymous:1996:KWIb

[Ano96-41] Anonymous. Key word index. Ecological Modelling, 85(2– 3):310–311, March 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096900599.

Anonymous:1996:KWIc

[Ano96-42] Anonymous. Key word index. Ecological Modelling, 86(2– 3):306–307, May 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096900253.

Anonymous:1996:KWId

[Ano96-43] Anonymous. Key word index. Ecological Modelling, 87(1–3): 303, June 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438009690040X.

Anonymous:1996:KIa

[Ano96-44] Anonymous. Keyword index. Ecological Modelling, 88(1–3): 309, July 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096900691.

Anonymous:1996:KIb

[Ano96-45] Anonymous. Keyword index. Ecological Modelling, 89(1– 3):291, August 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096900551.

Anonymous:1996:KIc

[Ano96-46] Anonymous. Keyword index. Ecological Modelling, 90(3): 293, November 1, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438009690099X.

Anonymous:1996:KId

[Ano96-47] Anonymous. Keyword index. Ecological Modelling, 91(1–3): 297–298, November 15, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096900459.

Anonymous:1996:KIe

[Ano96-48] Anonymous. Keyword index. Ecological Modelling, 92(2– 3):289–290, December 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096901076.

Anonymous:1996:KIf

[Ano96-49] Anonymous. Keyword index. Ecological Modelling, 93(1– 3):293–294, December 16, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096900915.

Anonymous:1996:PS

[Ano96-50] Anonymous. Pages 1–107 (September 1996). Ecological Modelling, 90(1):??, September 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1996:PNc

[Ano96-51] Anonymous. Pages 1–108 (21 November 1996). Ecological Modelling, 92(1):??, November 21, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1996:PAb

[Ano96-52] Anonymous. Pages 1–111 (April 1996). Ecological Modelling, 86(1):??, April 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1996:PAc

[Ano96-53] Anonymous. Pages 1–296 (August 1996). Ecological Modelling, 89(1–3):??, August 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1996:PD

[Ano96-54] Anonymous. Pages 1–298 (16 December 1996). Ecological Modelling, 93(1–3):??, December 16, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1996:PNb

[Ano96-55] Anonymous. Pages 1–302 (15 November 1996). Ecological Modelling, 91(1–3):??, November 15, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1996:PJb

[Ano96-56] Anonymous. Pages 1–308 (June 1996). Ecological Modelling, 87(1–3):??, June 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1996:PJc

[Ano96-57] Anonymous. Pages 1–313 (July 1996). Ecological Modelling, 88(1–3):??, July 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1996:PJa

[Ano96-58] Anonymous. Pages 1–332 (January 1996). Ecological Modelling, 84(1–3):??, January 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1996:PNa

[Ano96-59] Anonymous. Pages 187–298 (1 November 1996). Ecological Modelling, 90(3):??, November 1, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1996:PM

[Ano96-60] Anonymous. Pages 99–314 (March 1996). Ecological Modelling, 85(2–3):??, March 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1996:PAa

[Ano96-61] Anonymous. Publisher’s acknowledgement. Ecological Modelling, 85(2–3):309, March 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096900587.

Anonymous:1997:AIa

[Ano97a] Anonymous. Author index. Ecological Modelling, 94(2–3):317, January 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097839178.

Anonymous:1997:AIb

[Ano97b] Anonymous. Author index. Ecological Modelling, 95(2–3): 329, February 14, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097839142.

Anonymous:1997:AIc

[Ano97c] Anonymous. Author index. Ecological Modelling, 96(1–3):311, March 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097879665.

Anonymous:1997:AId

[Ano97d] Anonymous. Author index. Ecological Modelling, 99(2–3):293, June 30, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097839312.

Anonymous:1997:AIe

[Ano97e] Anonymous. Author index. Ecological Modelling, 103(2– 3):289, November 17, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438009788177X.

Anonymous:1997:AIVc

[Ano97f] Anonymous. Author index volume 101 (1997). Ecological Modelling, 101(2–3):373, August 15, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097868494.

Anonymous:1997:AIVa

[Ano97g] Anonymous. Author index volume 97 (1997). Ecological Modelling, 97(3):253, April 28, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097839348.

Anonymous:1997:AIVb

[Ano97h] Anonymous. Author index volume 98 (1997). Ecological Modelling, 98(2–3):249–251, May 30, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097839269.

Anonymous:1997:BR

[Ano97i] Anonymous. Book review. Ecological Modelling, 103(1): 103–104, November 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097001130.

Anonymous:1997:Ca

[Ano97j] Anonymous. Contents. Ecological Modelling, 98(1):ix, May 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097839300.

Anonymous:1997:CEMa

[Ano97k] Anonymous. Contents of ecological modelling, vol. 95. Ecological Modelling, 95(2–3):333–334, February 14, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097839166.

Anonymous:1997:CEMb

[Ano97l] Anonymous. Contents of ecological modelling, vol. 99. Ecological Modelling, 99(2–3):297, June 30, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097839336.

Anonymous:1997:CVc

[Ano97m] Anonymous. Contents of vol. 101 101 (1997). Ecological Modelling, 101(2–3):377–378, August 15, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097868512.

Anonymous:1997:CVd

[Ano97n] Anonymous. Contents of vol. 103. Ecological Modelling, 103(2–3):293, November 17, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097881793.

Anonymous:1997:CVa

[Ano97o] Anonymous. Contents of vol. 94. Ecological Modelling, 94 (2–3):321–322, January 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097839191.

Anonymous:1997:CVb

[Ano97p] Anonymous. Contents of vol. 96. Ecological Modelling, 96 (1–3):315–316, March 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097900077.

Anonymous:1997:Cb

[Ano97q] Anonymous. Correspondence. Ecological Modelling, 102(2– 3):375–377, October 29, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097000811.

Anonymous:1997:Ea

[Ano97r] Anonymous. Editorial. Ecological Modelling, 100(1–3):1–4, December 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001488.

Anonymous:1997:EBa

[Ano97s] Anonymous. Editorial Board. Ecological Modelling, 94(1):iii, January 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097900053.

Anonymous:1997:EBb

[Ano97t] Anonymous. Editorial Board. Ecological Modelling, 95(1):iii, February 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097900119.

Anonymous:1997:EBc

[Ano97u] Anonymous. Editorial Board. Ecological Modelling, 96(1–3):iii, March 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097900065.

Anonymous:1997:EBd

[Ano97v] Anonymous. Editorial Board. Ecological Modelling, 97(1–2):iii, April 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097900156.

Anonymous:1997:EBe

[Ano97w] Anonymous. Editorial Board. Ecological Modelling, 98(1): iii, May 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097900107.

Anonymous:1997:EBf

[Ano97x] Anonymous. Editorial Board. Ecological Modelling, 99(1):iii, June 16, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097900089.

Anonymous:1997:EBg

[Ano97y] Anonymous. Editorial Board. Ecological Modelling, 101(1):iii, August 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097900004.

Anonymous:1997:EBh

[Ano97z] Anonymous. Editorial Board. Ecological Modelling, 102(1):iii, October 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097900041.

Anonymous:1997:Ec

[Ano97-27] Anonymous. Erratum. Ecological Modelling, 102(2–3):379, October 29, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001166.

Anonymous:1997:Ed

[Ano97-28] Anonymous. Erratum. Ecological Modelling, 104(2–3):307– 308, December 8, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097001592.

Anonymous:1997:Ia

[Ano97-29] Anonymous. Index. Ecological Modelling, 100(1–3):171–197, December 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438009700207X.

Anonymous:1997:Ib

[Ano97-30] Anonymous. Index. Ecological Modelling, 100(1–3):199–223, December 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097002081.

Anonymous:1997:Ic

[Ano97-31] Anonymous. Index. Ecological Modelling, 100(1–3):225, December 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097002093.

Anonymous:1997:Id

[Ano97-32] Anonymous. Index. Ecological Modelling, 100(1–3):227, December 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438009700210X.

Anonymous:1997:Ie

[Ano97-33] Anonymous. Index. Ecological Modelling, 102(2–3):381, October 29, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001440.

Anonymous:1997:If

[Ano97-34] Anonymous. Index. Ecological Modelling, 102(2–3):383–384, October 29, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001452.

Anonymous:1997:Ig

[Ano97-35] Anonymous. Index. Ecological Modelling, 102(2–3):385–386, October 29, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001464.

Anonymous:1997:Ih

[Ano97-36] Anonymous. Index. Ecological Modelling, 104(2–3):313, December 8, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001932.

Anonymous:1997:Ii

[Ano97-37] Anonymous. Index. Ecological Modelling, 104(2–3):315– 316, December 8, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097001944.

Anonymous:1997:Ij

[Ano97-38] Anonymous. Index. Ecological Modelling, 104(2–3):317, December 8, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001956.

Anonymous:1997:PNa

[Ano97-39] Anonymous. Pages 1–104 (1 November 1997). Ecological Modelling, 103(1):??, November 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1997:PDb

[Ano97-40] Anonymous. Pages 1–112 (1 December 1997). Ecological Modelling, 104(1):??, December 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1997:PFa

[Ano97-41] Anonymous. Pages 1–118 (1 February 1997). Ecological Modelling, 95(1):??, February 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1997:PDd

[Ano97-42] Anonymous. Pages 1–128 (14 December 1997). Ecological Modelling, 105(1):??, December 14, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1997:PAc

[Ano97-43] Anonymous. Pages 1–137 (1 August 1997). Ecological Modelling, 101(1):??, August 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1997:PAa

[Ano97-44] Anonymous. Pages 1–151 (15 April 1997). Ecological Modelling, 97(1–2):??, April 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1997:PDa

[Ano97-45] Anonymous. Pages 1–228 (December 1997). Ecological Modelling, 100(1–3):??, December 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1997:PMa

[Ano97-46] Anonymous. Pages 1–317 (1 March 1997). Ecological Modelling, 96(1–3):??, March 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1997:PJb

[Ano97-47] Anonymous. Pages 1–97 (16 June 1997). Ecological Modelling, 99(1):??, June 16, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1997:PJc

[Ano97-48] Anonymous. Pages 101–297 (30 June 1997). Ecological Modelling, 99(2–3):??, June 30, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1997:PNb

[Ano97-49] Anonymous. Pages 105–293 (17 November 1997). Ecological Modelling, 103(2–3):??, November 17, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1997:PDc

[Ano97-50] Anonymous. Pages 113–318 (8 December 1997). Ecological Modelling, 104(2–3):??, December 8, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1997:PFb

[Ano97-51] Anonymous. Pages 119–334 (14 February 1997). Ecological Modelling, 95(2–3):??, February 14, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1997:PAe

[Ano97-52] Anonymous. Pages 139–378 (15 August 1997). Ecological Modelling, 101(2–3):??, August 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1997:PAb

[Ano97-53] Anonymous. Pages 153–258 (28 April 1997). Ecological Modelling, 97(3):??, April 28, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1997:PO

[Ano97-54] Anonymous. Pages 175–386 (29 October 1997). Ecological Modelling, 102(2–3):??, October 29, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1997:PMb

[Ano97-55] Anonymous. Pages 89–255 (30 May 1997). Ecological Modelling, 98(2–3):??, May 30, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1997:PJa

[Ano97-56] Anonymous. Pages 95–322 (15 January 1997). Ecological Modelling, 94(2–3):??, January 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1997:P

[Ano97-57] Anonymous. Preface. Ecological Modelling, 98(1):vii–viii, May 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097839294.

Anonymous:1997:PAd

[Ano97-58] Anonymous. Publisher’s acknowledgement. Ecological Modelling, 101(1):137, August 1, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097900016.

Anonymous:1997:SIa

[Ano97-59] Anonymous. Subject index. Ecological Modelling, 94(2– 3):319–320, January 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438009783918X.

Anonymous:1997:SIb

[Ano97-60] Anonymous. Subject index. Ecological Modelling, 95(2–3): 331–332, February 14, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097839154.

Anonymous:1997:SIc

[Ano97-61] Anonymous. Subject index. Ecological Modelling, 96(1– 3):313–314, March 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097879677.

Anonymous:1997:SId

[Ano97-62] Anonymous. Subject index. Ecological Modelling, 99(2–3): 295–296, June 30, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097839324.

Anonymous:1997:SIe

[Ano97-63] Anonymous. Subject index. Ecological Modelling, 103(2–3): 291–292, November 17, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097881781.

Anonymous:1997:SIVc

[Ano97-64] Anonymous. Subject index volume 101 (1997). Ecological Modelling, 101(2–3):375–376, August 15, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097868500.

Anonymous:1997:SIVa

[Ano97-65] Anonymous. Subject index volume 97 (1997). Ecological Modelling, 97(3):255–256, April 28, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438009783935X.

Anonymous:1997:SIVb

[Ano97-66] Anonymous. Subject index volume 98 (1997). Ecological Modelling, 98(2–3):253–254, May 30, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097839270.

Anonymous:1997:VCVa

[Ano97-67] Anonymous. Volume contents volume 97 (1997). Ecological Modelling, 97(3):257–258, April 28, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097839361.

Anonymous:1997:VCVb

[Ano97-68] Anonymous. Volume contents volume 98 (1997)DM98. Ecological Modelling, 98(2–3):255, May 30, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097839282.

Anonymous:1998:BR

[Ano98a] Anonymous. Book review. Ecological Modelling, 109(1): 115–119, June 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380098000465.

Anonymous:1998:Ca

[Ano98b] Anonymous. Correspondence. Ecological Modelling, 106 (1):93–95, February 16, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097001713.

Anonymous:1998:Cb

[Ano98c] Anonymous. Correspondence. Ecological Modelling, 106 (1):95–96, February 16, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097001725.

Anonymous:1998:Cc

[Ano98d] Anonymous. Correspondence. Ecological Modelling, 106 (1):96–98, February 16, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097001737.

Anonymous:1998:Cd

[Ano98e] Anonymous. Correspondence. Ecological Modelling, 106(1): 98, February 16, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097002068.

Anonymous:1998:Ea

[Ano98f] Anonymous. Editorial. Ecological Modelling, 108(1–3):1–2, May 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098000143.

Anonymous:1998:Eb

[Ano98g] Anonymous. Editorial. Ecological Modelling, 110(1):1–4, July 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098000994.

Anonymous:1998:Ed

[Ano98h] Anonymous. Editorial. Ecological Modelling, 113(1–3): 1, November 2, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438009800129X.

Anonymous:1998:Ec

[Ano98i] Anonymous. Erratum. Ecological Modelling, 111(1):99, August 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098001094.

Anonymous:1998:FLA

[Ano98j] Anonymous. Full length article. Ecological Modelling, 108 (1–3):53–65, May 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380098000180.

Anonymous:1998:Ia

[Ano98k] Anonymous. Index. Ecological Modelling, 105(2–3):353, January 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097101934.

Anonymous:1998:Ib

[Ano98l] Anonymous. Index. Ecological Modelling, 105(2–3):355–356, January 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097101946.

Anonymous:1998:Ic

[Ano98m] Anonymous. Index. Ecological Modelling, 106(2–3):313, March 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098000441.

Anonymous:1998:Id

[Ano98n] Anonymous. Index. Ecological Modelling, 106(2–3):315–316, March 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098000453.

Anonymous:1998:Ie

[Ano98o] Anonymous. Index. Ecological Modelling, 107(2–3):305, April 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098000556.

Anonymous:1998:If

[Ano98p] Anonymous. Index. Ecological Modelling, 107(2–3):307–308, April 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098000568.

Anonymous:1998:Ig

[Ano98q] Anonymous. Index. Ecological Modelling, 108(1–3):279, May 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098000702.

Anonymous:1998:Ih

[Ano98r] Anonymous. Index. Ecological Modelling, 108(1–3):281–282, May 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098000714.

Anonymous:1998:Ii

[Ano98s] Anonymous. Index. Ecological Modelling, 109(3):325, June 22, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098000970.

Anonymous:1998:Ij

[Ano98t] Anonymous. Index. Ecological Modelling, 109(3):327–328, June 22, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098000982.

Anonymous:1998:Ik

[Ano98u] Anonymous. Index. Ecological Modelling, 110(3):307, July 22, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098001070.

Anonymous:1998:Il

[Ano98v] Anonymous. Index. Ecological Modelling, 110(3):309–310, July 22, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098001082.

Anonymous:1998:Im

[Ano98w] Anonymous. Index. Ecological Modelling, 111(2–3):301, September 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380098001513.

Anonymous:1998:In

[Ano98x] Anonymous. Index. Ecological Modelling, 111(2–3):303– 304, September 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380098001525.

Anonymous:1998:Io

[Ano98y] Anonymous. Index. Ecological Modelling, 112(2–3):251, October 15, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098001665.

Anonymous:1998:Ip

[Ano98z] Anonymous. Index. Ecological Modelling, 112(2–3):253, October 15, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098001677.

Anonymous:1998:Iq

[Ano98-27] Anonymous. Index. Ecological Modelling, 113(1–3):241, November 2, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438009800177X.

Anonymous:1998:Ir

[Ano98-28] Anonymous. Index. Ecological Modelling, 113(1–3):243– 244, November 2, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380098001781.

Anonymous:1998:PJe

[Ano98-29] Anonymous. Pages 1–104 (1 July 1998). Ecological Modelling, 110(1):??, July 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1998:PMb

[Ano98-30] Anonymous. Pages 1–104 (16 March 1998). Ecological Modelling, 107(1):??, March 16, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1998:PAb

[Ano98-31] Anonymous. Pages 1–106 (1 August 1998). Ecological Modelling, 111(1):??, August 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1998:PF

[Ano98-32] Anonymous. Pages 1–106 (16 February 1998). Ecological Modelling, 106(1):??, February 16, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1998:PD

[Ano98-33] Anonymous. Pages 1–112 (1 December 1998). Ecological Modelling, 114(1):??, December 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1998:PJb

[Ano98-34] Anonymous. Pages 1–120 (1 June 1998). Ecological Modelling, 109(1):??, June 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1998:PN

[Ano98-35] Anonymous. Pages 1–252 (2 November 1998). Ecological Modelling, 113(1–3):??, November 2, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1998:PMc

[Ano98-36] Anonymous. Pages 1–290 (1 May 1998). Ecological Modelling, 108(1–3):??, May 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1998:POa

[Ano98-37] Anonymous. Pages 1–80 (1 October 1998). Ecological Modelling, 112(1):??, October 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1998:PJf

[Ano98-38] Anonymous. Pages 105–210 (11 July 1998). Ecological Modelling, 110(2):??, July 11, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1998:PAa

[Ano98-39] Anonymous. Pages 105–310 (1 April 1998). Ecological Modelling, 107(2–3):??, April 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1998:PS

[Ano98-40] Anonymous. Pages 107–306 (1 September 1998). Ecological Modelling, 111(2–3):??, September 1, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1998:PMa

[Ano98-41] Anonymous. Pages 107–317 (1 March 1998). Ecological Modelling, 106(2–3):??, March 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1998:PJc

[Ano98-42] Anonymous. Pages 121–224 (11 June 1998). Ecological Modelling, 109(2):??, June 11, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1998:PJa

[Ano98-43] Anonymous. Pages 129–358 (1 January 1998). Ecological Modelling, 105(2–3):??, January 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1998:PJg

[Ano98-44] Anonymous. Pages 211–312 (22 July 1998). Ecological Modelling, 110(3):??, July 22, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1998:PJd

[Ano98-45] Anonymous. Pages 225–330 (22 June 1998). Ecological Modelling, 109(3):??, June 22, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1998:POb

[Ano98-46] Anonymous. Pages 81–256 (15 October 1998). Ecological Modelling, 112(2–3):??, October 15, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1999:BR

[Ano99a] Anonymous. Book review. Ecological Modelling, 115(1):99– 100, February 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380098001720.

Anonymous:1999:ESA

[Ano99b] Anonymous. Earth System Analysis. Integrating Science for Sustainability, Edited by H. J. Schnellnhuber and V. Wenzel, Springer Verlag, 1998, 530 pages. ISBN. 3-540-58017-

4. 92 figures, several in colour, hard cover. DM 128, US $80.00. Ecological Modelling, 118(1):87, June 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000290.

Anonymous:1999:ENE

[Ano99c] Anonymous. Ecological Numeracy, Edited by Robert A. Herendeen, John Wiley, New York, 331 pages. ISBN 0-471-183 091. Soft bound. £25 British pounds = approximately US$55. Ecological Modelling, 118(1):89–90, June 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099000320.

Anonymous:1999:EST

[Ano99d] Anonymous. Ecological Scale, Theory and Application, Edited by David L. Peterson and V. Thomas Parker, Columbia University Press, 1998, 615 pages. ISBN 0-231-10502-9 cloth, ISBN 0-231-10503-7 paper. Domestic price cloth, $60.00; paper, $35.00. International price cloth, $69.00; paper, $40.50. Ecological Modelling, 118(1):88, June 1, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099000307.

Anonymous:1999:FSA

[Ano99e] Anonymous. Forestry Sciences: Assessment of Biodiversity for Improved Forest Planning. Proceedings of the Conference on Assessment of Biodiversity for Improved Forest Planning, 7–11 October 1996, held in Monte Verita, Switzerland. Edited by Peter Bachman, Michael Köhl and Risto Pälvinen. Kluwer Academic Publishers, 1998. ISBN: 0-7923-4872-9. 444 pages hardbound, $169. Ecological Modelling, 118(1):90, June 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099000332.

Anonymous:1999:TEC

[Ano99f] Anonymous. Terrestrial Ecosystems in Changing Environments, Edited by Herman H. Shugart, Cambridge University Press Cambridge, UK, 1998, 537 pages. ISBN 0-521-56523-5 (paper back, US$40), ISBN 0-521-56342-9 (hard back edition, US$90). Ecological Modelling, 118(1):89, June 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099000319.

Anonymous:1999:Ea

[Ano99g] Anonymous. Editorial. Ecological Modelling, 115(2–3):109– 110, February 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380098001847.

Anonymous:1999:Eb

[Ano99h] Anonymous. Editorial. Ecological Modelling, 122(3):135–137, October 20, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001337.

Anonymous:1999:Ec

[Ano99i] Anonymous. Editorial. Ecological Modelling, 123(1):1– 3, November 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380099001635.

Anonymous:1999:Ia

[Ano99j] Anonymous. Index. Ecological Modelling, 114(2–3):313, January 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098002075.

Anonymous:1999:Ib

[Ano99k] Anonymous. Index. Ecological Modelling, 114(2–3):315–316, January 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098002087.

Anonymous:1999:Ic

[Ano99l] Anonymous. Index. Ecological Modelling, 115(2–3):285, February 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099000022.

Anonymous:1999:Id

[Ano99m] Anonymous. Index. Ecological Modelling, 115(2–3):287– 288, February 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380099000034.

Anonymous:1999:Ie

[Ano99n] Anonymous. Index. Ecological Modelling, 116(2–3):305, March 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099000265.

Anonymous:1999:If

[Ano99o] Anonymous. Index. Ecological Modelling, 116(2–3):307–308, March 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099000277.

Anonymous:1999:Ig

[Ano99p] Anonymous. Index. Ecological Modelling, 117(2–3):373, May 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099000757.

Anonymous:1999:Ih

[Ano99q] Anonymous. Index. Ecological Modelling, 117(2–3):375–376, May 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099000769.

Anonymous:1999:Ii

[Ano99r] Anonymous. Index. Ecological Modelling, 118(2–3):279, June 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099000848.

Anonymous:1999:Ij

[Ano99s] Anonymous. Index. Ecological Modelling, 118(2–3):281–282, June 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438009900085X.

Anonymous:1999:Ik

[Ano99t] Anonymous. Index. Ecological Modelling, 119(2–3):277, July 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099001167.

Anonymous:1999:Il

[Ano99u] Anonymous. Index. Ecological Modelling, 119(2–3):279–280, July 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001179.

Anonymous:1999:Im

[Ano99v] Anonymous. Index. Ecological Modelling, 120(2–3):361, August 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001532.

Anonymous:1999:In

[Ano99w] Anonymous. Index. Ecological Modelling, 120(2–3):363–364, August 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001544.

Anonymous:1999:Io

[Ano99x] Anonymous. Index. Ecological Modelling, 121(2–3):297, September 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380099001672.

Anonymous:1999:Ip

[Ano99y] Anonymous. Index. Ecological Modelling, 121(2–3):299– 300, September 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380099001684.

Anonymous:1999:Iq

[Ano99z] Anonymous. Index. Ecological Modelling, 122(3):307, October 20, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001830.

Anonymous:1999:Ir

[Ano99-27] Anonymous. Index. Ecological Modelling, 122(3):309–310, October 20, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001842.

Anonymous:1999:Is

[Ano99-28] Anonymous. Index. Ecological Modelling, 123(2–3):243, November 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380099001970.

Anonymous:1999:It

[Ano99-29] Anonymous. Index. Ecological Modelling, 123(2–3):245– 246, November 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380099001982.

Anonymous:1999:Iu

[Ano99-30] Anonymous. Index. Ecological Modelling, 124(2–3):259, December 13, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099002033.

Anonymous:1999:Iv

[Ano99-31] Anonymous. Index. Ecological Modelling, 124(2–3):261– 262, December 13, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380099002045.

Anonymous:1999:PSa

[Ano99-32] Anonymous. Pages 1–102 (1 September 1999). Ecological Modelling, 121(1):??, September 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1999:PMa

[Ano99-33] Anonymous. Pages 1–106 (1 March 1999). Ecological Modelling, 116(1):??, March 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1999:PFa

[Ano99-34] Anonymous. Pages 1–108 (1 February 1999). Ecological Modelling, 115(1):??, February 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1999:POa

[Ano99-35] Anonymous. Pages 1–134 (1 October 1999). Ecological Modelling, 122(1–2):??, October 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1999:PAa

[Ano99-36] Anonymous. Pages 1–178 (1 April 1999). Ecological Modelling, 117(1):??, April 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1999:PNa

[Ano99-37] Anonymous. Pages 1–52 (1 November 1999). Ecological Modelling, 123(1):??, November 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1999:PAb

[Ano99-38] Anonymous. Pages 1–64 (3 August 1999). Ecological Modelling, 120(1):??, August 3, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1999:PJd

[Ano99-39] Anonymous. Pages 1–88 (1 July 1999). Ecological Modelling, 119(1):??, July 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1999:PJb

[Ano99-40] Anonymous. Pages 1–94 (1 June 1999). Ecological Modelling, 118(1):??, June 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1999:PDa

[Ano99-41] Anonymous. Pages 1–98 (1 December 1999). Ecological Modelling, 124(1):??, December 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1999:PSb

[Ano99-42] Anonymous. Pages 103–302 (15 September 1999). Ecological Modelling, 121(2–3):??, September 15, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1999:PMb

[Ano99-43] Anonymous. Pages 107–310 (15 March 1999). Ecological Modelling, 116(2–3):??, March 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1999:PFb

[Ano99-44] Anonymous. Pages 109–290 (15 February 1999). Ecological Modelling, 115(2–3):??, February 15, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1999:PJa

[Ano99-45] Anonymous. Pages 113–318 (1 January 1999). Ecological Modelling, 114(2–3):??, January 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1999:POb

[Ano99-46] Anonymous. Pages 135–312 (20 October 1999). Ecological Modelling, 122(3):??, October 20, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1999:PMc

[Ano99-47] Anonymous. Pages 179–378 (17 May 1999). Ecological Modelling, 117(2–3):??, May 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1999:PNb

[Ano99-48] Anonymous. Pages 53–248 (15 November 1999). Ecological Modelling, 123(2–3):??, November 15, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1999:PAc

[Ano99-49] Anonymous. Pages 65–368 (17 August 1999). Ecological Modelling, 120(2–3):??, August 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1999:PJe

[Ano99-50] Anonymous. Pages 89–282 (15 July 1999). Ecological Modelling, 119(2–3):??, July 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1999:PJc

[Ano99-51] Anonymous. Pages 95–284 (15 June 1999). Ecological Modelling, 118(2–3):??, June 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:1999:PDb

[Ano99-52] Anonymous. Pages 99–264 (13 December 1999). Ecological Modelling, 124(2–3):??, December 13, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Antonic:1998:MDT

[Ant98] Oleg Antonić. Modelling daily topographic solar radiation without site-specific hourly radiation data. Ecological Modelling, 113(1–3):31–40, November 2, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009800132X.

Alexandrov:1997:CVT

[AO97a] Georgii A. Alexandrov and Takehisa Oikawa. Contemporary variations of terrestrial net primary production: the use of satellite data in the light of an extremal principle. Ecological Modelling, 95(1):113–118, February 1, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096000312.

Arp:1997:FSV

[AO97b] Paul A. Arp and Tõnu Oja. A forest soil vegetation atmosphere model (ForSVA), i: Concepts. Ecological Modelling, 95(2–3):211–224, February 14, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096000361.

Aber:1997:MNS

[AOD97] John D. Aber, Scott V. Ollinger, and Charles T. Driscoll. Modeling nitrogen saturation in forest ecosystems in response to land use and atmospheric deposition. Ecological Modelling, 101(1):61–78, August 1, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097019534.

Alexandrov:1999:ETN

[AOE99] G. A. Alexandrov, T. Oikawa, and G. Esser. Estimating terrestrial NPP: what the data say and how they may be interpreted? Ecological Modelling, 117(2–3):361–369, May 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099000198.

Aoki:1990:MVE

[Aok90] Ichiro Aoki. Monthly variations of entropy production in Lake Biwa. Ecological Modelling, 51(3–4):227–232, June 1990. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009090067Q.

Aoki:1992:EAN

[Aok92] Ichiro Aoki. Exergy analysis of network systems at steady state. Ecological Modelling, 62(1–3):183–193, July 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380092900902.

Aoki:1993:IKI

[Aok93] Ichiro Aoki. Inclusive Kullback index — a macroscopic measure in ecological systems. Ecological Modelling, 66(3–4):289– 299, April 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390119D.

Aoki:1994:ITA

[Aok94] Ichiro Aoki. Information theoretical approach to comparative study of lakes. Ecological Modelling, 73(1–2):1–12, May 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094900949.

Aoki:1995:DRA

[Aok95a] Ichiro Aoki. Diversity and rank-abundance relationship concerning biotic compartments. Ecological Modelling, 82(1): 21–26, September 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438009400076T.

Aoki:1995:FIC

[Aok95b] Ichiro Aoki. Flow-indices characterizing eutrophication in lakeecosystems. Ecological Modelling, 82(3):225–232, November 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400093W.

Aoki:1997:CSF

[Aok97] Ichiro Aoki. Comparative study of flow-indices in lakeecosystems and the implication for maturation process. Ecological Modelling, 95(2–3):165–169, February 14, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096000476.

Abrecht:1996:TTD

[AR96] Douglas G. Abrecht and Stephen D. Robinson. TACT: a tactical decision aid using a CERES based wheat simulation model. Ecological Modelling, 86(2–3):241–244, May 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095000585.

Acock:1997:DOO

[AR97a] B. Acock and V. R. Reddy. Designing an object-oriented structure for crop models. Ecological Modelling, 94(1):33–44, January 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096019266.

Acock:1997:IMP

[AR97b] Basil Acock and James F. Reynolds. Introduction: modularity in plant models. Ecological Modelling, 94(1):1–6, January 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096019230.

Amacher:1994:MME

[AS94] Michael C. Amacher and H. M. Selim. Mathematical models to evaluate retention and transport of chromium (VI) in soil. Ecological Modelling, 74(3–4):205–230, August 1994. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094901201.

Agarwal:1993:GFG

[ASP93] M. Agarwal, A. Shukla, and V. N. Pal. Grazing of forested grassland and its conservation. Ecological Modelling, 69(1–2):57–62, September 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009390048W.

Allen:1991:PMM

[ASTL91] Linda J. S. Allen, Monty J. Strauss, Harlan G. Thorvilson, and William N. Lipe. A preliminary mathematical model of the apple twig borer (Coleoptera: Bostrichidae) and grapes on the Texas high plains. Ecological Modelling, 58(1–4):369–382, November 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380091900464.

Acevedo:1996:MFD

[AUS96] M. F. Acevedo, D. L. Urban, and H. H. Shugart. Models of forest dynamics based on roles of tree species. Ecological Modelling, 87(1–3):267–284, June 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094002088.

Asner:1997:SPA

[AW97] Gregory P. Asner and Carol A. Wessman. Scaling PAR absorption from the leaf to landscape level in spatially heterogeneous ecosystems. Ecological Modelling, 103(1):81–97, November 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438009700080X.

Alveteg:1998:ESU

[AWS98] M. Alveteg, C. Walse, and H. Sverdrup. Evaluating simplifications used in regional applications of the SAFE and MAKEDEP models. Ecological Modelling, 107(2–3):265–277, April 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098000039.

Axelsen:1993:APD

[Axe93] Jørgen Aagaard Axelsen. Analysis of the populations dynamics of the pod gall midge (Dasyneura brassicae Winn.) in winter rape and spring rape by computer simulation. Ecological Modelling, 69(1–2):43–55, September 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390047V.

Axelsen:1994:HPI

[Axe94] Jørgen Aagaard Axelsen. Host-parasitoid interactions in an agricultural ecosystem: a computer simulation. Ecological Modelling, 73(3–4):189–203, June 1994. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094900620.

Alexandrov:1999:TMP

[AYO99] G. A. Alexandrov, Y. Yamagata, and T. Oikawa. Towards a model for projecting net ecosystem production of the world forests. Ecological Modelling, 123(2–3):183–191, November 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099001283.

Bach:1993:DMD

[Bac93] Hanne K. Bach. A dynamic model describing the seasonal variations in growth and the distribution of eelgrass (Zostera marina L.) I. Model theory. Ecological Modelling, 65(1–2):31–50, January 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390125C.

Baker:1996:LPS

[Bak96] Bryan D. Baker. Landscape pattern, spatial behavior, and a dynamic state variable model. Ecological Modelling, 89 (1–3):147–160, August 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438009500128X.

Barreto:1991:SSP

[Bar91] Luís Soares Barreto. SPESS — a simulator for pure evenaged self-thinned stands. Ecological Modelling, 54(1–2):127– 132, May 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380091901027.

Bardossy:1996:UFR

[Bár96a] András Bárdossy. The use of fuzzy rules for the description of elements of the hydrological cycle. Ecological Modelling, 85(1):59–65, February 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095000119.

Bartholow:1996:SSP

[Bar96b] John M. Bartholow. Sensitivity of a salmon population model to alternative formulations and initial conditions. Ecological Modelling, 88(1–3):215–226, July 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001085.

Bartelink:1998:RIF

[Bar98] H. H. Bartelink. Radiation interception by forest trees: a simulation study on effects of stand density and foliage clustering on absorption and transmission. Ecological Modelling, 105(2–3):213–225, January 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001658.

Bastianoni:1998:DPB

[Bas98] Simone Bastianoni. A definition of ‘pollution’ based on thermodynamic goal functions. Ecological Modelling, 113(1–3): 163–166, November 2, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380098001410.

Batabyal:1996:SAM

[Bat96a] Amitrajeet A. Batabyal. On some aspects of the management of a stochastically developing forest. Ecological Modelling, 89(1–3):67–72, August 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001220.

Batabyal:1996:QTA

[Bat96b] Amitrajeet A. Batabyal. The queuing theoretic approach to groundwater management. Ecological Modelling, 85(2–3):219– 227, March 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094001782.

Batabyal:1998:OSA

[Bat98] Amitrajeet A. Batabyal. An optimal stopping approach to the conservation of biodiversity. Ecological Modelling, 105 (2–3):293–298, January 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097001646.

Baker:1992:FMF

[BBH92] B. B. Baker, R. M. Bourdon, and J. D. Hanson. FOR-AGE: a model of forage intake in beef cattle. Ecological Modelling, 60(3–4):257–279, April 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290036E.

Balzter:1998:CAM

[BBK98] Heiko Balzter, Paul W. Braun, and Wolfgang Köhler. Cellular automata models for vegetation dynamics. Ecological Modelling, 107(2–3):113–125, April 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097002020.

Belotelov:1996:MTD

[BBKV96] Nickolay V. Belotelov, Boris G. Bogatyrev, Andrew P. Kirilenko, and Sergey V. Venevsky. Modelling of timedependent biome shifts under global climate changes. Ecological Modelling, 87(1–3):29–40, June 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094002002.

Bolstad:1997:MMH

[BBL97] Paul V. Bolstad, Barbara J. Bentz, and Jesse A. Logan. Modelling micro-habitat temperature for Dendroctonus ponderosae (coleoptera: scolytidae). Ecological Modelling, 94(2–3):287–297, January 15, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438009600021X.

Bachelet:1998:SBM

[BBN98] D. Bachelet, M. Brugnach, and R. P. Neilson. Sensitivity of a biogeography model to soil properties. Ecological Modelling, 109(1):77–98, June 1, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098000106.

Bearlin:1999:SMS

[BBR99] Andrew R. Bearlin, Mark A. Burgman, and Helen M. Regan. A stochastic model for seagrass (Zostera muelleri) in Port Phillip Bay, Victoria, Australia. Ecological Modelling, 118(2–3):131–148, June 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000113.

Brunt:1990:BMA

[BC90] James W. Brunt and Walt Conley. Behavior of a multivariate algorithm for ecological edge detection. Ecological Modelling, 49(3–4):179–203, January 1990. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009090027E.

Boudjema:1996:RDE

[BC96] Gérard Boudjema and Nguyen Phong Chau. Revealing dynamics of ecological systems from natural recordings. Ecological Modelling, 91(1–3):15–23, November 15, 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001565.

Bocci:1997:MBN

[BCB97] M. Bocci, G. Coffaro, and G. Bendoricchio. Modelling biomass and nutrient dynamics in eelgrass (Zostera marina L.): applications to the Lagoon of Venice (Italy) and Øresund (Denmark). Ecological Modelling, 102(1):67–80, October 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097000951.

Bendoricchio:1993:MPE

[BCD93] G. Bendoricchio, G. Coffaro, and M. Di Luzio. Modelling the photosynthetic efficiency for Ulva rigida growth. Ecological Modelling, 67(2–4):221–232, June 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390006E.

Bai:1997:MDS

[BCH^{+}97] T. Jay Bai, Tom Cottrell, Dun-Yuan Hao, Tala Te, and Robert J. Brozka. Multi-dimensional sphere model and instantaneous vegetation trend analysis. Ecological Modelling, 97(1–2):75–86, April 15, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096000750.

Blaine:1997:ISS

[BD97] Tegan W. Blaine and Donald L. DeAngelis. The interaction of spatial scale and predator–prey functional response. Ecological Modelling, 95(2–3):319–328, February 14, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096000452.

Bergen:1999:IRS

[BD99] Kathleen M. Bergen and M. Craig Dobson. Integration of remotely sensed radar imagery in modeling and mapping of forest biomass and net primary production. Ecological Modelling, 122(3):257–274, October 20, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099001416.

Basset:1997:EFR

[BDD97] Alberto Basset, Donald L. DeAngelis, and James E. Diffendorfer. The effect of functional response on stability of a grazer population on a landscape. Ecological Modelling, 101(2–3):153–162, August 15, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097019753.

Bergez:1999:APB

[BEB99] Jacques-Eric Bergez, Michel Etienne, and Philippe Balandier. ALWAYS: a plot-based silvopastoral system model. Ecological Modelling, 115(1):1–17, February 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001537.

Belyaev:1992:MIT

[Bel92] Valeriy I. Belyaev. Modelling the influence of turbulence on phytoplankton photosynthesis. Ecological Modelling, 60(1):11– 29, January 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290010C.

Belov:1998:MPR

[Bel98] A. P. Belov. A model of phycotoxin release by cyanobacterial cells. Ecological Modelling, 110(2):105–117, July 11, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098000131.

Boren:1997:VCT

[BEM97] Jon C. Boren, David M. Engle, and Ronald E. Masters. Vegetation cover type and avian species changes on landscapes within a wildland-urban interface. Ecological Modelling, 103(2–3):251–266, November 17, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001087.

Benjamin:1999:CDR

[Ben99] L. R. Benjamin. A comparison of different rules of partitioning of crop growth between individual plants. Ecological Modelling, 115(2–3):111–118, February 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001859.

Beyer:1998:SST

[Bey98] Jan E. Beyer. Stochastic stomach theory of fish: an introduction. Ecological Modelling, 114(1):71–93, December 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098001288.

Beecham:1998:AFI

[BF98] J. A. Beecham and K. D. Farnsworth. Animal foraging from an individual perspective: an object orientated model. Ecological Modelling, 113(1–3):141–156, November 2, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001392.

Bugmann:1996:MCS

[BFK96] Harald Bugmann, Andreas Fischlin, and Felix Kienast. Model convergence and state variable update in forest gap models. Ecological Modelling, 89(1–3):197–208, August 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001352.

Ball:1992:SDE

[BG92] George L. Ball and Randy Gimblett. Spatial dynamic emergent hierarchies simulation and assessment system. Ecological Modelling, 62(1–3):107–121, July 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290084R.

Bachelet:1993:ICC

[BG93] Dominique Bachelet and Cheryl A. Gay. The impacts of climate change on rice yield: a comparison of four model performances. Ecological Modelling, 65(1–2):71–93, January 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009390127E.

Barciela:1999:MPP

[BGF99] Rosa M. Barciela, Emilio García, and Emilio Fernández. Modelling primary production in a coastal embayment affected by upwelling using dynamic ecosystem models and artificial neural networks. Ecological Modelling, 120(2–3):199–211, August 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001027.

Brown:1999:SVM

[BGL99] Martin Brown, Steve R. Gunn, and Hugh G. Lewis. Support vector machines for optimal classification and spectral unmixing. Ecological Modelling, 120(2–3):167–179, August 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099001003.

Brovkin:1997:CCV

[BGS97] Victor Brovkin, Andrei Ganopolski, and Yuri Svirezhev. A continuous climate-vegetation classification for use in climatebiosphere studies. Ecological Modelling, 101(2–3):251–261, August 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097000495.

Brosse:1999:UAN

[BGTL99] Sébastien Brosse, Jean-François Guegan, Jean-Nöel Tourenq, and Sovan Lek. The use of artificial neural networks to assess fish abundance and spatial occupancy in the littoral zone of a mesotrophic lake. Ecological Modelling, 120(2–3):299–311, August 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001106.

Brown:1991:MST

[BH91] Kendall P. Brown and Edward Z. Hosseinipour. Modeling speciation, transport and transformation of metals from mine wastes. Ecological Modelling, 57(1–2):65–89, October 1, 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380091900556.

Bowers:1994:LSM

[BH94] Michael A. Bowers and Leon C. Harris. A large-scale metapopulation model of interspecific competition and environmental change. Ecological Modelling, 72(3–4):251–273, April 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094900868.

Byers:1996:ISS

[BH96] R. E. Byers and R. I. C. Hansell. Implications of semi-stable attractors for ecological modelling. Ecological Modelling, 89(1–3):59–65, August 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001204.

Bailey:1997:WIS

[BH97] Mark C. Bailey and David P. Hamilton. Wind induced sediment resuspension: a lake-wide model. Ecological Modelling, 99(2–3):217–228, June 30, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097019558.

Byers:1999:EPS

[BH99] R. E. Byers and R. I. C. Hansell. Exploiting properties of semistable attractors to predict change in non-linear systems. Ecological Modelling, 121(2–3):255–260, September 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099000897.

Boyer:1996:OSM

[BHBM96] Elizabeth W. Boyer, George M. Hornberger, Kenneth E. Bencala, and Diane McKnight. Overview of a simple model describing variation of dissolved organic carbon in an upland catchment. Ecological Modelling, 86(2–3):183–188, May 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095000496.

Bhat:1996:CTW

[BHL96] Mahadev G. Bhat, Ray G. Huffaker, and Suzanne M. Lenhart. Controlling transboundary wildlife damage: modeling under alternative management scenarios. Ecological Modelling, 92(2–3):215–224, December 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001697.

Burns:1991:TUC

[BHWP91] Thomas P. Burns, Masahiko Higashi, Sam C. Wainright, and Bernard C. Patten. Trophic unfolding of a continental shelf energy-flow network. Ecological Modelling, 55(1–2):1–26, July 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380091900615.

Brisson:1998:PSW

[BILL98] Nadine Brisson, Bernard Itier, Jean Claude L’Hotel, and Jean Yves Lorendeau. Parameterisation of the shuttleworth-Wallace model to estimate daily maximum transpiration for use in crop models. Ecological Modelling, 107(2–3):159–169, April 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097002159.

Birkett:1994:IES

[Bir94] S. H. Birkett. On the interconnection of ecological system models. Ecological Modelling, 72(3–4):153–174, April 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094900817.

Bach:1994:MWQ

[BJ94] Hanne K. Bach and Jørgen Krogsgaard Jensen. Modelling of the water quality of a proposed impounded lake of a tidally influenced river. Ecological Modelling, 74(1–2):77–90, July 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094901120.

Bouten:1995:WBS

[BJ95] Willem Bouten and Per-Erik Jansson. Water balance of the Solling spruce stand as simulated with various forest-soil-atmosphere models. Ecological Modelling, 83(1–2):245–253, December 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095001022.

Bendoricchio:1997:EGF

[BJ97] G. Bendoricchio and S. E. Jørgensen. Exergy as goal function of ecosystems dynamic. Ecological Modelling, 102(1):5–15, October 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097000914.

Bai:1992:EDA

[BJM92] Jun Bai, Anthony J. Jakeman, and Michael McAleer. Estimation and discrimination of alternative air pollution models. Ecological Modelling, 64(2–3):89–124, October 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290111Q.

Blackburn:1991:MDS

[BK91a] H. D. Blackburn and M. M. Kothmann. Modelling diet selection and intake for grazing herbivores. Ecological Modelling, 57(1–2):145–163, October 1, 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009190059A.

Bossel:1991:SMN

[BK91b] Hartmut Bossel and Holger Krieger. Simulation model of natural tropical forest dynamics. Ecological Modelling, 59(1–2):37–71, December 1, 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009190127M.

Belyaev:1992:LIM

[BK92a] V. I. Belyaev and M. Yu. Khudoshina. Logical-information modelling of ecosystems. Ecological Modelling, 60(2):119–138, March 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290042D.

Belyaev:1992:MSE

[BK92b] V. I. Belyaev and N. V. Konduforova. Modelling of the shelf ecosystem. Ecological Modelling, 60(2):95–118, March 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009290041C.

Brown:1996:VMC

[BK96] T. N. Brown and D. Kulasiri. Validating models of complex, stochastic, biological systems. Ecological Modelling, 86(2–3):129–134, May 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095000399.

Benz:1997:CCM

[BK97] J. Benz and M. Knorrenschild. Call for a common model documentation etiquette. Ecological Modelling, 97(1–2):141–143, April 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096000555.

Barlow:1998:SMI

[BK98] N. D. Barlow and J. M. Kean. Simple models for the impact of rabbit calicivirus disease (RCD) on Australasian rabbits. Ecological Modelling, 109(3):225–241, June 22, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009800009X.

Brown:1997:RMB

[BKG97] T. N. Brown, D. Kulasiri, and R. E. Gaunt. A rootmorphology based simulation for plant/soil microbial ecosystem modelling. Ecological Modelling, 99(2–3):275–287, June 30, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097019613.

Bachelet:1995:BRC

[BKT95] Dominique Bachelet, Jeffrey Kern, and Michael Tölg. Balancing the rice carbon budget in China using spatiallydistributed data. Ecological Modelling, 79(1–3):167–177, May 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400031C.

Bobba:1990:HMA

[BL90] A. G. Bobba and D. C. L. Lam. Hydrological modelling of acidified Canadian watersheds. Ecological Modelling, 50(1–3):5–32, March 1990. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009090040N.

Baveco:1992:OOT

[BL92] J. M. Baveco and R. Lingeman. An object-oriented tool for individual-oriented simulation: host-parasitoid system application. Ecological Modelling, 61(3–4):267–286, June 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380092900227.

Burke:1993:WDL

[BL93] Ingrid C. Burke and William K. Lauenroth. What do LTER results mean? Extrapolating from site to region and decade to century. Ecological Modelling, 67(1):19–35, May 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009390097C.

Berck:1996:SSC

[BL96] Peter Berck and Ethan Ligon. The swamp and the shopping center: an interest rate parable. Ecological Modelling, 92(2–3):275, December 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001824.

Blanco:1995:RT

[Bla95] Jose Ma Blanco. A relationship is not an hat-trick. Ecological Modelling, 77(2–3):97–98, February 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400155B.

Blackwell:1997:RDM

[Bla97] P. G. Blackwell. Random diffusion models for animal movement. Ecological Modelling, 100(1–3):87–102, December 1997.

CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001531.

Bartell:1999:EMA

[BLK^{+}99] Steven M. Bartell, Guy Lefebvre, Grégoire Kaminski, Michel Carreau, and Kym Rouse Campbell. An ecosystem model for assessing ecological risks in Québec rivers, lakes, and reservoirs. Ecological Modelling, 124(1):43–67, December 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099001556.

Botterweg:1998:EGM

[BLRV98] Peter Botterweg, Rodney Leek, Eirik Romstad, and Arild Vatn. The EUROSEM-GRIDSEM modeling system for erosion analyses under different natural and economic conditions. Ecological Modelling, 108(1–3):115–129, May 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000234.

Bouloux:1998:MHP

[BLS98] Catherine Bouloux, Michel Langlais, and Patrick Silan. A marine host–parasite model with direct biological cycle and age structure. Ecological Modelling, 107(1):73–86, March 16, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097002123.

Boeve:1994:SVE

[BM94] Jean-Luc Boevé and Christoph Meier. Spider venom and eco-ethological implications: a simple mathematical model. Ecological Modelling, 73(1–2):149–157, May 1994. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094901031.

Brown:1996:EAP

[BM96] M. T. Brown and T. R. McClanahan. Emergy analysis perspectives of Thailand and Mekong River dam proposals. Ecological Modelling, 91(1–3):105–130, November 15, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095001832.

Bastianoni:1997:EER

[BM97a] Simone Bastianoni and Nadia Marchettini. Emergy/exergy ratio as a measure of the level of organization of systems. Ecological Modelling, 99(1):33–40, June 16, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096019205.

Beisel:1997:SFC

[BM97b] Jean-Nicolas Beisel and Jean-Claude Moreteau. A simple formula for calculating the lower limit of Shannon’s diversity index. Ecological Modelling, 99(2–3):289–292, June 30, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097019546.

Bancroft:1999:IBM

[BM99] J. S. Bancroft and D. C. Margolies. An individual-based model of an acarine tritrophic system: lima bean, Phaseolus lunatus L., twospotted spider mite, Tetranychus urticae (Acari: Tetranychidae), and Phytoseiulus persimilis (Acari: Phytoseiidae). Ecological Modelling, 123(2–3):161–181, November 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099001313.

Barak:1990:OEM

[BMHC90] Phillip Barak, J. A. E. Molina, Aviva Hadas, and C. E. Clapp. Optimization of an ecological model with the Marquardt algorithm. Ecological Modelling, 51(3–4):251–263, June 1990. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009090069S.

Becerra-Munoz:1999:SRD

[BMHT99] Salvador Becerra-Muñoz, Daniel B. Hayes, and William W. Taylor. Stationarity and rate of dampening of modeled indices of fish abundance in relation to their exploitation status in the Northwest Atlantic Ocean. Ecological Modelling, 117(2–3):225–238, May 17, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098002063.

Bonilla:1999:OSW

[BMV99] Carlos A. Bonilla, José F. Muñoz, and Michel Vauclin. Opus simulation of water dynamics and nitrate transport in a field plot. Ecological Modelling, 122(1–2):69–80, October 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099001192.

Bocci:1999:MGN

[Boc99] Martina Bocci. Modelling the growth of Nile tilapia (Oreochromis niloticus) feeding on natural resources in enclosures in Laguna de Bay (Philippines). Ecological Modelling, 119(2–3):135–148, July 15, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438009900040X.

Bolter:1990:ECA

[Böl90] Manfred Bölter. Evaluation — by cluster analysis — of descriptors for the establishment of significant subunits in antartic soils. Ecological Modelling, 50(1–3):79–94, March 1990. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009090043G.

Bonan:1993:ANC

[Bon93] Gordon B. Bonan. Analysis of neighborhood competition among annual plants: implications of a plant growth model. Ecological Modelling, 65(1–2):123–136, January 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390129G.

Bossel:1991:ASE

[Bos91a] Hartmut Bossel. Applied systems ecology: F. Recknagel. Akademie-Verlag, Berlin, 1989. 138 pp., DM 32.00. ISBN 3-05-500475-2. Ecological Modelling, 54(1–2):144, May 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009190107C.

Bossel:1991:SEP

[Bos91b] Hartmut Bossel. Simulation of ecophysiological processes of growth in several annual crops: F. W. T. Penning de Vries, D. M. Jansen, H. F. M. ten Berge, A. Bakema. Simulation Monographs, 29. Pudoc, Wageningen, Netherlands, 1989. Dfl.100/US$57. ISBN 90-220-0937-8. Ecological Modelling, 59(3–4):298–299, December 15, 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380091901865.

Bossel:1992:RSP

[Bos92] Hartmut Bossel. Real-structure process description as the basis of understanding ecosystems and their development. Ecological Modelling, 63(1–4):261–276, September 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290072M.

Bossel:1993:IEM

[Bos93a] Hartmut Bossel. Introduction to environmental management: P. E. Hansen and S. E. Jørgensen (Editors). Elsevier, Amsterdam, 1991, 403 pp., Dfl 265.00, ISBN 0-444-88469-6, paperback Dfl 135.00, ISBN 0-444-88532-3. Ecological Modelling, 65(3–4):297–298, February 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093900857.

Bossel:1993:TDF

[Bos93b] Hartmut Bossel. Temperate deciduous forests: E. Röhrig and B. Ulrich (Editors). Ecosystems of the World 7, Elsevier, Amsterdam, 1991, 635 pp., US $225.50/DFl 440.00, ISBN 0-444-88599-4. Ecological Modelling, 65(3–4):301–302, February 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009390088A.

Bossel:1996:TFS

[Bos96] Hartmut Bossel. treedyn 3 forest simulation model. Ecological Modelling, 90(3):187–227, November 1, 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001395.

Botterweg:1995:UIM

[Bot95] Peter Botterweg. The user’s influence on model calibration results: an example of the model SOIL, independently calibrated by two users. Ecological Modelling, 81(1–3):71–81, August 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400161A.

Bott:1998:NMC

[Bot98] Andreas Bott. A numerical model of the cloud-topped planetary boundary-layer: radiative forcing of aerosols in stratiform clouds. Ecological Modelling, 113(1–3):13–30, November 2, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098001318.

Bouten:1995:SWD

[Bou95] Willem Bouten. Soil water dynamics of the Solling spruce stand, calculated with the forhyd simulation package. Ecological Modelling, 83(1–2):67–75, December 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009500085A.

Boping:1993:MII

[BP93] Han Boping and Lin Peng. The mechanisms of interspecific and intraspecific competition for resources and the dynamics model of community. Ecological Modelling, 69(3–4):303–309, October 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390032N.

Bouman:1999:QEB

[BPN99] B. A. M. Bouman, R. A. J. Plant, and A. Nieuwenhuyse. Quantifying economic and biophysical sustainability trade-offs in tropical pastures. Ecological Modelling, 120(1):31–46, August 3, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099000666.

Boychuk:1997:MES

[BPTM^{+}97] Dennis Boychuk, Ajith H. Perera, Michael T. Ter-Mikaelian, David L. Martell, and Chao Li. Modelling the effect of spatial scale and correlated fire disturbances on forest age distribution. Ecological Modelling, 95(2–3):145–164, February 14, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096000427.

Breckling:1992:UEV

[Bre92] Broder Breckling. Uniqueness of ecosystems versus generalizability and predictability in ecology. Ecological Modelling, 63(1–4):13–27, September 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290059N.

Bretschko:1995:RWE

[Bre95] G. Bretschko. Running water ecosystems — a bare field for modelling? Ecological Modelling, 78(1–2):77–81, March 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400121W.

Berg:1995:TDSa

[BS95a] Mitchell T. Berg and Larry J. Shuman. A three-dimensional stochastic model of the behavior of radionuclides in forests I. Model structure. Ecological Modelling, 83(3):359–372, December 15, 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094001043.

Berg:1995:TDSb

[BS95b] Mitchell T. Berg and Larry J. Shuman. A three-dimensional stochastic model of the behavior of radionuclides in forests II. Cs-137 behavior in forest soils. Ecological Modelling, 83 (3):373–386, December 15, 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380094001067.

Berg:1995:TDSc

[BS95c] Mitchell T. Berg and Larry J. Shuman. A three-dimensional stochastic model of the behavior of radionuclides in forests III. Cs-137 uptake and release by vegetation. Ecological Modelling, 83(3):387–404, December 15, 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094001055.

Battaglia:1998:ASA

[BS98a] Michael Battaglia and Peter Sands. Application of sensitivity analysis to a model of Eucalyptus globulus plantation productivity. Ecological Modelling, 111(2–3):237–259, September 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098001148.

Bock:1998:FKB

[BS98b] Walter Bock and Arkadiusz Salski. A fuzzy knowledgebased model of population dynamics of the yellow-necked mouse (Apodemus flavicollis) in a beech forest. Ecological Modelling, 108(1–3):155–161, May 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438009800026X.

Bettinger:1997:UTS

[BSB97] Pete Bettinger, John Sessions, and Kevin Boston. Using tabu search to schedule timber harvests subject to spatial wildlife goals for big game. Ecological Modelling, 94(2–3):111–123, January 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096000075.

Bonato:1999:SMM

[BSB99] O. Bonato, F. Schulthess, and J. Baumgärtner. Simulation model for maize crop growth based on acquisition and allocation processes for carbohydrate and nitrogen. Ecological Modelling, 124(1):11–28, December 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099001465.

Blomback:1995:SWN

[BSE95] Karin Blombäck, Manfred Stähli, and Henrik Eckersten. Simulation of water and nitrogen flows and plant growth for a winter wheat stand in central Germany. Ecological Modelling, 81(1–3):157–167, August 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400168H.

Belyaev:1997:MHS

[BSL97] V. I. Belyaev, E. E. Sovga, and S. P. Lyubartseva. Modelling the hydrogen sulphide zone of the Black Sea. Ecological Modelling, 96(1–3):51–59, March 1, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096000567.

Bouman:1998:QEB

[BSN^{+}98] B. A. M. Bouman, R. A. Schipper, A. Nieuwenhuyse, H. Hengsdijk, and H. G. P. Jansen. Quantifying economic and biophysical sustainability trade-offs in land use exploration at the regional level: a case study for the Northern Atlantic Zone of Costa Rica. Ecological Modelling, 114(1):95–109, December 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098001215.

Badia:1991:ESH

[BSV91] Jacques Badia, François Spitz, and Gilbert Valet. Estimate of the size of a hunted population. Ecological Modelling, 55(1–2):113–122, July 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009190069D.

Bredemeier:1995:SNS

[BTvH95] Michael Bredemeier, Aaldrik Tiktak, and Kees van Heerden. The Solling Norway Spruce site. Ecological Modelling, 83(1–2):7–15, December 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438009500079B.

Bianciardi:1998:MEE

[BU98] C. Bianciardi and S. Ulgiati. Modelling entropy and exergy changes during a fluid self-organization process. Ecological Modelling, 110(3):255–267, July 22, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000696.

Buwalda:1991:MMC

[Buw91] J. G. Buwalda. A mathematical model of carbon acquisition and utilisation by kiwifruit vines. Ecological Modelling, 57(1–2):43–64, October 1, 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380091900545.

Boersma:1991:NGS

[BvSH91] Maarten Boersma, Carel P. van Schaik, and Pauline Hogeweg. Nutrient gradients and spatial structure in tropical forests: a model study. Ecological Modelling, 55(3–4):219–240, August 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190088I.

Boone:1994:MDH

[BW94] James L. Boone and Richard G. Wiegert. Modeling deer herd management: sterilization is a viable option. Ecological Modelling, 72(3–4):175–186, April 1994. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094900825.

Booty:1996:AFM

[BW96] W. G. Booty and I. W. S. Wong. Application of a fugacity model for assessing chemical fate in ecodistricts of southern Ontario. Ecological Modelling, 84(1–3):245–263, January 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094001375.

Berger:1999:VBO

[BWW99] Uta Berger, Gerd Wagner, and Wilfried F. Wolff. Virtual biologists observe virtual grasshoppers: an assessment of different mobility parameters for the analysis of movement patterns. Ecological Modelling, 115(2–3):119–127, February 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098001872.

Byers:1991:NAP

[Bye91] R. E. Byers. Non-autonomous properties of decoupled systems: implications for the use of simple autonomous models. Ecological Modelling, 54(1–2):59–71, May 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009190098L.

Byers:1996:ERM

[Bye96] John A. Byers. An encounter rate model of bark beetle populations searching at random for susceptible host trees. Ecological Modelling, 91(1–3):57–66, November 15, 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001662.

Bacsi:1995:VOT

[BZ95] Zsuzsanna Bacsi and Ferenc Zemankovics. Validation: an objective or a tool? Results on a winter wheat simulation model application. Ecological Modelling, 81(1–3):251–263, August 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400176I.

Chelle:1998:NRM

[CA98] Michael Chelle and Bruno Andrieu. The nested radiosity model for the distribution of light within plant canopies. Ecological Modelling, 111(1):75–91, August 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001008.

Cacho:1990:PFD

[Cac90]  Oscar J. Cacho.  Protein and fat dynamics in fish:

a bioenergetic model applied to aquaculture. Ecological Modelling, 50(1–3):33–56, March 1990. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009090041E.

Carter:1999:GPA

[CALW99] Jacoby Carter, Azmy S. Ackleh, Billy P. Leonard, and Haibin Wang. Giant panda (Ailuropoda melanoleuca) population dynamics and bamboo (subfamily Bambusoideae) life history: a structured population approach to examining carrying capacity when the prey are semelparous. Ecological Modelling, 123(2–3):207–223, November 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099001453.

Cao:1995:DNF

[Cao95] Guangxia Cao. The definition of the niche by fuzzy set theory. Ecological Modelling, 77(1):65–71, January 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093E0071A.

Caswell:1996:ALT

[Cas96] Hal Caswell. Analysis of life table response experiments II. Alternative parameterizations for size- and stage-structured models. Ecological Modelling, 88(1–3):73–82, July 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095000704.

Carlson:1996:WDA

[CB96] Toby N. Carlson and James A. Bunce. Will a doubling of atmospheric carbon dioxide concentration lead to an increase or a decrease in water consumption by crops? Ecological Modelling, 88(1–3):241–246, July 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001107.

Coffaro:1997:RCB

[CB97] G. Coffaro and M. Bocci. Resources competition between Ulva rigida and Zostera marina: a quantitative approach applied to the Lagoon of Venice. Ecological Modelling, 102(1):81–95, October 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097000963.

Coffaro:1997:ASD

[CBB97] G. Coffaro, M. Bocci, and G. Bendoricchio. Application of structural dynamic approach to estimate space variability of primary producers in shallow marine water. Ecological Modelling, 102(1):97–114, October 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097000975.

Cernusca:1998:ECA

[CBC^{+}98] Alexander Cernusca, Michael Bahn, Claudio Chemini, Werner Graber, Rolf Siegwolf, Ulrike Tappeiner, and John Tenhunen. ECOMONT: a combined approach of field measurements and process-based modelling for assessing effects of land-use changes in mountain landscapes. Ecological Modelling, 113(1–3):167–178, November 2, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001422.

Cyr:1995:VID

[CBP95] Linda Cyr, Ferdinand Bonn, and Alain Pesant. Vegetation indices derived from remote sensing for an estimation of soil protection against water erosion. Ecological Modelling, 79(1–3):277–285, May 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400182H.

Chiarello:1997:RCP

[CBS97] Ernest Chiarello and Marie-Hélène Barrat-Segretain. Recolonization of cleared patches by macrophytes: Modelling with point processes and random mosaics. Ecological Modelling, 96(1–3):61–73, March 1, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096000580.

Cloutman:1994:UMF

[CC94] Donald G. Cloutman and Lawrence D. Cloutman. A unified mathematical framework for population dynamics modeling. Ecological Modelling, 71(1–3):131–160, January 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094900795.

Chen:1996:MMS

[CC96] De-Xing Chen and M. B. Coughenour. A mechanistic model for submerged aquatic macrophyte photosynthesis: Hydrilla in ambient and elevated CO_{2}. Ecological Modelling, 89(1–3):133–146, August 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001271.

Chmura:1992:MCM

[CCK92] Gail L. Chmura, Robert Costanza, and Elisabeth C. Kosters. Modelling coastal marsh stability in response to sea level rise: a case study in coastal Louisiana, USA. Ecological Modelling, 64(1):47–64, October 15, 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290049K.

Chen:1994:MSC

[CCKO94] De-Xing Chen, M. B. Coughenour, A. K. Knapp, and C. E. Owensby. Mathematical simulation of C_{4} grass photosynthesis in ambient and elevated CO_{2}. Ecological Modelling, 73(1–2): 63–80, May 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094900981.

Childress:1998:CMM

[CCR98] W. Michael Childress, Charles M. Crisafulli, and Edward J. Rykiel, Jr. Comparison of Markovian matrix models of a primary successional plant community. Ecological Modelling, 107(1):93–102, March 16, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097002007.

Cropper:1999:SBB

[CD99] Wendell P. Cropper and Daniel DiResta. Simulation of a Biscayne Bay, Florida commercial sponge population: effects of harvesting after Hurricane Andrew. Ecological Modelling, 118(1):1–15, June 1, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099000393.

Cushing:1996:IAU

[CDDC96] J. M. Cushing, Brian Dennis, Robert A. Desharnais, and R. F. Costantino. An interdisciplinary approach to understanding nonlinear ecological dynamics. Ecological Modelling, 92(2–3):111–119, December 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001700.

Cescatti:1997:MRTa

[Ces97a] Alessandro Cescatti. Modelling the radiative transfer in discontinuous canopies of asymmetric crowns. I. Model structure and algorithms. Ecological Modelling, 101(2–3):263–274, August 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097000501.

Cescatti:1997:MRTb

[Ces97b] Alessandro Cescatti. Modelling the radiative transfer in discontinuous canopies of asymmetric crowns. II. Model testing and application in a Norway spruce stand. Ecological Modelling, 101(2–3):275–284, August 15, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097000550.

Carter:1999:MSE

[CF99] Jacoby Carter and John T. Finn. MOAB: a spatially explicit, individual-based expert system for creating animal foraging models. Ecological Modelling, 119(1):29–41, July 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099000447.

Cazelles:1991:SPL

[CFC91] B. Cazelles, D. Fontvieille, and N. P. Chau. Self-purification in a lotic ecosystem: a model of dissolved organic carbon and benthic microorganisms dynamics. Ecological Modelling, 58(1–4):91–117, November 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009190031U.

Christian:1996:NCN

[CFC^{+}96] Robert R. Christian, Elisenda Forés, Francisco Comin, Pierluigi Viaroli, Mariachiara Naldi, and Ireneo Ferrari. Nitrogen cycling networks of coastal ecosystems: influence of trophic status and primary producer form. Ecological Modelling, 87(1–3):111–129, June 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095000194.

Chen:1993:EMP

[CFS93] Jiquan Chen, Jerry F. Franklin, and Thomas A. Spies. An empirical model for predicting diurnal air-temperature gradients from edge into old-growth Douglas-fir forest. Ecological Modelling, 67(2–4):179–198, June 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009390004C.

Cropper:1993:SCD

[CG93] Wendell P. Cropper and Henry L. Gholz. Simulation of the carbon dynamics of a Florida slash pine plantation. Ecological Modelling, 66(3–4):231–249, April 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093901159.

Costanza:1998:MEE

[CG98] Robert Costanza and Sara Gottlieb. Modelling ecological and economic systems with STELLA: Part II. Ecological Modelling, 112(2–3):81–84, October 15, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000738.

Cooper:1997:LTB

[CH97] Kevin Cooper and Ray Huffaker. The long-term bioeconomic impacts of grazing on plant succession in a rangeland ecosystem. Ecological Modelling, 97(1–2):59–73, April 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096000725.

Chapelle:1995:PMN

[Cha95] A. Chapelle. A preliminary model of nutrient cycling in sediments of a Mediterranean lagoon. Ecological Modelling, 80(2–3):131–147, July 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438009400073Q.

Chattopadhyay:1996:ETS

[Cha96] J. Chattopadhyay. Effect of toxic substances on a two-species competitive system. Ecological Modelling, 84(1–3):287–289, January 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094001340.

Chave:1999:SSS

[Cha99] Jérôme Chave. Study of structural, successional and spatial patterns in tropical rain forests using TROLL, a spatially explicit forest model. Ecological Modelling, 124(2–3):233–254, December 13, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099001714.

Clements:1991:MSC

[CHC91] David R. Clements, Rudolf Harmsen, and Peter J. Clements. A mechanistic simulation to complement an empirical transition matrix model of acarine population dynamics. Ecological Modelling, 59(3–4):257–277, December 15, 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009190181Y.

Chertov:1990:SST

[Che90] O. G. Chertov. Specom — a single tree model of pine stand/ raw humus soil ecosystem. Ecological Modelling, 50(1–3):107– 132, March 1990. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009090045I.

Childress:1997:PDS

[Chi97] W. Michael Childress. Predicting dynamics of spatial automata models using Hamiltonian equations. Ecological Modelling, 96(1–3):293–303, March 1, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096000713.

Chiba:1998:AAR

[Chi98a] Yukihiro Chiba. Architectural analysis of relationship between biomass and basal area based on pipe model theory. Ecological Modelling, 108(1–3):219–225, May 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000301.

Chiba:1998:SCB

[Chi98b] Yukihiro Chiba. Simulation of CO_{2} budget and ecological implications of sugi (Cryptomeria japonica) man-made forests in Japan. Ecological Modelling, 111(2–3):269–281, September 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098001240.

Cheng:1996:FAG

[CHLT96] Z. Cheng, D. J. Horn, R. K. Lindquist, and R. A. J. Taylor. Fuzzy analysis for a greenhouse spider mite management system. Ecological Modelling, 90(2):111–121, October 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095001603.

Chen:1996:RCC

[CHM96] De-Xing Chen, H. W. Hunt, and J. A. Morgan. Responses of a C_{3} and C_{4} perennial grass to CO_{2} enrichment and climate change: Comparison between model predictions and experimental data. Ecological Modelling, 87(1–3):11–27, June 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094001995.

Christensen:1994:EBA

[Chr94] Villy Christensen. Emergy-based ascendency. Ecological Modelling, 72(1–2):129–144, March 1994. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094901481.

Christensen:1995:EMT

[Chr95] Villy Christensen. Ecosystem maturity — towards quantification. Ecological Modelling, 77(1):3–32, January 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093E0073C.

Chikumbo:1996:MSP

[CJMT96] O. Chikumbo, R. N. James, I. M. Y. Mareels, and B. J. Turner. Mortality simulations in Pinus radiata plantations in the Tarawera Valley regimes trial. Ecological Modelling, 86(2–3):253–258, May 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095000607.

Chertov:1997:SMS

[CK97] O. G. Chertov and A. S. Komarov. SOMM: a model of soil organic matter dynamics. Ecological Modelling, 94(2–3):177–189, January 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096000178.

Castrignano:1998:MVP

[CKK^{+}98] A. Castrignanò, N. Katerji, F. Karam, M. Mastrorilli, and A. Hamdy. A modified version of CERES-Maize model for predicting crop response to salinity stress. Ecological Modelling, 111(2–3):107–120, September 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000842.

Coffin:1990:GDS

[CL90] D. P. Coffin and W. K. Lauenroth. A gap dynamics simulation model of succession in a semiarid grassland. Ecological Modelling, 49(3–4):229–266, January 1990. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009090029G.

Christian:1999:OUW

[CL99] Robert R. Christian and Joseph J. Luczkovich. Organizing and understanding a winter’s seagrass foodweb network through effective trophic levels. Ecological Modelling, 117(1):99–124, April 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099000228.

Claereboudt:1999:FSS

[Cla99] Michel Claereboudt. Fertilization success in spatially distributed populations of benthic free-spawners: a simulation model. Ecological Modelling, 121(2–3):221–233, September 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099000800.

Chen:1999:DCP

[CLCG99] J. M. Chen, J. Liu, J. Cihlar, and M. L. Goulden. Daily canopy photosynthesis model through temporal and spatial scaling for remote sensing applications. Ecological Modelling, 124(2–3):99–119, December 13, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099001568.

Clemen:1998:USI

[Cle98] Thomas Clemen. The use of scale information for integrating simulation models into environmental information systems.

Ecological Modelling, 108(1–3):107–113, May 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000222.

Costanza:1991:SEM

[CM91] Robert Costanza and Thomas Maxwell. Spatial ecosystem modelling using parallel processors. Ecological Modelling, 58(1–4):159–183, November 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009190034X.

Corson:1998:SCI

[CMG98] Michael S. Corson, Miguel A. Mora, and William E. Grant. Simulating cholinesterase inhibition in birds caused by dietary insecticide exposure. Ecological Modelling, 105(2–3):299–323, January 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001749.

Choi:1999:MPC

[CMH99] Jae S. Choi, Asit Mazumder, and Roger I. C. Hansell. Measuring perturbation in a complicated, thermodynamic world. Ecological Modelling, 117(1):143–158, April 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000423.

Cronk:1990:EMM

[CMS90] Julie K. Cronk, William J. Mitsch, and Robert M. Sykes. Effective modelling of a major inland oil spill on the Ohio River. Ecological Modelling, 51(3–4):161–192, June 1990. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009090063M.

Carrer:1999:TNM

[CO99] Sebastiano Carrer and Silvia Opitz. Trophic network model of a shallow water area in the northern part of the Lagoon of Venice. Ecological Modelling, 124(2–3):193–219, December 13, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001611.

Collos:1997:PBE

[Col97] Yves Collos. A physiological basis for estimating inorganic carbon release during photosynthesis by natural phytoplankton. Ecological Modelling, 96(1–3):285–292, March 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096000701.

Colding:1998:AHO

[Col98] Johan Colding. Analysis of hunting options by the use of general food taboos. Ecological Modelling, 110(1):5–17, July 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098000386.

Comins:1997:ANC

[Com97] Hugh N. Comins. Analysis of nutrient-cycling dynamics, for predicting sustainability and CO_{2}-response of nutrient-limited forest ecosystems. Ecological Modelling, 99(1):51–69, June 16, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096019400.

Cook:1995:ODO

[Coo95] F. J. Cook. One-dimensional oxygen diffusion into soil with exponential respiration: analytical and numerical solutions. Ecological Modelling, 78(3):277–283, April 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009400179L.

Coquillard:1995:SCP

[Coq95] Patrick Coquillard. Simulation of the cyclical process of heathlands: induction of mosaic structures, evolution to irreversible states. Ecological Modelling, 80(2–3):97–111, July 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400057O.

Coughenour:1991:DSG

[Cou91] M. B. Coughenour. Dwarf shrub and graminoid responses to clipping, nitrogen and water: simplified simulations of biomass and nitrogen dynamics. Ecological Modelling, 54(1–2):81–110, May 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190100F.

Christensen:1992:EIS

[CP92] V. Christensen and D. Pauly. ECOPATH II — a software for balancing steady-state ecosystem models and calculating network characteristics. Ecological Modelling, 61(3–4):169–185, June 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380092900168.

Culberson:1996:APE

[CP96] Steven D. Culberson and Raul H. Piedrahita. Aquaculture pond ecosystem model: temperature and dissolved oxygen prediction — mechanism and application. Ecological Modelling, 89(1–3):231–258, August 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001409.

Congleton:1997:CII

[CPB97] William R. Congleton, Bryan R. Pearce, and Brian F. Beal. A C++ implementation of an individual/landscape model. Ecological Modelling, 103(1):1–17, November 1, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097000690.

Carlsson:1999:MMP

[CPH99] Lennart Carlsson, Johan Persson, and Lars Håkanson. A management model to predict seasonal variability in oxygen concentration and oxygen consumption in thermally stratified coastal waters. Ecological Modelling, 119(2–3):117–134, July 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099000137.

Cohen:1999:BCS

[CPM99] Yosef Cohen, John Pastor, and Ron Moen. Bite, chew, and swallow. Ecological Modelling, 116(1):1–14, March 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098001276.

Chon:1996:PCU

[CPMC96] Tae-Soo Chon, Young Seuk Park, Kyong Hi Moon, and Eui Young Cha. Patternizing communities by using an artificial neural network. Ecological Modelling, 90(1):69–78, September 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095001484.

Congleton:1999:MSG

[CPPB99] William R. Congleton, Bryan R. Pearce, Matthew R. Parker, and Brian F. Beal. Mariculture siting: a GIS description of intertidal areas. Ecological Modelling, 116(1):63–75, March 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098001586.

Chen:1997:GGP

[CR97a] Jia-Lin Chen and James F. Reynolds. GePSi: a generic plant simulator based on object-oriented principles. Ecological Modelling, 94(1):53–66, January 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438009601928X.

Clark:1997:IBM

[CR97b] Mark E. Clark and Kenneth A. Rose. Individual-based model of stream-resident rainbow trout and brook char: model description, corroboration, and effects of sympatry and spawning season duration. Ecological Modelling, 94(2–3):157–175, January 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096000105.

Crain:1998:SCL

[Cra98] Bradford R. Crain. Some comments on line transect grouped data analysis. Ecological Modelling, 109(3):243–249, June 22, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098000374.

Crittenden:1994:DMD

[Cri94a] Robert N. Crittenden. A diffusion model for the downstream migration of sockeye salmon smolts. Ecological Modelling, 71(1–3):69–84, January 1994. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094900760.

Crittenden:1994:MPR

[Cri94b] Robert N. Crittenden. A model for the processes regulating recruitment for a sockeye salmon stock. Ecological Modelling, 71(1–3):85–106, January 1994. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094900779.

Castella:1996:KRU

[CS96] E. Castella and M. C. D. Speight. Knowledge representation using fuzzy coded variables: an example based on the use of Syrphidae (Insecta, Diptera) in the assessment of riverine wetlands. Ecological Modelling, 85(1):13–25, February 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095000151.

Coffaro:1997:SMU

[CS97] Giovanni Coffaro and Adriano Sfriso. Simulation model of Ulva rigida growth in shallow water of the Lagoon of Venice. Ecological Modelling, 102(1):55–66, October 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009700094X.

Chaulya:1999:NMB

[CSCD99] S. K. Chaulya, R. S. Singh, M. K. Chakraborty, and B. B. Dhar. Numerical modelling of biostabilisation for a coal mine overburden dump slope. Ecological Modelling, 114(2–3):275–286, January 1, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098001574.

Carlson:1996:CEI

[CT96] D. H. Carlson and T. L. Thurow. Comprehensive evaluation of the improved SPUR model (SPUR-91). Ecological Modelling, 85(2–3):229–240, March 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094001901.

Chattopadhyay:1994:FDS

[CTDC94] J. Chattopadhyay, P. K. Tapaswi, D. Datta, and D. Chattopadhyay (Sarkar). Formation of a dissipative structure: a nonlinear analysis. Ecological Modelling, 73(3–4):205–214, June 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094900639.

Cook:1998:MOD

[CTKW98] F. J. Cook, S. M. Thomas, F. M. Kelliher, and D. Whitehead. A model of one-dimensional steady-state carbon dioxide diffusion from soil. Ecological Modelling, 109(2):155–164, June 11, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098000349.

Chen:1998:MAT

[CTL98] Carl W. Chen, W. T. Tsai, and Alan A. Lucier. A model of air-tree-soil system for ozone impact analysis. Ecological Modelling, 111(2–3):207–222, September 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001203.

Coffin:1993:INH

[CU93] Debra P. Coffin and Dean L. Urban. Implications of natural history traits to system-level dynamics: comparisons of a grassland and a forest. Ecological Modelling, 67(2–4):147– 178, June 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390003B.

CarbonelH:1999:NMP

[CV99] Carlos A. A. Carbonel H. and Jean Louis Valentin. Numerical modelling of phytoplankton bloom in the upwelling ecosystem of Cabo Frio (Brazil). Ecological Modelling, 116(2–3):135–148, March 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098002014.

Camilo:1995:DFC

[CW95] Gerardo R. Camilo and Michael R. Willig. Dynamics of a food chain model from an arthropod-dominated lotic community. Ecological Modelling, 79(1–3):121–129, May 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009400036H.

Christie:1995:MOD

[CWC95] I. Christie, J. W. Wilder, and J. J. Colbert. Modeling of one-dimensional spatial effects on the spread of gypsy moths. Ecological Modelling, 78(3):219–234, April 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093E0089L.

Claus:1995:DMD

[CWPD95] S. Claus, P. Wernecke, U. Pigla, and G. Dubsky. A dynamic model describing leaf temperature and transpiration of wheat plants. Ecological Modelling, 81(1–3):31–40, August 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400158E.

Zeng:1996:ADM

[cZdZ96] Qing cun Zeng and Xiao dong Zeng. An analytical dynamic model of grass field ecosystem with two variables. Ecological Modelling, 85(2–3):187–196, March 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094001863.

Diekkruger:1995:SWF

[DA95] B. Diekkrüger and M. Arning. Simulation of water fluxes using different methods for estimating soil parameters. Ecological Modelling, 81(1–3):83–95, August 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400162B.

Damgaard:1999:TAC

[Dam99] C. Damgaard. A test of asymmetric competition in plant monocultures using the maximum likelihood function of a simple growth model. Ecological Modelling, 116(2–3):285–292, March 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438009800204X.

Dayong:1990:DSP

[Day90] Zhang Dayong. Detection of spatial pattern in desert shrub populations: a comment. Ecological Modelling, 51(3–4):265– 271, June 1990. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009090070W.

Dayong:1991:NDL

[Day91] Zhang Dayong. A note on the detection of linear point patterns. Ecological Modelling, 55(1–2):151–152, July 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009190073A.

Dachs:1998:OMC

[DB98] Jordi Dachs and Josep M. Bayona. On the occurrence of microscale chemical patches in fractal aggregates. Ecological Modelling, 107(1):87–92, March 16, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097002160.

DeBerardinis:1994:EAD

[DBB94] E. De Berardinis, P. Baronio, and J. Baumgärtner. The effect of aphid (Dysaphis plantaginea pass., Hom., Aphididae) feeding on apple fruit growth. Ecological Modelling, 72(1–2):115– 127, March 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094901473.

deCastro:1995:CSD

[dC95] Francisco de Castro. Computer simulation of the dynamics of a dune system. Ecological Modelling, 78(3):205–217, April 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380093E0090P.

Dimopoulos:1999:NNM

[DCCSL99] Ioannis Dimopoulos, J. Chronopoulos, A. Chronopoulou-Sereli, and Sovan Lek. Neural network models to study relationships between lead concentration in grasses and permanent urban descriptors in Athens city (Greece). Ecological Modelling, 120(2–3):157–165, August 17, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009900099X.

deCastro:1999:ELC

[dCF99] Francisco de Castro and Ned Fetcher. The effect of leaf clustering in the interception of light in vegetal canopies: theoretical considerations. Ecological Modelling, 116(2–3):125–134, March 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098001707.

Dilks:1992:DBM

[DCM92] David W. Dilks, Raymond P. Canale, and Peter G. Meier. Development of Bayesian Monte Carlo techniques for water quality model uncertainty. Ecological Modelling, 62(1–3):149– 162, July 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290087U.

Davis:1999:MIT

[DCP99] S. A. Davis, E. A. Catchpole, and R. P. Pech. Models for the introgression of a transgene into a wild population within a stochastic environment, with applications to pest control. Ecological Modelling, 119(2–3):267–275, July 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099000630.

Dedik:1997:FRM

[DD̆Á97] Ladislav Dedík, Mária D̆urišová, and L̆ubica Ághová. The frequency response method used in modelling environmental systems: a working example. Ecological Modelling, 101(2–3):175–184, August 15, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097019807.

Debeljak:1999:IAR

[DDA99] Marko Debeljak, Sašo Džeroski, and Miha Adamič. Interactions among the red deer (Cervus elaphus, L.) population, meteorological parameters and new growth of the natural regenerated forest in Snežnik, Slovenia. Ecological Modelling, 121(1):51–61, September 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000721.

DeBlasio:1998:DVI

[De 98] F. V. De Blasio. Diversity variation in isolated environments: species-area effects from a stochastic model. Ecological Modelling, 111(1):93–98, August 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009800091X.

Deleersnijder:1992:RNM

[Del92] Eric Deleersnijder. Revisiting Nihoul’s model for oil slicks transport and spreading on the sea. Ecological Modelling, 64(1):71–75, October 15, 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290051F.

Derry:1998:MEI

[Der98] J. F. Derry. Modelling ecological interaction despite objectoriented modularity. Ecological Modelling, 107(2–3):145–158, April 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097002147.

Duarte:1997:MSM

[DF97] P. Duarte and J. G. Ferreira. A model for the simulation of macroalgal population dynamics and productivity. Ecological Modelling, 98(2–3):199–214, May 30, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096019151.

Dillon:1996:SMW

[DFHR96] Martin L. Dillon, Gary P. Fitt, J. Graeme Hamilton, and Wayne A. Rochester. A simulation model of wind-driven dispersal of Helicoverpa moths. Ecological Modelling, 86(2– 3):145–150, May 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095000429.

Dejak:1993:IAM

[DFP^{+}93] C. Dejak, Davide Franco, R. Pastres, G. Pecenik, and C. Solidoro. An informational approach to model time series of environmental data through negentropy estimation. Ecological Modelling, 67(2–4):199–220, June 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390005D.

Dale:1991:SEI

[DFPG91] V. H. Dale, R. L. A. Franklin, W. M. Post, and R. H. Gardner. Sampling ecological information: Choice of sample size. Ecological Modelling, 57(1–2):1–10, October 1, 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380091900512.

Dixon:1993:SEI

[DG93] Philip M. Dixon and Karen A. Garrett. Sampling ecological information: choice of sample size, reconsidered. Ecological Modelling, 68(1–2):67–73, July 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380093901085.

Darwen:1996:VPL

[DG96] P. J. Darwen and D. G. Green. Viability of populations in a landscape. Ecological Modelling, 85(2–3):165–171, March 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094001472.

deGee:1998:SYS

[dGG98] Maarten de Gee and Johan Grasman. Sustainable yields from seasonally fluctuating biological populations. Ecological Modelling, 109(2):203–212, June 11, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000520.

DeAngelis:1991:IBA

[DGS91] D. L. DeAngelis, L. Godbout, and B. J. Shuter. An individual-based approach to predicting density-dependent dynamics in smallmouth bass populations. Ecological Modelling, 57(1–2):91–115, October 1, 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380091900567.

Dzeroski:1997:UML

[DGWK97] Sašo Džeroski, Jasna Grbović, William J. Walley, and Boris Kompare. Using machine learning techniques in the construction of models. II. Data analysis with rule induction. Ecological Modelling, 95(1):95–111, February 1, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096000294.

Dodds:1995:SRC

[DH95] Walter K. Dodds and Geoffrey M. Henebry. Simulation of responses of community structure to species interactions driven by phenotypic change. Ecological Modelling, 79(1–3):85–94, May 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400030L.

Dodds:1996:EDD

[DH96] Walter K. Dodds and Geoffrey M. Henebry. The effect of density dependence on community structure. Ecological Modelling, 93(1–3):33–42, December 16, 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009500209X.

Dent:1999:MNP

[DH99] C. Lisa Dent and Julia Curro Henry. Modelling nutrientperiphyton dynamics in streams with surface–subsurface exchange. Ecological Modelling, 122(1–2):97–116, October 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099001210.

DeGrandi-Hoffman:1990:BHC

[DHRLE90] Gloria DeGrandi-Hoffman, Stephan A. Roth, Gerald M. Loper, and Eric H. Erickson. BEEAMATE: a honeybee colony genetics models. Ecological Modelling, 51(3–4):299–314, June 1990. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009090073P.

DiPasquale:1998:DPM

[DJ98] Cristina Di Pasquale and Claudia Maria Jacobi. Dynamics of pollination: a model of insect-mediated pollen transfer in self-incompatible plants. Ecological Modelling, 109(1):25–34, June 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097002238.

Dreyfus-Leon:1999:IBM

[DL99] Michel Jules Dreyfus-León. Individual-based modelling of fishermen search behaviour with neural networks and reinforcement learning. Ecological Modelling, 120(2–3):287–297, August 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438009900109X.

DeAngelis:1995:MNP

[DLN^{+}95] D. L. DeAngelis, M. Loreau, D. Neergaard, P. J. Mulholland, and E. R. Marzolf. Modelling nutrient-periphyton dynamics in streams: the importance of transient storage zones. Ecological Modelling, 80(2–3):149–160, July 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009400066Q.

Robles-Diaz-de-Leon:1998:PAT

[dLNT98] Luisa Fernanda Robles-Diaz de León and Alfredo Nava-Tudela. Playing with Asimina triloba (pawpaw): a species to consider when enhancing riparian forest buffer systems with nontimber products. Ecological Modelling, 112(2–3):169–193, October 15, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098000787.

Diekkruger:1995:MPD

[DNR95] B. Diekkrüger, P. Nörtersheuser, and O. Richter. Modeling pesticide dynamics of a loam site using HERBSIM and SIMULAT. Ecological Modelling, 81(1–3):111–119, August 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400164D.

Dowd:1997:PGC

[Dow97] Michael Dowd. On predicting the growth of cultured bivalves. Ecological Modelling, 104(2–3):113–131, December 8, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097001336.

Duffy:1999:RPA

[DPSB99] Kevin J. Duffy, Bruce R. Page, John H. Swart, and Vladimir B. Bajić. Realistic parameter assessment for a well known elephant–tree ecosystem model reveals that limit cycles are unlikely. Ecological Modelling, 121(2–3):115–125, September 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099000915.

Downing:1996:OOM

[DR96] Keith Downing and Mark Reed. Object-oriented migration modelling for biological impact assessment. Ecological Modelling, 93(1–3):203–219, December 16, 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096000038.

Droesen:1996:FEE

[Dro96] Wim J. Droesen. Formalisation of ecohydrological expert knowledge applying fuzzy techniques. Ecological Modelling, 85(1):75–81, February 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095000135.

Danilowicz:1999:MSO

[DS99] Bret S. Danilowicz and Peter F. Sale. Modelling settlement in open populations of reef fishes. Ecological Modelling, 121(2–3):261–276, September 15, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099000873.

Dixon:1999:SFS

[DSB^{+}99] Robert K. Dixon, Joel B. Smith, Sandra Brown, Omar Masera, Luis J. Mata, and Igor Buksha. Simulations of forest system response and feedbacks to global change: experiences and results from the U.S. Country Studies Program. Ecological Modelling, 122(3):289–305, October 20, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009900143X.

Diekkruger:1995:VAM

[DSKM95] B. Diekkrüger, D. Söndgerath, K. C. Kersebaum, and C. W. McVoy. Validity of agroecosystem models a comparison of results of different models applied to the same data set. Ecological Modelling, 81(1–3):3–29, August 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009400157D.

Daunicht:1996:FKB

[DSNN96] W. Daunicht, A. Salski, P. Nöhr, and C. Neubert. A fuzzy knowledge-based model of annual production of skylarks. Ecological Modelling, 85(1):67–73, February 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009500016X.

Dionne:1997:MEI

[DT97] D. Dionne and N. Thérien. Minimizing environmental impacts of hydroelectric reservoirs through operational control: a generic approach to reservoirs in northern Quebec. Ecological Modelling, 105(1):41–63, December 14, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001439.

Duarte:1995:MME

[Dua95] P. Duarte. A mechanistic model of the effects of light and temperature on algal primary productivity. Ecological Modelling, 82(2):151–160, October 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400088Y.

Dubey:1997:MDC

[Dub97a] B. Dubey. Modelling the depletion and conservation of resources: Effects of two interacting populations. Ecological Modelling, 101(1):123–136, August 1, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097019741.

DuBowy:1997:LTF

[DuB97b] Paul J. DuBowy. Long-term foraging optimization in northern shovelers. Ecological Modelling, 95(2–3):119–132, February 14, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096000403.

Dunkerley:1999:BCS

[Dun99] D. L. Dunkerley. Banded chenopod shrublands of arid Australia: modelling responses to interannual rainfall variability with cellular automata. Ecological Modelling, 121(2–3): 127–138, September 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380099000885.

Duplisea:1998:FBB

[Dup98] Daniel E. Duplisea. Feedbacks between benthic carbon mineralisation and community structure: a simulation-model analysis. Ecological Modelling, 110(1):19–43, July 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000398.

Durnota:1996:AOM

[Dur96] Bohdan Durnota. An abstract object model of an animal’s environment. Ecological Modelling, 86(2–3):119–123, May 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095000372.

deVries:1995:MIA

[dVKvdS95] W. de Vries, J. Kros, and C. van der Salm. Modelling the impact of acid deposition and nutrient cycling on forest soils. Ecological Modelling, 79(1–3):231–254, May 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093E0121I.

deVries:1995:MCL

[dVPO^{+}95] Wim de Vries, Maximilian Posch, Tonu Oja, Harmke van Oene, Hans Kros, Per Warfvinge, and Paul A. Arp. Modelling critical loads for the Solling spruce site. Ecological Modelling, 83(1–2):283–293, December 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001066.

deVasconcelos:1993:DES

[dVZ93] Maria J. Perestrello de Vasconcelos and Bernard P. Zeigler. Discrete-event simulation of forest landscape response to fire disturbances. Ecological Modelling, 65(3–4):177–198, February 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380093900798.

Dyer:1995:ACW

[Dye95] James M. Dyer. Assessment of climatic warming using a model of forest species migration. Ecological Modelling, 79(1–3):199– 219, May 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400038J.

Zeng:1996:TVD

[dZcZ96] Xiao dong Zeng and Qing cun Zeng. Two-variable dynamic model of grass field ecosystem with seasonal variation. Ecological Modelling, 85(2–3):197–202, March 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094001871.

Eiane:1997:SOP

[EAG97] Ketil Eiane, Dag L. Aksnes, and Jarl Giske. The significance of optical properties in competition among visual and tactile planktivores: a theoretical study. Ecological Modelling, 98(2–3):123–136, May 30, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096019096.

Elston:1993:SMR

[EB93] D. A. Elston and S. T. Buckland. Statistical modelling of regional GIS data: an overview. Ecological Modelling, 67(1):81– 102, May 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380093901007.

Espana:1999:MMC

[EBA^{+}99] María Luisa España, Frédéric Baret, Franck Aries, M. Chelle, B. Andrieu, and Laurent Prévot. Modeling maize canopy 3D architecture: Application to reflectance simulation. Ecological Modelling, 122(1–2):25–43, October 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000708.

Ertsen:1998:CPS

[EBBW98] A. C. D. Ertsen, A. M. F. Bio, W. Bleuten, and M. J. Wassen. Comparison of the performance of species response models in several landscape units in the province of Noord-Holland, The Netherlands. Ecological Modelling, 109(2):213–223, June 11, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098000532.

Ebenhoh:1993:CSS

[Ebe93] Wolfgang Ebenhöh. Coexistence of similar species in models with periodic environments. Ecological Modelling, 68 (3–4):227–247, August 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438009390019O.

ElJai:1999:PCA

[EFCM99] A. El Jai, M. Fournier, O. Cabrera, and Y. Maurissen. A prediction correction alternative for modelling spatiotemporal systems. Application to landscape modelling. Ecological Modelling, 116(2–3):237–251, March 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001951.

Esteve:1998:TDE

[EFGE98] Patrick Estève, Jacques Fontès, and Jean Philippe Gastellu-Etchegorry. Tropical dry ecosystems modelling and monitoring from space. Ecological Modelling, 108(1–3):175–188, May 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001099.

Eckersten:1995:MSN

[EGJ95] Henrik Eckersten, Annemieke Gärdenäs, and Per-Erik Jansson. Modelling seasonal nitrogen, carbon, water and heat dynamics of the Solling spruce stand. Ecological Modelling, 83(1–2):119–129, December 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095000919.

Elliott:1991:PDW

[EH91] N. C. Elliott and G. L. Hein. Population dynamics of the western corn rootworm: formulation, validation, and analysis of a simulation model. Ecological Modelling, 59(1–2):93–122, December 1, 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190129O.

Elston:1992:SAP

[Els92] D. A. Elston. Sensitivity analysis in the presence of correlated parameter estimates. Ecological Modelling, 64(1):11–22, October 15, 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290047I.

Emanuel:1996:MCC

[Ema96] William R. Emanuel. Modeling carbon cycling on disturbed landscapes. Ecological Modelling, 89(1–3):1–12, August 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009500114X.

Evans:1999:SSW

[EMHB99] Samuel P. Evans, Thomas R. Mayr, John M. Hollis, and Colin D. Brown. SWBCM: a soil water balance capacity model for environmental applications in the UK. Ecological Modelling, 121(1):17–49, September 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009900068X.

Eilers:1993:DBM

[EP93] P. H. C. Eilers and J. C. H. Peeters. Dynamic behaviour of a model for photosynthesis and photoinhibition. Ecological Modelling, 69(1–2):113–133, September 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390052T.

Ennola:1998:MZP

[ESD98] K. Ennola, J. Sarvala, and G. Dévai. Modelling zooplankton population dynamics with the extended Kalman filtering technique. Ecological Modelling, 110(2):135–149, July 11, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438009800057X.

Esser:1990:LE

[Ess90] G. Esser. Letter to the editor. Ecological Modelling, 50(1–3): 223, March 1990. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009090053J.

Evans:1995:SWBa

[ET95a] Samuel P. Evans and Marco Trevisan. A soil water-balance ‘bucket’ model for paleoclimatic purposes 1. Model structure and validation. Ecological Modelling, 82(2):109–129, October 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400084U.

Evans:1995:SWBb

[ET95b] Samuel P. Evans and Marco Trevisan. A soil water-balance ‘bucket’ model for paleoclimatic purposes 2. Model application to late Holocene climate. Ecological Modelling, 82(2):131–138, October 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400083T.

Fogel:1998:IES

[FAF98] Gary B. Fogel, Peter C. Andrews, and David B. Fogel. On the instability of evolutionary stable strategies in small populations. Ecological Modelling, 109(3):283–294, June 22, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098000684.

Fahrig:1998:WDF

[Fah98] Lenore Fahrig. When does fragmentation of breeding habitat affect population survival? Ecological Modelling, 105(2–3):273–292, January 1, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097001634.

Franklin:1993:I

[FB93] Jerry F. Franklin and Philip J. Bacon. Introduction. Ecological Modelling, 67(1):1–3, May 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009390095A.

Fitz:1996:DGE

[FDC^{+}96] H. C. Fitz, E. B. DeBellevue, R. Costanza, R. Boumans, T. Maxwell, L. Wainger, and F. H. Sklar. Development of a general ecosystem model for a range of scales and ecosystems. Ecological Modelling, 88(1–3):263–295, July 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001123.

Ferreira:1995:EOO

[Fer95] J. G. Ferreira. ECOWIN an object-oriented ecological model for aquatic ecosystems. Ecological Modelling, 79(1–3):21–34, May 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400033E.

Fisher:1991:MEP

[FF91] M. E. Fisher and H. I. Freedman. A model of environmental protection by a mutualist. Ecological Modelling, 58(1–4):119–139, November 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009190032V.

Flint:1999:MBM

[FFR99] Paul L. Flint, Ada C. Fowler, and Robert F. Rockwell. Modeling bird mortality associated with the M/V Citrus oil spill off St. Paul Island, Alaska. Ecological Modelling, 117(2–3):261–267, May 17, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438009900006X.

Friederichs:1996:FCE

[FFS96] M. Friederichs, O. Fränzle, and A. Salski. Fuzzy clustering of existing chemicals according to their ecotoxicological properties. Ecological Modelling, 85(1):27–40, February 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095000097.

Finzi:1992:OAP

[FG92] G. Finzi and G. Guariso. Optimal air pollution control strategies: a case study. Ecological Modelling, 64(2–3):221–239, October 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290116V.

Fischer:1994:UNP

[FG94] Gero W. Fischer and W. E. Grant. Use of a native predator to control overcrowding in warm-water polyculture ponds: simulation of a tucunare (Cichla monoculus)-tilapia (Oreochromis niloticus) system. Ecological Modelling, 72(3–4):205–227, April 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094900841.

Fomsgaard:1999:IMA

[FK99] Inge S. Fomsgaard and Kristian Kristensen. Influence of microbial activity, organic carbon content, soil texture and soil depth on mineralisation rates of low concentrations of 14C-mecoprop — development of a predictive model. Ecological Modelling, 122(1–2):45–68, October 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099001180.

Flindt:1997:MEE

[FKN97] Mogens Rene Flindt and Lars Kamp-Nielsen. Modelling of an estuarine eutrophication gradient. Ecological Modelling, 102(1):143–153, October 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097001014.

Flindt:1997:DTS

[FKNM^{+}97] M. R. Flindt, L. Kamp-Nielsen, J. C. Marques, M. A. Pardal, M. Bocci, G. Bendoricchio, J. Salomonsen, S. N. Nielsen, and S. E. Jørgensen. Description of the three shallow estuaries: Mondego River (Portugal), Roskilde Fjord (Denmark) and the Lagoon of Venice (Italy). Ecological Modelling, 102(1):17–31, October 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097000926.

Fleishman:1991:HLR

[Fle91] B. S. Fleishman. Hyperbolic law of reliability and its logarithmic effects in ecology. Ecological Modelling, 55(1–2):75– 88, July 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190066A.

Fomsgaard:1997:MMK

[Fom97] Inge S. Fomsgaard. Modelling the mineralization kinetics for low concentrations of pesticides in surface and subsurface soil. Ecological Modelling, 102(2–3):175–208, October 29, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097019820.

Foody:1996:FMV

[Foo96] G. M. Foody. Fuzzy modelling of vegetation from remotely sensed imagery. Ecological Modelling, 85(1):3–12, February 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095000127.

Foody:1999:ASO

[Foo99] Giles M. Foody. Applications of the self-organising feature map neural network in community data analysis. Ecological Modelling, 120(2–3):97–107, August 17, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000940.

Franko:1995:STW

[FOS95] U. Franko, B. Oelschlägel, and S. Schenk. Simulation of temperature-, water- and nitrogen dynamics using the model CANDY. Ecological Modelling, 81(1–3):213–222, August 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400172E.

Ferris:1993:MTF

[FP93] H. Ferris and J.-C. Prot. Masked trends in field data: nematode population assessments as an example. Ecological Modelling, 69(3–4):243–265, October 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390029R.

Fath:1998:NSE

[FP98] Brian D. Fath and Bernard C. Patten. Network synergism: Emergence of positive relations in ecological systems. Ecological Modelling, 107(2–3):127–143, April 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097002135.

Fath:1999:QRH

[FP99] Brian D. Fath and Bernard C. Patten. Quantifying resource homogenization using network flow analysis. Ecological Modelling, 123(2–3):193–205, November 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099001301.

Franzle:1992:P

[Frä92] Otto Fränzle. Preface. Ecological Modelling, 63(1–4):vii–ix, September 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290056K.

Friend:1995:PIM

[Fri95] A. D. Friend. PGEN: an integrated model of leaf photosynthesis, transpiration, and conductance. Ecological Modelling, 77(2–3):233–255, February 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380093E0082E.

Friend:1998:PGD

[Fri98] A. D. Friend. Parameterisation of a global daily weather generator for terrestrial ecosystem modelling. Ecological Modelling, 109(2):121–140, June 11, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098000362.

Forsythe:1995:MCD

[FRS^{+}95] William C. Forsythe, Edward J. Rykiel, Randal S. Stahl, Hsin i Wu, and Robert M. Schoolfield. A model comparison for daylength as a function of latitude and day of year. Ecological Modelling, 80(1):87–95, June 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009400034F.

Fang:1997:SCC

[FS97] Xing Fang and Heinz G. Stefan. Simulated climate change effects on dissolved oxygen characteristics in ice-covered lakes. Ecological Modelling, 103(2–3):209–229, November 17, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097000860.

Fang:1999:SVF

[FSA99] Xing Fang, Heinz G. Stefan, and Shoeb R. Alam. Simulation and validation of fish thermal DO habitat in northcentral US lakes under different climate scenarios. Ecological Modelling, 118(2–3):167–191, June 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000186.

Flindt:1997:LGT

[FSC^{+}97] Mogens Flindt, Jørgen Salomonsen, Marco Carrer, Martina Bocci, and Lars Kamp-Nielsen. Loss, growth and transport dynamics of Chaetomorpha aerea and Ulva rigida in the Lagoon of Venice during an early summer field campaign. Ecological Modelling, 102(1):133–141, October 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097000938.

Fuhlendorf:1996:SFS

[FSG96] Samuel D. Fuhlendorf, Fred E. Smeins, and William E. Grant. Simulation of a fire-sensitive ecological threshold: a case study of Ashe juniper on the Edwards Plateau of Texas, USA. Ecological Modelling, 90(3):245–255, November 1, 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095001514.

Friend:1997:PBT

[FSKC97] A. D. Friend, A. K. Stevens, R. G. Knox, and M. G. R. Cannell. A process-based, terrestrial biosphere model of ecosystem dynamics (Hybrid v3.0). Ecological Modelling, 95(2–3):249–287, February 14, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096000348.

Foy:1999:EUS

[FTH99] J. K. Foy, W. R. Teague, and J. D. Hanson. Evaluation of the upgraded SPUR model (spur2.4). Ecological Modelling, 118(2–3):149–165, June 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000162.

Fomichev:1997:RMS

[FV97] A. O. Fomichev and V. A. Vavilin. The reduced model of self-oscillating dynamics in an anaerobic system with sulfatereduction. Ecological Modelling, 95(2–3):133–144, February 14, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096000415.

Ferrier:1995:AMM

[FWCJ95] Robert C. Ferrier, Richard F. Wright, B. Jack Cosby, and Alan Jenkins. Application of the magic model to the Norway spruce stand at Solling, Germany. Ecological Modelling, 83(1–2):77–84, December 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009500086B.

Findlay:1999:EER

[FZ99] C. Scott Findlay and Ligang Zheng. Estimating ecosystem risks using cross-validated multiple regression and cross-validated holographic neural networks. Ecological Modelling, 119(1):57– 72, July 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099000551.

Gonzalez-Andujar:1996:HCM

[GA96] J. L. Gonzalez-Andujar. High control measures cannot produce extinction in weed populations. Ecological Modelling, 91(1–3):293–294, November 15, 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001948.

Garcia:1997:DSS

[GA97a] Luis A. Garcia and Michael Armbruster. A decision support system for evaluation of wildlife habitat. Ecological Modelling, 102(2–3):287–300, October 29, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097000641.

Gonzalez-Andujar:1997:MMP

[GA97b] J. L. Gonzalez-Andujar. A matrix model for the population dynamics and vertical distribution of weed seedbanks. Ecological Modelling, 97(1–2):117–120, April 15, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096000786.

Gaedke:1992:IEP

[Gae92] Ursula Gaedke. Identifying ecosystem properties: a case study using plankton biomass size distributions. Ecological Modelling, 63(1–4):277–298, September 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290073N.

Galitsky:1990:DDM

[Gal90] V. V. Galitsky. Dynamic 2-D model of plant communities. Ecological Modelling, 50(1–3):95–105, March 1990. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009090044H.

Gamito:1998:GMT

[Gam98] Sofia Gamito. Growth models and their use in ecological modelling: an application to a fish population. Ecological Modelling, 113(1–3):83–94, November 2, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001367.

Gao:1996:DME

[Gao96] Qiong Gao. Dynamic modeling of ecosystems with spatial heterogeneity. A structured approach implemented in Windows environment. Ecological Modelling, 85(2–3):241–252, March 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094001499.

Gonzalez-Andujar:1993:CMD

[GAP93] J. L. Gonzalez-Andujar and Joe N. Perry. Chaos, metapopulations and dispersal. Ecological Modelling, 65(3–4):255–263, February 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380093900824.

Gard:1990:SME

[Gar90] Thomas C. Gard. A stochastic model for the effects of toxicants on populations. Ecological Modelling, 51(3–4):273–280, June 1990. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009090071N.

Gavrikov:1995:MCG

[Gav95] Vladimir L. Gavrikov. A model of collisions of growing individuals: a further development. Ecological Modelling, 79(1–3): 59–66, May 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380093E0123K.

Granier:1999:LWB

[GBBV99] A. Granier, N. Bréda, P. Biron, and S. Villette. A lumped water balance model to evaluate duration and intensity of drought constraints in forest stands. Ecological Modelling, 116(2–3):269–283, March 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098002051.

Ganguly:1991:NES

[GC91] S. Ganguly and K. Chaudhuri. A note on an extension of Schaefer’s model. Ecological Modelling, 55(1–2):141–150, July 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380091900729.

Ganguly:1995:RSS

[GC95] S. Ganguly and K. S. Chaudhuri. Regulation of a singlespecies fishery by taxation. Ecological Modelling, 82(1):51–60, September 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400079W.

Ghosh:1996:IBL

[GCT96] Asit K. Ghosh, J. Chattopadhyay, and P. K. Tapaswi. Immunity boosted by low level of exposure to infection in an SIRS model. Ecological Modelling, 87(1–3):227–233, June 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095000208.

Grist:1999:SGI

[GdC99] Eric P. M. Grist and Sophie des Clers. Seasonal and genotypic influences on life cycle synchronisation: further insights from annual squid. Ecological Modelling, 115(2–3):149–163, February 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098001896.

Golubyatnikov:1998:MTE

[GDS98] Leonid L. Golubyatnikov, Elizaveta A. Denisenko, and Yuri M. Svirezhev. Model of the total exchange carbon flux for terrestrial ecosystems. Ecological Modelling, 108(1–3):265–276, May 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098000337.

Gastellu-Etchegorry:1998:MAP

[GET98] J. P. Gastellu-Etchegorry and V. Trichon. A modeling approach of PAR environment in a tropical rain forest in Sumatra: application to remote sensing. Ecological Modelling, 108(1–3):237–264, May 1, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098000325.

Grant:1990:RAT

[GF90] W. E. Grant and N. R. French. Response of alpine tundra to a changing climate: a hierarchical simulation model. Ecological Modelling, 49(3–4):205–227, January 1990. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009090028F.

Gertner:1999:BAE

[GFS99] George Z. Gertner, Shoufan Fang, and J. P. Skovsgaard. A Bayesian approach for estimating the parameters of a forest process model based on long-term growth data. Ecological Modelling, 119(2–3):249–265, July 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000678.

Gatto:1992:TOO

[GG92] Marino Gatto and Luca Luigi Ghezzi. Taxing overexploited open-access fisheries: the role of demand elasticity. Ecological Modelling, 60(3–4):185–198, April 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290033B.

Gertsev:1999:MSD

[GG99] V. I. Gertsev and V. V. Gertseva. A model of sturgeon distribution under a dam of a hydro-electric power plant. Ecological Modelling, 119(1):21–28, July 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000459.

Gin:1998:SBE

[GGC98] Karina Y. H. Gin, Jinghui Guo, and Hin-Fatt Cheong. A size-based ecosystem model for pelagic waters. Ecological Modelling, 112(1):53–72, October 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001264.

Gomboso:1996:MGF

[GGJ96] J. Gomboso, F. Ghassemi, and A. J. Jakeman. Modelling groundwater flow in the North Stirling Land Conservation District, Western Australia. Ecological Modelling, 86(2–3):169– 175, May 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009500047X.

Guan:1997:FUA

[GGP97] Biing T. Guan, George Z. Gertner, and Pablo Parysow. A framework for uncertainty assessment of mechanistic forest growth models: a neural network example. Ecological Modelling, 98(1):47–58, May 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096019369.

Gauthier:1999:GGO

[GGZ99] Laurent Gauthier, Christian Gary, and Hassan Zekki. GPSF: a generic and object-oriented framework for crop simulation. Ecological Modelling, 116(2–3):253–268, March 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098002038.

Ginot:1994:EPD

[GH94] Vincent Ginot and Jean-Christophe Hervé. Estimating the parameters of dissolved oxygen dynamics in shallow ponds. Ecological Modelling, 73(3–4):169–187, June 1994. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094900612.

Grasman:1997:LEM

[GH97] Johan Grasman and Reinier HilleRisLambers. On local extinction in a metapopulation. Ecological Modelling, 103 (1):71–80, November 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097000793.

Guenni:1996:MSV

[GHH^{+}96] Lelys Guenni, Michael F. Hutchinson, William Hogarth, Calvin W. Rose, and Roger Braddock. A model for seasonal variation of rainfall at Adelaide and Turen. Ecological Modelling, 85(2–3):203–217, March 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094001898.

Grant:1999:SAS

[GHQ99] W. E. Grant, Wayne T. Hamilton, and Esteban Quintanilla. Sustainability of agroecosystems in semi-arid grasslands: simulated management of woody vegetation in the Rio Grande Plains of southern Texas and northeastern Mexico. Ecological Modelling, 124(1):29–42, December 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099001477.

Groenenberg:1995:AMP

[GKvdSdV95] Bert-Jan Groenenberg, Hans Kros, Caroline van der Salm, and Wim de Vries. Application of the model NUCSAM to the Solling spruce site. Ecological Modelling, 83(1–2):97–107, December 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009500089E.

Gao:1997:MAD

[GLD97] Qiong Gao, Ning Liang, and Xuejun Dong. A modelling analysis on dynamics of hilly sandy grassland landscapes using spatial simulation. Ecological Modelling, 98(2–3):163–172, May 30, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096019126.

Gilruth:1995:DSM

[GMI95] Peter T. Gilruth, Stuart E. Marsh, and Robert Itami. A dynamic spatial model of shifting cultivation in the highlands of Guinea, West Africa. Ecological Modelling, 79(1–3):179– 197, May 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380093E0145S.

Grant:1991:SCM

[GMM91] W. E. Grant, J. H. Matis, and T. H. Miller. A stochastic compartmental model for migration of marine shrimp. Ecological Modelling, 54(1–2):1–15, May 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009190095I.

Gobas:1993:MPB

[Gob93] Frank A. P. C. Gobas. A model for predicting the bioaccumulation of hydrophobic organic chemicals in aquatic foodwebs: application to Lake Ontario. Ecological Modelling, 69(1–2):1–17, September 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438009390045T.

Gonzalez:1997:MWE

[Gon97] Ernesto J. González. Are models which explain the paradox of the plankton really different? Ecological Modelling, 97(3):247–251, April 28, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096019023.

Gorokhova:1998:EMG

[Gor98] Elena Gorokhova. Exploring and modeling the growth dynamics of Mysis mixta. Ecological Modelling, 110(1):45–54, July 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098000404.

Gottlieb:1998:NRA

[Got98] Sara J. Gottlieb. Nutrient removal by age-0 Atlantic menhaden (Brevoortia tyrranus) in Chesapeake Bay and implications for seasonal management of the fishery. Ecological Modelling, 112(2–3):111–130, October 15, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000751.

Goetz:1999:SRS

[GPG^{+}99] Scott J. Goetz, Stephen D. Prince, Samuel N. Goward, Michelle M. Thawley, and Jennifer Small. Satellite remote sensing of primary production: an improved production efficiency modeling approach. Ecological Modelling, 122(3):239–255, October 20, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001404.

Gilmanov:1997:TCE

[GPO97] Tagir G. Gilmanov, William J. Parton, and Dennis S. Ojima. Testing the ‘CENTURY’ ecosystem level model on data sets from eight grassland sites in the former USSR representing a wide climatic/soil gradient. Ecological Modelling, 96(1–3):191–210, March 1, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096000671.

Gove:1995:MPM

[GPT95] Jeffrey H. Gove, Ganapati P. Patil, and Charles Taillie. A mathematical programming model for maintaining structural diversity in uneven-aged forest stands with implications to other formulations. Ecological Modelling, 79(1–3):11–19, May 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400044I.

Goldberg:1997:ZFH

[GPW97] G. A. Goldberg, S. A. Piontkovski, and R. Williams. Zooplankton field heterogeneity formation on the mesoscale: elements of the theory and empirical characteristics. Ecological Modelling, 96(1–3):41–49, March 1, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009600052X.

Guowei:1991:NMM

[GQB91] Sun Guowei, Cui Qiwu, and Song Bo. A new mathematical model of interspecific competition an expansion of the classical Lotka–Volterra competition equations. Ecological Modelling, 58(1–4):273–284, November 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380091900408.

Grant:1994:SCB

[Gra94] Robert Grant. Simulation of competition between barley and wild oats under different managements and climates. Ecological Modelling, 71(4):269–287, February 1994. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094901384.

Grant:1995:DEW

[Gra95] R. F. Grant. Dynamics of energy, water, carbon and nitrogen in agricultural ecosystems: simulation and experimental validation. Ecological Modelling, 81(1–3):169–181, August 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400169I.

Grant:1998:PCC

[Gra98a] Alastair Grant. Population consequences of chronic toxicity: incorporating density dependence into the analysis of life table response experiments. Ecological Modelling, 105(2–3):325–335, January 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001762.

Grant:1998:SPE

[Gra98b] R. F. Grant. Simulation in ecosys of root growth response to contrasting soil water and nitrogen. Ecological Modelling, 107(2–3):237–264, April 1, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097002214.

Grant:1998:ENR

[Gra98c] W. E. Grant. Ecology and natural resource management: reflections from a systems perspective ^{1}. Ecological Modelling, 108(1–3):67–76, May 1, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098000192.

Grasso:1998:EEM

[Gra98d] Monica Grasso. Ecological–economic model for optimal mangrove trade off between forestry and fishery production: comparing a dynamic optimization and a simulation model. Ecological Modelling, 112(2–3):131–150, October 15, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000763.

Giske:1997:IFD

[GRBF97] Jarl Giske, Rune Rosland, Jarle Berntsen, and Øyvind Fiksen. Ideal free distribution of copepods under predation risk. Ecological Modelling, 95(1):45–59, February 1, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096000270.

Grimm:1999:TYI

[Gri99a] Volker Grimm. Ten years of individual-based modelling in ecology: what have we learned and what could we learn in the future? Ecological Modelling, 115(2–3):129–148, February 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098001884.

Grist:1999:SST

[Gri99b] Eric P. M. Grist. The significance of spatio-temporal neighbourhood on plant competition for light and space. Ecological Modelling, 121(1):63–78, September 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000733.

Grant:1996:ILN

[GRJP96] William E. Grant, Edward J. Rykiel, Sven E. Jørgensen, and Wolfgang Pittroff. ISEM launches a new era. Ecological Modelling, 93(1–3):vii, December 16, 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097827731.

Gatto:1992:ARD

[GRL92] Marino Gatto, Claudia Ricci, and Małgorzata Loga. Assessing the response of demographic parameters to density in a rotifer population. Ecological Modelling, 62(1–3):209–232, July 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290092S.

Gallerano:1993:AET

[GRV93] Francesco Gallerano, Ruggero Ricci, and Paolo Viotti. Analysis of the eutrophication trend in a deep lake. Ecological Modelling, 66(3–4):157–179, April 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380093901115.

Gavrikov:1996:SBT

[GS96] Vladimir L. Gavrikov and Olga P. Sekretenko. Shootbased three-dimensional model of young Scots pine growth. Ecological Modelling, 88(1–3):183–193, July 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095000879.

Giho:1997:TMM

[GS97] Hitomi Giho and Hiromi Seno. Transition matrix modelling on disturbance-controlled persistence of plant population. Ecological Modelling, 94(2–3):207–219, January 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096000208.

Ghosh:1998:SOR

[GS98a] Dipanwita Ghosh and A. K. Sarkar. Stability and oscillations in a resource-based model of two interacting species with nutrient cycling. Ecological Modelling, 107(1):25–33, March 16, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097002032.

Gronewold:1998:EBM

[GS98b] Anja Gronewold and Michael Sonnenschein. Event-based modelling of ecological systems with asynchronous cellular automata. Ecological Modelling, 108(1–3):37–52, May 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098000179.

Gersev:1995:PFO

[GSG95] V. I. Gersev, A. M. Safronov, and V. V. Gerseva. Population as a freely oscillating system. Ecological Modelling, 80(1):41– 45, June 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400047L.

Gertsev:1996:SUI

[GSG96] V. I. Gertsev, A. M. Safronov, and V. V. Gertseva. Structure of the unified ichthyological prognosis. Ecological Modelling, 85(2–3):145–149, March 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400188X.

Goto:1994:MSC

[GSS94] Naohiro Goto, Akiyoshi Sakoda, and Motoyuki Suzuki. Modelling of soil carbon dynamics as a part of carbon cycle in terrestrial ecosystems. Ecological Modelling, 74(3–4):183–204, August 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094901198.

Grimm:1992:ASC

[GSW92] Volker Grimm, Eric Schmidt, and Christian Wissel. On the application of stability concepts in ecology. Ecological Modelling, 63(1–4):143–161, September 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290067O.

Grant:1997:IEM

[GT97] W. E. Grant and Paul B. Thompson. Integrated ecological models: simulation of socio-cultural constraints on ecological dynamics. Ecological Modelling, 100(1–3):43–59, December 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001555.

Gross:1999:AAN

[GTF99] Lidwine Gross, Sylvie Thiria, and Robert Frouin. Applying artificial neural network methodology to ocean color remote sensing. Ecological Modelling, 120(2–3):237–246, August 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001052.

Gilbert:1997:ASP

[GU97] James R. Gilbert and Mark S. Udevitz. Adaptation of a stage-projection model for species with multiple year reproductive cycles. Ecological Modelling, 97(1–2):47–57, April 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096000737.

Guerrin:1991:QRA

[Gue91] François Guerrin. Qualitative reasoning about an ecological process: interpretation in hydroecology. Ecological Modelling, 59(3–4):165–201, December 15, 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380091901773.

Gathmann:1998:PIS

[GW98] F. Oliver Gathmann and D. Dudley Williams. Inter-site: a new tool for the simulation of spatially realistic population dynamics. Ecological Modelling, 113(1–3):125–139, November 2, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098001380.

Grimm:1999:IBM

[GWAU99] Volker Grimm, Tomasz Wyszomirski, David Aikman, and Janusz Uchmański. Individual-based modelling and ecological theory: synthesis of a workshop. Ecological Modelling, 115(2–3):275–282, February 15, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098001860.

Gutierrez:1999:PBT

[GYNE99] A. P. Gutierrez, J. S. Yaninek, P. Neuenschwander, and C. K. Ellis. A physiologically-based tritrophic metapopulation model of the African cassava food web. Ecological Modelling, 123(2–3):225–242, November 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099001441.

Gao:1996:MDM

[GYYL96] Qiong Gao, Xiusheng Yang, Rui Yun, and Chunping Li. MAGE, a dynamic model of alkaline grassland ecosystems with variable soil characteristics. Ecological Modelling, 93(1–3):19–32, December 16, 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095002081.

Haefner:1991:FWS

[Hae91] James W. Haefner. Food-web simulation on parallel computers: Inter-processor communication benchmarks. Ecological Modelling, 54(1–2):73–79, May 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009190099M.

Haakanson:1995:MPO

[Håk95a] Lars Håkanson. Models to predict organic content of lake sediments. Ecological Modelling, 82(3):233–245, November 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400100V.

Haakanson:1995:OSP

[Håk95b] Lars Håkanson. Optimal size of predictive models. Ecological Modelling, 78(3):195–204, April 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380093E0103A.

Haakanson:1996:GMD

[Håk96a] Lars Håkanson. A general method to define confidence limits for model predictions based on validations. Ecological Modelling, 91(1–3):153–168, November 15, 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001875.

Haakanson:1996:NSG

[Håk96b] Lars Håkanson. A new, simple, general technique to predict seasonal variability of river discharge and lake temperature for lake ecosystem models. Ecological Modelling, 88(1–3):157– 181, July 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095000836.

Haakanson:1996:SMP

[Håk96c] Lars Håkanson. A simple model to predict the duration of the mercury problem in Sweden. Ecological Modelling, 93(1–3):251–262, December 16, 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096000063.

Haakanson:1997:TDS

[Håk97] Lars Håkanson. Testing different sub-models for the partition coefficient and the retention rate for radiocesium in lake ecosystem modelling. Ecological Modelling, 101(2–3):229–250, August 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097000483.

Haakanson:1999:PFD

[Håk99] Lars Håkanson. On the principles and factors determining the predictive success of ecosystem models, with a focus on lake eutrophication models. Ecological Modelling, 121(2–3):139–160, September 15, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099000836.

Hanson:1991:ASR

[Han91] J. D. Hanson. Analytical solution of the rectangular hyperbola for estimating daily net photosynthesis. Ecological Modelling, 58(1–4):209–216, November 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380091900372.

Hansson:1995:EEF

[Han95] Sture Hansson. Effects of exploitative food competition on food niche dynamics — a simulation analysis. Ecological Modelling, 77(2–3):167–187, February 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093E0075E.

Han:1996:GFU

[Han96] Bo-Ping Han. A group of formulae to unfold econetworks. Ecological Modelling, 84(1–3):305–310, January 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001387.

Han:1997:SMC

[Han97a] Bo-Ping Han. On several measures concerning flow variables in ecosystems. Ecological Modelling, 104(2–3):289–302, December 8, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001373.

Han:1997:RTM

[Han97b] Bo-Ping Han. Residence time of matter and energy in econetworks at steady state. Ecological Modelling, 95(2–3): 301–310, February 14, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096000440.

Han:1998:DTS

[Han98] Bo-Ping Han. On the diversity of trophic structures and processes in ecosystems. Ecological Modelling, 107(1):51–62, March 16, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097002019.

Holst:1997:OOI

[HAOR97] N. Holst, J. A. Axelsen, J. E. Olesen, and P. Ruggle. Objectoriented implementation of the metabolic pool model. Ecological Modelling, 104(2–3):175–187, December 8, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001245.

Hill:1997:KBI

[HAW97] Michael J. Hill, Richard J. Aspinall, and Walter D. Willms. Knowledge-based and inductive modelling of rough fescue (Festuca altaica, F. campestris and F. hallii) distribution in Alberta, Canada. Ecological Modelling, 103(2–3):135– 150, November 17, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097000823.

Heil:1993:CSM

[HB93] Gerrit W. Heil and Roland Bobbink. “Calluna”, a simulation model for evaluation of impacts of atmospheric nitrogen deposition on dry heathlands. Ecological Modelling, 68(3–4):161–182, August 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390015K.

Hendry:1997:TDI

[HBM^{+}97] Ruth Hendry, P. J. Bacon, R. Moss, S. C. F. Palmer, and Jacqueline McGlade. A two-dimensional individualbased model of territorial behaviour: possible population consequences of kinship in red grouse. Ecological Modelling, 105(1):23–39, December 14, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097001361.

Higashi:1993:NTDb

[HBP93] Masahiko Higashi, Thomas P. Burns, and Bernard C. Patten. Network trophic dynamics: the tempo of energy movement and availability in ecosystems. Ecological Modelling, 66(1–2):43–64, March 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390038T.

Holt:1996:MDL

[HC96] Johnson Holt and Robert A. Cheke. Models of desert locust phase changes. Ecological Modelling, 91(1–3):131–137, November 15, 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095001840.

Harmon:1997:MCM

[HC97a] Robin Harmon and Peter Challenor. A Markov chain Monte Carlo method for estimation and assimilation into models. Ecological Modelling, 101(1):41–59, August 1, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097019479.

Holt:1997:DEM

[HC97b] Johnson Holt and John Colvin. A differential equation model of the interaction between the migration of the Senegalese grasshopper, Oedaleus senegalensis, its predators, and a seasonal habitat. Ecological Modelling, 101(2–3):185–193, August 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097019819.

Huntingford:1997:USN

[HC97c] C. Huntingford and P. M. Cox. Use of statistical and neural network techniques to detect how stomatal conductance responds to changes in the local environment. Ecological Modelling, 97(3):217–246, April 28, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096019059.

Hill:1998:AMI

[HCdVM98] David Hill, Patrick Coquillard, Jean de Vaugelas, and Alexandre Meinesz. An algorithmic model for invasive species: Application to Caulerpa taxifolia (Vahl) C. Agardh development in the North-Western Mediterranean Sea. Ecological Modelling, 109(3):251–266, June 22, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000581.

Honeycutt:1994:FAE

[HCL94] C. Wayne Honeycutt, William M. Clapham, and Simeon S. Leach. A functional approach to efficient nitrogen use in crop production. Ecological Modelling, 73(1–2):51–61, May 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094900973.

Humphries:1996:IBM

[HCL96] Hope C. Humphries, Debra P. Coffin, and William K. Lauenroth. An individual-based model of alpine plant distributions. Ecological Modelling, 84(1–3):99–126, January 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094001448.

Hendriks:1996:MRS

[HE96] A. Jan Hendriks and E. Lisette Enserink. Modelling response of single-species populations to microcontaminants as a function of species size with examples for waterfleas (Daphnia magna) and cormorants (Phalacrocorax carbo). Ecological Modelling, 88(1–3):247–262, July 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001115.

Heard:1995:STD

[Hea95] Stephen B. Heard. Short-term dynamics of processing chain systems. Ecological Modelling, 80(1):57–68, June 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009400050R.

Henebry:1995:SME

[Hen95] Geoffrey M. Henebry. Spatial model error analysis using autocorrelation indices. Ecological Modelling, 82(1):75–91, September 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400074R.

Haigen:1993:ACG

[HF93] Xu Haigen and Yang Fengxiang. Age classification and growth model of Phrynocephalus przewalskii. Ecological Modelling, 70(1–2):127–135, November 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380093900765.

Hof:1996:ACS

[HF96] John Hof and Curtis H. Flather. Accounting for connectivity and spatial correlation in the optimal placement of wildlife habitat. Ecological Modelling, 88(1–3):143–155, July 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095000828.

Hoch:1998:TSD

[HGB98] R. Hoch, T. Gabele, and J. Benz. Towards a standard for documentation of mathematical models in ecology. Ecological Modelling, 113(1–3):3–12, November 2, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001306.

Hikosaka:1998:LCP

[HH98] Kouki Hikosaka and Tadaki Hirose. Leaf and canopy photosynthesis of C_{3} plants at elevated CO_{2} in relation to optimal partitioning of nitrogen among photosynthetic components: theoretical prediction. Ecological Modelling, 106(2–3):247–259, March 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001981.

Halle:1999:MAS

[HH99] Stefan Halle and Barbara Halle. Modelling activity synchronisation in free-ranging microtine rodents. Ecological Modelling, 115(2–3):165–176, February 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001902.

Hill:1996:DWC

[Hil96] Michael J. Hill. Defining the white clover zone in eastern mainland Australia using a model and a geographic information system. Ecological Modelling, 86(2–3):245–252, May 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095000593.

Hjuler:1996:MSN

[Hju96] Henning Hjuler. A model of the simultaneous nitrogen and carbon removal in a biofilm. Ecological Modelling, 89(1–3): 269–290, August 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095001476.

Hilden:1991:CSA

[HK91] Mikael Hildén and Veijo Kaitala. Comprehensive sensitivity analysis of a bioeconomic stock-recruitment model. Ecological Modelling, 54(1–2):37–57, May 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009190097K.

Hauhs:1995:MRF

[HKMRS95] Michael Hauhs, Alois Kastner-Maresch, and Klaus Rost-Siebert. A model relating forest growth to ecosystemscale budgets of energy and nutrients. Ecological Modelling, 83(1–2):229–243, December 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009500101Z.

Hadfield:1995:APM

[HKV^{+}95] M. Hadfield, S. K. Kumar, W. F. Vincent, P. C. Austin, and G. C. Wake. Addendum to “Picoplankton and marine food chain dynamics in a variable mixed layer: a reactiondiffusion model”. Ecological Modelling, 82(1):105–108, September 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094000802.

Hanratty:1996:EMP

[HL96] Michael P. Hanratty and Karsten Liber. Evaluation of model predictions of the persistence and ecological effects of diflubenzuron in a littoral ecosystem. Ecological Modelling, 90(1):79–95, September 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001492.

Haakanson:1997:FDT

[HL97] Lars Håkanson and Martin Lindström. Frequency distributions and transformations of lake variables, catchment area and morphometric parameters in predictive regression models for small glacial lakes. Ecological Modelling, 99(2–3):171–201, June 30, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097019510.

Hunt:1999:FCD

[HLF99] E. Raymond Hunt, Michael B. Lavigne, and Steven E. Franklin. Factors controlling the decline of net primary production with stand age for balsam fir in Newfoundland assessed using an ecosystem simulation model. Ecological Modelling, 122(3):151–164, October 20, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099001350.

Hearne:1996:OOL

[HLG96] John Hearne, Roland Lamberson, and Peter Goodman. Optimising the offtake of large herbivores from a multi-species community. Ecological Modelling, 92(2–3):225–233, December 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095002030.

Hagiwara:1994:EMM

[HM94] Hiroyuki Hagiwara and William J. Mitsch. Ecosystem modeling of a multi-species integrated aquaculture pond in South China. Ecological Modelling, 72(1–2):41–73, March 1994. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094901457.

Hasenauer:1997:BPT

[HM97a] Hubert Hasenauer and Robert A. Monserud. Biased predictions for tree height increment models developed from smoothed ‘data’. Ecological Modelling, 98(1):13–22, May 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096019333.

Hraber:1997:CAM

[HM97b] Peter T. Hraber and Bruce T. Milne. Community assembly in a model ecosystem. Ecological Modelling, 103(2–3): 267–285, November 17, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097001117.

Hanson:1990:LFD

[HMA90] Jeffrey S. Hanson, George P. Malanson, and Marc P. Armstrong. Landscape fragmentation and dispersal in a model of riparian forest dynamics. Ecological Modelling, 49(3–4):277–296, January 1990. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009090031B.

He:1999:OOF

[HMB99] Hong S. He, David J. Mladenoff, and Joel Boeder. An object-oriented forest landscape model and its representation of tree species. Ecological Modelling, 119(1):1–19, July 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099000411.

He:1999:LEM

[HMC99] Hong S. He, David J. Mladenoff, and Thomas R. Crow. Linking an ecosystem model and a landscape model to study forest species response to climate warming. Ecological Modelling, 114(2–3):213–233, January 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001471.

Hendry:1996:CML

[HMW96] R. J. Hendry, J. M. McGlade, and J. Weiner. A coupled map lattice model of the growth of plant monocultures. Ecological Modelling, 84(1–3):81–90, January 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094001286.

Hogarth:1992:SPP

[HNCS92] W. L. Hogarth, J. Norbury, I. Cunning, and K. Sommers. Stability of a predator–prey model with harvesting. Ecological Modelling, 62(1–3):83–106, July 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290083Q.

Hoffmann:1995:FMG

[Hof95] Friedrich Hoffmann. FAGUS, a model for growth and development of beech. Ecological Modelling, 83(3):327–348, December 15, 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094001018.

Hogg:1999:SIR

[Hog99] E. H. Hogg. Simulation of interannual responses of trembling aspen stands to climatic variation and insect defoliation in western Canada. Ecological Modelling, 114(2–3):175–193, January 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098001501.

Hone:1992:MPV

[Hon92] J. Hone. Modelling of poisoning for vertebrate pest control, with emphasis on poisoning feral pigs. Ecological Modelling, 62(4):311–327, August 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290005Y.

Howarth:1996:DRSa

[How96a] Richard B. Howarth. Discount rates and sustainable development. Ecological Modelling, 92(2–3):263–270, December 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095001727.

Howarth:1996:DRSb

[How96b] Richard B. Howarth. Discount rates and sustainable development: reply. Ecological Modelling, 92(2–3):271–272, December 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380096000117.

Hocking:1994:MTD

[HP94] Graeme C. Hocking and John C. Patterson. Modelling tracer dispersal and residence time in a reservoir. Ecological Modelling, 74(1–2):63–75, July 1994. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094901112.

Higashi:1993:NTDa

[HPB93] Masahiko Higashi, Bernard C. Patten, and Thomas P. Burns. Network trophic dynamics: the modes of energy utilization in ecosystems. Ecological Modelling, 66(1–2):1–42, March 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009390037S.

Higgins:1996:RMA

[HR96] S. I. Higgins and D. M. Richardson. A review of models of alien plant spread. Ecological Modelling, 87(1–3):249–265, June 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095000224.

Hirvonen:1999:SMP

[HRRP99] Heikki Hirvonen, Esa Ranta, Hannu Rita, and Nina Peuhkuri. Significance of memory properties in prey choice decisions. Ecological Modelling, 115(2–3):177–189, February 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001914.

Hastings:1996:LPM

[HRV^{+}96] Harold M. Hastings, Donald J. Rumignani, Alfred G. Vassalotti, Marysia T. Weiss, and Yihren Wu. Lack of predictability in model ecosystems based on coupled logistic equations. Ecological Modelling, 92(2–3):209–214, December 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095001719.

Haness:1991:TUF

[HRWC91] Steven J. Haness, Linda A. Roberts, John J. Warwick, and William G. Cale. Testing the utility of first order uncertainty analysis. Ecological Modelling, 58(1–4):1–23, November 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009190028Y.

Hearne:1991:OTS

[HS91] J. W. Hearne and Johan Swart. Optimal translocation strategies for saving the black rhino. Ecological Modelling, 59(3–4):279–292, December 15, 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009190182Z.

Henderson-Sellers:1992:DSS

[HS92a] A. Henderson-Sellers. Decision support system prototypes for greenhouse advice. Ecological Modelling, 64(2–3):241–260, October 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290117W.

Henderson-Sellers:1992:E

[HS92b] B. Henderson-Sellers. Editorial. Ecological Modelling, 64(2– 3):87–88, October 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438009290110Z.

Henderson-Sellers:1992:SPR

[HS92c]  B. Henderson-Sellers.  Sensitivity of plume rise mod-

els in the context of decision support. Ecological Modelling, 64(2–3):159–183, October 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290113S.

Halling-Sorensen:1994:MCE

[HS94] Bent Halling-Sørensen. Modelling change in environmental systems: A. J. Jakeman, M. B. Beck and McAleer (Editors). John Wiley & Sons Ltd, England, 584 pp., hardbound, ISBN 0-471-94263-4. Ecological Modelling, 74(3–4):315, August 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094901295.

Henderson-Sellers:1996:TCM

[HS96] B. Henderson-Sellers. Towards a conceptual model of randomness. Ecological Modelling, 85(2–3):303–308, March 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094001480.

Hamilton:1997:PWQ

[HS97a] David P. Hamilton and S. Geoffrey Schladow. Prediction of water quality in lakes and reservoirs. Part I — model description. Ecological Modelling, 96(1–3):91–110, March 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096000622.

Hirata:1997:HDS

[HS97b] Hisaya Hirata and Hiromi Seno. How does the size distribution of male territories depend on the spatial distribution of females? Ecological Modelling, 103(2–3):193–207, November 17, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097000859.

Hilgerloh:1999:SMM

[HS99] Gudrun Hilgerloh and Holger Siemoneit. A simple mathematical model upon the effect of predation by birds on a blue mussel (Mytilus edulis) population. Ecological Modelling, 124(2–3):175–182, December 13, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099001623.

Hof:1999:STO

[HSB99] John Hof, Carolyn Hull Sieg, and Michael Bevers. Spatial and temporal optimization in habitat placement for a threatened plant: the case of the western prairie fringed orchid. Ecological Modelling, 115(1):61–75, February 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001768.

Henderson-Sellers:1996:SEE

[HSHS96] B. Henderson-Sellers and A. Henderson-Sellers. Sensitivity evaluation of environmental models using fractional factorial experimentation. Ecological Modelling, 86(2–3):291–295, May 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095000666.

Henderson-Sellers:1996:E

[HSJM96] B. Henderson-Sellers, A. J. Jakeman, and M. McAleer. Editorial. Ecological Modelling, 86(2–3):vii, May 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009690023X.

Halling-Sorensen:1996:MNR

[HSN96] Bent Halling-Sørensen and Søren Nors Nielsen. A model of nitrogen removal from waste water in a fixed bed reactor using simultaneous nitrification and denitrification (SND). Ecological Modelling, 87(1–3):131–141, June 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095000259.

Hary:1996:LDA

[HSPM96] Ingo Hary, Horst-Juergen Schwartz, Volker H. C. Pielert, and Christoph Mosler. Land degradation in African pastoral systems and the destocking controversy. Ecological Modelling, 86(2–3):227–233, May 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095000569.

Halfon:1996:EFT

[HSU96] Efraim Halfon, Natalie Schito, and Robert E. Ulanowicz. Energy flow through the Lake Ontario food web: conceptual model and an attempt at mass balance. Ecological Modelling, 86(1):1–36, April 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094001952.

Hu:1998:MEW

[HSXP98] Weiping Hu, Jørgen Salomonsen, Fu-Liu Xu, and Peiming Pu. A model for the effects of water hyacinths on water quality in an experiment of physico-biological engineering in Lake Taihu, China. Ecological Modelling, 107(2–3):171–188, April 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097002196.

Hallam:1996:MEC

[HTW96] Thomas G. Hallam, Tamara L. Trawick, and Wilfried F. Wolff. Modeling effects of chemicals on a population: application to a wading bird nesting colony. Ecological Modelling, 92(2–3):155–178, December 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095002049.

Huang:1998:MCG

[Hua98] Wending Huang. Modelling the coexistence gain and interactions of populations in Taxodium ascendens-intercrop systems. Ecological Modelling, 107(2–3):189–212, April 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097002172.

Hughes:1991:EPD

[Hug91] Gareth Hughes. Epidemics of plant diseases: Mathematical analysis and modeling: (2nd, completely revised edition). Jürgen Kranz (Editor). Ecological Studies, 13. Springer, Berlin, 1990. 268 pp., DM160.00 ISBN 3-540-52116-X. Ecological Modelling, 59(3–4):296–298, December 15, 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380091901854.

Hulburt:1992:EAP

[Hul92] Edward M. Hulburt. Equivalence and the adaptationist program. Ecological Modelling, 64(4):305–329, November 1992.

CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290029E.

Hulburt:1996:SAP

[Hul96] Edward M. Hulburt. The symmetry of adaptation in predominantly asymmetrical contexts. Ecological Modelling, 85 (2–3):173–185, March 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380094001502.

Hulburt:1998:TAA

[Hul98] Edward M. Hulburt. Theory of adaptation: application of symbolic logic. Ecological Modelling, 107(1):35–50, March 16, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097001993.

Hutchings:1991:MGS

[Hut91] N. J. Hutchings. A model of a grass sward continuously grazed to a constant mean height. Ecological Modelling, 59(1–2):73–91, December 1, 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009190128N.

Hackett:1998:MEK

[HV98] C. Hackett and J. K. Vanclay. Mobilizing expert knowledge of tree growth with the PLANTGRO and INFER systems. Ecological Modelling, 106(2–3):233–246, March 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001853.

Hance:1999:IIA

[HV99] Th. Hance and G. Van Impe. The influence of initial age structure on predator–prey interaction. Ecological Modelling, 114(2–3):195–211, January 1, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098001483.

Hearne:1994:DSB

[HvCC94] J. W. Hearne, L. M. van Coller, and D. E. Conlong. Determining strategies for the biological control of a sugarcane stalk borer. Ecological Modelling, 73(1–2):117–133, May 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094901015.

Hinrichsen:1998:BPC

[HW98] Ulf Hinrichsen and Fredrik Wulff. Biogeochemical and physical controls of nitrogen fluxes in a highly dynamic marine ecosystem — model and network flow analysis of the Baltic Sea. Ecological Modelling, 109(2):165–191, June 11, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098000519.

Howden:1996:MSG

[HWB96] S. M. Howden, D. H. White, and P. J. Bowman. Managing sheep grazing systems in southern Australia to minimise greenhouse gas emissions: adaptation of an existing simulation model. Ecological Modelling, 86(2–3):201–206, May 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095000526.

Iwasa:1995:FGD

[IK95] Yoh Iwasa and Takuya Kubo. Forest gap dynamics with partially synchronized disturbances and patch age distribution. Ecological Modelling, 77(2–3):257–271, February 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093E0084G.

Ilichev:1996:MPB

[Il’96] Vitaly G. Il’ichev. Mathematical problems of biological competition theory in changing environment. Ecological Modelling, 93(1–3):191–201, December 16, 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438009600004X.

Iverson:1999:MPF

[IPS99] Louis R. Iverson, Anantha Prasad, and Mark W. Schwartz. Modeling potential future individual tree-species distributions in the eastern United States under a climate change scenario: a case study with Pinus virginiana. Ecological Modelling, 115(1):77–93, February 1, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098002002.

Tainaka:1996:UEC

[iT96] Kei ichi Tainaka. Uncertainty in ecological catastrophe. Ecological Modelling, 86(2–3):125–128, May 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095000380.

Invernizzi:1997:GLM

[IT97] Sergio Invernizzi and Katia Terpin. A generalized logistic model for photosynthetic growth. Ecological Modelling, 94(2–3):231–242, January 15, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096000245.

Isabel:1991:SDC

[IV91] Denis Isabel and Jean-Pierre Villeneuve. Significance of the dispersion coefficient in the stochastic modelling of pesticides transport in the unsaturated zone. Ecological Modelling, 59(1–2):1–10, December 1, 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009190125K.

Wu:1997:UPI

[iWHS97] Hsin i Wu, Thomas J. Hatton, and Revin L. Stoker. Use of power index and two-phase density approach to study fine root dynamics. Ecological Modelling, 95(1):87–93, February 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096000300.

Wu:1996:SAG

[iWLSL96] Hsin i Wu, Bai-Lian Li, Revin Stoker, and Yang Li. A semiarid grazing ecosystem simulation model with probabilistic and fuzzy parameters. Ecological Modelling, 90(2):147–160, October 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009500162X.

Wu:1996:MLV

[iWSG96] Hsin i Wu, Revin L. Stoker, and Longchang Gao. A modified Lotka–Volterra simulation model to study the interaction between arrow bamboo (Sinarundinaria fangiana) and giant panda (Ailuropoda melanoleuca). Ecological Modelling, 84(1–3):11–17, January 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400109X.

Jayaweera:1996:MBS

[JA96] Mahesh Jayaweera and Takashi Asaeda. Modeling of biomanipulation in shallow, eutrophic lakes: an application to Lake Bleiswijkse Zoom, The Netherlands. Ecological Modelling, 85(2–3):113–127, March 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094001537.

Jackson:1996:SMP

[Jac96] Leland J. Jackson. A simulation model of PCB dynamics in the Lake Ontario pelagic food web. Ecological Modelling, 93(1–3):43–56, December 16, 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095002103.

James:1993:SAS

[Jam93] R. Thomas James. Sensitivity analysis of a simulation model of methane flux from the Florida Everglades. Ecological Modelling, 68(3–4):119–146, August 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009390013I.

Jarvis:1995:SSW

[Jar95] N. J. Jarvis. Simulation of soil water dynamics and herbicide persistence in a silt loam soil using the MACRO model. Ecological Modelling, 81(1–3):97–109, August 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009400163C.

Jiang:1999:MSP

[JAZ^{+}99] Hong Jiang, Michael J. Apps, Yanli Zhang, Changhui Peng, and Paul M. Woodard. Modelling the spatial pattern of net primary productivity in Chinese forests. Ecological Modelling, 122(3):275–288, October 20, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099001428.

Johst:1997:EDL

[JB97] Karin Johst and Roland Brandl. The effect of dispersal on local population dynamics. Ecological Modelling, 104(1):87–101, December 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001129.

Janovy:1995:SDD

[JCC^{+}95] J. Janovy, R. E. Clopton, D. A. Clopton, Scott D. Snyder, Aris Efting, and Laura Krebs. Species density distributions as null models for ecologically significant interactions of parasite species in an assemblage. Ecological Modelling, 77(2–3):189–196, February 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380093E0087J.

Jager:1997:MLB

[JCS^{+}97] Henriette I. Jager, Hal E. Cardwell, Michael J. Sale, Mark S. Bevelhimer, Charles C. Coutant, and Webb Van Winkle. Modelling the linkages between flow management and salmon recruitment in rivers. Ecological Modelling, 103(2–3):171–191, November 17, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097000847.

Jorgensen:1999:CMT

[JD99] Eric E. Jorgensen and Stephen Demarais. A comparison of modelling techniques for small mammal diversity. Ecological Modelling, 120(1):1–8, August 3, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009900054X.

Jensen:1991:SFP

[Jen91] A. L. Jensen. Simulation of fish population responses to exploitation. Ecological Modelling, 55(3–4):203–218, August 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009190087H.

Jensen:1993:DFP

[Jen93] A. L. Jensen. Dynamics of fish populations with different compensatory processes when subjected to random survival of eggs and larvae. Ecological Modelling, 68(3–4):249–256, August 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390020S.

Jensen:1994:DPN

[Jen94] A. L. Jensen. Dynamics of populations with nonoverlapping generations, continuous mortality, and discrete reproductive periods. Ecological Modelling, 74(3–4):305–309, August 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094901252.

Jensen:1995:SDD

[Jen95] A. L. Jensen. Simple density-dependent matrix model for population projection. Ecological Modelling, 77(1):43–48, January 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380093E0081D.

Jensen:1996:DDM

[Jen96a] A. L. Jensen. Density-dependent matrix yield equation for optimal harvest of age-structured wildlife populations.

Ecological Modelling, 88(1–3):125–132, July 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095000755.

Jensen:1996:PMF

[Jen96b] A. L. Jensen. A process model for food competition. Ecological Modelling, 87(1–3):1–9, June 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094001960.

Jensen:1997:LSC

[Jen97] A. L. Jensen. Limiting similarity for coexistence of lake herring (Coregonus artedii) and chubs (Coregonus hoyi). Ecological Modelling, 95(1):11–15, February 1, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095002146.

Jessup:1998:SSB

[Jes98] Benjamin K. Jessup. A strategy for simulating brown trout population dynamics and habitat quality in an urbanizing watershed. Ecological Modelling, 112(2–3):151–167, October 15, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098000775.

Jesse:1999:MDL

[Jes99] Klemens J. Jesse. Modelling of a diffusive Lotka–Volterrasystem: the climate-induced shifting of tundra and forest realms in North-America. Ecological Modelling, 123(2–3): 53–64, November 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438009900126X.

Jetschke:1992:SPM

[Jet92] Gottfried Jetschke. Stochastic population models and their relevance for the conservation of species. Ecological Modelling, 63(1–4):71–89, September 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290062J.

Janssen:1995:CPO

[JH95] P. H. M. Janssen and P. S. C. Heuberger. Calibration of process-oriented models. Ecological Modelling, 83(1–2):55–66, December 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095000849.

Johnson:1995:SDD

[JHM95] Alan R. Johnson, Colleen A. Hatfield, and Bruce T. Milne. Simulated diffusion dynamics in river networks. Ecological Modelling, 83(3):311–325, December 15, 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094001079.

Jans-Hammermeister:1997:ETS

[JHM97] Desirée C. Jans-Hammermeister and William B. McGill. Evaluation of three simulation models used to describe plant residue decomposition in soil. Ecological Modelling, 104(1):1–13, December 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097000720.

Janowitz:1999:EEM

[JK99] Gerald S. Janowitz and Daniel Kamykowski. An expanded Eulerian model of phytoplankton environmental response. Ecological Modelling, 118(2–3):237–247, June 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009900037X.

Jorgensen:1998:DME

[JLS98] Sven Erik Jørgensen, Hans Christian Lützhøft, and Bent Halling Sørensen. Development of a model for environmental risk assessment of growth promoters. Ecological Modelling, 107(1):63–72, March 16, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097002111.

Jorgensen:1997:MFS

[JMA97] S. E. Jørgensen, J. C. Marques, and P. M. Anastácio. Modelling the fate of surfactants and pesticides in a rice field. Ecological Modelling, 104(2–3):205–213, December 8, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001348.

Jorgensen:1998:ESO

[JMN98] S. E. Jørgensen, Henning Mejer, and Søren Nors Nielsen. Ecosystem as self-organizing critical systems. Ecological Modelling, 111(2–3):261–268, September 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001045.

Johnson:1999:MFP

[JMPT99] Glen D. Johnson, Wayne L. Myers, Ganapati P. Patil, and Charles Taillie. Multiresolution fragmentation profiles for assessing hierarchically structured landscape patterns. Ecological Modelling, 116(2–3):293–301, March 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098002026.

Jorgensen:1992:I

[JMW92] Sven Erik Jørgensen, Felix Müller, and Wilhelm Windhorst. Introduction. Ecological Modelling, 63(1–4):xi–xiii, September 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009290057L.

Jorgensen:1994:MSD

[JN94] S. E. Jørgensen and Søren Nors Nielsen. Models of the structural dynamics in lakes and reservoirs. Ecological Modelling, 74(1–2):39–46, July 1994. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094901090.

Jorgensen:1995:EEE

[JNM95] Sven Erik Jørgensen, Søren Nors Nielsen, and Henning Mejer. Emergy, environ, exergy and ecological modelling. Ecological Modelling, 77(2–3):99–109, February 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093E0080M.

Jolayemi:1995:MPP

[JO95] Joel K. Jolayemi and J. O. Olaomi. A mathematical programming procedure for selecting crops for mixed-cropping schemes. Ecological Modelling, 79(1–3):1–9, May 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009400041F.

Jogiste:1998:PMS

[Jõg98] Kalev Jõgiste. Productivity of mixed stands of Norway spruce and birch affected by population dynamics: a model analysis. Ecological Modelling, 106(1):77–91, February 16, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438009700183X.

Jolayemi:1996:IMP

[Jol96] Joel K. Jolayemi. An integrated model for planning and managing multi-regional mixed-crop farming schemes. Ecological Modelling, 84(1–3):63–74, January 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094001464.

Jorgensen:1990:AYJ

[Jør90] S. E. Jørgensen. About your journal: Ecological modelling. Ecological Modelling, 50(1–3):1–4, March 1990. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009090039J.

Jorgensen:1991:EIO

[Jør91a] S. E. Jørgensen. Ecological impacts of the oil industry: Brian Dicks (Editor). Wiley, 316 pp. Published on behalf of Petroleum, London, 1989. Proceedings of International Meeting organized by the Institute of Petroleum and held in London in November 1987. £80.00. ISBN 0-471-92193-9. Ecological Modelling, 54(1–2):146–147, May 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009190109E.

Jorgensen:1991:GAO

[Jør91b] S. E. Jørgensen. Global alert: the ozone pollution crisis: Jack Fishman and Robert Kalish. Plenum, New York, 1990. Hardbound, illustrated, 311 pp., US$24.50 ISBN 0-306-43455-5. Ecological Modelling, 59(1–2):139–140, December 1, 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190131J.

Jorgensen:1991:MME

[Jør91c] S. E. Jørgensen. Mathematical modeling in ecology — a workbook for students: Clark Jeffries. Birkhäuser, Boston/Basel/ Berlin, 1989. Hardbound, illustrated, 193 pp., US$29.95. ISBN 0-8176-3421-5. Ecological Modelling, 59(1–2):140–141, December 1, 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190132K.

Jorgensen:1992:ACS

[Jør92a] S. E. Jørgensen. The art of computer systems performance analysis: Techniques for Experimental Design, Measurement, Simulation and Modeling. Raj Jain. John Wiley, New York. Hardcover, 720 p. U. S. $52.95. Ecological Modelling, 64(1):81–82, October 15, 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290054I.

Jorgensen:1992:ETE

[Jør92b] S. E. Jørgensen. Ecosystem theory and ecological modelling. Ecological Modelling, 62(1–3):vii–ix, July 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290079T.

Jorgensen:1992:DMA

[Jør92c] Sven Erik Jørgensen. Development of models able to account for changes in species composition. Ecological Modelling, 62(1–3):195–208, July 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290091R.

Jorgensen:1992:EE

[Jør92d] Sven Erik Jørgensen. Exergy and ecology. Ecological Modelling, 63(1–4):185–214, September 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290069Q.

Jorgensen:1992:PEC

[Jør92e] Sven Erik Jørgensen. Parameters, ecological constraints and exergy. Ecological Modelling, 62(1–3):163–170, July 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290088V.

Jorgensen:1993:EES

[Jør93a] S. E. Jørgensen. Ecological economics: the science and management of sustainability: Robert Costanza (Editor), Columbia University Press, 1991. Softbound $21.50, hardbound $57.00, 525 pp. Proceedings of a workshop held May 1990 by ISEE (International Society of Ecological Economics). Ecological Modelling, 70(3–4):323–324, December 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390064Y.

Jorgensen:1993:SEE

[Jør93b] S. E. Jørgensen. Soil erosion — experiments and models: R. B. Brian (Editor). Catena Supplement No. 17, Catena Verlag, Brockenblick 8, W-3302 Cremlingen 4, Germany, 1990. 208 pp. DM 139/US $75. ISSN 0722-0723/ISBN 3-923381-22-0. Ecological Modelling, 65(1–2):148–149, January 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009390132C.

Jorgensen:1993:JRC

[Jør93c] S. E. Jørgensen. Jörg Richter (Coordinator), Models for processes in the soil — programs and exercises: Catena Verlag, Brockenbleck 8, W-3302 Cremlinger 4, Germany, 1990. 240 pp. DM 38.50/US $24. ISBN 3-923381-24-7. Ecological Modelling, 65(1–2):147–148, January 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390131B.

Jorgensen:1993:MTA

[Jør93d] S. E. Jørgensen. Model-theoretical aspects of ecological modelling. Ecological Modelling, 68(1–2):viii, July 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390103Y.

Jorgensen:1993:YJN

[Jør93e] Sven Erik Jørgensen. Your journal has a new name! Ecological Modelling, 65(1–2):v, January 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009390123A.

Jorgensen:1994:DHC

[Jør94a] S. E. Jørgensen. Development in hydrobiology comparative reservoir limnology and water quality management: M. Straškraba, J. G. Tundisi and A. Duncan (Editors). Kluwer Academic Publishers, Dordrecht, 291 pp. hardbound, ISBN 0-7923-1919-2. Ecological Modelling, 72(3–4):279–281, April 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094900906.

Jorgensen:1994:GWE

[Jør94b] S. E. Jørgensen. Global warming and economic development. A holistic approach to international policy. Co-operation and Co-ordination: Anantha K. Durraiappah. Kluwer Academic Publishers, Dordrecht, 224 pp., hardbound, ISBN 0-7923-2149-9. Ecological Modelling, 72(3–4):278–279, April 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094900892.

Jorgensen:1994:IEE

[Jør94c] S. E. Jørgensen. Integrating economics, ecology and thermodynamics: Matthias Ruth. Kluwer Academics Publishers, Dordrecht, The Netherlands, 1993, 251 pp. ISBN 0-7923-2377-7. Ecological Modelling, 72(3–4):275–276, April 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094900876.

Jorgensen:1994:TUE

[Jør94d] S. E. Jørgensen. Toward a unified ecology: Timothy F. H. Allen and Thomas W. Hoekstra. Columbia University Press, New York, 1992. Softbound edition, $32.50, 383 pp., ISBN 0-231-06919-7. Ecological Modelling, 72(3–4):276–277, April 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094900884.

Jorgensen:1994:HWI

[Jør94e] Sven Erik Jørgensen. Hazardous wastes and international law: Transboundary Movements and Disposal of Hazardous Wastes in International Law. Barbara Kwiatkowska and Alfred H. A. Soons (Editors). With a preface by Mostafa K. Tolba. Martinus Nijhoff /Graham Trotman, Dordrecht, 1390 pp., US$278, ISBN 0-7923-1667-3. Ecological Modelling, 74(3–4):311, August 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094901260.

Jorgensen:1994:ISS

[Jør94f] Sven Erik Jørgensen. Integrated Soil and Sediment Research: a Basis for Proper Protection. Selected Proceedings of the First European Conference on Integrated Research for Soil and Sediment Protection and Remediation (EUROSOL), Maastricht 6–12 September 1992. J. P. Eijsackers and Timo Hamers (Editors). Kluwer Academic Publ., Dordrecht, 765 pp., US $236, £160, ISBN 0-7923-2321-1. Ecological Modelling, 74(3–4):313–314, August 1994. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094901287.

Jorgensen:1994:RSI

[Jør94g] Sven Erik Jørgensen. Risk and society: the Interaction of Science, Technology and Public Policy. Marvin Waterstone (Editor), Kluwer Academic Publishers, Dordrecht, 180 pp., US$126, ISBN 0-7923-1370-4. Ecological Modelling, 74(3–4):312–313, August 1994. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094901279.

Jorgensen:1995:CEE

[Jør95a] Sven Erik Jørgensen. Concise encyclopedia of environmental systems: Peter C. Young (Editor), 1993. Pergamon, Oxford. 769 pp., £195, ISBN 0-08-036198. Ecological Modelling, 77(1):85–86, January 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095900330.

Jorgensen:1995:DEB

[Jør95b] Sven Erik Jørgensen. Dynamic energy budgets in biological systems. Theory and applications in ecotoxicology: S. A. L. M. Kooijman. Cambridge University Press, Cambridge, 1993. Hardbound, 350 pp., ISBN 0-521-452323-6. Ecological Modelling, 79(1–3):288–289, May 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095900675.

Jorgensen:1995:EAS

[Jør95c] Sven Erik Jørgensen. Evolution in age-structured populations: 2nd ed. Brian Charlesworth. Cambridge University Press, 1994. ISBN 0-521-45967-2, 306 pp., US $29.95. Ecological Modelling, 78(3):288, April 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095900799.

Jorgensen:1995:NSC

[Jør95d] Sven Erik Jørgensen. Natural sinks of CO_{2}: J. Wisniewski and A. E. Lugo (Editors). Kluwer Academic, Dordrecht, 1992. ISBN 0-7923-1805-6, 468 pp. Ecological Modelling, 79(1–3): 287, May 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900659.

Jorgensen:1995:QES

[Jør95e] Sven Erik Jørgensen. Quantitative ecology. Spatial and temporal scaling: David C. Schneider. Academic Press, San Diego, 1994. ISBN 12-627860-1, 395 pp., US $49.95. Ecological Modelling, 79(1–3):288, May 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095900667.

Jorgensen:1995:SDN

[Jør95f] Sven Erik Jørgensen. Sulphur dioxide and nitrogen oxides in industrial waste gases: Emission, legislation and abatement: D. van Velzen (Editor). Kluwer Academic, Dordrecht, 1991. Eurocourses. Chemical and Environmental Science, Vol. 3. ISBN 0-7923-1386-0, 281 pp. Ecological Modelling, 80(2–3): 299, July 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900411.

Jorgensen:1995:WHR

[Jør95g] Sven Erik Jørgensen. Wildlife habitat relationships in forested ecosystems: David R. Patton. Timber Press, Portland, OR, 1992. ISBN 0-88192-202-1, 394 pp. Ecological Modelling, 78(3): 287, April 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900780.

Jorgensen:1995:SAE

[Jør95h] Sven Erik Jørgensen. State of the art of ecological modelling in limnology. Ecological Modelling, 78(1–2):101–115, March 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094001207.

Jorgensen:1996:CMW

[Jør96a] Sven Erik Jørgensen. Computational methods in water resources X, volumes 1 and 2: A. Peters, G. Wittum, B. Herrling, U. Meissner, C. A. Brebbia, W. Gray and G. F. Pinder (Editors). Kluwer, Dordrecht, 1994. ISBN 0-7923-2935-X and 0-7923-2936-8, 1548 pp. Ecological Modelling, 87(1–3):300– 301, June 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380096881713.

Jorgensen:1996:E

[Jør96b] Sven Erik Jørgensen. Editorial. Ecological Modelling, 93 (1–3):v–vi, December 16, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096900897.

Jorgensen:1997:IDM

[Jør97a] S. E. Jørgensen. Introduction to the development of models with dynamic structure for marine ecosystems. Ecological Modelling, 102(1):1–3, October 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097000902.

Jorgensen:1997:IEM

[Jør97b] Sven Erik Jørgensen. An introduction to ecological modelling putting practice into theory. The methods in ecology series: Michael Gillman and Rosemary Hails (Eds.), Blackwell Science, Oxford; 1997, 202 pp.; Soft bound; Illustrated. Ecological Modelling, 103(2–3):287, November 17, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097001142.

Jorgensen:1997:DEI

[Jør97c] Sven Erik Jørgensen. Detecting ecological impacts, concepts and applications in coastal habitats: Edited by Russel J. Schmitt and Craig W. Osenberg. Academic Press, San Diego, New York, Boston, London, Sidney, Tokyo and Toronto, 1996. Hardbound, 401 pages. Ecological Modelling, 97(1–2):148–149, April 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097839221.

Jorgensen:1997:DEB

[Jør97d] Sven Erik Jørgensen. Dynamics of environmental bioprocesses, modelling and simulation: By J. B. Snarpe, I. J. Dunn, J. Ingham and J. E. Prenosil. VHC Weinheim, New York, Basel, Cambridge and Tokyo, 1995. Hardbound, Price DM 258, ISBN 3-527-28705-1, 492 pages and with a software diskette. Ecological Modelling, 97(1–2):148, April 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438009783921X.

Jorgensen:1997:FWI

[Jør97e] Sven Erik Jørgensen. Food webs, integration of patterns and dynamics: Edited by Gary A. Polis and Kirk O. Winemiller. Chapman and Hall, New York, Albany, Bonn, Boston, Cincinnati, Detroit, London, Mardi Melbourne, Mexico City, Pacific Grove, Paris, San Francisco, Singapore, Tokyo, Toronto, Washington, 1996. Hardbound, £45, ISBN 0-412-04051-4, 472 pages. Ecological Modelling, 97(1–2):147–148, April 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097839208.

Jorgensen:1997:ORE

[Jør97f] Sven Erik Jørgensen. Operations research and environmental management: By Carlo Carraro and Alain Haurie. Kluwer Academic Publishers, Dordrecht, Boston and London, 1996. Hardbound, US $129.50 or £83.00, ISBN 0-7923-3767-0, 267 pages. Ecological Modelling, 97(1–2):149, April 15, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097839233.

Jorgensen:1997:RSE

[Jør97g] Sven Erik Jørgensen. River and stream ecosystems, ecosystems edited of the world 22: Edited by C. E. Cushing, K. W. Cummins and G. W. Minshall. Elsevier, Amsterdam, Lausanne, New York, Oxford, Shannon and Tokyo. Hardbound, 817 pages. Ecological Modelling, 97(1–2):150–151, April 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097839257.

Jorgensen:1997:FLM

[Jør97h] Sven Erik Jørgensen. The future of the land, mobilising and integrating knowledge for land use options: Edited by L. O. Fresco, L. Stroosnijder, J. Bouma and H. van Keulen. John Wiley and Sons, Chichester, New York, Brisbane, Toronto and Singapore, 1994. Hardbound, ISBN: 0-47-1-95017, 3–65§, 409 pages. Ecological Modelling, 97(1–2):149–150, April 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097839245.

Jorgensen:1997:EME

[Jør97i] Sven Erik Jørgensen. Ecological modelling by ‘Ecological Modelling’. Ecological Modelling, 100(1–3):5–10, December 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097001580.

Jorgensen:1999:CSI

[Jør99a] Sven Erik Jørgensen. Contaminated Soils 3rd International Conference on the Biogeochemistry of Trace Elements, Paris 15–19 May 1995. R. Prost (Ed.) An INRA Edition, 1997. 528 pp. + CD Rom, ISBN 2-7380-0775-9; 380 ff. Ecological Modelling, 117(1):165, April 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000289.

Jorgensen:1999:FSI

[Jør99b] Sven Erik Jørgensen. Forest Sciences. Impact of Global Change on Tree Physiology Forest Ecosystems. Proceedings of the International Conference on Impacts of Global Change on tree physiology and Forest Ecosystems, 26–29 November 1996 in Wageningen, The Netherlands. Edited by G. M. J. Mohren, K. Kramer and S. Sabate. Kluwer, Dordrecht, 1997. ISBN: 0-7923-4921-0. 372 pages hardbound, US$169. Ecological Modelling, 118(1):90–91, June 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000344.

Jorgensen:1999:EMS

[Jør99c] Sven Erik Jørgensen. Ecological modelling and systems ecology. Ecological Modelling, 117(1):1–2, April 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000253.

Jorgensen:1999:SAE

[Jør99d] Sven Erik Jørgensen. State-of-the-art of ecological modelling with emphasis on development of structural dynamic models. Ecological Modelling, 120(2–3):75–96, August 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099000939.

Jiang:1999:MNP

[JPA^{+}99] Hong Jiang, Changhui Peng, Michael J. Apps, Yanli Zhang, Paul M. Woodard, and Zhiming Wang. Modelling the net primary productivity of temperate forest ecosystems in China with a GAP model. Ecological Modelling, 122(3):225–238, October 20, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001398.

Jorgensen:1992:EET

[JPS92] Sven E. Jørgensen, Bernard C. Patten, and Milan Straškraba. Ecosystems emerging: toward an ecology of complex systems in a complex future. Ecological Modelling, 62(1–3):1– 27, July 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290080X.

Jorgensen:1999:EEO

[JPS99] Sven E. Jørgensen, Bernard C. Patten, and Milan Straškraba. Ecosystems emerging: 3. Openness. Ecological Modelling, 117(1):41–64, April 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380098001963.

Jaworska:1997:IBM

[JRB97] Joanna S. Jaworska, Kenneth A. Rose, and Antoinette L. Brenkert. Individual-based modeling of PCBs effects on youngof-the-year largemouth bass in southeastern USA reservoira. Ecological Modelling, 99(2–3):113–135, June 30, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096019436.

Jizhong:1991:IED

[JSC91] Zhou Jizhong, Ma Shijun, and Chen Changming. An index of ecosystem diversity. Ecological Modelling, 59(3–4): 151–163, December 15, 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380091901762.

Jakeman:1992:ETA

[JTS92] A. J. Jakeman, J. A. Taylor, and H. D. Symons. Efficient tools for analysing the influence of sources and meteorology on urban ambient concentration trends illustrated for Canberra, Australia. Ecological Modelling, 64(2–3):125–157, October 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290112R.

Jeltsch:1992:ODP

[JWEB92] F. Jeltsch, Ch. Wissel, S. Eber, and R. Brandl. Oscillating dispersal patterns of tephritid fly populations. Ecological Modelling, 60(1):63–75, January 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380092900135.

Jonsson:1995:ASM

[JWS95] Charlotta Jönsson, Per Warfvinge, and Harald Sverdrup. Application of the safe model to the Solling spruce site. Ecological Modelling, 83(1–2):85–96, December 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009500088D.

Jax:1992:RTD

[JZV92] Kurt Jax, Gerd-Peter Zauke, and Ekkehard Vareschi. Remarks on terminology and the description of ecological systems. Ecological Modelling, 63(1–4):133–141, September 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290066N.

Keane:1990:MSD

[KABT90] Robert E. Keane, Stephen F. Arno, James K. Brown, and Diana F. Tomback. Modelling stand dynamics in whitebark pine (Pinus albicaulis) forests. Ecological Modelling, 51(1–2):73– 95, May 1990. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009090059P.

Kaiser:1996:MCM

[Kai96] J. Kaiser. Modelling composting as a microbial ecosystem: a simulation approach. Ecological Modelling, 91(1– 3):25–37, November 15, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095001573.

Kalita:1999:TFE

[Kal99] Prasanta K. Kalita. Transient finite element method solution of oxygen diffusion in soil. Ecological Modelling, 118(2–3):227–236, June 15, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099000381.

Kangas:1990:DDL

[Kan90] Patrick Kangas. Deforestation and diversity of life zones in the Brazilian Amazon: a map analysis. Ecological Modelling, 49(3–4):267–275, January 1990. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009090030K.

Kangas:1991:MMD

[Kan91] Patrick Kangas. Macroscopic minimodels of deforestation and diversity. Ecological Modelling, 57(3–4):277–294, October 15, 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190117J.

Karev:1993:ADP

[Kar93] G. P. Karev. Age-dependent population dynamics with several interior variables and spatial spread. Ecological Modelling, 70(3–4):277–288, December 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009390061V.

Kasimovsky:1993:ASM

[Kas93] A. A. Kasimovsky. Analytical solution for the model of soil radionuclide migration with fixation-leaching reaction. Ecological Modelling, 66(3–4):217–229, April 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093901148.

Katerji:1990:USM

[Kat90] Nader Katerji. Use of simulation methods for determining critical leaf water potential for stomatal closure in field conditions. Ecological Modelling, 50(1–3):133–144, March 1990. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009090046J.

Kawata:1997:ECE

[Kaw97] Masakado Kawata. Exploitative competition and ecological effective abundance. Ecological Modelling, 94(2–3):125–137, January 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096000087.

Klepper:1997:NPE

[KB97a] Olivier Klepper and Jacques J. M. Bedaux. Nonlinear parameter estimation for toxicological threshold models. Ecological Modelling, 102(2–3):315–324, October 29, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097000665.

Krebs:1997:EVO

[KB97b] Friedrich Krebs and Hartmut Bossel. Emergent value orientation in self-organization of an animat. Ecological Modelling, 96(1–3):143–164, March 1, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096000506.

Kindlmann:1999:EPT

[KB99] Pavel Kindlmann and Zuzana Balounová. Energy partitioning in terrestrial orchids a model for assessing their performance. Ecological Modelling, 119(2–3):167–176, July 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099000472.

Kawashima:1996:GBN

[KBL96] H. Kawashima, M. J. Bazin, and J. M. Lynch. Global N_{2}O balance and nitrogen fertilizer. Ecological Modelling, 87(1–3): 51–57, June 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094002045.

Krivtsov:1998:ASX

[KBSC98] V. Krivtsov, E. Bellinger, D. Sigee, and J. Corliss. Application of SEM XRMA data to lake ecosystem modelling. Ecological Modelling, 113(1–3):95–123, November 2, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001379.

Kolbe:1999:GEU

[KBV99] Audra E. Kolbe, Joseph Buongiorno, and Michael Vasievich. Geographic extension of an uneven-aged, multi-species matrix growth model for northern hardwood forests. Ecological Modelling, 121(2–3):235–253, September 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000794.

Krapivin:1998:AMT

[KCP^{+}98] V. F. Krapivin, V. A. Cherepenin, G. W. Phillips, R. A. August, A. Yu. Pautkin, M. J. Harper, and F. Y. Tsang. An application of modelling technology to the study of radionuclear pollutants and heavy metals dynamics in the Angara–Yenisey river system. Ecological Modelling, 111(2–3):121–134, September 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098000908.

Kindlmann:1996:PDT

[KD96] P. Kindlmann and A. F. G. Dixon. Population dynamics of a tree-dwelling aphid: individuals to populations. Ecological Modelling, 89(1–3):23–30, August 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001166.

Keller:1999:MMK

[KD99] Richard M. Keller and Jennifer L. Dungan. Meta-modeling: a knowledge-based approach to facilitating process model construction and reuse. Ecological Modelling, 119(2–3):89–116, July 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098001975.

Klepper:1999:CSE

[KdH99] Olivier Klepper and Henri A. den Hollander. A comparison of spatially explicit and box models for the fate of chemicals in water, air and soil in Europe. Ecological Modelling, 116(2–3):183–202, March 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001616.

Katterer:1997:WWB

[KEAP97] Thomas Kätterer, Henrik Eckersten, Olof Andrén, and Roger Pettersson. Winter wheat biomass and nitrogen dynamics under different fertilization and water regimes: application of a crop growth model. Ecological Modelling, 102(2–3):301–314, October 29, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097000653.

Keitt:1997:SCL

[Kei97] Timothy H. Keitt. Stability and complexity on a lattice: coexistence of species in an individual-based food web model. Ecological Modelling, 102(2–3):243–258, October 29, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097000598.

Kendal:1992:FSG

[Ken92] Wayne S. Kendal. Fractal scaling in the geographic distribution of populations. Ecological Modelling, 64(1):65–69, October 15, 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290050O.

Kendal:1995:PMV

[Ken95] Wayne S. Kendal. A probabilistic model for the variance to mean power law in ecology. Ecological Modelling, 80(2–3):293– 297, July 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400053K.

Kersebaum:1995:ASM

[Ker95] K. C. Kersebaum. Application of a simple management model to simulate water and nitrogen dynamics. Ecological Modelling, 81(1–3):145–156, August 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400167G.

Killion:1995:CGM

[KG95] Michael J. Killion and William E. Grant. A colony-growth model for the imported fire ant: potential geographic range of an invading species. Ecological Modelling, 77(1):73–84, January 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400029H.

King:1996:SMI

[KG96] Sammy L. King and William E. Grant. A simulation model of the impacts of green-tree reservoir management on bottomland hardwood seedling growth and survival. Ecological Modelling, 87(1–3):69–82, June 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094002061.

Kerr:1997:SMI

[KGD97] Sara F. Kerr, William E. Grant, and Norman O. Dronen. A simulation model of the infection cycle of Leishmania mexicana in Neotoma micropus. Ecological Modelling, 98(2–3):187–197, May 30, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438009601914X.

Klepper:1994:MRC

[KH94a] Olivier Klepper and Eligius M. T. Hendrix. A method for robust calibration of ecological models under different types of uncertainty. Ecological Modelling, 74(3–4):161–182, August 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009490118X.

Kmet:1994:DLV

[KH94b] T. Kmet’ and J. Holčík. The diffusive Lotka–Volterra model as applied to the population dynamics of the German carp and predator and prey species in the Danube River basin. Ecological Modelling, 74(3–4):277–285, August 1994. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094901236.

Kaitaniemi:1996:SSS

[KH96] Pekka Kaitaniemi and Tuija Honkanen. Simulating sourcesink control of carbon and nutrient translocation in a modular plant. Ecological Modelling, 88(1–3):227–240, July 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001093.

Kohler:1998:ETS

[KH98] Peter Köhler and Andreas Huth. The effects of tree species grouping in tropical rainforest modelling: Simulations with the individual-based model Formind. Ecological Modelling, 109(3):301–321, June 22, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000660.

Koh:1997:CEE

[KHL97] Hock Lye Koh, Thomas G. Hallam, and Hooi Ling Lee. Combined effects of environmental and chemical stressors on a model Daphnia population. Ecological Modelling, 103(1):19–32, November 1, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097000732.

Kawano:1993:LSM

[KI93] Kohji Kawano and Yoh Iwasa. A lattice-structured model for beech forest dynamics: the effect of understory dwarf bamboo. Ecological Modelling, 66(3–4):261–275, April 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390117B.

Kubo:1998:SIB

[KI98] Takuya Kubo and Hideyuki Ida. Sustainability of an isolated beech-dwarf bamboo stand: analysis of forest dynamics with individual based model. Ecological Modelling, 111(2–3):223–235, September 1, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438009800115X.

Kim:1994:ASS

[Kim94] Seungdo Kim. Analytical solution schemes for phosphorus transport equations of a steady state in a stream. Ecological Modelling, 71(4):221–243, February 1994. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009490135X.

King:1995:EAD

[Kin95] David A. King. Equilibrium analysis of a decomposition and yield model applied to Pinus radiata plantations on sites of contrasting fertility. Ecological Modelling, 83(3):349–358, December 15, 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094001031.

King:1996:MEF

[Kin96] David A. King. A model to evaluate factors controlling growth in Eucalyptus plantations of southeastern Australia. Ecological Modelling, 87(1–3):181–203, June 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380096881658.

Kellomaki:1999:MSG

[KIPK99] Seppo Kellomäki, Veli-Pekka Ikonen, Heli Peltola, and Taneli Kolström. Modelling the structural growth of Scots pine with implications for wood quality. Ecological Modelling, 122(1–2):117–134, October 1, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099000861.

Kirsta:1992:TDQ

[Kir92] Yu. B. Kirsta. Time-dynamic quantization of moleculargenetic, photosynthesis and ecosystem hierarchical levels of the biosphere. Ecological Modelling, 62(4):259–274, August 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009290002V.

Kirsta:1994:EIN

[Kir94] Yu. B. Kirsta. Exchange of information in natural hierarchical systems. Ecological Modelling, 73(3–4):269–280, June 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094900655.

Kirschbaum:1999:CFG

[Kir99] Miko U. F. Kirschbaum. CenW, a forest growth model with linked carbon, energy, nutrient and water cycles. Ecological Modelling, 118(1):17–59, June 1, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099000204.

Kittlein:1997:AIO

[Kit97] Marcelo J. Kittlein. Assessing the impact of owl predation on the growth rate of a rodent prey population. Ecological Modelling, 103(2–3):123–134, November 17, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097000781.

Kozlowski:1994:DDR

[KJ94] Jan Kozłowski and Mariusz Janczur. Density-dependent regulation of population number and life-history evolution: Optimization of age at maturity in a simple allocation model for annuals and biennials. Ecological Modelling, 73(1–2):81– 96, May 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009490099X.

Kropp:1997:PPR

[KK97] J. Kropp and Th. Klenke. Phenomenological pattern recognition in the dynamical structures of tidal sediments from the German Wadden Sea. Ecological Modelling, 103(2–3):151–170, November 17, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097000835.

Kishi:1991:BBM

[KKNY91] Michio J. Kishi, Shingo Kimura, Hideaki Nakata, and Yoh Yamashita. A biomass-based model for the sand lance (Ammodytes personatus) in Seto Inland Sea, Japan. Ecological Modelling, 54(3–4):247–263, June 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009190078F.

Kokko:1998:SDD

[KL98a] Hanna Kokko and Jan Lindström. Seasonal density dependence, timing of mortality, and sustainable harvesting. Ecological Modelling, 110(3):293–304, July 22, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000891.

Kuuluvainen:1998:EST

[KL98b] Timo Kuuluvainen and Tapio Linkosalo. Estimation of a spatial tree-influence model using iterative optimization. Ecological Modelling, 106(1):63–75, February 16, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001828.

Kolstrom:1999:DSS

[KL99] Marja Kolström and Jaana Lumatjärvi. Decision support system for studying effect of forest management on species richness in boreal forests. Ecological Modelling, 119(1):43–55, July 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099000605.

Klepper:1995:MOF

[Kle95] Olivier Klepper. Modelling the oceanic food web using a quasi steady-state approach. Ecological Modelling, 77(1):33–41, January 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380093E0079I.

Klepper:1997:MAM

[Kle97] Olivier Klepper. Multivariate aspects of model uncertainty analysis: tools for sensitivity analysis and calibration. Ecological Modelling, 101(1):1–13, August 1, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096019229.

Knorrenschild:1996:UDE

[KLFH96] M. Knorrenschild, R. Lenz, E. Forster, and C. Herderich. UFIS: a database of ecological models. Ecological Modelling, 86(2–3):141–144, May 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095000410.

Kimmins:1999:MFE

[KMS99] J. P. Kimmins, Daniel Mailly, and Brad Seely. Modelling forest ecosystem net primary production: the hybrid simulation approach used in forecast. Ecological Modelling, 122(3):195–224, October 20, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001386.

Krysanova:1998:DTS

[KMWB98] Valentina Krysanova, Dirk-Ingmar Müller-Wohlfeil, and Alfred Becker. Development and test of a spatially distributed hydrological/water quality model for mesoscale watersheds. Ecological Modelling, 106(2–3):261–289, March 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097002044.

Kokorin:1995:AGP

[KN95] A. O. Kokorin and I. M. Nazarov. The analysis of growth parameters of Russian boreal forests in warming, and its use in carbon budget model. Ecological Modelling, 82(2):139–150, October 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400086W.

Kalabokidis:1995:IAF

[KO95] Kostas D. Kalabokidis and Philip N. Omi. Isarithmic analysis of forest fire fuelbed arrays. Ecological Modelling, 80(1):47– 55, June 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400045J.

Kohlmaier:1990:LE

[Koh90] G. H. Kohlmaier. Letter to the editor. Ecological Modelling, 50(1–3):221–223, March 1990. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009090052I.

Kohlmaier:1991:EME

[Koh91] G. H. Kohlmaier. Environmental models: Emission and consequences: J. Fenhann, H. Larsen, G. A. Mackenzie and B. Rasmussen (Editors). Developments in Environmental Modelling,

15. Elsevier, Amsterdam, 1990. 489 pp., Dfl. 195.00. ISBN 0-444-88609-5. Ecological Modelling, 59(1–2):146–147, December 1, 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190135N.

Kolstrom:1998:ESM

[Kol98] Marja Kolström. Ecological simulation model for studying diversity of stand structure in boreal forests. Ecological Modelling, 111(1):17–36, August 1, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098001021.

Kompare:1994:WQM

[Kom94] Boris Kompare. Water quality modelling: Roger A. Falconer (Editor). Ashgate and The Institution of Water and Environmental Management, UK. 1992, 139 pp., hard cover. ISBN 1-85742-034-9. Ecological Modelling, 72(1–2):145–149, March 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009490149X.

Kompare:1997:MSH

[Kom97] Boris Kompare. Modelling and simulation: H. Bossel, Modelling and Simulation, A. K. Peters, Wellesley, MA, USA (1994), ISBN 1-56881-033-4 and Verlag Vieweg, Wiesbaden, D (1994), ISBN 3-528-05419-0; pp. 484, hardback, accompanied by a 3.5 in. HD PC format diskette with simpas model building shell and simzoo collection of 50 simulation models. Price: DM98. Ecological Modelling, 98(2–3):245–247, May 30, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096019163.

Kindlmann:1992:PCS

[KR92] Pavel Kindlmann and Zdeněk Ro̊žička. Possible consequences of a specific interaction between predators and parasites of aphids. Ecological Modelling, 61(3–4):253–265, June 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380092900216.

Kremer:1996:SSS

[KR96] Robert G. Kremer and Steven W. Running. Simulating seasonal soil water balance in contrasting semi-arid vegetation communities. Ecological Modelling, 84(1–3):151–162, January 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094001405.

Krapivin:1993:MMG

[Kra93] V. F. Krapivin. Mathematical model for global ecological investigations. Ecological Modelling, 67(2–4):103–127, June 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380093900019.

Krambeck:1995:AAS

[Kra95a] H.-J. Krambeck. Application and abuse of statistical methods in mathematical modelling in limnology. Ecological Modelling, 78(1–2):7–15, March 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400113V.

Krauchi:1995:AMP

[Krä95b] Norbert Kräuchi. Application of the model FORSUM to the Solling spruce site. Ecological Modelling, 83(1–2):219–228, December 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009500100A.

Krapivin:1996:EPC

[Kra96] V. F. Krapivin. The estimation of the Peruvian current ecosystem by a mathematical model of biosphere. Ecological Modelling, 91(1–3):1–14, November 15, 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001557.

Kolev:1996:DSW

[KRD96] Boyko Kolev, Svetla Rousseva, and Dimitar Dimitrov. Derivation of soil water capacity parameters from standard soil texture information for Bulgarian soils. Ecological Modelling, 84(1–3):315–319, January 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001344.

Kremer:1991:SFR

[Kre91] Robert G. Kremer. Simulating forest response to air pollution: Integrating physiological responses to sulphur dioxide with climate-dependent growth processes. Ecological Modelling, 54(1–2):111–126, May 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380091901016.

Knudsen:1991:SRW

[KS91] Guy R. Knudsen and Dennis J. Schotzko. Simulation of Russian wheat aphid movement and population dynamics on preferred and non-preferred host plants. Ecological Modelling, 57(1–2):117–131, October 1, 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380091900578.

Kellomaki:1995:MSG

[KS95] Seppo Kellomäki and Harri Strandman. A model for the structural growth of young Scots pine crowns based on light interception by shoots. Ecological Modelling, 80(2–3):237–250, July 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400065P.

Kulasiri:1996:MHM

[KS96] D. Kulasiri and S. Samarasinghe. Modelling heat and mass transfer in drying of biological materials: a simplified approach to materials with small dimensions. Ecological Modelling, 86(2–3):163–167, May 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095000461.

Kram:1999:AFS

[KSD^{+}99] Pavel Krám, Robert C. Santore, Charles T. Driscoll, John D. Aber, and Jakub Hruška. Application of the forest–soil– water model (PnET-BGC/CHESS) to the Lysina catchment, Czech Republic. Ecological Modelling, 120(1):9–30, August 3, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099000642.

Kesh:1997:STS

[KSR97] Dipak Kesh, A. K. Sarkar, and A. B. Roy. Succession in a three-species food-chain model. Ecological Modelling, 96(1–3):211–219, March 1, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096000683.

Kirsta:1994:RBT

[KT94] Yu. B. Kirsta and V. A. Tarabrin. Real biological time and its calculation in wheat. Ecological Modelling, 71(4):259–267, February 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094901376.

Karafyllidis:1997:MPF

[KT97] Ioannis Karafyllidis and Adonios Thanailakis. A model for predicting forest fire spreading using cellular automata. Ecological Modelling, 99(1):87–97, June 16, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096019424.

Korzukhin:1995:ITB

[KTM95] Michael D. Korzukhin and Michael T. Ter-Mikaelian. An individual tree-based model of competition for light. Ecological Modelling, 79(1–3):221–229, May 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400039K.

Kishi:1994:NSM

[KUI94] Michio J. Kishi, Masato Uchiyama, and Yoshiyasu Iwata. Numerical simulation model for quantitative management of aquaculture. Ecological Modelling, 72(1–2):21–40, March 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094901449.

Krapivin:1990:MEE

[KV90] V. F. Krapivin and L. P. Vilkova. Model estimation of excess CO_{2}, distribution in biosphere structure. Ecological Modelling, 50(1–3):57–78, March 1990. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009090042F.

Kellomaki:1997:MDF

[KV97] Seppo Kellomäki and Hannu Väisänen. Modelling the dynamics of the forest ecosystem for climate change studies in the boreal conditions. Ecological Modelling, 97(1–2):121–140, April 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096000816.

Kumar:1991:PMF

[KVAW91] S. Kishore Kumar, Warwick F. Vincent, Paul C. Austin, and Graeme C. Wake. Picoplankton and marine food chain dynamics in a variable mixed-layer: a reaction-diffusion model. Ecological Modelling, 57(3–4):193–219, October 15, 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009190113F.

Kros:1995:EMB

[KW95] Hans Kros and Per Warfvinge. Evaluation of model behaviour with respect to the biogeochemistry at the Solling spruce site. Ecological Modelling, 83(1–2):255–262, December 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095001033.

Kellomaki:1999:STE

[KW99] Seppo Kellomäki and Kai-Yun Wang. Short-term environmental controls of heat and water vapour fluxes above a boreal coniferous forest: model computations compared with measurements by eddy correlation. Ecological Modelling, 124(2–3):145–173, December 13, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099001593.

Lee:1991:SMS

[LA91] C. S. Lee and P. Ang. A simple model for seaweed growth and optimal harvesting strategy. Ecological Modelling, 55(1–2):67– 74, July 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380091900659.

Li:1996:ECD

[LA96] Chao Li and Michael J. Apps. Effects of contagious disturbance on forest temporal dynamics. Ecological Modelling, 87(1–3):143–151, June 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095000232.

Li:1997:WMP

[LA97] Bai-Lian Li and Steve Archer. Weighted mean patch size: a robust index for quantifying landscape structure. Ecological Modelling, 102(2–3):353–361, October 29, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097000719.

Liu:1998:FIB

[LA98] Jianguo Liu and Peter S. Ashton. FORMOSAIC: an individualbased spatially explicit model for simulating forest dynamics in landscape mosaics. Ecological Modelling, 106(2–3):177–200, March 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001919.

Lathrop:1995:SVD

[LAB95] Richard G. Lathrop, John D. Aber, and John A. Bognar. Spatial variability of digital soil maps and its impact on regional ecosystem modeling. Ecological Modelling, 82(1):1–10, September 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400068S.

Legaspi:1998:AMC

[LAB^{+}98] B. C. Legaspi, Jr., J. C. Allen, C. C. Brewster, J. A. Morales-Ramos, and E. G. King. Areawide management of the cotton boll weevil: use of a spatio–temporal model in augmentative biological control. Ecological Modelling, 110(2):151–164, July 11, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438009800060X.

Lek-Ang:1999:PMC

[LADL99] Sithan Lek-Ang, Louis Deharveng, and Sovan Lek. Predictive models of collembolan diversity and abundance in a riparian habitat. Ecological Modelling, 120(2–3):247–260, August 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001064.

Laval:1995:HOO

[Lav95] Ph. Laval. Hierarchical object-oriented design of a concurrent, individual-based, model of a pelagic tunicate bloom. Ecological Modelling, 82(3):265–276, November 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009400092V.

Laval:1996:RSO

[Lav96] Ph. Laval. The representation of space in an object-oriented computational pelagic ecosystem. Ecological Modelling, 88 (1–3):113–124, July 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095000747.

Lin:1997:FVV

[LB97a] Ching-Rong Lin and Joseph Buongiorno. Fixed versus variableparameter matrix models of forest growth: the case of maple– birch forests. Ecological Modelling, 99(2–3):263–274, June 30, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097019601.

Luk:1997:APB

[LB97b] G. K. Luk and F. Brockway. Application of a polychlorinated biphenyls bioaccumulation model to Lake Ontario lake trout. Ecological Modelling, 101(1):97–111, August 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438009701956X.

Lindenmayer:1995:RGC

[LBA^{+}95] D. B. Lindenmayer, M. A. Burgman, H. R. Akçakaya, R. C. Lacy, and H. P. Possingham. A review of the generic computer programs ALEX, RAMAS/space and VORTEX for modelling the viability of wildlife metapopulations. Ecological Modelling, 82(2):161–174, October 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400085V.

Lin:1996:MSD

[LBV96] Ching Rong Lin, Joseph Buongiorno, and Mike Vasievich. A multi-species, density-dependent matrix growth model to predict tree diversity and income in northern hardwood stands. Ecological Modelling, 91(1–3):193–211, November 15, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001905.

Lynn:1991:STP

[LC91] Barry H. Lynn and Toby N. Carlson. Simulating transpiration plateaus: the importance of leaf water potential. Ecological Modelling, 58(1–4):199–208, November 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009190036Z.

Legovic:1997:MPG

[LC97a] T. Legović and A. Cruzado. A model of phytoplankton growth on multiple nutrients based on the Michaelis– Menten-Monod uptake, Droop’s growth and Liebig’s law. Ecological Modelling, 99(1):19–31, June 16, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096019199.

Lemmon:1997:OOD

[LC97b] Hal Lemmon and Ning Chuk. Object-oriented design of a cotton crop model. Ecological Modelling, 94(1):45–51, January 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096019278.

Luckyanov:1995:SAE

[LCH95] N. K. Luckyanov, W. P. Cropper, and M. A. Harwell. State analysis of ecological models: model reactions to parameter change. Ecological Modelling, 82(1):99–104, September 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400077U.

Li:1993:EPC

[LCL93] Bai-Lian Li, Yi-Te Chu, and Douglas K. Loh. Event probability correlation analysis for comparison of two-phase ecological maps. Ecological Modelling, 69(3–4):287–302, October 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009390031M.

Legaspi:1996:FRC

[LCMR96] B. C. Legaspi, R. I. Carruthers, and J. A. Morales-Ramos. Functional response as a component of dynamic simulation models in biological control: the Catolaccus-boll weevil system. Ecological Modelling, 89(1–3):43–57, August 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001190.

Lee:1997:SSS

[LD97] Hooi Ling Lee and Donald L. DeAngelis. A simulation study of the spatio-temporal dynamics of the unionid mussels. Ecological Modelling, 95(2–3):171–180, February 14, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096000397.

Lek:1996:ANN

[LDB^{+}96] Sovan Lek, Marc Delacoste, Philippe Baran, Ioannis Dimopoulos, Jacques Lauga, and Stéphane Aulagnier. Application of neural networks to modelling nonlinear relationships in ecology. Ecological Modelling, 90(1):39–52, September 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095001425.

Leary:1997:TMU

[Lea97] Rolfe A. Leary. Testing models of unthinned red pine plantation dynamics using a modified bakuzis matrix of stand properties. Ecological Modelling, 98(1):35–46, May 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096019357.

Lee:1991:SCM

[Lee91] Danny C. Lee. A stochastic, compartmental model of the migration of juvenile anadromous salmonids in the Columbia River Basin. Ecological Modelling, 54(3–4):227–245, June 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009190077E.

Legovic:1991:CFW

[Leg91a] Tarzan Legovic. Community food webs: J. E. Cohen, F. Briand and C. M. Newman. Biomathematics, 20. Springer, Berlin, 1990. 308 pp., DM148.00. ISBN 3-540-51129-6. Ecological Modelling, 59(3–4):294–296, December 15, 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380091901843.

Legovic:1991:SEP

[Leg91b] Tarzan Legovic. Statistical ecology. A primer on methods and computing: J. A. Ludwig and J. F. Reynolds. Wiley-Interscience Publication, New York, 1988. 337 pp. + MS-DOS diskette of 21 Basic Programs, paperback, £16.75. ISBN 0-471-61315-0. Ecological Modelling, 54(1–2):143–144, May 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009190106B.

Legovic:1992:TSE

[Leg92] T. Legović. Theoretical studies of ecosystems: the network perspective: M. Higashi and T. P. Burns (Editors). Cambridge University Press, New York, 1991, 364 pp. U. S. $74.00. ISBN 521-36138-9. Ecological Modelling, 64(1):79–80, October 15, 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290053H.

Legovic:1993:DNC

[Leg93] Tarzan Legovic. Dynamics of nutrient cycling and food webs: D. L. DeAngelis. Chapman and Hall, London, 1992, £17.95, XV + 270 pp., ISBN 0-412-29840-6 paper cover, ISBN 0-412-29830-9 hard cover. Ecological Modelling, 65(3–4):299–301, February 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380093900879.

Legovic:1997:EMI

[Leg97a] Tarzan Legović. Ecological modelling Internet resources. Ecological Modelling, 100(1–3):163–169, December 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009700152X.

Legovic:1997:TMA

[Leg97b] Tarzan Legović. Toxicity may affect predictability of eutrophication models in the coastal sea. Ecological Modelling, 99(1):1– 6, June 16, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096019394.

Luo:1997:AGG

[LFM97] Y. Luo, C. B. Field, and H. A. Mooney. Adapting GePSi (Generic Plant Simulator) for modeling studies in the Jasper Ridge CO_{2} project. Ecological Modelling, 94(1):81–88, January 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096019308.

Lewis:1991:PCC

[LG91] David L. Lewis and David K. Gattie. Predicting chemical concentration effects on transformation rates of dissolved organics by complex microbial assemblages. Ecological Modelling, 55(1–2):27–46, July 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380091900626.

Lek:1999:ANN

[LG99] Sovan Lek and J. F. Guégan. Artificial neural networks as a tool in ecological modelling, an introduction. Ecological Modelling, 120(2–3):65–73, August 17, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000927.

Latore:1999:EHH

[LGM99] J. Latore, P. Gould, and A. M. Mortimer. Effects of habitat heterogeneity and dispersal strategies on population persistence in annual plants. Ecological Modelling, 123(2–3):127–139, November 15, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099001325.

Loza:1992:PBL

[LGSF92] H. J. Loza, W. E. Grant, J. W. Stuth, and T. D. A. Forbes. Physiologically based landscape use model for large herbivores. Ecological Modelling, 61(3–4):227–252, June 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290020F.

Lenz:1992:ARF

[LH92] R. Lenz and W. Haber. Approaches for the restoration of forest ecosystems in northeastern Bavaria. Ecological Modelling, 63(1–4):299–317, September 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290074O.

Lindstrom:1999:EMP

[LHAJ99] Martin Lindström, Lars Håkanson, Otto Abrahamsson, and Håkan Johansson. An empirical model for prediction of lake water suspended particulate matter. Ecological Modelling, 121(2–3):185–198, September 15, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099000812.

Lhotka:1994:IIO

[Lho94] Ladislav Lhotka. Implementation of individual-oriented models in aquatic ecology. Ecological Modelling, 74(1–2):47–62, July 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094901104.

Li:1995:SAN

[Li95] Bai-Lian Li. Stability analysis of a nonhomogeneous Markovian landscape model. Ecological Modelling, 82(3):247–256, November 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400091U.

Lindgren:1998:CCT

[Lin98] Elisabet Lindgren. Climate change, tick-borne encephalitis and vaccination needs in Sweden — a prediction model.

Ecological Modelling, 110(1):55–63, July 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000416.

Liu:1993:DIP

[Liu93a] Jianguo Liu. Discounting initial population sizes for prediction of extinction probabilities in patchy environments. Ecological Modelling, 70(1–2):51–61, November 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390072Z.

Liu:1993:EEE

[Liu93b] Jianguo Liu. ECOLECON: an ECOLogical-ECONomic model for species conservation in complex forest landscapes. Ecological Modelling, 70(1–2):63–87, November 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093900732.

Liu:1996:IDL

[Liu96] De Li Liu. Incorporating diurnal light variation and canopy light attenuation into analytical equations for calculating daily gross photosynthesis. Ecological Modelling, 93(1–3): 175–189, December 16, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095002235.

Liu:1997:NMP

[Liu97] Shuguang Liu. A new model for the prediction of rainfall interception in forest canopies. Ecological Modelling, 99 (2–3):151–159, June 30, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097019480.

Loehle:1994:FMM

[LJ94] Craig Loehle and Paul Johnson. A framework for modeling microbial transport and dynamics in the subsurface. Ecological Modelling, 73(1–2):31–49, May 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094900965.

Lhomme:1991:SMC

[LK91] Jean-Paul Lhomme and Nader Katerji. A simple modelling of crop water balance for agrometeorological applications. Ecological Modelling, 57(1–2):11–25, October 1, 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380091900523.

Link:1999:MSP

[LK99] Jason S. Link and Robert E. Keen. A model of salmonid planktivory: field test of a mechanistic approach to sizeselection. Ecological Modelling, 117(2–3):269–283, May 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099000356.

Levine:1996:CSS

[LKS96] E. R. Levine, D. S. Kimes, and V. G. Sigillito. Classifying soil structure using neural networks. Ecological Modelling, 92(1):101–108, November 21, 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001999.

Loehle:1996:MBA

[LL96a] Craig Loehle and David LeBlanc. Model-based assessments of climate change effects on forests: a critical review. Ecological Modelling, 90(1):1–31, September 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380096837094.

Loehle:1996:SPE

[LL96b] Craig Loehle and Bai-Lian Li. Statistical properties of ecological and geologic fractals. Ecological Modelling, 85(2–3):271– 284, March 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094001774.

Lung:1996:MCR

[LL96c] Wu-Seng Lung and Ronald N. Light. Modelling copper removal in wetland ecosystems. Ecological Modelling, 93(1–3):89–100, December 16, 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095002162.

Logofet:1997:WMP

[LL97] Dmitrii O. Logofet and Ekaterina V. Lesnaya. Why are the middle points the most ‘sensitive’ in the sensitivity experiments? Ecological Modelling, 104(2–3):303–306, December 8, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097001257.

Lischke:1997:CTD

[LLF97] Heike Lischke, Thomas J. Löffler, and Andreas Fischlin. Calculating temperature dependence over long time periods: derivation of methods. Ecological Modelling, 98(2–3):105–122, May 30, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096019072.

Li:1996:MTT

[LLM96] Bai-Lian Li, Craig Loehle, and David Malon. Microbial transport through heterogeneous porous media: random walk, fractal, and percolation approaches. Ecological Modelling, 85(2–3):285–302, March 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094001987.

Lae:1999:PFY

[LLM99] Raymond Laë, Sovan Lek, and Jacques Moreau. Predicting fish yield of African lakes using neural networks. Ecological Modelling, 120(2–3):325–335, August 17, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009900112X.

Liu:1995:UMD

[LLWP95] Duo-Sen Liu, Zhen-Gao Li, Zong-Sheng Wang, and Ying-Hua Pan. An unified model for the different phases of growth of microbial populations in a closed system. Ecological Modelling, 82(2):193–198, October 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400089Z.

Luan:1996:HAF

[LMG96] Jingsheng Luan, Robert I. Muetzelfeldt, and John Grace. Hierarchical approach to forest ecosystem simulation. Ecological Modelling, 86(1):37–50, April 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001212.

Loehle:1997:HTF

[Loe97a] Craig Loehle. A hypothesis testing framework for evaluating ecosystem model performance. Ecological Modelling, 97(3):153–165, April 28, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438009601900X.

Loehle:1997:PTA

[Loe97b] Craig Loehle. The pathogen transmission avoidance theory of sexual selection. Ecological Modelling, 103(2–3):231– 250, November 17, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380097001063.

Lofgren:1996:BEC

[Löf96] Karl-Gustaf Löfgren. A back of the envelope calculation method for calculations of the gains from genetic progress in forestry with some theoretical underpinning. Ecological Modelling, 92(2–3):245–252, December 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001743.

Logofet:1997:SCM

[Log97] D. O. Logofet. Svicobians of the compartment models and DaD-stability of the Svicobians: aggregating ‘0dimensional’ models of global biogeochemical cycles. Ecological Modelling, 104(1):39–49, December 1, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001075.

Lomnicki:1999:IBM

[Łom99] Adam Łomnicki. Individual-based models and the individualbased approach to population ecology. Ecological Modelling, 115(2–3):191–198, February 15, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098001926.

Losordo:1991:MTV

[LP91] Thomas M. Losordo and Raul H. Piedrahita. Modelling temperature variation and thermal stratification in shallow aquaculture ponds. Ecological Modelling, 54(3–4):189–226, June 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009190076D.

Li:1996:I

[LR96] Bai-Lian Li and Edward J. Rykiel. Introduction. Ecological Modelling, 90(2):109–110, October 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380096900721.

Letcher:1997:PPL

[LR97] Benjamin H. Letcher and James A. Rice. Prey patchiness and larval fish growth and survival: inferences from an individualbased model. Ecological Modelling, 95(1):29–43, February 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096000154.

Liu:1998:SEF

[LRG98] Shuguang Liu, Hans Riekerk, and Henry L. Gholz. Simulation of evapotranspiration from Florida pine flatwoods. Ecological Modelling, 114(1):19–34, December 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001033.

Ludeke:1996:USN

[LRK96] Matthias K. B. Lüdeke, Peter H. Ramage, and G. H. Kohlmaier. The use of satellite NDVI data for the validation of global vegetation phenology models: application to the Frankfurt Biosphere model. Ecological Modelling, 91(1–3):255–270, November 15, 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001921.

Li:1993:PRT

[LRM93] C. Li, B. D. Roitberg, and M. Mackauer. Patch residence time and parasitism of Aphelinus asychis: a simulation model. Ecological Modelling, 69(3–4):227–241, October 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009390028Q.

Levine:1993:FED

[LRS^{+}93] E. R. Levine, K. J. Ranson, J. A. Smith, D. L. Williams, R. G. Knox, H. H. Shugart, D. L. Urban, and W. T. Lawrence. Forest ecosystem dynamics: linking forest succession, soil process and radiation models. Ecological Modelling, 65(3–4):199–219, February 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390080C.

Lappi:1998:JEA

[LS98a] Juha Lappi and Pauline Stenberg. Joint effect of angular distribution of radiation and spatial pattern of trees on radiation interception. Ecological Modelling, 112(1):45–51, October 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098001124.

Lorek:1998:OOS

[LS98b] Helmut Lorek and Michael Sonnenschein. Object-oriented support for modelling and simulation of individual-oriented ecological models. Ecological Modelling, 108(1–3):77–96, May 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098000209.

Lorek:1999:MSS

[LS99] Helmut Lorek and Michael Sonnenschein. Modelling and simulation software to support individual-based ecological modelling. Ecological Modelling, 115(2–3):199–216, February 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098001938.

Littleboy:1996:SMD

[LSB96] M. Littleboy, D. M. Smith, and M. J. Bryant. Simulation modelling to determine suitability of agricultural land. Ecological Modelling, 86(2–3):219–225, May 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095000550.

Lett:1999:CCA

[LSB99] Christophe Lett, Catherine Silber, and Nicolas Barret. Comparison of a cellular automata network and an individualbased model for the simulation of forest dynamics. Ecological Modelling, 121(2–3):277–293, September 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000903.

Larkin:1995:SDM

[LSC95] Timothy S. Larkin, Anthony W. Sweeney, and Raymond I. Carruthers. Simulation of the dynamics of a microsporidian pathogen of mosquitoes. Ecological Modelling, 77(2–3): 143–165, February 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380093E0076F.

Lehn:1996:FCS

[LT96] Karsten Lehn and Karl-Heinz Temme. Fuzzy classification of sites suspected of being contaminated. Ecological Modelling, 85(1):51–58, February 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095000143.

Li:1997:TFD

[LTMP97] Chao Li, Michael Ter-Mikaelian, and Ajith Perera. Temporal fire disturbance patterns on a forest landscape. Ecological Modelling, 99(2–3):137–150, June 30, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096019448.

Lauenroth:1993:MVS

[LUC^{+}93a] W. K. Lauenroth, D. L. Urban, D. P. Coffin, W. J. Parton, H. H. Shugart, T. B. Kirchner, and T. M. Smith. Modeling vegetation structure-ecosystem process interactions across sites and ecosystems. Ecological Modelling, 67(1):49–80, May 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009390099E.

Luckyanov:1993:AEM

[Luc93b] N. K. Luckyanov. On adequacy of ecological models. Ecological Modelling, 68(1–2):51–65, July 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380093901074.

Lokshina:1999:KAK

[LV99] L. Ya. Lokshina and V. A. Vavilin. Kinetic analysis of the key stages of low temperature methanogenesis. Ecological Modelling, 117(2–3):285–303, May 17, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000083.

Li:1996:MND

[LVH96] Jian Li, Magda Vincx, and Peter M. J. Herman. A model of nematode dynamics in the Westerschelde Estuary. Ecological Modelling, 90(3):271–284, November 1, 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001549.

Liden:1999:NSA

[LVS^{+}99] R. Lidén, A. Vasilyev, P. Stålnacke, E. Loigu, and H. B. Wittgren. Nitrogen source apportionment — a comparison between a dynamic and a statistical model. Ecological Modelling, 114(2–3):235–250, January 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009800146X.

Loehle:1994:LHD

[LW94] Craig Loehle and Gary Wein. Landscape habitat diversity: a multiscale information theory approach. Ecological Modelling, 73(3–4):311–329, June 1994. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009490068X.

Levine:1996:MRH

[LW96] Stephen H. Levine and Richard E. Wetzler. Modelling the role of host plant dispersion in the search success of herbivorous insects: Implications for ecological pest management. Ecological Modelling, 89(1–3):183–196, August 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009500131X.

Li:1997:RMS

[LWMT97] Wenjun Li, Zijian Wang, Zhijun Ma, and Hongxiao Tang. A regression model for the spatial distribution of red-crown crane in Yancheng Biosphere Reserve, China. Ecological Modelling, 103(2–3):115–121, November 17, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009700077X.

Lyon:1996:DRS

[Lyo96a] Kenneth S. Lyon. Discount rates and sustainable development: comment. Ecological Modelling, 92(2–3):273–274, December 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095001816.

Lyon:1996:WED

[Lyo96b] Kenneth S. Lyon. Why economists discount future benefits. Ecological Modelling, 92(2–3):253–262, December 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001735.

MaBlanco:1993:RBL

[M.a93a] Jose M.a Blanco. Relationship between the logistic equation and the Lotka–Volterra models. Ecological Modelling, 66(3–4):301–303, April 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438009390120H.

Mossman:1993:AFH

[MA93b] Deborah J. Mossman and Ronald B. Achelpohl. Added flexibility for Holly-preissmann advection operator. Ecological Modelling, 69(3–4):311–317, October 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390033O.

Malanson:1996:DPF

[MA96a] George P. Malanson and Marc P. Armstrong. Dispersal probability and forest diversity in a fragmented landscape. Ecological Modelling, 87(1–3):91–102, June 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094002029.

Mossman:1996:ODU

[MA96b] Deborah J. Mossman and Nael Al Mulki. One-dimensional unsteady flow and unsteady pesticide transport in a reservoir. Ecological Modelling, 89(1–3):259–267, August 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001468.

MaRey:1999:MPE

[M.a99] José M.a Rey. Modelling potential evapotranspiration of potential vegetation. Ecological Modelling, 123(2–3):141– 159, November 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380099001295.

Mahamah:1998:SAW

[Mah98a] D. S. Mahamah. Simplifying assumptions in water quality modeling. Ecological Modelling, 109(3):295–300, June 22, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098000635.

Mahmood:1998:ATV

[Mah98b] Rezaul Mahmood. Air temperature variations and rice productivity in Bangladesh: a comparative study of the performance of the YIELD and the CERES-Rice models. Ecological Modelling, 106(2–3):201–212, March 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001920.

Malchow:1995:FLI

[Mal95] H. Malchow. Flow- and locomotion-induced pattern formation in nonlinear population dynamics. Ecological Modelling, 82(3):257–264, November 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438009400095Y.

Malanson:1996:EDM

[Mal96] George P. Malanson. Effects of dispersal and mortality on diversity in a forest stand model. Ecological Modelling, 87(1–3):103–110, June 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001441.

Marino:1991:IVD

[Mar91] Paul C. Marino. The influence of varying degress of spore aggregation on the coexistence of the mosses Splachnum ampullacuem and S. luteum: a simulation study. Ecological Modelling, 58(1–4):333–345, November 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380091900442.

Markus:1992:ODM

[Mar92] Mario Markus. Are one-dimensional maps of any use in ecology? Ecological Modelling, 63(1–4):243–259, September 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009290071L.

Marani:1993:RSH

[Mar93] A. Marani. Remote sensing in hydrology: E. T. Engman and R. J. Gurney. Chapman & Hall, London, 1991, 225, pp., £38.00. Ecological Modelling, 65(3–4):304–305, February 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009390090F.

Marin:1997:SBS

[Mar97a] Victor H. Marín. A simple-biology, stage-structured population model of the spring dynamics of Calanus chilensis at Mejillones del Sur Bay, Chile. Ecological Modelling, 105(1):65–82, December 14, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001397.

Martin:1997:UCS

[Mar97b] Paul-Gerhard Martin. The use of canonical S-system modelling for condensation of complex dynamic models. Ecological Modelling, 103(1):43–70, November 1, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097000756.

Martin:1998:PME

[Mar98] Stephen Martin. A population model for the ectoparasitic mite Varroa jacobsoni in honey bee (Apis mellifera) colonies. Ecological Modelling, 109(3):267–281, June 22, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000593.

Marques:1999:IET

[Mar99a] João Carlos Marques. Integration of Ecosystem Theories: a Pattern, by Sven Erik Jørgensen, Kluwer Academic Publishers, Dordrecht, Boston, London, Second Revised Edition, 1997. 388 pp. ISBN Hb: 0-7923-4523-1; Pb: 0-7923-4524-X. Ecological Modelling, 117(1):165–167, April 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000241.

Martin:1999:EBS

[Mar99b] Philippe H. Martin. Exchanges between structured canopies and their physical environment: a simple analytical solution for a generic configuration. Ecological Modelling, 122(1–2):1–24, October 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001209.

Matter:1999:PDA

[Mat99] Stephen F. Matter. Population density and area: the role of within- and between-generation processes over time. Ecological Modelling, 118(2–3):261–275, June 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000514.

Maurer:1990:DPE

[Mau90] Brian A. Maurer. Dipodomys populations as energy-processing systems: Regulation, competition, and hierarchical organization. Ecological Modelling, 50(1–3):157–176, March 1990. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009090048L.

Mauersberger:1993:ILE

[Mau93] Peter Mauersberger. Intertidal and littoral ecosystems: A. C. Mathieson and P. H. Nienhuis (Editors). Ecosystems of the World 24, Elsevier, Amsterdam, 1991, XIII + 564 pp., US $195.00/Dfl 380.00, ISBN 0-444-87409-7 (Vol. 24). Ecological Modelling, 65(3–4):306–307, February 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093900916.

Maunder:1997:IDD

[Mau97] M. N. Maunder. Investigation of density dependence in salmon spawner–egg relationships using queueing theory. Ecological Modelling, 104(2–3):189–197, December 8, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001269.

Menshutkin:1998:MME

[MAY^{+}98] V. V. Menshutkin, G. P. Astrakhantsev, N. B. Yegorova, L. A. Rukhovets, T. L. Simo, and N. A. Petrova. Mathematical modeling of the evolution and current conditions of the Ladoga Lake ecosystem. Ecological Modelling, 107(1):1–24, March 16, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001841.

Mayer:1993:SV

[MB93] D. G. Mayer and D. G. Butler. Statistical validation. Ecological Modelling, 68(1–2):21–32, July 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380093901052.

Mooij:1996:OOS

[MB96] Wolf M. Mooij and Maarten Boersma. An object-oriented simulation framework for individual-based simulations (OSIRIS): Daphnia population dynamics as an example. Ecological Modelling, 93(1–3):139–153, December 16, 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095002200.

McCarthy:1996:LSB

[MBF96] M. A. McCarthy, M. A. Burgman, and S. Ferson. Logistic sensitivity and bounds for extinction risks. Ecological Modelling, 86(2–3):297–303, May 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095000674.

Maley:1993:ISC

[MC93a] Carlo C. Maley and Hal Caswell. Implementing i-state configuration models for population dynamics: an object-oriented programming approach. Ecological Modelling, 68(1–2):75–89, July 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380093901096.

Maxwell:1993:AMD

[MC93b] Thomas Maxwell and Robert Costanza. An approach to modelling the dynamics of evolutionary self-organization. Ecological Modelling, 69(1–2):149–161, September 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390054V.

Maxwell:1997:LMS

[MC97] Tom Maxwell and Robert Costanza. A language for modular spatio-temporal simulation. Ecological Modelling, 103(2–3):105–113, November 17, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097001038.

McClanahan:1992:RUC

[McC92] T. R. McClanahan. Resource utilization, competition, and predation: a model and example from coral reef grazers. Ecological Modelling, 61(3–4):195–215, June 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290018A.

McClanahan:1995:CRE

[McC95a] T. R. McClanahan. A coral reef ecosystem-fisheries model: impacts of fishing intensity and catch selection on reef structure and processes. Ecological Modelling, 80(1):1–19, June 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400042G.

McClanahan:1995:HUW

[McC95b] T. R. McClanahan. Harvesting in an uncertain world: impact of resource competition on harvesting dynamics. Ecological Modelling, 80(1):21–26, June 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400035G.

McCarthy:1996:EDH

[McC96] Michael A. McCarthy. Extinction dynamics of the helmeted honeyeater: effects of demography, stochasticity, inbreeding and spatial structure. Ecological Modelling, 85(2–3):151–163, March 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400191X.

McCarthy:1997:AEF

[McC97] M. A. McCarthy. The Allee effect, finding mates and theoretical models. Ecological Modelling, 103(1):99–102, November 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438009700104X.

McCarthy:1999:ECN

[McC99] M. A. McCarthy. Effects of competition on natal dispersal distance. Ecological Modelling, 114(2–3):305–310, January 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098001628.

McDonnell:1993:WPM

[McD93] R. McDonnell. Water pollution: Modelling, measuring and prediction: L. C. Wrobel and C. A. Brebbia (Editors). Computational Mechanics Publications, Elsevier Applied Science, 1991, 762 pp., £99.00, ISBN 1-85312-146-0. Ecological Modelling, 65(3–4):302–303, February 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390089B.

Masson:1999:SSD

[MCGLJ99] Marie H. Masson, Stéphane Canu, Yves Grandvalet, and Anders Lynggaard-Jensen. Software sensor design based on empirical data. Ecological Modelling, 120(2–3):131–139, August 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099000976.

McKelvey:1996:VAE

[McK96] Robert McKelvey. Viability analysis of endangered species: a decision-theoretic perspective. Ecological Modelling, 92 (2–3):193–207, December 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095001751.

Mukhopadhyay:1997:SHT

[MCT97] A. Mukhopadhyay, J. Chattopadhyay, and P. K. Tapaswi. Selective harvesting in a two species fishery model. Ecological Modelling, 94(2–3):243–253, January 15, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096000191.

Mesple:1995:SDM

[MCTB95] Fabrice Mesple, Claude Casellas, Marc Troussellier, and Jean Bontoux. Some difficulties in modelling chlorophyll a evolution in a high rate algal pond ecosystem. Ecological Modelling, 78(1–2):25–36, March 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400115X.

Mesple:1996:MOE

[MCTB96] Fabrice Mesplé, Claude Casellas, Marc Troussellier, and Jean Bontoux. Modelling orthophosphate evolution in a high rate algal pond. Ecological Modelling, 89(1–3):13–21, August 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095001158.

Mageau:1998:QTE

[MCU98] Michael T. Mageau, Robert Costanza, and Robert E. Ulanowicz. Quantifying the trends expected in developing ecosystems. Ecological Modelling, 112(1):1–22, October 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000921.

Manel:1997:LRC

[MD97] Stéphanie Manel and Domitien Debouzie. Logistic regression and continuation ratio models to estimate insect development under variable temperatures. Ecological Modelling, 98(2–3):237–243, May 30, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096019084.

Maier:1998:UAN

[MDB98] Holger R. Maier, Graeme C. Dandy, and Michael D. Burch. Use of artificial neural networks for modelling cyanobacteria Anabaena spp. in the River Murray, South Australia. Ecological Modelling, 105(2–3):257–272, January 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001610.

Myerscough:1996:SPS

[MDH96] M. R. Myerscough, M. J. Darwen, and W. L. Hogarth. Stability, persistence and structural stability in a classical predator– prey model. Ecological Modelling, 89(1–3):31–42, August 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095001174.

Manel:1999:CDA

[MDO99] Stéphanie Manel, Jean-Marie Dias, and Steve J. Ormerod. Comparing discriminant analysis, neural networks and logistic regression for predicting species distributions: a case study with a Himalayan river bird. Ecological Modelling, 120(2–3):337–347, August 17, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099001131.

Menard:1993:SMO

[MDT93] F. Ménard, S. Dallot, and G. Thomas. A stochastic model for ordered categorical time series. Application to planktonic abundance data. Ecological Modelling, 66(1–2):101–112, March 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009390041P.

Momen:1996:AMS

[MEBZ96] B. Momen, L. W. Eichler, C. W. Boylen, and J. P. Zehr. Application of multivariate statistics in detecting temporal and spatial patterns of water chemistry in Lake George, New York. Ecological Modelling, 91(1–3):183–192, November 15, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095001891.

Messer:1992:SMT

[Mes92] J. Messer. Statistical mechanics of terrestrial ecosystems. Ecological Modelling, 63(1–4):319–323, September 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290075P.

Metternicht:1998:FCJ

[Met98] G. I. Metternicht. Fuzzy classification of JERS-1 SAR data: an evaluation of its performance for soil salinity mapping. Ecological Modelling, 111(1):61–74, August 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000957.

Moatar:1999:PMN

[MFP99] Florentina Moatar, Françoise Fessant, and Alain Poirel. pH modelling by neural networks. Application of control and validation data series in the Middle Loire River. Ecological Modelling, 120(2–3):141–156, August 17, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000988.

Mesterton-Gibbons:1996:TFO

[MG96a] Michael Mesterton-Gibbons. A technique for finding optimal two-species harvesting policies. Ecological Modelling, 92(2–3):235–244, December 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009500176X.

Mills:1996:MBC

[MG96b] N. J. Mills and W. M. Getz. Modelling the biological control of insect pests: a review of host-parasitoid models. Ecological Modelling, 92(2–3):121–143, December 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001778.

Morand:1997:PMI

[MG97] Serge Morand and Ernesto Arias Gonzalez. Is parasitism a missing ingredient in model ecosystems? Ecological Modelling, 95(1):61–74, February 1, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096000282.

Martinez-Garmendia:1998:SAE

[MG98] Josué Martínez-Garmendia. Simulation analysis of evolutionary response of fish populations to size-selective harvesting with the use of an individual-based model. Ecological Modelling, 111(1):37–60, August 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098000933.

Marcum:1998:SBE

[MGCR98] H. A. Marcum, W. E. Grant, and F. Chavez-Ramirez. Simulated behavioral energetics of nonbreeding American robins: the influence of foraging time, intake rate and flying time on weight dynamics. Ecological Modelling, 106(2–3):161–175, March 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001907.

Marin:1998:SPD

[MGD98] S. L. Marín, W. E. Grant, and N. O. Dronen. Simulation of population dynamics of the parasite Haematoloechus coloradensis in its three host species: effects of environmental temperature and precipitation. Ecological Modelling, 105(2–3):185–211, January 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001609.

Moreno-Grau:1996:MMW

[MGGSMC^{+}96] S. Moreno-Grau, A. García-Sánchez, J. Moreno-Clavel, J. Serrano-Aniorte, and M. D. Moreno-Grau. A mathematical model for waste water stabilization ponds with macrophytes and microphytes. Ecological Modelling, 91(1–3):77– 103, November 15, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095001689.

Matis:1992:SMP

[MGM92] J. H. Matis, W. E. Grant, and T. H. Miller. A semi-Markov process model for migration of marine shrimp. Ecological Modelling, 60(3–4):167–184, April 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290032A.

Masser:1991:SMR

[MGNR91] Michael P. Masser, W. E. Grant, W. H. Neill, and E. H. Robinson. A simulation model representing effects on dietary energy/ protein ratio and water temperature on growth of channel catfish (Ictalurus punctatus). Ecological Modelling, 54(1–2):17– 35, May 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190096J.

McKelvey:1996:I

[MH96] Robert McKelvey and Vincent Hull. Introduction. Ecological Modelling, 92(2–3):109, December 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096901064.

Madden:1995:RUH

[MHA95] A. D. Madden, J. Holt, and N. J. Armes. The role of uncultivated hosts in the spread of pyrethroid resistance in Helicoverpa armigera populations in Andhra Pradesh, India: a simulation approach. Ecological Modelling, 82(1):61–74, September 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400071O.

Monte:1996:UAV

[MHB^{+}96] Luigi Monte, Lars Håkanson, Ulla Bergström, John Brittain, and Rudie Heling. Uncertainty analysis and validation of environmental models: the empirically based uncertainty analysis. Ecological Modelling, 91(1–3):139–152, November 15, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095001859.

Maddison:1996:SDM

[MHJ96] A. C. Maddison, J. Holt, and M. J. Jeger. Spatial dynamics of a monocyclic disease in a perennial crop. Ecological Modelling, 88(1–3):45–52, July 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095000682.

Mohren:1995:MES

[MI95] Godefridus M. J. Mohren and Hannu Ilvesniemi. Modelling effects of soil acidification on tree growth and nutrient status. Ecological Modelling, 83(1–2):263–272, December 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095001044.

Miller:1992:SAT

[Mil92] Norman L. Miller. Stability analysis of toxic substances within aquatic ecosystems and their effect on aquatic populations. Ecological Modelling, 60(2):151–165, March 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290044F.

Mitsch:1994:IET

[Mit94] William J. Mitsch. Integration of ecosystem theories: a Pattern. S. E. Jørgensen. 1992. Kluwer Academic Publishers, Dordrecht, the Netherlands, 383 pp., US$165.00, UK£87.50, ISBN 0-7923-1985-0. Ecological Modelling, 74(3–4):316–317, August 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094901309.

Medlyn:1999:DUD

[MJ99] Belinda E. Medlyn and Paul G. Jarvis. Design and use of a database of model parameters from elevated [CO_{2}] experiments. Ecological Modelling, 124(1):69–83, December 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099001489.

Miller:1994:SMH

[MJGW94] M. W. Miller, K. C. Jensen, W. E. Grant, and M. W. Weller. A simulation model of helicopter disturbance of molting Pacific black brant. Ecological Modelling, 73(3–4):293–309, June 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094900671.

Mueller:1997:TVC

[MJMN97] T. Mueller, L. S. Jensen, J. Magid, and N. E. Nielsen. Temporal variation of C and N turnover in soil after oilseed rape straw incorporation in the field: simulations with the soil-plant-atmosphere model DAISY. Ecological Modelling, 99(2–3):247–262, June 30, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097019595.

Mahendrarajah:1996:WSM

[MJY96] S. Mahendrarajah, A. J. Jakeman, and P. C. Young. Water supply in monsoonal Asia: modelling and predicting small tank storage. Ecological Modelling, 84(1–3):127–137, January 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094001278.

Morgenstern:1995:SSW

[MK95] M. Morgenstern and R. Kloss. Simulation of the soil water balance on the “Intensive Loam Site”. Ecological Modelling, 81(1–3):41–52, August 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400159F.

Magnusson:1997:MSW

[MK97] Kjartan G. Magnusson and Toshio Kasuya. Mating strategies in whale populations: searching strategy vs. harem strategy. Ecological Modelling, 102(2–3):225–242, October 29, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097000586.

McVoy:1995:DSN

[MKA^{+}95] C. W. McVoy, K. C. Kersebaum, M. Arning, P. Kleeberg, H. Othmer, and U. Schröder. A data set from north Germany for the validation of agroecosystem models: documentation and evaluation. Ecological Modelling, 81(1–3):265–300, August 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400197P.

Marsili-Libelli:1992:PEE

[ML92a] S. Marsili-Libelli. Parameter estimation of ecological models. Ecological Modelling, 62(4):233–258, August 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290001U.

Monserud:1992:CGV

[ML92b] Robert A. Monserud and Rik Leemans. Comparing global vegetation maps with the kappa statistic. Ecological Modelling, 62(4):275–293, August 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290003W.

Marsili-Libelli:1996:SKT

[ML96] Stefano Marsili-Libelli. Simplified kinetics of tropospheric ozone. Ecological Modelling, 84(1–3):233–244, January 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094001294.

Montagna:1997:MCE

[ML97] Paul A. Montagna and Jian Li. Modelling contaminant effects on deposit feeding nematodes near Gulf of Mexico production platforms. Ecological Modelling, 98(2–3):151–162, May 30, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096019114.

McCarthy:1998:PDM

[ML98] Michael A. McCarthy and David B. Lindenmayer. Population density and movement data for predicting mating systems of arboreal marsupials. Ecological Modelling, 109(2):193–202, June 11, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098000544.

Marsili-Libelli:1996:SLM

[MLB96] Stefano Marsili-Libelli and Simone Beni. Shock load modelling in the anaerobic digestion process. Ecological Modelling, 84(1–3):215–232, January 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094001251.

Montague:1990:SCR

[MLDH90] Clay L. Montague, Lynn W. Lefebvre, David G. Decker, and Nicholas R. Holler. Simulation of cotton rat population dynamics and response to rodenticide applications in Florida sugarcane. Ecological Modelling, 50(1–3):177–203, March 1990. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009090049M.

Montoya:1999:SND

[MLGV99] Ruby A. Montoya, A. L. Lawrence, W. E. Grant, and M. Velaso. Simulation of nitrogen dynamics and shrimp growth in an intensive shrimp culture system: effects of feed and feeding parameters. Ecological Modelling, 122(1–2):81–95, October 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001234.

Melcher:1996:AFC

[MM96] Dirk Melcher and Michael Matthies. Application of fuzzy clustering to data dealing with phytotoxicity. Ecological Modelling, 85(1):41–49, February 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095000100.

Metzker:1997:MSD

[MM97] Kathleen D. Metzker and William J. Mitsch. Modelling selfdesign of the aquatic community in a newly created freshwater wetland. Ecological Modelling, 100(1–3):61–86, December 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097001579.

Mueller:1998:SCT

[MMJ^{+}98] T. Mueller, J. Magid, L. S. Jensen, H. Svendsen, and N. E. Nielsen. Soil C and N turnover after incorporation of chopped maize, barley straw and blue grass in the field: Evaluation of the DAISY soil–organic-matter submodel. Ecological Modelling, 111(1):1–15, August 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000945.

Martins:1997:MEG

[MMJN97] Irene Martins, J. C. Marques, S. E. Jørgensen, and S. N. Nielsen. Modelling the effects of green macroalgae blooms on the population dynamics of Cyathura carinata (Crustacea: Isopoda) in an eutrophied estuary. Ecological Modelling, 102(1):33–53, October 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097000987.

Mitra:1992:PCR

[MMRR92] D. Mitra, Debasis Mukherjee, A. B. Roy, and S. Ray. Permanent coexistence in a resource-based competition system. Ecological Modelling, 60(1):77–85, January 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380092900146.

Moore:1993:AMS

[MN93] A. D. Moore and I. R. Noble. Automatic model simplification: the generation of replacement sequences and their use in vegetation modelling. Ecological Modelling, 70(1–2):137–157, November 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380093900776.

Monte:1998:PLT

[Mon98a] Luigi Monte. Predicting the long term behaviour of ^{90}Sr in lacustrine systems by a collective model. Ecological Modelling, 106(2–3):141–159, March 1, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097001890.

Monte:1998:PMD

[Mon98b] Luigi Monte. Predicting the migration of dissolved toxic substances from catchments by a collective model. Ecological Modelling, 110(3):269–279, July 22, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000817.

Moreira:1992:LTE

[Mor92] Helmar Nunes Moreira. Liénard-type equations and the epidemiology of malaria. Ecological Modelling, 60(2):139–150, March 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290043E.

Morlini:1999:RBF

[Mor99] Isabella Morlini. Radial basis function networks with partially classified data. Ecological Modelling, 120(2–3):109–118, August 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099000952.

Mosetti:1992:CPG

[Mos92] Renzo Mosetti. On the control of phytoplankton growth via a decoupling feedback law. Ecological Modelling, 62(1–3):71– 81, July 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290082P.

Mowrer:1997:PUT

[Mow97] H. Todd Mowrer. Propagating uncertainty through spatial estimation processes for old-growth subalpine forests using sequential Gaussian simulation in GIS. Ecological Modelling, 98(1): 73–86, May 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096019382.

Murray:1999:AAF

[MP99] A. G. Murray and J. S. Parslow. The analysis of alternative formulations in a simple model of a coastal ecosystem. Ecological Modelling, 119(2–3):149–166, July 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000460.

Morell:1996:MLC

[MPA96] Ignacio Morell, Francisco Padilla, and Juan M. Alberto. Modelling of leaching of chloride and nitrogen species in an experimental citrus grove. Ecological Modelling, 87(1–3):235– 247, June 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095000216.

Moravie:1997:ICR

[MPA97] M.-A. Moravie, J.-P. Pascal, and P. Auger. Investigating canopy regeneration processes through individual-based spatial models: application to a tropical rain forest. Ecological Modelling, 104(2–3):241–260, December 8, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001282.

Marques:1997:APE

[MPNJ97] J. C. Marques, M. Â. Pardal, S. N. Nielsen, and S. E. Jørgensen. Analysis of the properties of exergy and biodiversity along an estuarine gradient of eutrophication. Ecological Modelling, 102(1):155–167, October 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097000999.

Malkina-Pykh:1998:SEP

[MPP98] I. G. Malkina-Pykh and Yu. A. Pykh. Simulation of ecological processes using response functions method. Ecological Modelling, 108(1–3):199–218, May 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000295.

Mitsch:1991:MNR

[MR91a] William J. Mitsch and Brian C. Reeder. Modelling nutrient retention of a freshwater coastal wetland: estimating the roles of primary productivity, sedimentation, resuspension and hydrology. Ecological Modelling, 54(3–4):151–187, June 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190075C.

Mukherjee:1991:PMC

[MR91b] Debasis Mukherjee and A. B. Roy. Predator mediated coexistence of three competing species with L-V type interactions. Ecological Modelling, 58(1–4):285–301, November 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009190041X.

Martin:1997:ISD

[MR97] Jay F. Martin and K. R. Reddy. Interaction and spatial distribution of wetland nitrogen processes. Ecological Modelling, 105(1):1–21, December 14, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097001221.

Macchiato:1992:MMS

[MRCL92] M. F. Macchiato, M. Ragosta, C. Cosmi, and A. Lo Porto. A method in multivariate statistics to analyze ecosystems starting from their species composition. Ecological Modelling, 62(4):295–310, August 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290004X.

Mauchamp:1994:SDV

[MRL94] A. Mauchamp, S. Rambal, and J. Lepart. Simulating the dynamics of a vegetation mosaic: a spatialized functional model. Ecological Modelling, 71(1–3):107–130, January 1994. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094900787.

Morales-Ramos:1996:MRS

[MRLC96] J. A. Morales-Ramos, B. C. Legaspi, and R. I. Carruthers. Modification of the random-search type II functional response equation for incorporation into simulation models. Ecological Modelling, 91(1–3):249–253, November 15, 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095002006.

Morales-Ramos:1996:AME

[MRSK96] Juan A. Morales-Ramos, Kenneth R. Summy, and Edgar G. King. ARCASIM, a model to evaluate augmentation strategies of the parasitoid Catolaccus grandis against boll weevil populations. Ecological Modelling, 93(1–3):221–235, December 16, 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096000026.

Makela:1993:OOI

[MRSW93] M. E. Makela, G. A. Rowell, W. J. Sames, and L. T. Wilson. An object-oriented intracolonial and population level model of honey bees based on behaviors of European and Africanized subspecies. Ecological Modelling, 67(2–4):259–284, June 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390009H.

Matsumoto:1995:PIM

[MS95] Haruo Matsumoto and Hiromi Seno. On predator invasion into a multi-patchy environment of two kinds of patches. Ecological Modelling, 79(1–3):131–147, May 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093E0144R.

McClanahan:1997:MRB

[MS97] T. R. McClanahan and E. Sala. A Mediterranean rockybottom ecosystem fisheries model. Ecological Modelling, 104 (2–3):145–164, December 8, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438009700121X.

Mangel:1998:MLF

[MS98] Marc Mangel and P. V. Switzer. A model at the level of the foraging trip for the indirect effects of krill (Euphausia superba) fisheries on krill predators. Ecological Modelling, 105(2–3):235–256, January 1, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097001671.

Miao:1993:MED

[MT93] Sha Miao and Shunchi Tu. Modeling the effect of daily ration and feeding frequency on growth of redtail shrimp Penaeus penicillatus (Alock) at controlled temperatures. Ecological Modelling, 70(3–4):305–321, December 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390063X.

Miao:1995:MTE

[MT95] Sha Miao and Shunchi Tu. Modeling thermal effect on growth of Chinese shrimp, Penaeus chinensis (Osbeck). Ecological Modelling, 80(2–3):187–196, July 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009400048M.

Miao:1996:MET

[MT96] Sha Miao and Shunchi Tu. Modeling effect of thermic amplitude on growing Chinese shrimp, Penaeus chinensis (Osbeck). Ecological Modelling, 88(1–3):93–100, July 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095000720.

Mesple:1996:ESS

[MTCL96] Fabrice Mesplé, Marc Troussellier, Claude Casellas, and Pierre Legendre. Evaluation of simple statistical criteria to qualify a simulation. Ecological Modelling, 88(1–3):9–18, July 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009500033X.

Miller:1999:MSF

[MU99] Carol Miller and Dean L. Urban. A model of surface fire, climate and forest pattern in the Sierra Nevada, California. Ecological Modelling, 114(2–3):113–135, January 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098001197.

Muller:1992:HAE

[Mül92] Felix Müller. Hierarchical approaches to ecosystem theory. Ecological Modelling, 63(1–4):215–242, September 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290070U.

Muller:1997:SAE

[Mül97] Felix Müller. State-of-the-art in ecosystem theory. Ecological Modelling, 100(1–3):135–161, December 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001567.

Muller:1998:GES

[Mül98] Felix Müller. Gradients in ecological systems. Ecological Modelling, 108(1–3):3–21, May 1, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098000155.

Mateu:1998:SPF

[MUM98] J. Mateu, J. L. Usó, and F. Montes. The spatial pattern of a forest ecosystem. Ecological Modelling, 108(1–3):163–174, May 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098000271.

Mastala:1993:MHM

[MVBP93] Zoltán Mastala, Katalin V.-Balogh, and Miklós Perényi. A model for heavy-metal pollution of fish populations and exploited fish. Ecological Modelling, 70(3–4):263–276, December 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380093900606.

Mohren:1995:FGR

[MvdV95] Godefridus M. J. Mohren and Jan Renger van de Veen. Forest growth in relation to site conditions. Application of the model FORGRO to the Solling spruce site. Ecological Modelling, 83(1–2):173–183, December 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009500096E.

Marinussen:1996:CAE

[MvdZ96] Mari P. J. C. Marinussen and Sjoerd E. A. T. M. van der Zee. Conceptual approach to estimating the effect of homerange size on the exposure of organisms to spatially variable soil contamination. Ecological Modelling, 87(1–3):83–89, June 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400207X.

Montealegre:1995:DSM

[MVS^{+}95] Ricardo Jiménez Montealegre, Johan Verreth, Kees Steenbergen, Jan Moed, and Marcel Machiels. A dynamic simulation model for the blooming of Oscillatoria agardhii in a monomictic lake. Ecological Modelling, 78(1–2):17–24, March 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400114W.

Muller:1992:RPE

[MWJ92] F. Müller, W. Windhorst, and S. E. Jørgensen. Recent problems in ecosystem theory — conclusions of the workshop. Ecological Modelling, 63(1–4):325–331, September 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290076Q.

Mirschel:1995:SSW

[MWK95] W. Mirschel, K.-O. Wenkel, and R. Koitzsch. Simulation of soil water and evapotranspiration using the model BOWET and data sets from krummbach and eisenbach, two research catchments in North Germany. Ecological Modelling, 81(1–3):53–69, August 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400160J.

Malanson:1992:RVF

[MWY92] George P. Malanson, Walter E. Westman, and Yeuh-Lih Yan. Realized versus fundamental niche functions in a model of chaparral response to climatic change. Ecological Modelling, 64(4):261–277, November 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290026B.

Moreira:1997:GSC

[MY97] Helmar Nunes Moreira and Wang Yuquan. Global stability in a class of competitive cubic systems. Ecological Modelling, 102(2–3):273–285, October 29, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097000616.

Nielsen:1999:CCR

[NAFM99] S. N. Nielsen, P. M. Anastácio, A. F. Frias, and J. C. Marques. CRISP-crayfish rice integrated system of production. 5. Simulation of nitrogen dynamics. Ecological Modelling, 123(1):41–52, November 1, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099001660.

Nakajima:1992:SSF

[Nak92] Hisao Nakajima. Sensitivity and stability of flow networks. Ecological Modelling, 62(1–3):123–133, July 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290085S.

Nally:1997:SAC

[Nal97] R. Mac Nally. Scaling artefacts in confinement experiments: a simulation model. Ecological Modelling, 99(2–3):229–245, June 30, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097019583.

Neave:1997:PES

[NCNN97] H. M. Neave, R. B. Cunningham, T. W. Norton, and H. A. Nix. Preliminary evaluation of sampling strategies to estimate the species richness of diurnal, terrestrial birds using Monte Carlo simulation. Ecological Modelling, 95(1):17–27, February 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096000166.

Norberg:1997:TES

[ND97] J. Norberg and D. DeAngelis. Temperature effects on stocks and stability of a phytoplankton-zooplankton model and the dependence on light and nutrients. Ecological Modelling, 95(1):75–86, February 1, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096000336.

Nedorezov:1997:EEP

[Ned97] Lev. V. Nedorezov. Escape effect and population outbreaks. Ecological Modelling, 94(2–3):95–110, January 15, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095002227.

Nedorezov:1999:RPP

[Ned99] L. V. Nedorezov. Restoration of phase portrait structure for the dynamics of a forest pest, the pine moth (Dendrolimus pini L.). Ecological Modelling, 115(1):35–44, February 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098001689.

Norro:1996:BBM

[NF96] A. Norro and M. Frankignoulle. Biogeochemical box modelling at small scale. Application to the inorganic carbon cycle in the Bay of Calvi. Ecological Modelling, 88(1–3):101– 112, July 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095000739.

Nevo:1996:SOW

[NG96] Amnon Nevo and Luis Garcia. Spatial optimization of wildlife habitat. Ecological Modelling, 91(1–3):271–281, November 15, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009500193X.

Nuessly:1991:DHZ

[NHWS91] Gregg S. Nuessly, Albert W. Hartstack, John A. Witz, and Winfield L. Sterling. Dislodgement of Heliothis zea (Lepidoptera: Noctuidae) eggs from cotton due to rain and wind: a predictive model. Ecological Modelling, 55(1–2):89–102, July 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009190067B.

Nielsen:1992:SSD

[Nie92] Søren Nors Nielsen. Strategies for structural-dynamic modelling. Ecological Modelling, 63(1–4):91–101, September 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290063K.

Nielsen:1994:MSD

[Nie94] Søren Nors Nielsen. Modelling structural dynamical changes in a Danish shallow lake. Ecological Modelling, 73(1–2):13– 30, May 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094900957.

Nielsen:1995:MGS

[Nie95a] S. N. Nielsen. Matrices and graphs — Stability problems in mathematical ecology: Dmitrii O. Logofet. 1993, CRC Press, Boca Raton, FL, ISBN 0-8493-4246-5, 308 pp., US $69.96. Ecological Modelling, 77(1):86–87, January 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095900349.

Nielsen:1995:GWO

[Nie95b] Søren Nors Nielsen. Global wetlands: Old world and new: W. J. Mitsch (Editor). Elsevier Science, Amsterdam, 1994. 992 pp., Dfl. 500.00, US$285.50, ISBN 0-444-81478-7. Ecological Modelling, 82(2):207–208, October 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095900381.

Nielsen:1995:WEW

[Nie95c] Søren Nors Nielsen. Wetlands: 2nd ed. W. J. Mitsch and J. G. Gosselink. Van Nostrand Reinhold, 1993. 722 pp., £23.95, ISBN 0-442-00805-8. Ecological Modelling, 80(2–3):300, July 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009590042X.

Nielsen:1995:OES

[Nie95d] Søren Nors Nielsen. Optimization of exergy in a structural dynamic model. Ecological Modelling, 77(2–3):111–122, February 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380093E0088K.

Nielsen:1997:EOD

[Nie97] S. N. Nielsen. Examination and optimization of different exergy forms in macrophyte societies. Ecological Modelling, 102(1):115–127, October 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097001026.

Nihoul:1991:MPS

[Nih91] J. C. J. Nihoul. Modelling for population and sustainable development: A. J. Gilbert and L. C. Braat (Editors). Routledge, London, 1991. xvii + 261 pp., 35 tables, 37 figures, £45.00. ISBN 0-415-06187-3. Ecological Modelling, 59(3–4):293–294, December 15, 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380091901832.

Nihoul:1993:CAE

[Nih93] Jacques C. J. Nihoul. Comparative analysis of ecosystems. Patterns, mechanisms, and theories: Jonathan Cole, Gary Lovett and Stuart Findlay (Editors). Springer, 1991, 375 pp., 75 illustrations, 1 in full color, DM 148.00, ISBN 3-540-97488-1. Ecological Modelling, 65(3–4):298–299, February 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380093900868.

Newman:1996:BMT

[NJ96] Michael C. Newman and Rosemary H. Jagoe. Bioaccumulation models with time lags: Dynamics and stability criteria. Ecological Modelling, 84(1–3):281–286, January 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094001391.

Nikolov:1995:CBB

[NMS95] Nedialko T. Nikolov, William J. Massman, and Anna W. Schoettle. Coupling biochemical and biophysical processes at the leaf level: an equilibrium photosynthesis model for leaves of C_{3} plants. Ecological Modelling, 80(2–3):205–235, July 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400072P.

Nedorezov:1995:CBM

[NN95] L. V. Nedorezov and B. N. Nedorezova. Correlation between models of population dynamics in continuous and discrete time. Ecological Modelling, 82(1):93–97, September 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400069T.

Nunnari:1998:ANT

[NNR98] G. Nunnari, A. F. M. Nucifora, and C. Randieri. The application of neural techniques to the modelling of timeseries of atmospheric pollution data. Ecological Modelling, 111(2–3):187–205, September 1, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098001185.

Norton:1996:RDB

[Nor96] J. P. Norton. Roles for deterministic bounding in environmental modelling. Ecological Modelling, 86(2–3):157–161, May 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095000453.

Nikolov:1992:SRA

[NZ92] Nedialko T. Nikolov and Karl F. Zeller. A solar radiation algorithm for ecosystem dynamic models. Ecological Modelling, 61(3–4):149–168, June 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380092900157.

Oja:1997:FSV

[OA97] Tõnu Oja and Paul A. Arp. A forest soil vegetation atmosphere model (ForSVA), II: Application to northern tolerant hardwoods. Ecological Modelling, 95(2–3):225–247, February 14, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438009600035X.

Ottosson:1998:PAM

[OA98] Fredrik Ottosson and Otto Abrahamsson. Presentation and analysis of a model simulating epilimnetic and hypolimnetic temperatures in lakes. Ecological Modelling, 110(3):233–253, July 22, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098000672.

Oelschlagel:1995:MDG

[Oel95] Burkhard Oelschlägel. A method for downscaling global climate model calculations by a statistical weather generator. Ecological Modelling, 82(2):199–204, October 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094000905.

Ogana:1996:MVD

[Oga96] W. Ogana. Modelling the vertical distribution of insects. Ecological Modelling, 89(1–3):225–230, August 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001379.

Ottosson:1997:PAM

[OH97] Fredrik Ottosson and Lars Håkanson. Presentation and analysis of a model simulating the pH response of lake liming. Ecological Modelling, 105(1):89–111, December 14, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001415.

Oshima:1999:ENB

[OKS99] Yuko Oshima, Michio J. Kishi, and Takashige Sugimoto. Evaluation of the nutrient budget in a seagrass bed. Ecological Modelling, 115(1):19–33, February 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001550.

Onstad:1990:SMT

[OM90] David W. Onstad and Joseph V. Maddox. Simulation model of Tribolium confusum and its pathogen, Nosema whitei. Ecological Modelling, 51(1–2):143–160, May 1990. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009090062L.

Odulaja:1997:SBM

[OM97a] Adedapo Odulaja and L. Chuka Madubunyi. A sampling bias model for odour-baited traps in relation to tsetse hunger cycle and population suppression. Ecological Modelling, 104(2–3):165–173, December 8, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097001233.

Ozesmi:1997:SHM

[ÖM97b] Uygar Özesmi and William J. Mitsch. A spatial habitat model for the marsh-breeding red-winged blackbird (Agelaius phoeniceus L.) in coastal Lake Erie wetlands. Ecological Modelling, 101(2–3):139–152, August 15, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097019832.

Omori:1995:ASL

[Omo95] Koji Omori. The adaptive significance of a lunar or semi-lunar reproductive cycle in marine animals. Ecological Modelling, 82(1):41–49, September 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400082S.

Omori:1997:MSD

[Omo97] Koji Omori. Mature size determination in copepods the adaptive significance of mature size in copepods: output or efficiency selection? Ecological Modelling, 99(2–3):203–215, June 30, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097019522.

Ozesmi:1999:ANN

[ÖÖ99] Stacy L. Özesmi and Uygar Özesmi. An artificial neural network approach to spatial habitat modelling with interspecific interaction. Ecological Modelling, 116(1):15–31, March 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098001495.

Odum:1996:SEE

[OP96] Howard T. Odum and Nils Peterson. Simulation and evaluation with energy systems blocks. Ecological Modelling, 93(1–3):155–173, December 16, 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095002219.

Omlin:1999:CTE

[OR99] Martin Omlin and Peter Reichert. A comparison of techniques for the estimation of model prediction uncertainty. Ecological Modelling, 115(1):45–59, February 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001744.

Olson:1995:ECA

[OS95] Richard L. Olson and Ronaldo A. Sequeira. An emergent computational approach to the study of ecosystem dynamics. Ecological Modelling, 79(1–3):95–120, May 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093E0124L.

Osho:1991:MMT

[Osh91] J. S. A. Osho. Matrix model for tree population projection in a tropical rain forest of south-western Nigeria. Ecological Modelling, 59(3–4):247–255, December 15, 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380091901809.

Osho:1996:MTP

[Osh96] J. S. A. Osho. Modelling the tree population dynamics of the most abundant species in a Nigerian tropical rain forest. Ecological Modelling, 89(1–3):175–181, August 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001301.

Oja:1995:FMS

[OYA95] Tõnu Oja, Xiwei Yin, and Paul A. Arp. The forest modelling series ForM-S: applications to the Solling spruce site. Ecological Modelling, 83(1–2):207–217, December 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009500099H.

Ogihara:1996:PNM

[OZIA96] Yousou Ogihara, Kresimir Zic, Jörg Imberger, and Steve Armfield. A parametric numerical model for lake hydrodynamics. Ecological Modelling, 86(2–3):271–276, May 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095000631.

Pachepsky:1996:MLG

[PA96] L. B. Pachepsky and B. Acock. A model 2DLEAF of leaf gas exchange: development, validation, and ecological application. Ecological Modelling, 93(1–3):1–18, December 16, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095002073.

Peng:1999:MRN

[PA99] Changhui Peng and Michael J. Apps. Modelling the response of net primary productivity (NPP) of boreal forest ecosystems to changes in climate and fire disturbance regimes. Ecological Modelling, 122(3):175–193, October 20, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099001374.

Padilla:1996:PCD

[PACS96] Dianna K. Padilla, Stephen C. Adolph, Kathryn L. Cottingham, and Daniel W. Schneider. Predicting the consequences of dreissenid mussels on a pelagic food web. Ecological Modelling, 85(2–3):129–144, March 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094001855.

Perianez:1996:MSM

[PAGL96] R. Periáñez, J. M. Abril, and M. García-León. Modelling the suspended matter distribution in an estuarine system. Application to the Odiel River in southwest Spain. Ecological Modelling, 87(1–3):169–179, June 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095000267.

Parker:1993:DMA

[Par93a] R. A. Parker. Dynamic models for ammonium inhibition of nitrate uptake by phytoplankton. Ecological Modelling, 66(1–2):113–120, March 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009390042Q.

Parker:1993:FCB

[Par93b] R. A. Parker. Feedback control of birth and death rates for optimal population density. Ecological Modelling, 65(1–2):137–146, January 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009390130K.

Parshotam:1996:RSC

[Par96] Aroon Parshotam. The Rothamsted soil-carbon turnover model — discrete to continuous form. Ecological Modelling, 86(2–3):283–289, May 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095000658.

Priyantha:1997:MEC

[PASG97] D. G. Nimal Priyantha, Takashi Asaeda, Satoki Saitoh, and Kohichi Gotoh. Modelling effects of curtain method on algal blooming in reservoirs. Ecological Modelling, 98(2–3):89–104, May 30, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096019060.

Patten:1992:EEE

[Pat92] Bernard C. Patten. Energy, emergy and environs. Ecological Modelling, 62(1–3):29–69, July 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290081O.

Patten:1995:NIE

[Pat95] Bernard C. Patten. Network integration of ecological extremal principles: exergy, emergy, power, ascendency, and indirect effects. Ecological Modelling, 79(1–3):75–84, May 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400037I.

Patten:1997:SCS

[Pat97] Bernard C. Patten. Synthesis of chaos and sustainability in a nonstationary linear dynamic model of the American black bear (Ursus americanus Pallas) in the Adirondack Mountains of New York. Ecological Modelling, 100(1–3):11–42, December 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001543.

Patten:1998:EAT

[Pat98] Bernard C. Patten. Ecology’s AWFUL theorem: sustaining sustainability. Ecological Modelling, 108(1–3):97–105, May 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098000210.

Payne:1999:VDE

[Pay99] Karen Payne. The VC dimension and ecological modelling. Ecological Modelling, 118(2–3):249–259, June 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000484.

Pfeifer:1996:MSD

[PBO^{+}96] D. Pfeifer, H.-P. Bäumer, H. Ortleb, G. Sach, and U. Schleier. Modeling spatial distributional patterns of benthic meiofauna species by Thomas and related processes. Ecological Modelling, 87(1–3):285–294, June 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009688166X.

Peterson:1999:SDP

[PC99a] A. Townsend Peterson and Kevin P. Cohoon. Sensitivity of distributional prediction algorithms to geographic data completeness. Ecological Modelling, 117(1):159–164, April 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438009900023X.

Pradhan:1999:DRM

[PC99b] T. Pradhan and K. S. Chaudhuri. A dynamic reaction model of a two-species fishery with taxation as a control instrument: a capital theoretic analysis. Ecological Modelling, 121(1):1–16, September 1, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099000629.

Polakow:1999:MFR

[PD99] Daniel A. Polakow and Timothy T. Dunne. Modelling firereturn interval T: stochasticity and censoring in the twoparameter Weibull model. Ecological Modelling, 121(1):79– 102, September 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380099000745.

Postek:1995:APP

[PDAS95] Kimberley M. Postek, Charles T. Driscoll, John D. Aber, and Robert C. Santore. Application of PNET-CN/CHESS to a spruce stand in Solling, Germany. Ecological Modelling, 83(1–2):163–172, December 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009500095D.

Perez-Espana:1999:MEM

[PEAS99] Horacio Pérez-España and Francisco Arreguín-Sánchez. A measure of ecosystem maturity. Ecological Modelling, 119 (1):79–85, July 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380099000587.

Pedersen:1998:MTM

[Ped98] Brian S. Pedersen. Modeling tree mortality in response to short- and long-term environmental stresses. Ecological Modelling, 105(2–3):347–351, January 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001622.

Persicani:1993:ALG

[Per93] Danilo Persicani. Atrazine leaching into groundwater: comparison of five simulation models. Ecological Modelling, 70(3–4):239–261, December 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380093900592.

Persicani:1996:PLF

[Per96] Danilo Persicani. Pesticide leaching into field soils: sensitivity analysis of four mathematical models. Ecological Modelling, 84(1–3):265–280, January 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094001367.

Perianez:1998:TDC

[Per98] R. Periáñez. A three dimensional σ-coordinate model to simulate the dispersion of radionuclides in the marine environment: application to the Irish Sea. Ecological Modelling, 114(1):59–70, December 1, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098001161.

Pastres:1995:UPC

[PFP^{+}95] R. Pastres, Davide Franco, G. Pecenik, C. Solidoro, and C. Dejak. Using parallel computers in environmental modelling: a working example. Ecological Modelling, 80(1):69–85, June 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400049N.

Parysow:1997:VEC

[PG97] Pablo Parysow and George Gertner. Virtual experimentation: conceptual models and hypothesis testing of ecological scenarios. Ecological Modelling, 98(1):59–71, May 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096019370.

Parthasarathy:1998:SCM

[PG98] S. Parthasarathy and J. Güémez. Synchronisation of chaotic metapopulations in a cascade of coupled logistic map models. Ecological Modelling, 106(1):17–25, February 16, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001750.

Parkhill:1999:MEL

[PG99a] Kenneth L. Parkhill and John S. Gulliver. Modeling the effect of light on whole-stream respiration. Ecological Modelling, 117(2–3):333–342, May 17, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099000174.

Parysow:1999:RIH

[PG99b] Pablo Parysow and George Gertner. The role of interactions in hypothesis testing of ecological scenarios with process models. Ecological Modelling, 116(2–3):107–124, March 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098001239.

Peterson:1991:SHP

[PGD91] M. J. Peterson, W. E. Grant, and D. S. Davis. Simulation of host — parasite interactions within a resource management framework: impact of brucellosis on bison population dynamics. Ecological Modelling, 54(3–4):299–320, June 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009190081B.

Peterson:1998:SRS

[PGS98] Markus J. Peterson, William E. Grant, and Nova J. Silvy. Simulation of reproductive stages limiting productivity of the endangered Attwater’s prairie chicken. Ecological Modelling, 111(2–3):283–295, September 1, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098001112.

Patten:1995:FPF

[PH95] Bernard C. Patten and Masahiko Higashi. First passage flows in ecological networks: measurement by input-output flow analysis. Ecological Modelling, 79(1–3):67–74, May 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009400040O.

Patten:1990:TDE

[PHB90] Bernard C. Patten, Masahiko Higashi, and Thomas P. Burns. Trophic dynamics in ecosystem networks: Significance of cycles and storage. Ecological Modelling, 51(1–2):1–28, May 1990. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009090055L.

Phillips:1995:TLE

[Phi95] Jonathan D. Phillips. Time lags and emergent stability in morphogenic/pedogenic system models. Ecological Modelling, 78(3):267–276, April 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400180P.

Piepho:1995:ISC

[Pie95] Hans-Peter Piepho. Implications of a simple competition model for the stability of an intercropping system.

Ecological Modelling, 80(2–3):251–256, July 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009400051I.

Park:1996:DSR

[PJ96] Seok S. Park and Peter R. Jaffé. Development of a sediment redox potential model for the assessment of postdepositional metal mobility. Ecological Modelling, 91(1–3):169–181, November 15, 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095001883.

Post:1999:PDS

[PJ99] David A. Post and Anthony J. Jakeman. Predicting the daily streamflow of ungauged catchments in S.E. Australia by regionalising the parameters of a lumped conceptual rainfallrunoff model. Ecological Modelling, 123(2–3):91–104, November 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001258.

Post:1996:MLC

[PJL^{+}96] D. A. Post, A. J. Jakeman, I. G. Littlewood, P. G. Whitehead, and M. D. A. Jayasuriya. Modelling land-cover-induced variations in hydrologic response: Picaninny Creek, Victoria. Ecological Modelling, 86(2–3):177–182, May 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095000488.

Pilette:1992:FFT

[PK92] Ron Pilette and Dwight T. Kincaid. First flow-thru analysis in ecosystem studies. Ecological Modelling, 64(1):1, 4–10, October 15, 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290046H.

Polymenopoulos:1990:EEM

[PL90] Alex D. Polymenopoulos and Garrel Long. Estimation and evaluation methods for population growth models with spatial diffusion: Dynamics of mountain pine beetle. Ecological Modelling, 51(1–2):97–121, May 1990. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009090060T.

Portielje:1995:ERB

[PL95] Robert Portielje and Lambertus Lijklema. The effect of reaeration and benthic algae on the oxygen balance of an artificial ditch. Ecological Modelling, 79(1–3):35–48, May 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380093E0122J.

Park:1996:MMS

[PL96] Seok Soon Park and Yong Seok Lee. A multiconstituent moving segment model for water quality predictions in steep and shallow streams. Ecological Modelling, 89(1–3):121–131, August 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095001263.

Phillips:1996:SUA

[PM96] Donald L. Phillips and Danny G. Marks. Spatial uncertainty analysis: propagation of interpolation errors in spatially distributed models. Ecological Modelling, 91(1–3): 213–229, November 15, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095001913.

Pykh:1997:PPM

[PMP97] Yu. A. Pykh and I. G. Malkina-Pykh. POLMOD.PEST — the model of pesticides dynamics in the elementary ecosystems. Ecological Modelling, 98(2–3):215–236, May 30, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096019175.

Pet:1996:SSS

[PMV96] J. S. Pet, M. A. M. Machiels, and W. L. T. Van Densen. A size-structured simulation model for evaluating management strategies in gillnet fisheries exploiting spatially differentiated populations. Ecological Modelling, 88(1–3):195–214, July 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095001077.

Pollett:1996:MLT

[Pol96] Philip K. Pollett. Modelling the long-term behaviour of evanescent ecological systems. Ecological Modelling, 86(2–3):135– 139, May 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095000402.

Porch:1998:NSL

[Por98] C. E. Porch. A numerical study of larval fish retention along the southeast Florida coast. Ecological Modelling, 109(1):35– 59, June 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098000052.

Power:1993:PVE

[Pow93] M. Power. The predictive validation of ecological and environmental models. Ecological Modelling, 68(1–2):33–50, July 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380093901063.

Power:1996:CCI

[Pow96] M. Power. Characterizing cumulative impacts using a brook trout population dynamics model. Ecological Modelling, 90(3):257–270, November 1, 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001530.

Power:1995:MFA

[PP95] M. Power and G. Power. A modelling framework for analyzing anthropogenic stresses on brook trout (Salvelinus fontinalis) populations. Ecological Modelling, 80(2–3):171–185, July 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400058P.

Pan:1995:DTA

[PR95a] Yude Pan and Dudley J. Raynal. Decomposing tree annual volume increments and constructing a system dynamic model of tree growth. Ecological Modelling, 82(3):299–312, November 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400096Z.

Pipp:1995:PCM

[PR95b] Eveline Pipp and Eugen Rott. A phytoplankton compartment model for a small meromictic lake with special reference to species-specific niches and long-term changes. Ecological Modelling, 78(1–2):129–148, March 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009400123Y.

Provencher:1995:TCI

[PR95c] Louis Provencher and Susan E. Riechert. Theoretical comparisons of individual success between phenotypically pure and mixed generalist predator populations. Ecological Modelling, 82(2):175–191, October 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009592851M.

Pacheco:1997:HBP

[PRF97] José M. Pacheco, César Rodríguez, and Isabel Fernández. Hopf bifurcations in predator–prey systems with social predator behaviour. Ecological Modelling, 105(1):83–87, December 14, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001403.

Potter:1997:SMN

[PRK97] Christopher S. Potter, Ralph H. Riley, and Steven A. Klooster. Simulation modeling of nitrogen trace gas emissions along an age gradient of tropical forest soils. Ecological Modelling, 97(3):179–196, April 28, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096019035.

Parrish:1990:MTW

[PS90] Rudolph S. Parrish and Charles N. Smith. A method for testing whether model predictions fall within a prescribed factor of true values, with an application to pesticide leaching. Ecological Modelling, 51(1–2):59–72, May 1990. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009090058O.

Parrish:1991:AR

[PS91] Rudolph S. Parrish and Charles N. Smith. Author’s response. Ecological Modelling, 54(1–2):139–141, May 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009190105A.

Patoucheas:1993:NHM

[PS93a] Dimitris P. Patoucheas and George Stamou. Non homogeneous Markovian models in ecological modelling: a study of zoobenthos dynamics in Thermaikos Gulf, Greece. Ecological Modelling, 66(3–4):197–215, April 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093901137.

Ping:1993:SMS

[PS93b] Sun Ping and Zeng Shimai. Statistical model to study the relationship between the wheat variety composition and the race composition of wheat stripe rust. Ecological Modelling, 66(3–4):277–287, April 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009390118C.

Prentice:1993:SMT

[PSC93] I. Colin Prentice, Martin T. Sykes, and Wolfgang Cramer. A simulation model for the transient effects of climate change on forest landscapes. Ecological Modelling, 65(1–2):51–70, January 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390126D.

Patten:1997:EEC

[PSJ97] Bernard C. Patten, Milan Straškraba, and Sven E. Jørgensen. Ecosystems emerging: 1. Conservation. Ecological Modelling, 96(1–3):221–284, March 1, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096000695.

Park:1998:AMS

[PSK98] Kyeong Park, Jian Shen, and Albert Y. Kuo. Application of a multi-step computation scheme to an intratidal estuarine water quality model. Ecological Modelling, 110(3):281–292, July 22, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098000829.

Pielke:1993:ATE

[PSL^{+}93] R. A. Pielke, D. S. Schimel, T. J. Lee, T. G. F. Kittel, and X. Zeng. Atmosphere-terrestrial ecosystem interactions: implications for coupled modeling. Ecological Modelling, 67(1): 5–18, May 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390096B.

Perttunen:1998:LMC

[PSN98] J. Perttunen, R. Sievänen, and E. Nikinmaa. LIGNUM: a model combining the structure and the functioning of trees. Ecological Modelling, 108(1–3):189–198, May 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000283.

Papadopoulos:1994:BAB

[PT94] Alex S. Papadopoulos and Ram C. Tiwari. Bayesian approach for BOD and DO when the discharged pollutants are random. Ecological Modelling, 71(4):245–257, February 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094901368.

Panizzuti:1995:PSM

[PT95] Monica Panizzuti and Gianni Tartari. pH simulation model in a mixed layer of a lacustrine environment (the case of Lake Orta). Ecological Modelling, 78(1–2):37–49, March 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400116Y.

Paruelo:1997:PFC

[PT97] JoséM. Paruelo and Fernando Tomasel. Prediction of functional characteristics of ecosystems: a comparison of artificial neural networks and regression models. Ecological Modelling, 98(2–3):173–186, May 30, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096019138.

Purger:1999:PAR

[PT99] Jenő J. Purger and Andreja Tepavčević. Pattern analysis of red-footed falcon (Falco vespertinus) nests in the rook (Corvus frugilegus) colony near torda (Voivodina, Yugoslavia), using fuzzy correspondences and entropy. Ecological Modelling, 117(1):91–97, April 1, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099000125.

Papadopoulos:1991:BPT

[PTM91] Alex S. Papadopoulos, Ram C. Tiwari, and Michael J. Muha. Bootstrap procedures for time series analysis of BOD data. Ecological Modelling, 55(1–2):57–65, July 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380091900648.

Park:1997:SMW

[PU97] Seok Soon Park and Christopher G. Uchrin. A stoichiometric model for water quality interactions in macrophyte dominated water bodies. Ecological Modelling, 96(1–3):165–174, March 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096000646.

Pahl-Wostl:1991:PST

[PW91] Claudia Pahl-Wostl. Patterns in space and time — a new method for their characterization. Ecological Modelling, 58(1–4):141–157, November 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009190033W.

Pahl-Wostl:1993:HOA

[PW93] Claudia Pahl-Wostl. The hierarchical organization of the aquatic ecosystem: an outline how reductionism and holism may be reconciled. Ecological Modelling, 66(1–2):81–100, March 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390040Y.

Pahl-Wostl:1997:DSF

[PW97] Claudia Pahl-Wostl. Dynamic structure of a food web model: comparison with a food chain model. Ecological Modelling, 100(1–3):103–123, December 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001518.

Pahl-Wostl:1993:QSF

[PWU93] Claudia Pahl-Wostl and Robert E. Ulanowicz. Quantification of species as functional units within an ecological network. Ecological Modelling, 66(1–2):65–79, March 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390039U.

Papadopoulos:1994:MAM

[PYE94] Basil K. Papadopoulos, Trisevgeni Yiannakopoulou, and Nicholas Elias. A model for analysis of multivariable systems with an application to ecology. Ecological Modelling, 74(3–4):139–160, August 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380094901171.

Rothschild:1996:PDI

[RA96] Brian J. Rothschild and Jerald S. Ault. Population-dynamic instability as a cause of patch structure. Ecological Modelling, 93(1–3):237–249, December 16, 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096000051.

Reynolds:1997:MGP

[RA97] James F. Reynolds and Basil Acock. Modularity and genericness in plant and ecosystem models. Ecological Modelling, 94(1):7–16, January 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096019242.

Rolff:1999:PED

[RÅ99] Carl Rolff and Göran I. Ågren. Predicting effects of different harvesting intensities with a model of nitrogen limited forest growth. Ecological Modelling, 118(2–3):193–211, June 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099000435.

Ricotta:1999:FSS

[RAM99] Carlo Ricotta, Giancarlo Avena, and Marco Marchetti. The flaming sandpile: self-organized criticality and wildfires. Ecological Modelling, 119(1):73–77, July 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000575.

Radford:1996:IMS

[RB96] P. J. Radford and J. C. Blackford. Interdisciplinary methods for successful ecological simulation. Ecological Modelling, 86(2–3):265–270, May 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095000623.

Roux:1998:EMF

[RB98] Olivier Roux and Johann Baumgärtner. Evaluation of mortality factors and risk analysis for the design of an integrated pest management system. Ecological Modelling, 109(1):61–75, June 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098000350.

Reuter:1999:EPI

[RB99] Hauke Reuter and Broder Breckling. Emerging properties on the individual level: modelling the reproduction phase of the European robin Erithacus rubecula. Ecological Modelling, 121(2–3):199–219, September 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000782.

Reich:1997:MSS

[RBM97] R. M. Reich, C. D. Bonham, and K. L. Metzger. Modeling small-scale spatial interaction of shortgrass prairie species. Ecological Modelling, 101(2–3):163–174, August 15, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097019765.

Rowell:1993:SAM

[RC93] Gareth A. Rowell and William H. Cade. Simulation of alternative male reproductive behavior: calling and satellite behavior in field crickets. Ecological Modelling, 65(3–4):265–280, February 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380093900835.

Rajar:1997:HWQa

[RC97] Rudi Rajar and Matjaz Cetina. Hydrodynamic and water quality modelling: an experience. Ecological Modelling, 101(2–3):195–207, August 15, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097000471.

Rose:1993:IBM

[RCD93] Kenneth A. Rose, Sigurd W. Christensen, and Donald L. DeAngelis. Individual-based modeling of populations with high mortality: a new method based on following a fixed number of model individuals. Ecological Modelling, 68(3–4):273–292, August 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390022K.

Rybczyk:1998:REM

[RCD98] John M. Rybczyk, J. C. Callaway, and J. W. Day, Jr. A relative elevation model for a subsiding coastal forested wetland receiving wastewater effluent. Ecological Modelling, 112(1):23–44, October 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098001252.

Rajar:1997:HWQb

[RCS97] Rudi Rajar, Matjaz Cetina, and Andrej Sirca. Hydrodynamic and water quality modelling: case studies. Ecological Modelling, 101(2–3):209–228, August 15, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097000525.

Rochester:1996:SML

[RDFZ96] W. A. Rochester, M. L. Dillon, G. P. Fitt, and M. P. Zalucki. A simulation model of the long-distance migration of Helicoverpa spp. moths. Ecological Modelling, 86(2–3):151– 156, May 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095000437.

Reckhow:1993:RCM

[Rec93] Kenneth H. Reckhow. A random coefficient model for chlorophyll-nutrient relationships in lakes. Ecological Modelling, 70(1–2):35–50, November 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://

www.sciencedirect.com/science/article/pii/030438009390071Y.

Reckhow:1994:WQS

[Rec94] Kenneth H. Reckhow. Water quality simulation modeling and uncertainty analysis for risk assessment and decision making. Ecological Modelling, 72(1–2):1–20, March 1994. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094901430.

Rizzotto:1997:PBM

[RF97] M. Rizzotto and S. Focardi. A physiologically-based model of a self-motivated hare in relation to its ecology. Ecological Modelling, 95(2–3):191–209, February 14, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096000373.

Richter:1999:RFF

[RFB99] Hartmut Richter, Ulfert Focken, and Klaus Becker. A review of the fish feeding model MAXIMS. Ecological Modelling, 120(1):47–64, August 3, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438009900071X.

Recknagel:1997:ANN

[RFHY97] Friedrich Recknagel, Mark French, Pia Harkonen, and Ken-Ichi Yabunaka. Artificial neural network approach for modelling and prediction of algal blooms. Ecological Modelling, 96(1–3):11–28, March 1, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438009600049X.

Reichelt:1990:APO

[RGBM90] R. E. Reichelt, W. Greve, R. H. Bradbury, and P. J. Moran. Acanthaster planci outbreak initiation: a starfishcoral site model. Ecological Modelling, 49(3–4):153–177, January 1990. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009090026D.

Riedo:1998:PSM

[RGRF98] Marcel Riedo, Anton Grub, Marc Rosset, and Jürg Fuhrer. A pasture simulation model for dry matter production, and fluxes of carbon, nitrogen, water and energy. Ecological Modelling, 105(2–3):141–183, January 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001105.

Rey:1996:SAU

[RH96a] T. D. Rey and W. A. Havranek. Some aspects of using the SimuSolv program for environmental, pharmacokinetics and toxicological applications. Ecological Modelling, 86(2–3):277– 282, May 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009500064X.

Robinson:1996:SIA

[RH96b] J. Brett Robinson and Keith R. Helyar. Simulating the impact of acidifying farming systems on Australian soils.

Ecological Modelling, 86(2–3):207–211, May 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095000534.

Richter:1995:E

[Ric95] Otto Richter. Editorial. Ecological Modelling, 77(1):1–2, January 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900322.

Ripl:1995:MWC

[Rip95] W. Ripl. Management of water cycle and energy flow for ecosystem control: the energy-transport-reaction (ETR) model. Ecological Modelling, 78(1–2):61–76, March 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094001182.

Roche:1999:CDM

[RJN99] Romain Roche, Marie-Hélène Jeuffroy, and Bertrand Ney. Comparison of different models predicting the date of beginning of flowering in pea (Pisum sativum L.). Ecological Modelling, 118(2–3):213–226, June 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000368.

Rana:1997:ESP

[RKM97] Gianfranco Rana, Nader Katerji, and Marcello Mastrorilli. Environmental and soil-plant parameters for modelling actual crop evapotranspiration under water stress conditions. Ecological Modelling, 101(2–3):363–371, August 15, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009700063X.

Ro:1999:GPS

[RL99] Tae Ho Ro and Garrell E. Long. GPA-Phenodynamics, a simulation model for the population dynamics and phenology of green peach aphid in potato: formulation, validation, and analysis. Ecological Modelling, 119(2–3):197–209, July 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099000538.

Railsback:1999:MRI

[RLHD99] Steven F. Railsback, Roland H. Lamberson, Bret C. Harvey, and Walter E. Duffy. Movement rules for individualbased models of stream fish. Ecological Modelling, 123(2– 3):73–89, November 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380099001246.

Rinaldi:1992:SFL

[RM92] S. Rinaldi and S. Muratori. Slow-fast limit cycles in predator– prey models. Ecological Modelling, 61(3–4):287–308, June 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380092900238.

Rinaldi:1993:CCS

[RM93] Sergio Rinaldi and Simona Muratori. Conditioned chaos in seasonally perturbed predator–prey models. Ecological Modelling, 69(1–2):79–97, September 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380093900503.

Ritchie:1995:SPB

[RM95] Scott A. Ritchie and Clay L. Montague. Simulated populations of the black salt march mosquito (Aedes taeniorhynchus) in a Florida mangrove forest. Ecological Modelling, 77(2–3):123–141, February 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380093E0083F.

Rossi:1992:ECC

[RMT92] G. E. Rossi, H. Muntau, and G. Tartari. The evolution of copper concentrations in Lake Orta, Italy. Ecological Modelling, 64(1):23–45, October 15, 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290048J.

Reichert:1997:UOE

[RO97] Peter Reichert and Martin Omlin. On the usefulness of overparameterized ecological models. Ecological Modelling, 95(2–3):289–299, February 14, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096000439.

Roberts:1996:LVM

[Rob96a] David W. Roberts. Landscape vegetation modelling with vital attributes and fuzzy systems theory. Ecological Modelling, 90(2):175–184, October 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001646.

Roberts:1996:MFD

[Rob96b] David W. Roberts. Modelling forest dynamics with vital attributes and fuzzy systems theory. Ecological Modelling, 90(2):161–173, October 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095001638.

Rodriguez:1999:MDH

[Rod99] Diego J. Rodríguez. A method to detect higher order interactions in ecological communities. Ecological Modelling, 117(1):81–89, April 1, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099000095.

Romey:1996:IDM

[Rom96] William L. Romey. Individual differences make a difference in the trajectories of simulated schools of fish. Ecological Modelling, 92(1):65–77, November 21, 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095002022.

Rott:1995:I

[Rot95] Eugen Rott. Introduction. Ecological Modelling, 78(1–2):1–2, March 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095900187.

Radeloff:1999:HPM

[RPH99] Volker C. Radeloff, Anna M. Pidgeon, and Patrick Hostert. Habitat and population modelling of roe deer using an interactive geographic information system. Ecological Modelling, 114(2–3):287–304, January 1, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098001641.

Recknagel:1994:HES

[RPJK94] Frieder Recknagel, Thomas Petzoldt, Olaf Jaeke, and Falk Krusche. Hybrid expert system DELAQUA — a toolkit for water quality control of lakes and reservoirs. Ecological Modelling, 71(1–3):17–36, January 1994. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094900744.

Rossi:1998:MAA

[RPM^{+}98] Claudio Rossi, Marcello Porcelli, Chiara Mocenni, Nadia Marchettini, Steven Loiselle, and Simone Bastianoni. A modelling approach for the analysis of xylose–ethanol bioconversion. Ecological Modelling, 113(1–3):157–162, November 2, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098001409.

Rego:1993:MCD

[RPT93] F. Rego, J. Pereira, and L. Trabaud. Modelling community dynamics of a Quercus coccifera L. garrigue in relation to fire using Markov chains. Ecological Modelling, 66(3–4):251–260, April 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390116A.

Roese:1991:HHF

[RRF91] John H. Roese, Ken L. Risenhoover, and L. Joseph Folse. Habitat heterogeneity and foraging efficiency: an individualbased model. Ecological Modelling, 57(1–2):133–143, October 1, 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380091900589.

Reddy:1990:TTA

[RRJ90] K. R. Reddy, P. S. C. Rao, and R. E. Jessup. Transformation and transport of ammonium nitrogen in a flooded organic soil. Ecological Modelling, 51(3–4):205–216, June 1990. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009090065O.

Rodrigo:1997:MTM

[RRNM97] Alfredo Rodrigo, Sylvie Recous, Catherine Neel, and Bruno Mary. Modelling temperature and moisture effects on C–N transformations in soils: comparison of nine models. Ecological Modelling, 102(2–3):325–339, October 29, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097000677.

Rotmans:1991:MTD

[RS91] Jan Rotmans and Robert J. Swart. Modelling tropical deforestation and its consequences for global climate. Ecological Modelling, 58(1–4):217–247, November 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380091900383.

Rai:1993:PDB

[RS93] Vikas Rai and R. Sreenivasan. Period-doubling bifurcations leading to chaos in a model food chain. Ecological Modelling, 69(1–2):63–77, September 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009390049X.

Reed:1996:CMF

[RS96] William J. Reed and Clement M. Simons. A contagion model of a fishery and its use in analyzing catch-effort data. Ecological Modelling, 92(2–3):179–191, December 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001786.

Rose:1998:SAM

[RS98a] Kenneth A. Rose and Eric P. Smith. Statistical assessment of model goodness-of-fit using permutation tests. Ecological Modelling, 106(2–3):129–139, March 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001889.

Ruxton:1998:NBR

[RS98b] Graeme D. Ruxton and Leonardo A. Saravia. The need for biological realism in the updating of cellular automata models. Ecological Modelling, 107(2–3):105–112, April 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097001798.

Raychaudhuri:1996:ETV

[RSC96] S. Raychaudhuri, D. K. Sinha, and J. Chattopadhyay. Effect of time-varying cross-diffusivity in a two-species Lotka–Volterra competitive system. Ecological Modelling, 92(1):55–64, November 21, 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095002014.

Reyes:1994:ROE

[RSD94] E. Reyes, F. H. Sklar, and J. W. Day. A regional organism exchange model for simulating fish migration. Ecological Modelling, 74(3–4):255–276, August 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094901228.

Rosado-Solorzano:1998:PTS

[RSdP98] R. Rosado-Solórzano and Sergio A. Guzmán del Próo. Preliminary trophic structure model for Tampamachoco lagoon, Veracruz, Mexico. Ecological Modelling, 109(2):141–154, June 11, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098000118.

Radke:1991:ANT

[RSKS91] J. K. Radke, M. J. Shaffer, K. S. Kroll, and J. Saponara. Application of the nitrogen-tillage-residue-management (NTRM) model for corn grown in low-input and conventional agricultural systems. Ecological Modelling, 55(3–4):241–255, August 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009190089J.

Racsko:1991:SAL

[RSS91] P. Racsko, L. Szeidl, and M. Semenov. A serial approach to local stochastic weather models. Ecological Modelling, 57(1–2):27–41, October 1, 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380091900534.

Raulier:1997:ISR

[RU97] Frédéric Raulier and Chhun-Huor Ung. Influence of shading on the relationship between leaf area and crown surface area in sugar maple stands. Ecological Modelling, 104(1):51–69, December 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438009700118X.

Rutherford:1993:EPP

[Rut93] Michael C. Rutherford. Empiricism and the prediction of primary production at the mesoscale: a savanna example. Ecological Modelling, 67(2–4):129–146, June 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390002A.

Ruxton:1995:FFN

[Rux95] G. D. Ruxton. Foraging in flocks: non-spatial models may neglect important costs. Ecological Modelling, 82(3):277–285, November 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094000983.

Ruxton:1996:EST

[Rux96] Graeme D. Ruxton. Effects of the spatial and temporal ordering of events on the behaviour of a simple cellular automaton. Ecological Modelling, 84(1–3):311–314, January 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094001456.

Regniere:1991:SMS

[RY91] Jacques Régnière and Minsheng You. A simulation model of spruce budworm (lepidoptera: Tortricidae) feeding on balsam fir and white spruce. Ecological Modelling, 54(3–4):277–297, June 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190080K.

Ryanzhin:1994:LAI

[Rya94] Sergei V. Ryanzhin. Latitudinal-altitudinal inter-relationships for the surface temperatures of the Northern Hemisphere freshwater lakes. Ecological Modelling, 74(3–4):231–253, August 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009490121X.

Rykiel:1996:TEM

[Ryk96] Edward J. Rykiel. Testing ecological models: the meaning of validation. Ecological Modelling, 90(3):229–244, November 1, 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095001522.

Sogn:1997:SED

[SA97] T. A. Sogn and G. Abrahamsen. Simulating effects of S and N deposition on soil water chemistry by the nutrient cycling model NuCM. Ecological Modelling, 99(2–3):101–111, June 30, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438009601945X.

Storey:1993:SSD

[SABG93] Michell L. Storey, Martin T. Auer, Anita K. Barth, and James M. Graham. Site-specific determination of kinetic coefficients for modeling algal growth. Ecological Modelling, 66(3–4):181–196, April 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380093901126.

Sheppard:1995:CGN

[SADZ95] Marsha I. Sheppard, B. D. Amiro, P. A. Davis, and R. Zach. Continental glaciation and nuclear fuel waste disposal: Canada’s approach and assessment of the impact on nuclide transport through the biosphere. Ecological Modelling, 78(3):249–265, April 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400181G.

Salomonsen:1992:EPE

[Sal92a] Jørgen Salomonsen. Examination of properties of exergy, power and ascendency along a eutrophication gradient. Ecological Modelling, 62(1–3):171–181, July 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290089W.

Salski:1992:FKB

[Sal92b] A. Salski. Fuzzy knowledge-based models in ecological research. Ecological Modelling, 63(1–4):103–112, September 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009290064L.

Salencon:1997:STD

[Sal97] Marie-José Salençon. Study of the thermal dynamics of two dammed lakes (Pareloup and Rochebut, France), using the EOLE model. Ecological Modelling, 104(1):15–38, December 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001051.

Samanta:1996:IEN

[Sam96] G. P. Samanta. Influence of environmental noises on the gomatam model of interacting species. Ecological Modelling, 91(1–3):283–291, November 15, 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001956.

Smith:1997:STO

[SB97] G. C. Smith and D. S. Bull. Spatial and temporal ordering of events in discrete time cellular automata — an overview. Ecological Modelling, 96(1–3):305–307, March 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096000531.

Scheffer:1995:SIS

[SBD^{+}95] M. Scheffer, J. M. Baveco, D. L. DeAngelis, K. A. Rose, and E. H. van Nes. Super-individuals a simple solution for modelling large populations on an individual basis. Ecological Modelling, 80(2–3):161–170, July 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009400055M.

Solidoro:1997:LTS

[SBD^{+}97] C. Solidoro, V. E. Brando, C. Dejak, D. Franco, R. Pastres, and G. Pecenik. Long term simulations of population dynamics of Ulva r. in the lagoon of Venice. Ecological Modelling, 102(2–3):259–272, October 29, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097000604.

Soudant:1997:DLB

[SBT97] D. Soudant, B. Beliaeff, and G. Thomas. Dynamic linear Bayesian models in phytoplankton ecology. Ecological Modelling, 99(2–3):161–169, June 30, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097019492.

Schleiter:1999:MWQ

[SBW^{+}99] Ingrid M. Schleiter, Dietrich Borchardt, Rüdiger Wagner, Thomas Dapper, Klaus-Dieter Schmidt, Hans-Heinrich Schmidt, and Heinrich Werner. Modelling water quality, bioindication and population dynamics in lotic ecosystems using neural networks. Ecological Modelling, 120(2–3):271–286, August 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001088.

Scala:1992:MMS

[SC92] C. Scala and P. Cappelini. A mathematical model to study leaf fall in apricot and peach trees. Ecological Modelling, 64(4):279–288, November 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290027C.

Sharov:1996:MTH

[SC96] Alexei A. Sharov and J. J. Colbert. A model for testing hypotheses of gypsy moth, Lymantria dispar L., population dynamics. Ecological Modelling, 84(1–3):31–51, January 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094001103.

Sequeira:1994:IPS

[SCEZ^{+}94] Ronaldo Antonio Sequeira, Mark Cochran, Kamal M. El-Zik, Nicholas D. Stone, and Merry E. Makela. Inclusion of plant structure and fiber quality into a distributed delay cotton model to improve management and optimize profit. Ecological Modelling, 71(1–3):161–186, January 1994. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094900809.

Seligman:1992:SDE

[SCH92] No’am G. Seligman, Juan B. Cavagnaro, and Manuel E. Horno. Simulation of defoliation effects on primary production of a warm-season, semiarid perennial-species grassland. Ecological Modelling, 60(1):45–61, January 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380092900124.

Schall:1995:AMP

[Sch95] Peter Schall. Application of the model FIWALD to the Solling spruce site. Ecological Modelling, 83(1–2):151–162, December 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009500094C.

Schenk:1996:MET

[Sch96a] H. Jochen Schenk. Modeling the effects of temperature on growth and persistence of tree species: a critical review of tree population models. Ecological Modelling, 92(1):1–32, November 21, 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009500212X.

Scherm:1996:VEW

[Sch96b] H. Scherm. On the velocity of epidemic waves in model plant disease epidemics. Ecological Modelling, 87(1–3):217– 222, June 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095000305.

Stage:1993:AAI

[SCM93] Albert R. Stage, Nicholas L. Crookston, and Robert A. Monserud. An aggregation algorithm for increasing the efficiency of population models. Ecological Modelling, 68(3–4):257–271, August 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390021J.

Shukla:1996:ECHa

[SD96] J. B. Shukla and B. Dubey. Effect of changing habitat on species: application to Keoladeo National Park, India. Ecological Modelling, 86(1):91–99, April 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094001944.

Semadeni-Davies:1997:MSM

[SD97] Annette Semádeni-Davies. Monthly snowmelt modelling for large-scale climate change studies using the degree day approach. Ecological Modelling, 101(2–3):303–323, August 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097000549.

Stankovski:1998:MPD

[SDBA98] Vlado Stankovski, Marko Debeljak, Ivan Bratko, and Miha Adamič. Modelling the population dynamics of red deer (Cervus elaphus L.) with regard to forest development. Ecological Modelling, 108(1–3):145–153, May 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000258.

Shukla:1996:ECHb

[SDF96] J. B. Shukla, B. Dubey, and H. I. Freedman. Effect of changing habitat on survival of species. Ecological Modelling, 87(1–3):205–216, June 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095000291.

Shukla:1996:EED

[SDS96] A. Shukla, B. Dubey, and J. B. Shukla. Effect of environmentally degraded soil on crop yield: the role of conservation. Ecological Modelling, 86(2–3):235–239, May 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095000577.

Stuefer:1998:MOR

[SDS98] Josef F. Stuefer, Heinjo J. During, and Feike Schieving. A model on optimal root–shoot allocation and water transport in clonal plants. Ecological Modelling, 111(2–3):171–186, September 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438009800101X.

Smith:1993:CMP

[SEJ93] Lynn C. Smith, D. J. Elliot, and A. James. Characterisation of mixing patterns in an anaerobic digester by means of tracer curve analysis. Ecological Modelling, 69(3–4):267–285, October 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390030V.

Selhorst:1995:MTE

[Sel95] Th. Selhorst. Modelling the temperature effect on bimodally distributed life times. Ecological Modelling, 80(1):27–34, June 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400043H.

Semovski:1999:BSL

[Sem99] S. V. Semovski. The Baltic Sea and Lake Baikal underwater bio-optical fields simulation using ecodynamical model. Ecological Modelling, 116(2–3):149–163, March 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001719.

Seno:1990:DDD

[Sen90] Hiromi Seno. A density-dependent diffusion model of shoaling of nesting fish. Ecological Modelling, 51(3–4):217–226, June 1990. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009090066P.

Seppelt:1999:AOC

[Sep99] Ralf Seppelt. Applications of optimum control theory to agroecosystem modelling. Ecological Modelling, 121(2–3): 161–183, September 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380099000824.

Seymour:1992:SIV

[Sey92] R. M. Seymour. A study of the interaction of virulence, resistance and resource limitation in a model of myxomatosis mediated by the European rabbit flea Spilopsyllus cuniculi (Dale). Ecological Modelling, 60(3–4):281–308, April 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009290037F.

Stefan:1994:DOM

[SF94] Heinz G. Stefan and Xing Fang. Dissolved oxygen model for regional lake analysis. Ecological Modelling, 71(1–3):37–68, January 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094900752.

Stoker:1994:SCE

[SFGF94] Revin L. Stoker, David K. Ferris, William E. Grant, and L. Joseph Folse. Simulating colonization by exotic species: a model of the red imported fire ant (Solenopsis invicta) in North America. Ecological Modelling, 73(3–4):281–292, June 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094900663.

Salomonsen:1997:SAT

[SFGH97] Jørgen Salomonsen, Mogens Rene Flindt, and Ole Geertz-Hansen. Significance of advective transport of Ulva lactuca for a biomass budget on a shallow water location. Ecological Modelling, 102(1):129–132, October 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001002.

Salski:1996:I

[SFK96] A. Salski, O. Fränzle, and P. Kandzia. Introduction. Ecological Modelling, 85(1):1–2, February 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380096900538.

Sparrow:1997:LSM

[SFS97] Ashley D. Sparrow, Margaret H. Friedel, and D. Mark Stafford Smith. A landscape-scale model of shrub and herbage dynamics in Central Australia, validated by satellite data. Ecological Modelling, 97(3):197–216, April 28, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096019047.

Schirripa:1997:SAA

[SG97] Michael J. Schirripa and C. Phillip Goodyear. Simulation of alternative assumptions of fish otolith–somatic growth with a bioenergetics model. Ecological Modelling, 102(2–3):209–223, October 29, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097000574.

Schreiber:1998:SDP

[SG98] Sebastian J. Schreiber and Andrew P. Gutierrez. A supply/demand perspective of species invasions and coexistence: applications to biological control. Ecological Modelling, 106(1):27–45, February 16, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097001786.

Schnase:1991:TEB

[SGML91] John L. Schnase, William E. Grant, Terry C. Maxwell, and John J. Leggett. Time and energy budgets of Cassin’s sparrow (Aimophila cassinii) during the breeding season: evaluation through modelling. Ecological Modelling, 55(3–4):285–319, August 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190091E.

Santha:1991:BCA

[SGNS91] Calista Rohini Santha, W. E. Grant, William H. Neill, and R. Kirk Strawn. Biological control of aquatic vegetation using grass carp: simulation of alternative strategies. Ecological Modelling, 59(3–4):229–245, December 15, 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380091901795.

Suzuki:1993:SDM

[SGS93] Motoyuki Suzuki, Naohiro Goto, and Akiyoshi Sakoda. Simplified dynamic model on carbon exchange between atmosphere and terrestrial ecosystems. Ecological Modelling, 70(3–4):161–194, December 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009390056X.

Silow:1995:MML

[SGS^{+}95] E. A. Silow, V. J. Gurman, D. J. Stom, D. M. Rosenraukh, and V. I. Baturin. Mathematical models of Lake Baikal ecosystem. Ecological Modelling, 82(1):27–39, September 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400078V.

Sheppard:1991:LSD

[SH91] Marsha I. Sheppard and J. L. Hawkins. A linear sorption/ dynamic water flow model applied to the results of a fouryear soil core study. Ecological Modelling, 55(3–4):175–201, August 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190086G.

Schulman:1992:DSS

[SH92] Lloyd L. Schulman and Steven R. Hanna. A decision system for selecting a site-specific air quality dispersion model. Ecological Modelling, 64(2–3):205–219, October 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290115U.

Schladow:1997:PWQ

[SH97] S. Geoffrey Schladow and David P. Hamilton. Prediction of water quality in lakes and reservoirs: Part II — model calibration, sensitivity analysis and application. Ecological Modelling, 96(1–3):111–123, March 1, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096000634.

Scardi:1999:DEM

[SH99] Michele Scardi and Lawrence W. Harding. Developing an empirical model of phytoplankton primary production: a neural network case study. Ecological Modelling, 120(2–3):213–223, August 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001039.

Stefan:1995:VFH

[SHEM95] Heinz G. Stefan, Midhat Hondzo, John G. Eaton, and J. Howard McCormick. Validation of a fish habitat model for lakes. Ecological Modelling, 82(3):211–224, November 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094000994.

Svendsen:1995:SCP

[SHJ95] H. Svendsen, S. Hansen, and H. E. Jensen. Simulation of crop production, water and nitrogen balances in two German agro-ecosystems using the DAISY model. Ecological Modelling, 81(1–3):197–212, August 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400171D.

Shukla:1998:MDW

[Shu98] V. P. Shukla. Modelling the dynamics of wetland macrophytes: Keoladeo National Park wetland, India. Ecological Modelling, 109(1):99–114, June 1, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438009800012X.

Silvert:1993:DDA

[Sil93a] William Silvert. The distributed derivative: an aid to modular modelling. Ecological Modelling, 68(3–4):293–302, August 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390023L.

Silvert:1993:OOE

[Sil93b] William Silvert. Object-oriented ecosystem modelling. Ecological Modelling, 68(1–2):91–118, July 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009390110E.

Silow:1995:CEP

[Sil95a] Eugene A. Silow. Complex ecology: the part-whole relation in ecosystem: Bernard C. Patten and Sven E. Jørgensen (Editors), with a foreword by Stanley I. Auerbach. Prentice Hall, Englewood Cliffs, NJ, 705 pp., ISBN 0-13-161506-8. Ecological Modelling, 80(2–3):300–301, July 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095900438.

Silvert:1995:LEL

[Sil95b] William Silvert. Is the logistic equation a Lotka–Volterra model? Ecological Modelling, 77(2–3):95–96, February 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009400154A.

Silvert:1997:EIC

[Sil97] William Silvert. Ecological impact classification with fuzzy sets. Ecological Modelling, 96(1–3):1–10, March 1, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096000518.

Silow:1999:UTL

[Sil99] Eugene A. Silow. The use of two lumped models for the analysis of consequences of external influences on the Lake Baikal ecosystem. Ecological Modelling, 121(2–3):103–113, September 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098001690.

Sekine:1997:MFD

[SIU97] Masahiko Sekine, Tsuyoshi Imai, and Masao Ukita. A model of fish distribution in rivers according to their preference for environmental factors. Ecological Modelling, 104(2–3):215–230, December 8, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001154.

Salomonsen:1996:ULM

[SJ96] Jørgen Salomonsen and Jan Juul Jensen. Use of a lake model to examine exergy response to changes in phytoplankton growth parameters and species composition. Ecological Modelling, 87(1–3):41–49, June 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094002037.

Straskraba:1999:EED

[SJP99] Milan Straškraba, Sven E. Jørgensen, and Bernard C. Patten. Ecosystems emerging: 2. Dissipation. Ecological Modelling, 117(1):3–39, April 1, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438009800194X.

Somlyody:1991:ISR

[SK91] L. Somlyódy and L. Koncsos. Influence of sediment resuspension on the light conditions and algal growth in Lake Balaton. Ecological Modelling, 57(3–4):173–192, October 15, 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009190112E.

Selhorst:1995:MET

[SKŞ95] Th. Selhorst, B. Kleinhenz, and Ç. Şengonca. Modelling the effect of temperature on the longevity of the grain aphid, Sitobion avenae (F.) (Hom., Aphididae). Ecological Modelling, 80(2–3):197–203, July 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400070X.

Summers:1991:WQM

[SKW91] J. Kevin Summers, Paul F. Kazyak, and Stephen B. Weisberg. A water quality model for a river receiving paper mill effluents and conventional sewage. Ecological Modelling, 58(1–4):25–54, November 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190029Z.

Sun:1991:CRM

[SL91a] Ping Sun and Changshan Lin. Classified and representative method — a useful way to select variables in ecology. Ecological Modelling, 55(1–2):123–131, July 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009190070H.

Sun:1991:CGT

[SL91b] Ping Sun and Changshan Lin. Composite growth type of population growth. Ecological Modelling, 55(1–2):133–139, July 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380091900718.

Swart:1996:EHP

[SL96] Johan Swart and M. J. Lawes. The effect of habitat patch connectivity on samango monkey (Cercopithecus mitis) metapopulation persistence. Ecological Modelling, 93(1–3):57–74, December 16, 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095002111.

Swart:1993:MMI

[SLP93] J. Swart, M. J. Lawes, and M. R. Perrin. A mathematical model to investigate the demographic viability of low-density samango monkey (Cercopithecus mitis) populations in Natal, South Africa. Ecological Modelling, 70(3–4):289–303, December 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390062W.

Schultz:1995:SSW

[SM95] A. Schultz and W. Mirschel. Simulating soil water balance, nitrogen behaviour and biomass components, using the agroecosystem model AGROSIM-winter-wheat and data from the north German krummbach catchment. Ecological Modelling, 81(1–3):133–144, August 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400166F.

Sterba:1997:AFS

[SM97] Hubert Sterba and Robert A. Monserud. Applicability of the forest stand growth simulator prognaus for the Austrian part of the Bohemian massif. Ecological Modelling, 98(1):23– 34, May 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096019345.

Selhorst:1999:EER

[SM99] T. Selhorst and T. Müller. An evaluation of the efficiency of rabies control strategies in fox (Vulpes vulpes) populations using a computer simulation program. Ecological Modelling, 124(2–3):221–232, December 13, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099001702.

Shui-Ming:1993:KMD

[SMDSZSXF93] Zhang Shui-Ming, Liu Duo-Sen, Wang Zong-Sheng, and Ma Xing-Fa. A kinetic model describing the effect of temperature on the loss rate of pesticides in soil. Ecological Modelling, 70(1–2):115–125, November 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093900754.

Smith:1991:UMP

[Smi91] Mark Smith. Using massively-parallel supercomputers to model stochastic spatial predator–prey systems. Ecological Modelling, 58(1–4):347–367, November 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380091900453.

Smith:1995:OSW

[Smi95a] Roger E. Smith. Opus simulation of a wheat/sugarbeet plot near Neuenkirchen, Germany. Ecological Modelling, 81 (1–3):121–132, August 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438009400165E.

Smith:1995:UES

[Smi95b] Roger E. Smith. Unbias the evaluation of soil hysteresis effects on transport. Ecological Modelling, 82(2):205–206, October 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009500133G.

Smith:1997:AFD

[Smi97a] G. C. Smith. An analysis of the form of density dependence in a simulation model of a seasonal breeder undergoing control. Ecological Modelling, 95(2–3):181–189, February 14, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096000385.

Smith:1997:ECP

[Smi97b] Jeff P. Smith. An energy-circuit population model for Great Egrets (Ardea alba) at Lake Okeechobee, Florida, U.S.A. Ecological Modelling, 97(1–2):1–21, April 15, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096000610.

Saravia:1998:MPD

[SML98] Leonardo A. Saravia, Fernando Momo, and Lucila D. Boffi Lissin. Modelling periphyton dynamics in running water. Ecological Modelling, 114(1):35–47, December 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001136.

Sondgerath:1996:MDC

[SMP96] Dagmar Söndgerath and Wolfgang Müller-Pietralla. A model for the development of the cabbage root fly (Delia radicum L.) based on the extended Leslie model. Ecological Modelling, 91(1–3):67–76, November 15, 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001670.

Sherratt:1998:MAE

[SMWF98] T. N. Sherratt, A. D. Macdougall, S. D. Wratten, and A. B. Forbes. Models to assist the evaluation of the impact of avermectins on dung insect populations. Ecological Modelling, 110(2):165–173, July 11, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000647.

Soyupak:1997:EEC

[SMY^{+}97] S. Soyupak, L. Mukhallalati, D. Yemişen, A. Bayar, and C. Yurteri. Evaluation of eutrophication control strategies for the Keban Dam reservoir. Ecological Modelling, 97 (1–2):99–110, April 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380096000774.

Seno:1995:MAF

[SN95] Hiromi Seno and Katsuki Nakai. Mathematical analysis on fish shoaling by a density-dependent diffusion model. Ecological Modelling, 79(1–3):149–157, May 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093E0143Q.

Seno:1999:TMM

[SN99a] Hiromi Seno and Hisao Nakajima. Transition matrix model for persistence of monocarpic perennial plant population under periodically ecological disturbance. Ecological Modelling, 117(1):65–80, April 1, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099000101.

Shiklomanov:1999:ARA

[SN99b] N. I. Shiklomanov and F. E. Nelson. Analytic representation of the active layer thickness field, Kuparuk River Basin, Alaska. Ecological Modelling, 123(2–3):105–125, November 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099001271.

Sekine:1996:SFM

[SNU96] M. Sekine, H. Nakanishi, and M. Ukita. Study on fish mortality caused by the combined effects of pesticides and changes in environmental conditions. Ecological Modelling, 86(2–3):259– 264, May 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095000615.

Sekine:1991:SSE

[SNUM91] Masahiko Sekine, Hiroshi Nakanishi, Masao Ukita, and Sadaaki Murakami. A shallow-sea ecological model using an object-oriented programming language. Ecological Modelling, 57(3–4):221–236, October 15, 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009190114G.

Sogn:1993:TCE

[Sog93] T. A. Sogn. A test of chemical equilibrium equations and assumptions commonly used in soil-oriented charge balance models for soil and freshwater acidification. Ecological Modelling, 70(3–4):221–238, December 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390058Z.

Sogn:1994:AMM

[Sog94] Trine A. Sogn. Application of the MAGIC model to lysimeters with Cambic Arenosol, Nordmoen, Norway. Ecological Modelling, 71(1–3):1–16, January 1994. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094900736.

Solovjova:1999:SEE

[Sol99] N. V. Solovjova. Synthesis of ecosystemic and ecoscreening modelling in solving problems of ecological safety. Ecological Modelling, 124(1):1–10, December 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099001222.

Sequeira:1997:IGO

[SOM97] R. A. Sequeira, R. L. Olson, and J. M. McKinion. Implementing generic, object-oriented models in biology. Ecological Modelling, 94(1):17–31, January 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096019254.

South:1999:EIM

[Sou99] Andy South. Extrapolating from individual movement behaviour to population spacing patterns in a ranging mammal. Ecological Modelling, 117(2–3):343–360, May 17, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000150.

Spencer:1994:EIT

[Spe94] David F. Spencer. Estimating the impact of triploid grass carp on sago pondweed in the Byrnes Canal: Implications for biological control in northern California irrigation systems. Ecological Modelling, 72(3–4):187–204, April 1994. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094900833.

Schaaf:1995:CFD

[SPE95] Th. Schaaf, E. Priesack, and Th. Engel. Comparing field data from north Germany with simulations of the nitrogen model N-SIM. Ecological Modelling, 81(1–3):223–232, August 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400173F.

Spencer:1997:EHS

[Spe97] Matthew Spencer. The effects of habitat size and energy on food web structure: an individual-based cellular automata model. Ecological Modelling, 94(2–3):299–316, January 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096000269.

Skirvin:1997:MDP

[SPH97] D. J. Skirvin, J. N. Perry, and R. Harrington. A model describing the population dynamics of Sitobion avenae and Coccinella septempunctata. Ecological Modelling, 96(1–3):29–39, March 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096000488.

Solidoro:1997:MMU

[SPP^{+}97] C. Solidoro, G. Pecenik, R. Pastres, Davide Franco, and C. Dejak. Modelling macroalgae (Ulva rigida) in the Venice Lagoon: Model structure identification and first parameters estimation. Ecological Modelling, 94(2–3):191–206, January 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096000257.

Svirezhev:1992:MRC

[SR92] Ju. M. Svirezhev and P. Racsko. Minimum risk criterion for optimal allocation of crops. Ecological Modelling, 61(3–4):217– 225, June 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290019B.

Sarkar:1993:OBR

[SR93] A. K. Sarkar and A. B. Roy. Oscillatory behaviour in a resource-based plant-herbivore model with random herbivore attack. Ecological Modelling, 68(3–4):213–226, August 1993. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009390018N.

Spivak:1994:MSM

[SR94] I. R. Spivak and J. S. Rokem. Mathematical simulations of man-made microbial mixture grown on natural gas. Ecological Modelling, 74(3–4):287–304, August 1994. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094901244.

Smith:1995:MGF

[SR95a] Eric P. Smith and Kenneth A. Rose. Model goodness-of-fit analysis using regression and related techniques. Ecological Modelling, 77(1):49–64, January 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380093E0074D.

Spivak:1995:APS

[SR95b] I. R. Spivak and J. S. Rokem. Analyses of possible steady states for mixed culture grown on natural gas. Ecological Modelling, 80(2–3):257–278, July 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400056N.

Skiles:1998:SWG

[SR98] J. W. Skiles and C. W. Richardson. A stochastic weather generation model for Alaska. Ecological Modelling, 110(3):211–232, July 22, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098000611.

Sands:1999:USP

[SRB99] P. J. Sands, W. Rawlins, and M. Battaglia. Use of a simple plantation productivity model to study the profitability of irrigated Eucalyptus globulus. Ecological Modelling, 117(1):125–141, April 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380099000216.

Schroder:1995:PPG

[SRV95] U. Schröder, O. Richter, and K. Velten. Performance of the plant growth models of the special collaborative project 179 with respect to winter wheat. Ecological Modelling, 81(1–3):243–250, August 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400175H.

Steinberg:1997:CPM

[SRW97] Laura J. Steinberg, Kenneth H. Reckhow, and Robert L. Wolpert. Characterization of parameters in mechanistic models: a case study of a PCB fate and transport model. Ecological Modelling, 97(1–2):35–46, April 15, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096000658.

Schroder:1996:CBT

[SS96] U. Schröder and D. Söndgerath. The concept of biological time for computing the switching points of a growth model for winter wheat. Ecological Modelling, 88(1–3):1–8, July 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095000321.

Skop:1998:GBM

[SS98] Eli Skop and Peter B. Sørensen. GIS-based modelling of solute fluxes at the catchment scale: a case study of the agricultural contribution to the riverine nitrogen loading in the Vejle Fjord catchment, Denmark. Ecological Modelling, 106(2–3):291–310, March 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097002056.

Svirezhev:1997:DB

[SSH97] Yu. Svirezhev and A. Svirejeva-Hopkins. Diversity of the biosphere. Ecological Modelling, 97(1–2):145–146, April 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096000798.

Svirezhev:1998:SBC

[SSH98] Yu. M. Svirezhev and A. Svirejeva-Hopkins. Sustainable biosphere: critical overview of basic concept of sustainability. Ecological Modelling, 106(1):47–61, February 16, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001774.

Streit:1991:MVE

[SSK^{+}91] Bruno Streit, Ernst-Olof Siré, Gundolf H. Kohlmaier, Franz W. Badeck, and Stephan Winter. Modelling ventilation efficiency of teleost fish gills for pollutants with high affinity to plasma proteins. Ecological Modelling, 57(3–4):237–262, October 15, 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009190115H.

Sequeira:1993:GMV

[SSM^{+}93] Ronaldo Antonio Sequeira, Nicholas D. Stone, Merry E. Makela, Kamal M. El-Zik, and Peter J. H. Sharpe. Generation of mechanistic variability in a process-based objectoriented plant model. Ecological Modelling, 67(2–4):285–306, June 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390010P.

Sequeira:1991:OOS

[SSS^{+}91] R. A. Sequeira, P. J. H. Sharpe, N. D. Stone, K. M. El-Zik, and M. E. Makela. Object-oriented simulation: plant growth and discrete organ to organ interactions. Ecological Modelling, 58(1–4):55–89, November 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380091900305.

Selhorst:1991:MDP

[SSW91] Thomas Selhorst, Dagmar Söndgerath, and Susanne Weigand. A model describing the predator–prey interaction between Scolothrips longicornis and Tetranychus cinnabarinus based upon the Leslie theory. Ecological Modelling, 59(1–2):123– 138, December 1, 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030438009190130S.

Shao:1995:STS

[SSY95] Guofan Shao, Herman H. Shugart, and Donald R. Young. Simulation of transpiration sensitivity to environmental changes for shrub (Myrica cerifera) thickets on a Virginia barrier island. Ecological Modelling, 78(3):235–248, April 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093E0091G.

Sterbacek:1990:CLE

[ŠŠZ90] Zdeněk Štěrbáček, Václav Škopek, and Vladimír Zavázal. A composite landscape ecology prognostic expert system — colepes. Part I. System philosophy and design. Ecological Modelling, 50(1–3):145–156, March 1990. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009090047K.

Skipper:1992:CMP

[ST92] Donna Skipper and Marty Tittlebaum. A computer model for predicting effluent BOD concentrations from a rock-plant filter. Ecological Modelling, 64(4):289–304, November 1992. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290028D.

Salencon:1996:SMM

[ST96] Marie-José Salençon and Jean-Marc Thébault. Simulation model of a mesotrophic reservoir (Lac de Pareloup, France): melodia, an ecosystem reservoir management model. Ecological Modelling, 84(1–3):163–187, January 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094001413.

Shiyomi:1999:MSP

[ST99] Masae Shiyomi and Mikinori Tsuiki. Model for the spatial pattern formed by a small herd in grazing cattle. Ecological Modelling, 119(2–3):231–238, July 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000599.

Stephenson:1993:BEB

[Ste93] D. Stephenson. Borrowed earth, borrowed time: G. E. Schweitzer. 1991, 350 pp. US $27.50. 0-306-43766-X. Ecological Modelling, 65(1–2):150–151, January 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390133D.

Steel:1995:MAP

[Ste95a] J. A. Steel. Modelling adaptive phytoplankton in a variable environment. Ecological Modelling, 78(1–2):117–127, March 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400122X.

Stenberg:1995:PWS

[Ste95b] Pauline Stenberg. Penumbra in within-shoot and betweenshoot shading in conifers and its significance for photosynthesis. Ecological Modelling, 77(2–3):215–231, February 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380093E0086I.

Stockle:1992:CPT

[Sto92] Claudio O. Stockle. Canopy photosynthesis and transpiration estimates using radiation interception models with different levels of detail. Ecological Modelling, 60(1):31–44, January 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380092900113.

Straskraba:1991:MSW

[Str91] Milan Straškraba. Mathematical submodels in water quality systems: S. E. Jørgensen and M. J. Gromiec (Editors). Development in Environmental Modeling, 14. Elsevier, Amsterdam, 1989. 408 pp., Dfl. 240.00. ISBN 0-444-88030-5. Ecological Modelling, 59(1–2):143–146, December 1, 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009190134M.

Straskraba:1992:TAE

[Str92] M. Straškraba. Terrestrial and aquatic ecosystem. Perturbation and recovery: Oscar Ravera (Editor). Ellis Horwood, New York, 1991, 613 pp. ISBN 0-13-928201-7. Price: £159.95. Ecological Modelling, 64(1):77–79, October 15, 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290052G.

Straskraba:1994:MRE

[Str94a] M. Straškraba. Modelling reservoir ecosystems. Ecological Modelling, 74(1–2):vi, July 1994. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094901074.

Straskraba:1994:EMR

[Str94b] Milan Straškraba. Ecotechnological models for reservoir water quality management. Ecological Modelling, 74(1–2):1–38, July 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094901082.

Straskraba:1995:MMM

[Str95] Milan Straškraba. Messages from the meeting “Mathematical Modelling in Limnology” Innsbruck. Ecological Modelling, 78(1–2):3–5, March 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400112U.

Stroosnijder:1996:MEG

[Str96] Leo Stroosnijder. Modelling the effect of grazing on infiltration, runoff and primary production in the Sahel. Ecological Modelling, 92(1):79–88, November 21, 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001972.

Sawyer:1993:GML

[STTR93] A. J. Sawyer, M. J. Tauber, C. A. Tauber, and J. R. Ruberson. Gypsy moth (Lepidoptera: Lymantriidae) egg development: a simulation analysis of laboratory and field data. Ecological Modelling, 66(1–2):121–155, March 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390043R.

Sturtevant:1998:MWV

[Stu98] Brian R. Sturtevant. A model of wetland vegetation dynamics in simulated beaver impoundments. Ecological Modelling, 112(2–3):195–225, October 15, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098000799.

Suarez:1999:MCE

[Suá99] Inti Suárez. Mastering chaos in ecology. Ecological Modelling, 117(2–3):305–314, May 17, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099000071.

Sands:1996:FBE

[SV96] P. J. Sands and E. O. Voit. Flux-based estimation of parameters in S-systems. Ecological Modelling, 93(1–3):75–88, December 16, 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095002154.

Svirezhev:1998:CIM

[SV98] Yu. M. Svirezhev and S. V. Venevsky. Constraints on information measures of embodied solar energy fluxes in the biosphere. Ecological Modelling, 105(2–3):227–234, January 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438009700166X.

Scherm:1993:RSM

[SvB93] H. Scherm and A. H. C. van Bruggen. Response surface models for germination and infection of Bremia lactucae, the fungus causing downy mildew of lettuce. Ecological Modelling, 65(3–4):281–296, February 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093900846.

Svirezhev:1996:MMI

[SvB96] Yuri M. Svirezhev and Werner von Bloh. A minimal model of interaction between climate and vegetation: qualitative approach. Ecological Modelling, 92(1):89–99, November 21, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095001980.

Svirezhev:1997:CVG

[SvB97] Yuri M. Svirezhev and Werner von Bloh. Climate, vegetation, and global carbon cycle: the simplest zero-dimensional model. Ecological Modelling, 101(1):79–95, August 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009701973X.

Svirezhev:1998:ZDC

[SvB98] Yuri M. Svirezhev and Werner von Bloh. A zero-dimensional climate-vegetation model containing global carbon and hydrological cycle. Ecological Modelling, 106(2–3):119–127, March 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097001877.

Supit:1999:NWY

[SvdG99] I. Supit and E. van der Goot. National wheat yield prediction of France as affected by the prediction level. Ecological Modelling, 116(2–3):203–223, March 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001756.

Svirezhev:1991:ASM

[Svi91a] Y. M. Svirezhev. Applied simulation of marine communities with reference to the fouling community: Mikhail Vladimirovich Gal’perin. Akademie-Verlag, Berlin, 1990. 205 pp., DM42.00. ISBN 3-05-500761-1. Ecological Modelling, 59(1–2):141–142, December 1, 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009190133L.

Svirezhev:1991:AME

[Svi91b] Yu. M. Svirezhev. Applied mathematical ecology. Simon A. Levin, Thomas G. Hallam, Louis J. Gross (Editors). Biomathematics, Vol. 18. Springer, Berlin, 1989, 505 pp. ISBN 3-540-19465-7. Ecological Modelling, 57(3–4):295–297, October 15, 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009190118K.

Svirezhev:1997:SGP

[Svi97a] Yu. M. Svirezhev. On some general properties of trophic networks. Ecological Modelling, 99(1):7–17, June 16, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096019187.

Svirezhev:1997:EB

[Svi97b] Yuri Svirezhev. Exergy of the Biosphere. Ecological Modelling, 96(1–3):309–310, March 1, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380096000543.

Svirezhev:1999:SDM

[Svi99] Yuri M. Svirezhev. Simplest dynamic models of the global vegetation pattern. Ecological Modelling, 124(2–3):131–144, December 13, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001581.

Strandman:1993:PGS

[SVK93] Harri Strandman, Hannu Väisänen, and Seppo Kellomäki. A procedure for generating synthetic weather records in conjunction of climatic scenario for modelling of ecological impacts of changing climate in boreal conditions. Ecological Modelling, 70(3–4):195–220, December 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390057Y.

Stephan:1999:ERP

[SW99] Thomas Stephan and Christian Wissel. The extinction risk of a population exploiting a resource. Ecological Modelling, 115(2–3):217–225, February 15, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098001811.

Swartzman:1996:RMM

[Swa96] Gordon Swartzman. Resource modeling moves into the courtroom. Ecological Modelling, 92(2–3):277–288, December 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095001794.

Summers:1993:MQP

[SWK93] J. Kevin Summers, Harold T. Wilson, and Jingyee Kou. A method for quantifying the prediction uncertainties associated with water quality models. Ecological Modelling, 65(3–4):161–176, February 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380093900787.

Smith:1990:LSS

[SY90] S. M. Smith and M. You. A life system simulation model for improving inundative releases of the egg parasite, Trichogramma minutum against the spruce budworm. Ecological Modelling, 51(1–2):123–142, May 1990. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009090061K.

Sun:1999:DPC

[SY99] P. Sun and X. B. Yang. Deterministic property changes in population models under random error perturbations. Ecological Modelling, 119(2–3):239–247, July 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000526.

Szyper:1996:OMP

[Szy96] James P. Szyper. Observations and model predictions of daily areal primary production in a eutrophic brackish water culture pond. Ecological Modelling, 88(1–3):83–92, July 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095000712.

Tyutyunov:1993:MFO

[TAB^{+}93] Yuri Tyutyunov, Roger Arditi, Bernard Büttiker, Yuri Dombrovsky, and Erich Staub. Modelling fluctuations and optimal harvesting in perch populations. Ecological Modelling, 69(1–2):19–42, September 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009390046U.

Tourenq:1999:UAN

[TAM^{+}99] Christophe Tourenq, Stéphane Aulagnier, François Mesléard, Laurent Durieux, Alan Johnson, Georges Gonzalez, and Sovan Lek. Use of artificial neural networks for predicting rice crop damage by greater flamingos in the Camargue, France. Ecological Modelling, 120(2–3):349–358, August 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099001143.

Taplin:1993:SVF

[Tap93] Ross H. Taplin. Sources of variation for fire spread rate in non-homogeneous fuel. Ecological Modelling, 68(3–4):205–211, August 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390017M.

Taylor:1996:FFE

[Tay96] John A. Taylor. Fossil fuel emissions required to achieve atmospheric CO_{2} stabilisation using ANU-BACE: a box-diffusion carbon cycle model. Ecological Modelling, 86(2–3):195–199, May 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095000518.

Tamo:1993:ACAa

[TB93] M. Tamò and J. Baumgärtner. Analysis of the cowpea agroecosystem in West Africa. I. A demographic model for carbon acquisition and allocation in cowpea, Vigna unguiculata (L.) Walp. Ecological Modelling, 65(1–2):95–121, January 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390128F.

Tamo:1993:ACAb

[TBG93] M. Tamò, J. Baumgärtner, and A. P. Gutierrez. Analysis of the cowpea agro-ecosystem in West Africa. II. Modelling the interactions between cowpea and the bean flower thrips Megalurothrips sjostedti (Trybom) (Thysanoptera, Thripidae). Ecological Modelling, 70(1–2):89–113, November 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093900743.

Tiezzi:1996:ECS

[TBM96] Enzo Tiezzi, Simone Bastianoni, and Nadia Marchettini. Environmental cost and steady state: the problem of adiabaticity in the emergy value. Ecological Modelling, 90(1):33–37, September 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095001417.

Tiktak:1995:SDS

[TBvH95] Aaldrik Tiktak, Michael Bredemeier, and Kees van Heerden. The Solling dataset. Site characteristics, monitoring data and deposition scenarios. Ecological Modelling, 83(1–2):17–34, December 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009500080F.

Tapaswi:1996:GSR

[TC96] P. K. Tapaswi and J. Chattopadhyay. Global stability results of a “susceptible-infective-immune-susceptible” (SIRS) epidemic model. Ecological Modelling, 87(1–3):223–226, June 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095000313.

Tappeiner:1998:MSS

[TC98] Ulrike Tappeiner and Alexander Cernusca. Model simulation of spatial distribution of photosynthesis in structurally differing plant communities in the Central Caucasus. Ecological Modelling, 113(1–3):201–223, November 2, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001446.

Trebitz:1997:MBL

[TCC^{+}97] Anett Trebitz, Stephen Carpenter, Paul Cunningham, Brett Johnson, Richard Lillie, David Marshall, Thomas Martin, Richard Narf, Thomas Pellett, Scott Stewart, Christine Storlie, and Jean Unmuth. A model of bluegill-largemouth bass interactions in relation to aquatic vegetation and its management. Ecological Modelling, 94(2–3):139–156, January 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096000099.

Tiwari:1994:KFM

[TD94] Ram C. Tiwari and Timothy P. Dienes. The Kalman filter model and Bayesian outlier detection for time series analysis of BOD data. Ecological Modelling, 73(1–2):159–165, May 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009490104X.

Todorovski:1998:MPP

[TDK98] Ljupčo Todorovski, Sašo Džeroski, and Boris Kompare. Modelling and prediction of phytoplankton growth with equation discovery. Ecological Modelling, 113(1–3):71–81, November 2, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098001355.

Turner:1996:CAS

[TDM96] David P. Turner, Rusty Dodson, and Danny Marks. Comparison of alternative spatial resolutions in the application of a spatially distributed biogeochemical model over complex terrain. Ecological Modelling, 90(1):53–67, September 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095001433.

Teramoto:1993:DSE

[Ter93] Ei Teramoto. Dynamical structure of energy trophic levels. Ecological Modelling, 69(1–2):135–147, September 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390053U.

Tapaswi:1995:TJE

[TGM95] P. K. Tapaswi, Asit K. Ghosh, and B. B. Mukhopadhyay. Transmission of Japanese encephalitis in a 3-population model. Ecological Modelling, 83(3):295–309, December 15, 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009400102X.

Thulke:1999:PPS

[TGM^{+}99] H.-H. Thulke, V. Grimm, M. S. Müller, C. Staubach, L. Tischendorf, C. Wissel, and F. Jeltsch. From pattern to practice: a scaling-down strategy for spatially explicit modelling illustrated by the spread and control of rabies. Ecological Modelling, 117(2–3):179–202, May 17, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001987.

Tung:1990:SWL

[TH90] Yeou-Koung Tung and Wade E. Hathhorn. Stochastic waste load allocation. Ecological Modelling, 51(1–2):29–46, May 1990. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009090056M.

Thorpe:1997:MEV

[Tho97] G. R. Thorpe. Modelling ecosystems in ventilated conical bottomed farm grain silos. Ecological Modelling, 94(2–3):255–286, January 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096000221.

Tuma:1996:CFE

[THR96] A. Tuma, H.-D. Haasis, and O. Rentz. A comparison of fuzzy expert systems, neural networks and neuro-fuzzy approaches. Controlling energy and material flows. Ecological Modelling, 85(1):93–98, February 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095000186.

Tischendorf:1998:SEP

[TIH98] Lutz Tischendorf, Ulrich Irmler, and Rainer Hingst. A simulation experiment on the potential of hedgerows as movement corridors for forest carabids. Ecological Modelling, 106(2–3):107–118, March 1, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097001865.

Tischendorf:1997:MIM

[Tis97] Lutz Tischendorf. Modelling individual movements in heterogeneous landscapes: potentials of a new approach. Ecological Modelling, 103(1):33–42, November 1, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097000744.

Trouslard-Kerdiles:1997:CCU

[TKG97] V. Trouslard-Kerdiles and M. O. Grondona. A case of combined use of crop simulation models and general linear models. Ecological Modelling, 99(1):71–85, June 16, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096019412.

Timm:1991:SMS

[TKO91] Uwe Timm, John M. Klinck, and Akira Okubo. Self- and mutual shading and competition effect on competing algal distributions: Biological implications of the model. Ecological Modelling, 59(1–2):11–36, December 1, 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009190126L.

Tufford:1999:STH

[TM99] Daniel L. Tufford and Hank N. McKellar. Spatial and temporal hydrodynamic and water quality modeling analysis of a large reservoir on the South Carolina (USA) coastal plain. Ecological Modelling, 114(2–3):137–173, January 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098001227.

Teel:1996:SHP

[TMG96] P. D. Teel, S. L. Marin, and W. E. Grant. Simulation of host-parasite-landscape interactions: influence of season and habitat on cattle fever tick (Boophilus sp.) population dynamics. Ecological Modelling, 84(1–3):19–30, January 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094001421.

Teel:1997:SHP

[TMGS97] P. D. Teel, S. Marin, W. E. Grant, and J. W. Stuth. Simulation of host-parasite-landscape interactions: influence of season and habitat on cattle fever tick (Boophilus sp.) population dynamics in rotational grazing systems. Ecological Modelling, 97(1–2):87–97, April 15, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096000762.

Troell:1998:MOR

[TN98] Max Troell and Jon Norberg. Modelling output and retention of suspended solids in an integrated salmon–mussel culture. Ecological Modelling, 110(1):65–77, July 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000428.

Tomasino:1996:IFP

[Tom96] M. G. Tomasino. Is it feasible to predict “slime blooms” or “mucilage” in the northern Adriatic Sea? Ecological Modelling, 84(1–3):189–198, January 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094001081.

Torres:1996:SMA

[Tor96] Néstor V. Torres. S-system modelling approach to ecosystem: Application to a study of magnesium flow in a tropical forest. Ecological Modelling, 89(1–3):109–120, August 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095001255.

Timlin:1997:MSR

[TP97] Dennis J. Timlin and Yakov A. Pachepsky. A modular soil and root process simulator. Ecological Modelling, 94(1):67–80, January 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096019291.

Tsanis:1998:MPS

[TPS98] I. K. Tsanis, K. L. Prescott, and H. Shen. Modelling of phosphorus and suspended solids in cootes paradise marsh. Ecological Modelling, 114(1):1–17, December 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000830.

Trebitz:1991:TSL

[Tre91] Anett S. Trebitz. Timing of spawning in largemouth bass: implications of an individual-based model. Ecological Modelling, 59(3–4):203–227, December 15, 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380091901784.

Trost:1990:AFD

[Tro90] Norbert Trost. An approximate formula for the daily photoproduction of forest tree canopies. Ecological Modelling, 49(3–4):297–309, January 1990. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009090032C.

Tiwari:1996:ACR

[TRR96] Shrish Tiwari, Ramakrishna Ramaswamy, and J. Subba Rao. Adaptive control in a resource management model. Ecological Modelling, 84(1–3):53–62, January 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009400143X.

Thebault:1993:SMM

[TS93] Jean-Marc Thébault and Marie-José Salençon. Simulation model of a mesotrophic reservoir (Lac de Pareloup, France): biological model. Ecological Modelling, 65(1–2):1–30, January 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390124B.

Tan:1996:PGC

[TS96] Sen S. Tan and Fred E. Smeins. Predicting grassland community changes with an artificial neural network model. Ecological Modelling, 84(1–3):91–97, January 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094001316.

Tayasu:1996:PMC

[TSMC96] Ichiro Tayasu, Nanako Shigesada, Hiroshi Mukai, and Hal Caswell. Predator-mediated coexistence of epiphytic grass shrimps that compete for refuges. Ecological Modelling, 84 (1–3):1–10, January 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380094001308.

Torres-Sorando:1997:MST

[TSR97] Lourdes Torres-Sorando and Diego J. Rodríguez. Models of spatio-temporal dynamics in malaria ^{1}. Ecological Modelling, 104(2–3):231–240, December 8, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009700135X.

Tappeiner:1998:MVP

[TTT98] Ulrike Tappeiner, Erich Tasser, and Gottfried Tappeiner. Modelling vegetation patterns using natural and anthropogenic influence factors: preliminary experience with a GIS based model applied to an alpine area. Ecological Modelling, 113(1–3):225–237, November 2, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001458.

Tung:1990:EPV

[Tun90] Yeou-Koung Tung. Evaluating the probability of violating dissolved oxygen standard. Ecological Modelling, 51(3–4):193– 204, June 1990. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009090064N.

Tuzinkevich:1991:MST

[Tuz91] A. V. Tuzinkevich. Modelling of spatial-temporal dynamics of the bisexual population of Tribolium confusum. Ecological Modelling, 58(1–4):185–198, November 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009190035Y.

Tiktak:1995:RSF

[TvG95] Aaldrik Tiktak and Hans J. M. van Grinsven. Review of sixteen forest-soil-atmosphere models. Ecological Modelling, 83(1–2):35–53, December 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095000816.

Taskinen:1994:PUA

[TVS^{+}94] A. Taskinen, O. Varis, H. Sirviö, J. Mutanen, and P. Vakkilainen. Probabilistic uncertainty assessment of phosphorus balance calculations in a watershed. Ecological Modelling, 74(1–2):125–135, July 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380094901155.

Turner:1993:LSM

[TWRW93] Monica G. Turner, Yegang Wu, William H. Romme, and Linda L. Wallace. A landscape simulation model of winter foraging by large ungulates. Ecological Modelling, 69(3–4):163–184, October 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009390026O.

Taylor:1996:DMS

[TZE96] John A. Taylor, Patrick R. Zimmerman, and David J. Erickson. A 3-D modelling study of the sources and sinks of atmospheric carbon monoxide. Ecological Modelling, 88(1–3):53– 71, July 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095000690.

Ulanowicz:1997:ISE

[UAA97] Robert E. Ulanowicz and Luis G. Abarca-Arenas. An informational synthesis of ecosystem structure and function. Ecological Modelling, 95(1):1–10, February 1, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096000324.

Ulgiati:1997:DSD

[UB97] S. Ulgiati and C. Bianciardi. Describing states and dynamics in far from equilibrium systems. Needed a metric within a system state space. Ecological Modelling, 96(1–3):75–89, March 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096000609.

Ulgiati:1998:MPS

[UB98] Sergio Ulgiati and Mark T. Brown. Monitoring patterns of sustainability in natural and man-made ecosystems. Ecological Modelling, 108(1–3):23–36, May 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098000167.

Uchmanski:1999:WPP

[Uch99] Janusz Uchmański. What promotes persistence of a single population: an individual-based model. Ecological Modelling, 115(2–3):227–241, February 15, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098001793.

Uso-Domenech:1997:MSF

[UDML97] J. L. Usó-Domènech, J. Mateu, and J. A. Lopez. Mathematical and statistical formulation of an ecological model with applications. Ecological Modelling, 101(1):27–40, August 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097019467.

Uso-Domenech:1995:MMC

[UDVESM^{+}95] J. L. Usó-Domenech, Y. Villacampa-Esteve, G. Stübing-Martinez, T. Karjalainen, and M. P. Ramo. MARI-OLA: a model for calculating the response of Mediterranean bush ecosystem to climatic variations. Ecological Modelling, 80(2–3):113–129, July 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400052J.

Ulbrich:1999:ICS

[UH99] Karin Ulbrich and JohannesR. Henschel. Intraspecific competition in a social spider. Ecological Modelling, 115(2–3): 243–251, February 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438009800180X.

Ulanowicz:1995:USA

[Ula95] Robert E. Ulanowicz. Utricularia’s secret: the advantage of positive feedback in oligotrophic environments. Ecological Modelling, 79(1–3):49–57, May 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400032D.

Ulgiati:1995:FEM

[Ulg95] Sergio Ulgiati. Fundamentals of ecological modelling: Developments in Environmental Modelling, Vol. 19. S. E. Jørgensen. Elsevier, Amsterdam, 1994. 632 pp., HB ISBN 0-444-81572-4, Dfl. 410.00, PB ISBN 0-444-81578-3, Dfl. 240.00. Ecological Modelling, 78(3):285–286, April 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095900772.

Ulrich:1992:FET

[Ulr92] B. Ulrich. Forest ecosystem theory based on material balance. Ecological Modelling, 63(1–4):163–183, September 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290068P.

Umeki:1995:MRB

[Ume95] Kiyoshi Umeki. Modeling the relationship between the asymmetry in crown display and local environment. Ecological Modelling, 82(1):11–20, September 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400081R.

Ulbrich:1997:MEI

[UMJ^{+}97] K. Ulbrich, R. Marsula, F. Jeltsch, H. Hofmann, and C. Wissel. Modelling the ecological impact of contaminated river sediments on wetlands. Ecological Modelling, 94(2–3):221–230, January 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096000233.

Ulgiati:1994:EUE

[UOB94] S. Ulgiati, H. T. Odum, and S. Bastianoni. Emergy use, environmental loading and sustainability an emergy analysis of Italy. Ecological Modelling, 73(3–4):215–268, June 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094900647.

Ulvestad:1997:MHA

[US97] Kåre B. Ulvestad and Øivind Strand. Modelling hydrography and algal production in a poll in western Norway. Ecological Modelling, 101(2–3):285–301, August 15, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009700046X.

Usher:1991:EKE

[Ush91] Michael B. Usher. Ecological knowledge and environmental problem-solving: Concepts and case studies: Committee on the Applications of Ecological Theory to Environmental Problems, Commission on Life Sciences, National Research Council. National Academy Press, Washington, DC, 1986. Paperback, 388 pp. ISBN 0-309-03645-3. Ecological Modelling, 54(1–2):145– 146, May 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190108D.

Valentine:1999:ENP

[Val99] Harry T. Valentine. Estimation of the net primary productivity of even-aged stands with a carbon-allocation model. Ecological Modelling, 122(3):139–149, October 20, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099001349.

Vandermeer:1994:QBC

[Van94] John Vandermeer. The qualitative behavior of coupled predator–prey oscillations as deduced from simple circle maps. Ecological Modelling, 73(1–2):135–148, May 1994. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094901023.

Vandermeer:1996:DNC

[Van96] John Vandermeer. Disturbance and neutral competition theory in rain forest dynamics. Ecological Modelling, 85(2–3):99–111, March 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094001529.

Vandermeer:1997:PBS

[Van97] John Vandermeer. Period ‘bubbling’ in simple ecological models: Pattern and chaos formation in a quartic model. Ecological Modelling, 95(2–3):311–317, February 14, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096000464.

Vardavas:1993:SGR

[Var93] Ilias Mihail Vardavas. A simple groundwater rechargedepletion model for the tropical Magela Creek catchment. Ecological Modelling, 68(3–4):147–159, August 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009390014J.

Vasiliev:1994:BIE

[Vas94] Oleg F. Vasiliev. Biological indicators in environmental protection: Margit Kovács (Editor), Ellis Horwood and Académiai Kiadó, 1992. 207 pp. $67.50, ISBN 0-13-084989-8, hard cover. Ecological Modelling, 72(1–2):149–152, March 1994. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094901503.

Vose:1999:CMN

[VB99] James M. Vose and Paul V. Bolstad. Challenges to modelling NPP in diverse eastern deciduous forests: species-level comparisons of foliar respiration responses to temperature and nitrogen. Ecological Modelling, 122(3):165–174, October 20, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099001362.

Villeneuve:1990:PAG

[VBL90] Jean-Pierre Villeneuve, Olivier Banton, and Pierre Lafrance. A probabilistic approach for the groundwater vulnerability to contamination by pesticides: the VULPEST model. Ecological Modelling, 51(1–2):47–58, May 1990. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009090057N.

Vila:1996:MSI

[VCGG96] X. Vila, J. Colomer, and L. J. Garcia-Gil. Modelling spectral irradiance in freshwater in relation to phytoplankton and solar radiation. Ecological Modelling, 87(1–3):59–68, June 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094002053.

vanDam:1995:AMP

[vD95] D. van Dam. Application of the model NICCCE to the Solling spruce site. Ecological Modelling, 83(1–2):131–138, December 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009500092A.

vandenBerg:1998:PPP

[vdB98] Hugo Antontus van den Berg. Propagation of permanent perturbations in food chains and food webs. Ecological Modelling, 107(2–3):225–235, April 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097002202.

vandenBelt:1998:CBS

[vdBDJ98] Marjan van den Belt, Lisa Deutsch, and Åsa Jansson. A consensus-based simulation model for management in the Patagonia coastal zone. Ecological Modelling, 110(1):79–103, July 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438009800043X.

vanderHoeven:1996:MBS

[vdHvEH96] N. van der Hoeven, J. D. van Elsas, and C. E. Heijnen. A model based on soil structural aspects describing the fate of genetically modified bacteria in soil. Ecological Modelling, 89(1–3):161–173, August 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001298.

Verburg:1999:SEA

[VdKK^{+}99] P. H. Verburg, G. H. J. de Koning, K. Kok, A. Veldkamp, and J. Bouma. A spatial explicit allocation procedure for modelling the pattern of land use change based upon actual land use. Ecological Modelling, 116(1):45–61, March 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098001562.

vanderPeijl:1999:MCN

[vdPV99] M. J. van der Peijl and J. T. A. Verhoeven. A model of carbon, nitrogen and phosphorus dynamics and their interactions in river marginal wetlands. Ecological Modelling, 118(2–3):95–130, June 15, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099000149.

vanderSalm:1996:MTS

[vdSdVK96] C. van der Salm, W. de Vries, and J. Kros. Modelling trends in soil solution concentrations under five forest-soil combinations in the Netherlands. Ecological Modelling, 88(1–3):19– 37, July 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095000348.

Vereecken:1995:ESH

[VDV95] H. Vereecken, J. Diels, and P. Viaene. The effect of soil heterogeneity and hysteresis on solute transport: a numerical experiment. Ecological Modelling, 77(2–3):273–288, February 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400183I.

vanDam:1995:NMC

[vDvB95] D. van Dam and N. van Breemen. NICCE: a model for cycling of nitrogen and carbon isotopes in coniferous forest ecosystems. Ecological Modelling, 79(1–3):255–275, May 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400184J.

Valkov:1999:OGM

[VEA99] N. G. Valkov, J. Eitzinger, and V. Alexandrov. An organism growth model optimised for soybean under arbitrary function of the external conditions. Ecological Modelling, 123(2–3):65–72, November 15, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099000770.

Vereecken:1996:RUE

[Ver96] H. Vereecken. Reply to: Unbias the evaluation of soil hysteresis effects on transport, by R. Smith. Ecological Modelling, 87(1–3):295–296, June 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095002138.

Vardavas:1996:ELE

[VF96a] Ilias Mihail Vardavas and Antonis Fountoulakis. Estimation of lake evaporation from standard meteorological measurements: application to four Australian lakes in different climatic regions. Ecological Modelling, 84(1–3):139–150, January 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009400126X.

Veldkamp:1996:CCM

[VF96b] A. Veldkamp and L. O. Fresco. CLUE: a conceptual model to study the conversion of land use and its effects. Ecological Modelling, 85(2–3):253–270, March 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094001510.

Veldkamp:1996:CCI

[VF96c] A. Veldkamp and L. O. Fresco. CLUE-CR: an integrated multi-scale model to simulate land use change scenarios in Costa Rica. Ecological Modelling, 91(1–3):231–248, November 15, 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095001581.

Voinov:1998:SWF

[VFC98] Alexey A. Voinov, H. Carl Fitz, and Robert Costanza. Surface water flow in landscape models: 1. Everglades case study. Ecological Modelling, 108(1–3):131–144, May 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000246.

vanGrinsven:1995:A

[vG95] Hans van Grinsven. Acknowledgements. Ecological Modelling, 83(1–2):viii, December 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380095900373.

Varriale:1998:STS

[VG98] M. C. Varriale and A. A. Gomes. A study of a three species food chain. Ecological Modelling, 110(2):119–133, July 11, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098000623.

vanGrinsven:1995:WCF

[vGDT95] Hans J. M. van Grinsven, Charles T. Driscoll, and Aaldrik Tiktak. Workshop on Comparison of Forest-Soil-Atmosphere Models: Preface. Ecological Modelling, 83(1–2):1–6, December 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009500078A.

Vink:1997:SMC

[VGDvdZ97] Jos P. M. Vink, Bernhard Gottesbüren, Bernd Diekkrüger, and Sjoerd E. A. T. M. van der Zee. Simulation and model comparison of unsaturated movement of pesticides from a large clay lysimeter. Ecological Modelling, 105(1):113–127, December 14, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097001476.

vanHeerden:1995:SFS

[vHvGT95] Kees van Heerden, Hans J. M. van Grinsven, and Aaldrik Tiktak. Simulation of forest stress for the Solling spruce site with SOILVEG. Ecological Modelling, 83(1–2):185–195, December 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095983423.

vanHeerden:1995:ESF

[vHY95] Kees van Heerden and Ruth D. Yanai. Effects of stresses on forest growth in models applied to the Solling spruce site. Ecological Modelling, 83(1–2):273–282, December 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001055.

Vincent:1996:MME

[Vin96] Thomas L. Vincent. Modeling and managing the evolutionary component of biological systems. Ecological Modelling, 92(2–3):145–153, December 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001808.

Virtala:1992:OHP

[Vir92] Matti Virtala. Optimal harvesting of a plant-hervibore system: lichen and reindeer in northern Finland. Ecological Modelling, 60(3–4):233–255, April 1992. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009290035D.

Virtala:1996:HLR

[Vir96] Matti Virtala. Harvesting a lichen-reindeer system in an uncertain environment. Ecological Modelling, 89(1–3):209–224, August 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095001360.

VanWinkle:1998:IBM

[VJR^{+}98] W. Van Winkle, H. I. Jager, S. F. Railsback, B. D. Holcomb, T. K. Studley, and J. E. Baldrige. Individual-based model of sympatric populations of brown and rainbow trout for instream flow assessment: model description and calibration. Ecological Modelling, 110(2):175–207, July 11, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000659.

Varis:1997:JUM

[VK97] Olli Varis and Sakari Kuikka. Joint use of multiple environmental assessment models by a Bayesian meta-model: the Baltic salmon case. Ecological Modelling, 102(2–3):341–351, October 29, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097000689.

Varis:1999:LBD

[VK99] Olli Varis and Sakari Kuikka. Learning Bayesian decision analysis by doing: lessons from environmental and natural resources management. Ecological Modelling, 119(2–3):177–195, July 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099000617.

Virtanen:1994:PCT

[VKH^{+}94] Markku Virtanen, Jorma Koponen, Seppo Hellsten, Olli Nenonen, and Kari Kinnunen. Principles for calculation of transport and water quality in strongly regulated reservoirs. Ecological Modelling, 74(1–2):103–123, July 1994. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094901147.

Varis:1994:MWQ

[VKT94] Olli Varis, Sakari Kuikka, and Antti Taskinen. Modeling for water quality decisions: uncertainty and subjectivity in information, in objectives, and in model structure. Ecological Modelling, 74(1–2):91–101, July 1994. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380094901139.

VanMinnen:1995:APF

[VMB95] Jelle G. Van Minnen, Rob Meijers, and Leon C. Braat. Application of the FORSOL model to the spruce site at Solling, Germany. Ecological Modelling, 83(1–2):197–205, December 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009500098G.

Vasconcellos:1997:STM

[VMSP97] Marcelo Vasconcellos, Steven Mackinson, Katherine Sloman, and Daniel Pauly. The stability of trophic mass-balance models of marine ecosystems: a comparative analysis. Ecological Modelling, 100(1–3):125–134, December 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001506.

Vereecken:1996:SDD

[VNLB96] H. Vereecken, O. Neuendorf, G. Lindenmayr, and A. Basermann. A Schwarz domain decomposition method for solution of transient unsaturated water flow on parallel computers. Ecological Modelling, 93(1–3):275–289, December 16, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095002243.

vanOene:1995:AMP

[vOÅ95] Harmke van Oene and Göran I. Ågren. The application of the model NAP to the Solling spruce site. Ecological Modelling, 83(1–2):139–149, December 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009500093B.

Vold:1998:GOY

[Vol98] Arild Vold. A generalization of ordinary yield response functions. Ecological Modelling, 108(1–3):227–236, May 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098000313.

VanderMolen:1993:EMC

[VP93] Diederik T. Van der Molen and János Pintér. Environmental model calibration under different specifications: an application to the model SED. Ecological Modelling, 68(1–2):1– 19, July 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390104Z.

Vardavas:1997:WRD

[VPFM97] I. M. Vardavas, J. Papamastorakis, A. Fountoulakis, and M. Manousakis. Water resources in the desertificationthreatened Messara Valley of Crete: estimation of potential lake evaporation. Ecological Modelling, 102(2–3):363–374, October 29, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097000707.

Viney:1996:HRC

[VS96a] Neil R. Viney and Murugesu Sivapalan. The hydrological response of catchments to simulated changes in climate. Ecological Modelling, 86(2–3):189–193, May 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009500050X.

Voit:1996:MFGa

[VS96b] Eberhard O. Voit and Peter J. Sands. Modeling forest growth I. Canonical approach. Ecological Modelling, 86(1):51–71, April 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094001928.

Voit:1996:MFGb

[VS96c] Eberhard O. Voit and Peter J. Sands. Modeling forest growth II. Biomass partitioning in Scots pine. Ecological Modelling, 86(1):73–89, April 1996. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094001936.

Vanclay:1997:EFG

[VS97] J. K. Vanclay and J. P. Skovsgaard. Evaluating forest growth models. Ecological Modelling, 98(1):1–12, May 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096019321.

Volohonsky:1992:IWI

[VSG92] H. Volohonsky, G. Shaham, and M. Gophen. The impact of water inflow reduction on the trophic status of lakes. Ecological Modelling, 62(1–3):135–147, July 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290086T.

vanTongeren:1995:DAS

[vT95] Onno F. R. van Tongeren. Data analysis or simulation model: a critical evaluation of some methods. Ecological Modelling, 78(1–2):51–60, March 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009400117Z.

VanOrden:1993:DOD

[VU93a] George N. Van Orden and Christopher G. Uchrin. Dissolved oxygen dynamics in the Whippany river, New Jersey: deterministic/stochastic time-variable model. Ecological Modelling, 70(1–2):19–34, November 1993. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380093900709.

VanOrden:1993:SDO

[VU93b] George N. Van Orden and Christopher G. Uchrin. The study of dissolved oxygen dynamics in the Whippany river, New Jersey using the QUAL2E model. Ecological Modelling, 70(1–2):1–17, November 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380093900695.

Villacampa:1999:GRG

[VUDM^{+}99] Y. Villacampa, J. L. Usó-Domènech, J. Mateu, F. Vives, and P. Sastre. Generative and recognoscitive grammars of ecological models. Ecological Modelling, 117(2–3):315–332, May 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099000046.

Vasiliev:1991:GSA

[VV91] V. B. Vasiliev and V. A. Vavilin. Generalists or specialists in activated sludge microorganisms community. Ecological Modelling, 55(1–2):47–55, July 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380091900637.

Vasiliev:1992:SCA

[VV92] V. B. Vasiliev and V. A. Vavilin. Substrate consumption by an activated sludge with changing bacterial size and form. Ecological Modelling, 60(1):1–9, January 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380092900094.

Voinov:1999:SWF

[VVC99] Alexey A. Voinov, Helena Voinov, and Robert Costanza. Surface water flow in landscape models: 2. Patuxent watershed case study. Ecological Modelling, 119(2–3):211–230, July 15, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099000563.

Vanclooster:1995:DEA

[VVDF95] M. Vanclooster, P. Viaene, J. Diels, and J. Feyen. A deterministic evaluation analysis applied to an integrated soilcrop model. Ecological Modelling, 81(1–3):183–195, August 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400170M.

VanWensem:1997:CNF

[VVK97] Joke Van Wensem, Nico M. Van Straalen, and Sebastiaan A. L. M. Kooijman. Carbon and nitrogen fluxes in decomposing leaf litter with microbial-detritivore interactions: model simulations compared to microcosm ecotoxicity tests. Ecological Modelling, 96(1–3):175–189, March 1, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009600066X.

Venterink:1997:CSM

[VW97] Harry Olde Venterink and Martin J. Wassen. A comparison of six models predicting vegetation response to hydrological habitat change. Ecological Modelling, 101(2–3):347–361, August 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380097000628.

vanWijk:1999:WCF

[vWB99] M. T. van Wijk and W. Bouten. Water and carbon fluxes above European coniferous forests modelled with artificial neural networks. Ecological Modelling, 120(2–3):181–197, August 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001015.

Vila:1999:NNA

[VWN^{+}99] Jean-Pierre Vila, Vérène Wagner, Pascal Neveu, Marc Voltz, and Philippe Lagacherie. Neural network architecture selection: new Bayesian perspectives in predictive modelling: Application to a soil hydrology problem. Ecological Modelling, 120(2–3):119–130, August 17, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099000964.

Veldkamp:1996:MAT

[VZF96] A. Veldkamp, G. Zuidema, and L. O. Fresco. A model analysis of the terrestrial vegetation model of Image 2.0 for Costa Rica. Ecological Modelling, 93(1–3):263–273, December 16, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095002065.

Waelbroeck:1993:CSP

[Wae93] C. Waelbroeck. Climate-soil processes in the presence of permafrost: a systems modelling approach. Ecological Modelling, 69(3–4):185–225, October 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009390027P.

Wang:1995:MNT

[Wan95] Jihuai Wang. The models of niche and their application. Ecological Modelling, 80(2–3):279–291, July 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009400067R.

Washida:1995:ECC

[Was95] Toyoaki Washida. Ecosystem configurations consequent on the maximum respiration hypothesis. Ecological Modelling, 78 (3):173–193, April 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304380093E01029.

Wohlfahrt:1998:NSM

[WBH^{+}98] Georg Wohlfahrt, Michael Bahn, Ingrid Horak, Ulrike Tappeiner, and Alexander Cernusca. A nitrogen sensitive model of leaf carbon dioxide and water vapour gas exchange: application to 13 key species from differently managed mountain grassland ecosystems. Ecological Modelling, 113(1–3):179–199, November 2, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098001434.

With:1996:TAS

[WC96] Kimberly A. With and Thomas O. Crist. Translating across scales: Simulating species distributions as the aggregate response of individuals to heterogeneity. Ecological Modelling, 93(1–3):125–137, December 16, 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095002197.

Wilder:1995:MTD

[WCC95] J. W. Wilder, I. Christie, and J. J. Colbert. Modelling of two-dimensional spatial effects on the spread of forest pests and their management. Ecological Modelling, 82(3):287–298, November 1995. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009500097F.

Welsh:1996:MAR

[WCDL96] A. H. Welsh, R. B. Cunningham, C. F. Donnelly, and D. B. Lindenmayer. Modelling the abundance of rare species: statistical models for counts with extra zeros. Ecological Modelling, 88(1–3):297–308, July 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095001131.

Watson:1995:BIN

[WCS95] R. A. Watson, G. M. Carlos, and M. A. Samoilys. Bias introduced by the non-random movement of fish in visual transect surveys. Ecological Modelling, 77(2–3):205–214, February 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380093E0085H.

Wirtz:1996:EVE

[WE96] Kai-W. Wirtz and Bruno Eckhardt. Effective variables in ecosystem models with an application to phytoplankton succession. Ecological Modelling, 92(1):33–53, November 21, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095001964.

Webb:1995:HVP

[Web95] James N. Webb. Hamilton’s variational principle and ecological models. Ecological Modelling, 80(1):35–40, June 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009400046K.

Wenzel:1992:SSE

[Wen92] Volker Wenzel. Semantics and syntax elements of a unique calculus for modelling of complex ecological systems. Ecological Modelling, 63(1–4):113–131, September 1992. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009290065M.

Ward:1999:NDR

[WF99] Jeffrey S. Ward and Francis J. Ferrandino. New derivation reduces bias and increases power of Ripley’s L index. Ecological Modelling, 116(2–3):225–236, March 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001732.

Ware:1996:PCB

[WFB96] John R. Ware, Daphne Gail Fautin, and Robert W. Buddemeier. Patterns of coral bleaching: modeling the adaptive bleaching hypothesis. Ecological Modelling, 84(1–3):199–214, January 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094001324.

Weber:1996:SMA

[WFJ96] Gerhard E. Weber, Karin Furch, and Wolfgang J. Junk. A simple modelling approach towards hydrochemical seasonality of major cations in a Central Amazonian floodplain lake. Ecological Modelling, 91(1–3):39–56, November 15, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095001654.

Wilson:1991:CDM

[WFK^{+}91] James A. Wilson, John French, Peter Kleban, Susan R. McKay, and Ralph Townsend. Chaotic dynamics in a multiple species fishery: a model of community predation. Ecological Modelling, 58(1–4):303–322, November 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009190042Y.

Wallach:1991:LE

[WG91a] D. Wallach and B. Goffinet. Letter to the editor. Ecological Modelling, 54(1–2):137–139, May 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380091901049.

Wennekers:1991:SAS

[WG91b]  Thomas Wennekers and Christoph Giersch.  Sensitiv-

ity analysis of a simple model food chain. Ecological Modelling, 54(3–4):265–276, June 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009190079G.

Wallach:1998:EUI

[WG98] D. Wallach and M. Genard. Effect of uncertainty in input and parameter values on model prediction error. Ecological Modelling, 105(2–3):337–345, January 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001804.

Wiehn:1996:CMW

[WGP96] Georg Wiehn, Andrzej Górak, and Witold Pedrycz. Cost modelling of waste in incineration plants: an application of fuzzy sets toward decision making under uncertainty. Ecological Modelling, 85(1):83–91, February 1996. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0304380095000178.

Woolhouse:1991:PDA

[WH91] M. E. J. Woolhouse and R. Harmsen. Population dynamics of Aphis pomi: a transition matrix approach. Ecological Modelling, 55(1–2):103–111, July 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009190068C.

Wortmann:1997:MME

[WHA97] Joanne Wortmann, John W. Hearne, and Janine B. Adams. A mathematical model of an estuarine seagrass. Ecological Modelling, 98(2–3):137–149, May 30, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096019102.

Wortmann:1998:EEF

[WHA98] J. Wortmann, J. W. Hearne, and J. B. Adams. Evaluating the effects of freshwater inflow on the distribution of estuarine macrophytes. Ecological Modelling, 106(2–3):213–232, March 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438009700197X.

Whitmore:1995:MML

[Whi95] A. P. Whitmore. Modelling the mineralization and leaching of nitrogen from crop residues during three successive growing seasons. Ecological Modelling, 81(1–3):233–241, August 1995. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030438009400174G.

Whitehead:1998:MMS

[Whi98] Hal Whitehead. Male mating strategies: models of roving and residence. Ecological Modelling, 111(2–3):297–298, September 1, 1998. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098001100.

Whipple:1999:AES

[Whi99] Stuart J. Whipple. Analysis of ecosystem structure and function: extended path and flow analysis of a steady-state oyster reef model. Ecological Modelling, 114(2–3):251–274, January 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380098001549.

Wratt:1992:PSO

[WHJ^{+}92] D. S. Wratt, M. G. Hadfield, M. T. Jones, G. M. Johnson, and I. McBurney. Power stations, oxides of nitrogen emissions, and photochemical smog: a modelling approach to guide decision makers. Ecological Modelling, 64(2–3):185–203, October 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290114T.

White:1996:MAP

[WHN96] David H. White, S. Mark Howden, and Henry A. Nix. Modelling agricultural and pastoral systems under environmental change. Ecological Modelling, 86(2–3):213–217, May 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095000542.

Wilhoit:1991:MPD

[Wil91] L. R. Wilhoit. Modelling the population dynamics of different aphid genotypes in plant variety mixtures. Ecological Modelling, 55(3–4):257–283, August 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009190090N.

Williams:1996:AOR

[Wil96a] Byron K. Williams. Adaptive optimization of renewable natural resources: Solution algorithms and a computer program. Ecological Modelling, 93(1–3):101–111, December 16, 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095002170.

Williams:1996:TDM

[Wil96b] Mathew Williams. A three-dimensional model of forest development and competition. Ecological Modelling, 89(1–3):73–98, August 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095001239.

Wilder:1999:PMG

[Wil99] J. W. Wilder. A predictive model for gypsy moth population dynamics with model validation. Ecological Modelling, 116(2–3):165–181, March 15, 1999. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438009800163X.

Wissel:1992:ALE

[Wis92a] Christian Wissel. Aims and limits of ecological modelling exemplified by island theory. Ecological Modelling, 63(1–4):1–12, September 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290058M.

Wissel:1992:MMC

[Wis92b] Christian Wissel. Modelling the mosaic cycle of a Middle European beech forest. Ecological Modelling, 63(1–4):29–43, September 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290060R.

Wiegand:1999:APD

[WJW99] Kerstin Wiegand, Florian Jeltsch, and David Ward. Analysis of the population dynamics of Acacia trees in the Negev desert, Israel with a spatially-explicit computer simulation model. Ecological Modelling, 117(2–3):203–224, May 17, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380098001999.

Winkler:1997:CPS

[WK97a] Eckart Winkler and Stefan Klotz. Clonal plant species in a dry-grassland community: a simulation study of long-term population dynamics. Ecological Modelling, 96(1–3):125–141, March 1, 1997. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380096000579.

Wook:1997:MNP

[WK97b] Kim Jong Wook and Joon-Ho Kim. Modelling the net photosynthetic rate of Quercus mongolica stands affected by ambient ozone. Ecological Modelling, 97(3):167–177, April 28, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380096019011.

Walker:1996:MPS

[WL96] Brian H. Walker and Jenny L. Langridge. Modelling plant and soil water dynamics in semi-arid ecosystems with limited site data. Ecological Modelling, 87(1–3):153–167, June 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095000240.

Wu:1997:PBS

[WL97] Jianguo Wu and Simon A. Levin. A patch-based spatial modeling approach: conceptual framework and simulation scheme. Ecological Modelling, 101(2–3):325–346, August 15, 1997. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097000562.

Wang:1999:CDP

[WLW99] Gang Wang, Xue-Gong Liang, and Fei-Zhi Wang. The competitive dynamics of populations subject to an Allee effect. Ecological Modelling, 124(2–3):183–192, December 13, 1999. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438009900160X.

Wesselink:1995:MAS

[WM95] Lambert G. Wesselink and Jan Mulder. Modelling Alsolubility controls in an acid forest soil, Solling, Germany.

Ecological Modelling, 83(1–2):109–117, December 1995. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009500090I.

Wilhelm:1993:AGV

[WMRK93] W. W. Wilhelm, Gregory S. McMaster, R. W. Rickman, and Betty Klepper. Above-ground vegetative development and growth of winter wheat as influenced by nitrogen and water availability. Ecological Modelling, 68(3–4):183–203, August 1993. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009390016L.

Wolff:1994:IOM

[Wol94] W. F. Wolff. An individual-oriented model of a wading bird nesting colony. Ecological Modelling, 72(1–2):75–114, March 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094901465.

Woodwell:1998:SMI

[Woo98] John C. Woodwell. A simulation model to illustrate feedbacks among resource consumption, production, and factors of production in ecological-economic systems. Ecological Modelling, 112(2–3):227–248, October 15, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000805.

Watson:1997:PTP

[WQ97] R. A. Watson and T. J. Quinn II. Performance of transect and point count underwater visual census methods. Ecological Modelling, 104(1):103–112, December 1, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097001178.

Wu:1994:IRM

[WRHW94] H. Wu, E. J. Rykiel, T. Hatton, and J. Walker. An integrated rate methodology (IRM) for multi-factor growth rate modelling. Ecological Modelling, 73(1–2):97–116, May 1994. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380094901007.

Wilber:1997:ALA

[WS97] Patricia G. Wilber and Henry D. Shapiro. An artificial life approach to host-parasite interactions. Ecological Modelling, 101(1):113–122, August 1, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097019571.

Wu:1996:FSE

[WSGR96] Yegang Wu, Fred H. Sklar, Kishore Gopu, and Ken Rutchey. Fire simulations in the Everglades landscape using parallel programming. Ecological Modelling, 93(1–3):113–124, December 16, 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380095002189.

Wuttke:1991:SNT

[WTL91] G. Wuttke, B. Thober, and H. Lieth. Simulation of nitrate transport in groundwater with a three-dimensional groundwater model run as a subroutine in an agroecosystem model. Ecological Modelling, 57(3–4):263–276, October 15, 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009190116I.

Wu:1990:MEE

[Wu90] Jianguo Wu. Modelling the energy exchange processes between plant communities and environment. Ecological Modelling, 51(3–4):233–250, June 1990. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009090068R.

Wu:1995:DMM

[Wu95] Jianguo (Jingle) Wu. Dynamic modeling (Mac version): B. Hannon and M. Ruth. Springer-Verlag, New York, 1994. 248 pp., US$59.95, ISBN 0-387-94309-9 (Mac), ISBN 3-540-94287-4 (Windows), two 3 1/2in diskettes. Ecological Modelling, 82(2):208–210, October 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009590039X.

Weishampel:1996:CSE

[WU96] J. F. Weishampel and D. L. Urban. Coupling a spatiallyexplicit forest gap model with a 3-D solar routine to simulate latitudinal effects. Ecological Modelling, 86(1):101–111, April 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094002010.

Wu:1991:ABM

[WV91] Jianguo Wu and John L. Vankat. An area-based model of species richness dynamics of forest islands. Ecological Modelling, 58(1–4):249–271, November 1991. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380091900394.

Wu:1993:EPC

[WVB93] Jianguo Wu, John L. Vankat, and Yaman Barlas. Effects of patch connectivity and arrangement on animal metapopulation dynamics: a simulation study. Ecological Modelling, 65(3–4):221–254, February 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380093900813.

Wilder:1994:TVD

[WVCS94] J. W. Wilder, N. Voorhis, J. J. Colbert, and A. Sharov. A three variable differential equation model for gypsy moth population dynamics. Ecological Modelling, 72(3–4):229–250, April 1994. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009490085X.

West:1992:MAM

[WW92] P. W. West and K. F. Wells. Method of application of a model to predict the light environment of individual tree crowns and its use in a eucalypt forest. Ecological Modelling, 60(3–4):199– 231, April 1992. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009290034C.

Wyszomirski:1999:SMS

[WWJ99] Tomasz Wyszomirski, Izabela Wyszomirska, and Ingeborga Jarzyna. Simple mechanisms of size distribution dynamics in crowded and uncrowded virtual monocultures. Ecological Modelling, 115(2–3):253–273, February 15, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001823.

Xie:1999:WTD

[XHW^{+}99] Jialong Xie, Harry R. Hill, Scott R. Winterstein, Henry Campa, Robert V. Doepker, Timothy R. Van Deelen, and Jianguo Liu. White-tailed deer management options model (DeerMOM): design, quantification, and application. Ecological Modelling, 124(2–3):121–130, December 13, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438009900157X.

Xu:1999:EIA

[XJT99] Fu-Liu Xu, Sven Erik Jørgensen, and Shu Tao. Ecological indicators for assessing freshwater ecosystem health. Ecological Modelling, 116(1):77–106, March 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001604.

Xu:1999:MEE

[XJTL99] Fu-Liu Xu, Sven Erik Jørgensen, Shu Tao, and Ben-Gang Li. Modeling the effects of ecological engineering on ecosystem health of a shallow eutrophic Chinese lake (Lake Chao). Ecological Modelling, 117(2–3):239–260, May 17, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099000058.

Xu:1997:ESE

[Xu97] Fuliu Xu. Exergy and structural exergy as ecological indicators for the development state of the Lake Chaohu ecosystem. Ecological Modelling, 99(1):41–49, June 16, 1997. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380096019217.

Xu:1995:SMA

[XY95] Haigen Xu and Fengxiang Yang. Simulation model of activity of Phrynocephalus przewalskii. Ecological Modelling, 77(2–3):197–204, February 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380093E0092H.

Yang:1997:ASP

[YAKS97] Y. Yang, J. C. Allen, J. L. Knapp, and P. A. Stansly. An age-structured population model of citrus rust mite: a fruit–mite–fungal pathogen system. Ecological Modelling, 104(1):71–85, December 1, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097001191.

Yamamura:1997:OSL

[Yam97] Kohji Yamamura. Optimality in the spatial leaf distribution of the weed Portulaca oleracea L. Ecological Modelling, 104(2–3):133–143, December 8, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097001208.

Yamazaki:1993:NLM

[YH93] Hidekatsu Yamazaki and Loren R. Haury. A new Lagrangian model to study animal aggregation. Ecological Modelling, 69(1–2):99–111, September 1993. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/030438009390051S.

Yin:1996:MLD

[YH96] Chengjun Yin and Dehua Huang. A model of litter decomposition and accumulation in grassland ecosystems. Ecological Modelling, 84(1–3):75–80, January 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380094001332.

Yokozawa:1999:GVL

[YH99] M. Yokozawa and T. Hara. Global versus local coupling models and theoretical stability analysis of size-structure dynamics in plant populations. Ecological Modelling, 118(1):61–72, June 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099000496.

Yin:1996:RMG

[Yin96] Xiwei Yin. Reconstructing monthly global solar radiation from air temperature and precipitation records: a general algorithm for Canada. Ecological Modelling, 88(1–3):39–44, July 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304380095000356.

Yokozawa:1998:ECM

[YKH98] M. Yokozawa, Y. Kubota, and T. Hara. Effects of competition mode on spatial pattern dynamics in plant communities. Ecological Modelling, 106(1):1–16, February 16, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380097001816.

Yokozawa:1999:ECM

[YKH99] M. Yokozawa, Y. Kubota, and T. Hara. Effects of competition mode on the spatial pattern dynamics of wave regeneration in subalpine tree stands. Ecological Modelling, 118(1):73–86, June 1, 1999. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099000502.

Yamazaki:1995:SGA

[YO95] Hidekatsu Yamazaki and Akira Okubo. A simulation of grouping: an aggregating random walk. Ecological Modelling, 79(1–3):159–165, May 1995. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380093E0142P.

Yool:1999:CPR

[Yoo99] Andrew Yool. Comments on the paper: on repeated parameter sampling in Monte Carlo simulations. Ecological Modelling, 115(1):95–98, February 1, 1999. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098001653.

Yang:1998:TDM

[YS98] M. D. Yang and R. M. Sykes. Trophic-dynamic modeling in a shallow eutrophic river ecosystem. Ecological Modelling, 105(2–3):129–139, January 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097000768.

Yang:1998:ANM

[YWMM98] Y. Yang, L. T. Wilson, M. E. Makela, and M. A. Marchetti. Accuracy of numerical methods for solving the advection– diffusion equation as applied to spore and insect dispersal. Ecological Modelling, 109(1):1–24, June 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380097002226.

Zhu:1996:ASI

[ZBDN96] A-Xing Zhu, Lawrence E. Band, Barry Dutton, and Thomas J. Nimlos. Automated soil inference under fuzzy logic. Ecological Modelling, 90(2):123–145, October 1996. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0304380095001611.

Zeide:1991:QCE

[Zei91] Boris Zeide. Quality as a characteristic of ecological models. Ecological Modelling, 55(3–4):161–174, August 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/030438009190085F.

Zizhen:1997:NFM

[ZH97] Li Zizhen and Lin Hong. The niche-fitness model of crop population and its application. Ecological Modelling, 104(2–3):199–203, December 8, 1997. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380097001270.

Zizhen:1998:RRW

[ZH98] Li Zizhen and Lin Hong. Research on the regulation of water and fertilizers and a crop growth model of spring wheat in farmland of semi-arid regions. Ecological Modelling, 107(2–3):279–287, April 1, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098000040.

Ziv:1998:EHH

[Ziv98] Yaron Ziv. The effect of habitat heterogeneity on species diversity patterns: a community-level approach using an objectoriented landscape simulation model (SHALOM). Ecological Modelling, 111(2–3):135–170, September 1, 1998. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380098000969.

Zhou:1996:EEA

[ZMH96] Jizhong Zhou, Shijun Ma, and George W. Hinman. Ecological exergy analysis: a new method for ecological energetics research. Ecological Modelling, 84(1–3):291–303, January 1996. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0304380094001359.

Zonneveld:1998:CBM

[Zon98a] C. Zonneveld. A cell-based model for the chlorophyll a to carbon ratio in phytoplankton. Ecological Modelling, 113(1–3):55–70, November 2, 1998. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380098001343.

Zonneveld:1998:LLM

[Zon98b] Cor Zonneveld. Light-limited microalgal growth: a comparison of modelling approaches. Ecological Modelling, 113 (1–3):41–54, November 2, 1998. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380098001331.

Zucchetto:1991:EEI

[Zuc91] James J. Zucchetto. Ecological engineering: an introduction to ecotechnology: W. J. Mitsch and S. E. Jørgensen (Editors). Wiley, Chichester, UK, 1989. 472 pp., £43.15. ISBN 0-471-62559-0. Ecological Modelling, 59(1–2):148–150, December 1, 1991. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/030438009190136O.